### v1

# ADer: A Comprehensive Benchmark for Multi-class Visual Anomaly Detection

Abstract
Visual anomaly detection aims to identify anomalous regions in images through unsupervised learning paradigms, with increasing application demand and value in fields such as industrial inspection and medical lesion detection. Despite significant progress in recent years, there is a lack of comprehensive benchmarks to adequately evaluate the performance of various mainstream methods across different datasets under the practical multi-class setting. The absence of standardized experimental setups can lead to potential biases in training epochs, resolution, and metric results, resulting in erroneous conclusions. This paper addresses this issue by proposing a comprehensive visual anomaly detection benchmark, ADer, which is a modular framework that is highly extensible for new methods. The benchmark includes multiple datasets from industrial and medical domains, implementing fifteen state-of-the-art methods and nine comprehensive metrics. Additionally, we have open-sourced the GPU-assisted ADEval package to address the slow evaluation problem of metrics like time-consuming mAU-PRO on large-scale data, significantly reducing evaluation time by more than 1000-fold. Through extensive experimental results, we objectively reveal the strengths and weaknesses of different methods and provide insights into the challenges and future directions of multi-class visual anomaly detection. We hope that ADer will become a valuable resource for researchers and practitioners in the field, promoting the development of more robust and generalizable anomaly detection systems. Full codes have been attached in Appendix and open-sourced at https://github.com/zhangzjn/ader.
1 Introduction
In recent years, with the rapid advancement in model iteration and computational power, Visual Anomaly Detection (VAD) has made significant progress across various fields [25, 7]. It has become a crucial component in key tasks such as industrial quality inspection and medical lesion detection. Due to its unsupervised experimental setup, VAD demonstrates immense application value in real-world scenarios where the yield rate is high, defect samples are difficult to obtain, and potential defect patterns are diverse. However, the field faces challenges such as small dataset sizes and insufficient evaluation metrics, resulting in potentially unfair comparison outcomes due to differing training recipes among methods. Moreover, most methods have not been compared on the latest large-scale datasets (e.g., Real-IAD  [32] and COCO-AD [42]) and new evaluation metrics (e.g., mAD [40] and mIoU-max [42]). The fundamental issue lies in the absence of standardized training strategies, akin to those in object detection, to evaluate different algorithms. Factors such as training epoch and resolution can potentially affect evaluation results, leading to erroneous conclusions.
To address this pressing issue, we believe that establishing a comprehensive and fair benchmark is crucial for the sustained and healthy development of this field. Therefore, we have constructed an integrated ADer benchmark, utilizing a unified evaluation interface under the more practical multi-class setting. This benchmark is designed as a highly extensible modular framework (see Sec. 3), allowing for the easy implementation of new methods. Specifically, the framework integrates multiple datasets from industrial, medical, and general-purpose domains (see Sec. 3.2), and implements fifteen state-of-the-art methods (including augmentation-based, embedding-based, reconstruction-based, and hybrid methods, see Sec. 3.1) and nine comprehensive evaluation metrics (see Sec. 3.3), ensuring thorough and unbiased performance evaluation for each method. Additionally, to address the efficiency issue of evaluating time-consuming metrics like mAU-PRO on large-scale data, we have developed and open-sourced the GPU-assisted ADEval package (see Sec. 3.5), significantly reducing evaluation time by over 1000 times, making previously impractical extensive detailed evaluations feasible on large-scale datasets.
Through extensive and fair experiments, we objectively reveal the strengths and weaknesses of different visual anomaly detection methods, comparing their efficiency (i.e., model parameter count and FLOPs) and training resource consumption across different datasets, as shown in Fig. 1. Detailed results and analyses (see Sec. 4) elucidate the challenges of multi-class visual anomaly detection and provide valuable insights for future research directions.
In summary, the contributions of this paper are as follows: 1) Comprehensive benchmark: We introduce a modular and extensible benchmark termed ADer for visual anomaly detection, which evaluates various representative methods on multiple mainstream datasets. 2) Implementation of state-of-the-Art methods: We implement fifteen state-of-the-art anomaly detection methods and nine comprehensive evaluation metrics within the ADer framework. 3) GPU-assisted evaluation package: We develop and open-source the ADEval package, significantly reducing the evaluation time of complex metrics by over 1000 times. 4) Extensive experimental analysis: We conduct extensive experiments to objectively evaluate the performance of different methods, providing insights into their strengths, weaknesses, and potential areas for improvement. 5) Open-source resources: We open-source the complete ADer code, making it a valuable resource for the research community and promoting further advancements in the field.
Figure 1: Intuitive benchmarked results comparison on MVTec AD [3] (Left), VisA [47] (Middle), and Real-IAD [32] (Right) datasets among mainstream methods. For each dataset, the horizontal axis represents the training time for different methods, the left vertical axis represents mAD [40] (marked as circles, with radius indicating model parameter count), and the right vertical axis represents mIoU-max [42] (marked as squares, with side length indicating model FLOPs).
2 Background and Related Work
2.1 Problem Definition and Objective
Visual anomaly detection (VAD) is a critical task in computer vision, aiming at identifying patterns or instances in visual data that deviate significantly from the norm. These anomalies can manifest as industrial defects, medical lesion, or rare objects that are not typically present in the training data. The primary objective of VAD is to develop algorithms capable of discerning these irregularities with high accuracy and reliability. This task is particularly challenging due to the inherent variability and complexity of visual data, the scarcity of anomalous examples, and the need for robust generalization across diverse scenarios. In a formal context, multi-class VAD can be defined as follows: Given a training dataset $D_{train}=\{x_{1},x_{2},...,x_{n}\}$ with $C$ categories and each visual image $x_{i}$ belonging to a specific category, the goal is to learn a unified AD model $M$ that can predict an anomaly score $s_{i}=M(x_{i})$ for each image. This score reflects the likelihood of each pixel in $x_{i}$ being an anomaly. The model $M$ is typically trained on $D_{train}$ that predominantly contains normal instances, with the assumption that anomalies are rare and not well-represented in the training set. Inevitably, there are some mislabeled or inaccurately labeled noisy samples, which constitute inherent biases within the dataset. These are typically disregarded under standard settings. The performance of the model is then evaluated based on its ability to correctly identify anomalous images and their defect regions in a unified testset that contains normal and anomalous images.
2.2 Challenges in Multi-class VAD
The complexity of VAD arises from several factors: 1) Data Imbalance: Anomalies are rare, leading to highly imbalanced datasets where normal instances and region areas vastly outnumber anomalous ones. 2) Variability of Anomalies: Anomalies can vary widely in appearance, making it difficult to capture all possible variations during training. 3) Context Sensitivity: The definition of what constitutes an anomaly can be context-dependent, requiring models to understand the broader context in which the visual data is situated. 4) Efficiency Requirements: Many applications of VAD require real-time processing and limited GPU memory, necessitating efficient and scalable algorithms. 5) Comprehensive and Fair Evaluation: Current methods exhibit significant differences in training hyperparameters and insufficient evaluation of performance metrics, so it is necessary to benchmark them using fair and standardized criteria. In this benchmark study, we systematically evaluate a range of state-of-the-art VAD methods (Sec. 3.1) across multiple datasets (Sec. 3.2) and comprehensive metrics (Sec. 3.3). Our goal is to provide a comprehensive assessment of current capabilities, identify key challenges, and suggest directions for future research in visual anomaly detection.
Figure 2: A comparative diagram of different frameworks for the benchmarked methods in Sec. 3.1.
2.3 Visual Anomaly Detection
Visual anomaly detection methods can generally be categorized into three types: 1) Augmentation-based methods generate pseudo-supervised information for anomalies by creating abnormal regions [23, 37], constructing anomalous data [38, 19], or adding feature perturbations [27, 31]. This enables the model to learn the differences between normal and abnormal distributions. 2) Embedding-based methods leverage pretrained models to extract powerful feature representations and judge anomalies in high-dimensional space. Typical approaches include distribution-map-based methods [14, 22], teacher-student frameworks [4, 33], and memory-bank techniques [29, 13]. 3) Reconstruction-based methods use encoder-decoder architectures to locate anomalies by analyzing the reconstruction error. They typically include both image-level [2, 24, 16] and feature-level approaches [10, 40, 42]. There are also some hybrid methods [35, 31, 44] that attempt to integrate multiple techniques to further enhance model performance.
Basic network structures of VAD. Early visual anomaly detection methods typically employ UNet-based autoencoder architectures [1, 24, 37]. With advancements in foundational visual model structures [17, 34, 26, 39, 43, 41] and pretraining techniques [18, 8], more recent methods often utilize models pretrained on ImageNet-1K [11] as feature extractors, such as the ResNet [17] series, Wide ResNet-50 [36], and EfficientNet-b4 [30]. Recently, benefiting from the dynamic modeling capabilities of Vision Transformers (ViT) [12], some studies [9, 35, 40] have attempted to incorporate this architecture into the design of anomaly detection models.
3 Methodology: ADer Benchmark
Table 1: Attribute comparison for mainstream representative methods. Notations: Augmentation-based (Aug.), Embedding-based (Emb.), Reconstruction-based (Rec.), Parameters (Params), Memory (Mem.), Batch Size (BS), Optimizer (Optim.), Time (T.), ResNet (RN), Wide-ResNet (WRN), EfficientNet (EN), hours (h), minutes (m), seconds (s), unavailable (-), out-of-memory (OOM). Train and test time are evaluated under the standard setting described in Sec. 4.1 in one L40S GPU. Memory is tested under the standard setting with a batch size of 8, and the results for different methods are presented in Sec. 4.2. Bold, underline, and wavy-line represent the best, second-best, and third-best results, respectively.
Method Hyper Params. Efficiency Train Mem. (M) Test Mem. (M) MVTec AD VisA Real-IAD
BS Optim. LR Params. FLOPs Backbone Train T. Test T. Train T. Test T. Train T. Test T.
Aug. DRAEM [37] 8 Adam 1e-4 97.4M 198G UNet 12,602 2,858 3.4h 35s 8.0h 36s 39.2 18m41s
SimpleNet [27] 32 AdamW 1e-4 72.8M 17.715G WRN50 2,266 4,946 1.7h 5m50s 3.9h 7m21s 15.9h 4h51m
RealNet [45] 16 Adam 1e-4 591M 115G WRN50 14,004 3,794 2.6h 41s 5.4h 41s - -
Emb. CFA [21] 4 AdamW 1e-3 38.6M 55.3G WRN50 4,364 2,826 1.2h 18s 2.7h 17s 10.9h 14m20s
PatchCore [29] 8 - - - - WRN50 - - - 9h22m - OOM - OOM
CFLOW-AD [14] 32 Adam 2e-4 237M 28.7G WRN50 3,048 1,892 5.2h 56s 10.4h 1m15s 40.9h 22m49s
PyramidalFlow [22] 2 Adam 2e-4 34.3M 962G RN18 3,904 2,836 7.9h 1m30s 10.5h 2m43s 45h 38m15s
Rec. RD [10] 16 Adam 5e-3 80.6M 28.4G WRN50 3,286 1,464 0.8h 13s 1.1h 18s 4.1h 7m48s
DiAD [16] 12 Adam 1e-5 1331M 451.5G RN50 26,146 20,306 9.1h 16m 21.6h 19m 90h 16h20m
ViTAD [40] 8 AdamW 1e-4 39.0M 9.7G ViT-S 1,470 800 0.8h 15s 1.1h 15s 5.2h 10m2s
InvAD [42] 32 Adam 1e-3 95.6M 45.4G WRN50 5,920 3,398 1.0h 31s 2.3h 33s 9.2h 21m
InvAD-lite [42] 32 Adam 1e-3 17.1M 9.3G RN34 1,846 1,100 0.8h 20s 1.1h 31s 3.4h 9m27s
MambaAD [15] 16 AdamW 5e-3 25.7M 8.3G RN34 6,542 1,484 2.4h 34s 5.6h 23s 23.6h 24m6s
Hybrid UniAD [35] 8 AdamW 1e-4 24.5M 3.4G EN-b4 1,856 968 0.8h 22s 1.0h 18s 4.1h 7m2s
RD++ [31] 16 Adam 1e-3 96.1M 37.5G WRN50 4,772 1,480 3.3h 28s 7.8h 33s 42.4h 15m17s
DesTSeg [44] 32 SGD 0.4 35.2M 30.7G RN18 3,446 1,240 1.0h 19s 1.7h 16s 6.8h 8m13s
3.1 Supported VAD Methods
Following the categories of current VAD methods in Sec. 2.3, we choose representative models for each category. The selection criteria are based on the method’s popularity, effectiveness, and ease of use. 1) For Augmentation-based methods, we choose DRAEM [37], SimpleNet [27], and RealNet [45]. 2) For Embedding-based methods, we select CFA [21], PatchCore [29], CFLOW-AD [14], and PyramidalFlow [22]. 3) For Reconstruction-based methods, we include RD [10], DiAD [16], ViTAD [40], InvAD [42], InvAD-lite [42], and MambaAD [15]. Additionally, UniAD [35], RD++ [31], and DesTSeg [44] are categorized as hybrid methods due to their use of multiple techniques. Fig. 2 presents schematic diagrams and comparisons of the frameworks for each method belonging to different types, facilitating a better understanding of the differences among these methods. Tab. 1 provides a direct comparison of the hyperparameters, efficiency, and training time on three mainstream datasets for different methods, using one L40S GPU. Note that different methods may yield varying results when tested on different hardware, but the overall relative trends remain largely unchanged.
3.2 VAD Datasets
Table 2: Comparison of representative VAD datasets, i.e., industrial, medical, and general-purpose fields, respectively. Large-scale Real-IAD and COCO-AD only employ the 100 epoch setting.
Dataset Category Number Image Quantity Epoch Setting in ADer
Train Test Train Test
Normal Anomaly Normal
MVTec AD [3] 15 15 03,629 01,258 00467 100 300
MVTec AD 3D [6] 10 10 02,950 00249 00948 100 300
MVTec LOCO-AD [5] 05 05 01,772 00993 00575 100 300
VisA [47] 12 12 08,659 00962 01,200 100 300
BTAD [28] 03 03 01,799 00580 00451 100 300
MPDD [20] 06 06 00888 00282 00176 100 300
MAD_Real [46] 10 10 00490 00221 00050 100 300
MAD_Sim [46] 20 20 04,200 04,951 00638 100 300
Real-IAD [32] 30 30 36,465 51,329 63,256 100 -
Uni-Medical [40] 03 03 13,339 04,499 02,514 100 300
COCO-AD [42] 61 81 30,438 01,291 03,661 100 -
65,133 02,785 02,167 100 -
79,083 03,328 01,624 100 -
77,580 03,253 01,699 100 -
To comprehensively evaluate the effectiveness, stability, and generalization of different methods, we benchmark extensive and fair experiments on three types of datasets: 1) Real and synthetic industrial anomaly detection (AD) datasets, i.e., MVTec AD [3], MVTec AD 3D [6], MVTec LOCO-AD [5], VisA [47], BTAD [28], MPDD [20], MAD_Real [46], MAD_Sim [46], and Real-IAD [32]. 2) The medical Uni-Medical [40] dataset. 3) The general-purpose COCO-AD [42] dataset. Detailed descriptions of the datasets are provided in Tab. 2, including the categories and scales of the datasets. Note that COCO-AD is inherently a multi-class dataset with four splits, and the average is taken when evaluating the comprehensive results.
3.3 Evaluation Metrics
[…]
Figure 3: Core sub-modules of the training framework in ADer. The blue area represents standard components, while the red area indicates that a new method requires only three corresponding files.
3.4 Simplify Implementation by Structured ADer Codebase
To ensure fair comparison among different methods, we construct a standardized ADer framework. As shown in Fig. 3, it includes shared foundational training/testing components and implements various metric calculations (compatible with our ADEval). The standardized dataset allows for easy comparison, eliminating potential unfair settings from different codebases. Additionally, ADer is highly extensible for new methods, requiring only compliant model, trainer, and configuration files.
Feature comparison with anomalib. Our ADer framework supports more recent and powerful representative methods, datasets, and comprehensive metrics. The code organization is straightforward and clear, making it easy for researches on new methods and practical deployments.
3.5 ADEval: Fast and Memory-efficient Routines for mAU-ROC/mAP/mAU-PRO
The speed of metric evaluation is crucial for the iterative process of model algorithms. When the number of test images increases and the resolution becomes higher, the pixel-level evaluation algorithms implemented naively using sklearn and skimage packages become time-consuming. This is particularly evident with large-scale datasets such as Real-IAD [32] and COCO-AD [42], where evaluation times can exceed one hour. To address this issue, we have released the GPU-assisted ADEval library, which employs an iterative-accumulating algorithm with optional CUDA acceleration. For instance, in the case of the most time-consuming mAU-PRO metric on the multi-class MVTec AD dataset, as demonstrated by UniAD [35], the naive implementation requires 242.7 seconds, whereas the optimized version reduces this to less than 0.1 second, achieving a more than 1000-fold speedup.
Table 3: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3 under 100/300 epochs. Bold, underline, and wavy-line represent the best, second-best, and third-best results, respectively. Patchcore requires no training that shares results under different epoch settings.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Aug. DRAEM [37] 54.5/55.2 76.3/77.0 83.6/83.9 47.6/48.7 03.2/03.1 06.7/06.3 14.3/15.8 03.5/03.3 44.7/45.3
SimpleNet [27] 95.4/79.2 98.3/90.8 95.7/87.6 96.8/82.4 48.8/24.0 51.9/29.0 86.9/62.0 36.4/17.8 83.8/67.6
RealNet [45] 84.8/82.9 94.1/93.3 90.9/90.9 72.6/69.8 48.2/50.0 41.4/40.4 56.8/51.2 28.8/28.5 72.3/70.9
Emb. CFA [21] 57.6/55.8 78.3/78.8 84.7/84.5 54.8/43.9 11.9/04.8 14.7/08.9 25.3/19.3 08.9/04.7 50.1/46.1
PatchCore [29] 98.8/0-00 99.5/0-00 98.4/0-00 98.3/0-00 59.9/0-00 61.0/0-00 94.2/0-00 44.9/0-00 88.6/0-00
CFLOW-AD [14] 91.6/92.7 96.7/97.2 93.4/94.0 95.7/95.8 45.9/46.8 48.6/49.6 88.3/89.0 33.2/34.0 81.8/82.5
PyramidalFlow [22] 70.2/66.2 85.5/84.3 85.5/85.1 80.0/74.2 22.3/17.2 22.0/19.6 47.5/40.0 12.8/11.4 61.7/58.1
Rec. RD [10] 93.6/90.5 97.2/95.0 95.6/95.1 95.8/95.9 48.2/47.1 53.6/52.1 91.2/91.2 37.0/35.8 83.8/82.6
DiAD [16] 88.9/92.0 95.8/96.8 93.5/94.4 89.3/89.3 27.0/27.3 32.5/32.7 63.9/64.4 21.1/21.3 70.1/71.0
ViTAD [40] 98.3/98.4 99.3/99.4 97.3/97.5 97.6/97.5 55.2/55.2 58.4/58.1 92.0/91.7 42.3/42.0 87.1/87.0
InvAD [42] 98.1/98.9 99.0/99.6 97.6/98.2 98.0/98.1 56.3/57.1 59.2/59.6 94.4/94.4 42.8/43.1 87.6/88.1
InvAD-lite [42] 97.9/98.1 99.2/99.1 96.8/96.8 97.3/97.3 54.4/55.0 57.8/58.1 93.3/93.1 41.4/41.7 86.8/86.9
MambaAD [15] 97.8/98.5 99.3/99.5 97.3/97.7 97.4/97.6 55.1/56.1 57.6/58.7 93.4/93.6 41.2/42.3 87.0/87.5
Hybrid UniAD [35] 92.5/96.8 97.3/98.9 95.4/97.0 95.8/96.8 42.7/45.0 48.0/50.2 89.3/91.0 32.5/34.2 82.0/84.2
RD++ [31] 97.9/95.8 98.8/98.0 96.4/96.6 97.3/97.3 54.7/53.0 58.0/57.0 93.2/92.9 41.5/40.5 86.8/85.9
DesTSeg [44] 96.4/96.3 98.6/98.8 96.2/96.1 92.0/92.6 71.1/75.8 68.2/71.3 83.4/82.6 52.8/56.6 87.9/88.8
Table 4: Benchmarked results on VisA dataset [47] by the suggested metrics under 100/300 epochs.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Aug. DRAEM [37] 55.1/56.2 62.4/64.6 72.9/74.9 37.5/45.0 00.6/00.7 01.7/01.8 10.0/16.0 00.9/00.9 38.0/40.6
SimpleNet [27] 86.4/80.7 89.1/83.8 82.8/79.3 96.6/94.4 34.0/29.2 37.8/33.1 79.2/74.2 25.7/22.1 74.0/69.5
RealNet [45] 71.4/79.2 79.5/84.8 74.7/78.3 61.0/65.4 25.7/29.2 22.6/27.9 27.4/33.9 13.5/17.4 54.7/59.9
Emb. CFA [21] 66.3/67.1 74.3/73.8 74.2/75.3 81.3/83.0 22.1/13.7 26.2/18.7 50.8/48.7 17.0/11.3 58.4/56.5
CFLOW-AD [14] 86.5/87.2 88.8/89.5 84.9/85.1 97.7/97.8 33.9/34.2 37.2/37.2 86.8/87.3 24.9/24.9 75.3/75.7
PyramidalFlow [22] 58.2/69.0 66.3/72.9 74.4/75.8 77.0/79.1 07.2/07.9 09.6/08.7 42.8/52.6 05.6/04.7 50.2/54.8
Rec. RD [10] 90.6/93.9 90.9/94.8 89.3/90.4 98.0/98.1 35.4/38.4 42.5/43.7 91.9/91.9 27.9/29.0 78.6/80.5
DiAD [16] 84.8/90.5 88.5/91.4 86.9/90.4 82.5/83.4 17.9/19.2 23.2/25.0 44.5/44.3 14.9/16.2 61.2/63.5
ViTAD [40] 90.4/90.3 91.1/91.2 86.0/86.4 98.2/98.2 36.4/36.4 41.0/40.9 85.7/85.8 27.5/27.5 77.2/77.3
InvAD [42] 95.4/95.6 95.7/96.0 91.6/92.3 98.9/99.0 43.3/43.7 46.8/46.9 93.1/93.0 32.5/32.6 82.4/82.6
InvAD-lite [42] 94.9/95.3 95.2/95.8 90.7/91.0 98.6/98.7 40.2/41.2 44.0/44.9 93.1/93.2 29.8/30.6 81.3/81.8
MambaAD [15] 94.5/93.6 94.9/93.9 90.2/89.8 98.4/98.2 39.3/34.0 43.7/39.3 92.1/90.5 29.5/25.9 80.8/79.0
Hybrid UniAD [35] 89.0/91.4 91.0/93.3 85.8/87.5 98.3/98.5 34.5/35.3 39.6/40.2 86.5/89.0 26.4/26.5 76.7/78.2
RD++ [31] 93.9/93.1 94.7/94.1 90.2/90.0 98.4/98.4 42.3/40.4 46.3/44.8 91.9/91.4 31.2/29.9 81.3/80.6
DesTSeg [44] 89.9/89.0 91.4/90.3 86.7/85.9 86.7/84.8 46.6/43.3 47.2/44.4 61.1/57.5 32.7/30.1 74.9/73.0
4 Results and Analysis
4.1 Experimental Setup
Different methods potentially introduce various factors that can impact model performance. To ensure a fair and comprehensive evaluation of the effectiveness and convergence of different methods, we fix the most influential parameters, i.e., resolution (256$\times$256) and training epochs (100 and 300). The reason lies in the fact that tasks such as classification, detection, and segmentation typically set specific resolutions and standard training epochs. We observe that for most methods, 100 epochs generally suffice to reach saturation [40, 42], with only a few methods [35] requiring more epochs for training. Therefore, we also establish a setting with 300 epochs. Meanwhile, we maintain consistency with the original papers for batch size, optimizer, learning rate, and data augmentation. We report the evaluation results corresponding to the final epoch at the end of training to ensure fairness. All experiments are conducted on one L40S GPU.
Table 5: Benchmarked results on Real-IAD dataset [32] by the suggested metrics under 100 epoch.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Aug. DRAEM [37] 50.9 45.9 61.3 44.0 00.2 00.4 13.6 00.2 33.6
SimpleNet [27] 54.9 50.6 61.5 76.1 01.9 04.9 42.4 02.5 43.5
Emb. CFA [21] 55.7 50.5 61.9 81.3 01.6 03.8 48.8 02.0 45.0
CFLOW-AD [14] 77.0 75.8 69.9 94.8 17.6 21.7 80.4 12.4 63.9
PyramidalFlow [22] 54.4 48.0 62.0 71.1 01.2 01.1 34.9 00.5 40.9
Rec. RD [10] 82.7 79.3 74.1 97.2 25.2 32.8 90.0 20.0 70.0
DiAD [16] 75.6 66.4 69.9 88.0 02.9 07.1 58.1 03.7 52.6
ViTAD [40] 82.7 80.2 73.7 97.2 24.3 32.3 84.8 19.6 69.3
InvAD [42] 89.4 87.0 80.2 98.4 32.6 38.9 92.7 24.6 75.6
InvAD-lite [42] 87.2 85.2 77.8 98.0 31.7 37.9 92.0 23.8 74.2
MambaAD [15] 87.0 85.3 77.6 98.6 32.4 38.1 91.2 23.9 74.2
Hybrid UniAD [35] 83.1 81.2 74.5 97.4 23.3 30.9 87.1 18.6 69.6
RD++ [31] 83.6 80.6 74.8 97.7 25.9 33.6 90.7 20.5 70.8
DesTSeg [44] 79.3 76.7 70.7 80.3 36.9 40.3 56.1 26.2 64.5
Table 6: Benchmarked results on all other datasets by the mIoU/mAD metrics under 100 epoch.
Method MVTec 3D MVTec LOCO BTAD MPDD MAD_Real MAD_Sim Uni-Medical COCO-AD
Aug. DRAEM [37] 01.0/46.6 05.6/41.8 03.4/47.4 02.5/35.4 00.8/48.9 00.7/48.0 03.0/37.0 08.0/38.3
SimpleNet [27] 13.9/70.1 21.2/67.4 28.6/78.8 24.5/76.2 06.3/56.7 04.2/60.9 23.3/68.9 11.5/42.9
RealNet [45] 0-00/0-00 0-00/0-00 36.6/76.1 28.2/72.3 0-00/0-00 0-00/0-00 0-00/0-00 0-00/0-00
Emb. CFA [21] 09.3/58.6 09.1/52.2 33.6/80.2 16.6/65.1 08.7/58.7 04.6/58.2 14.7/57.2 08.9/41.0
PatchCore [29] 24.5/78.4 20.4/67.7 38.0/83.3 35.0/83.5 16.6/68.2 0-00/0-00 0-00/0-00 0-00/0-00
CFLOW-AD [14] 15.8/71.6 17.3/64.0 33.8/78.9 20.1/69.7 08.8/63.4 02.7/60.0 17.7/69.9 16.0/53.0
PyramidalFlow [22] 06.4/61.9 08.0/47.5 18.3/68.3 10.4/64.6 05.1/59.0 02.5/57.4 09.4/47.6 08.0/38.4
Rec. RD [10] 22.2/75.3 15.8/63.0 42.1/84.7 31.4/80.6 07.2/60.6 04.5/62.6 26.9/71.1 11.5/45.8
DiAD [16] 05.4/62.8 14.9/56.5 15.7/68.5 08.2/58.1 03.6/55.8 04.2/58.9 23.2/69.1 11.6/44.1
ViTAD [40] 20.4/75.4 19.8/64.8 40.1/83.1 27.7/77.3 05.0/57.4 05.0/62.8 33.7/75.5 20.1/53.5
InvAD [42] 27.4/80.3 23.1/69.2 44.3/85.9 34.0/81.9 16.8/67.6 08.6/67.4 32.6/75.4 14.3/50.5
InvAD-lite [42] 26.9/79.6 20.6/67.1 42.6/84.4 30.9/79.8 08.1/62.5 06.0/64.9 26.5/71.6 13.7/49.2
MambaAD [15] 25.9/79.0 20.6/66.4 39.0/82.5 26.8/78.1 07.2/60.5 05.0/63.6 33.5/75.9 12.9/48.8
Hybrid UniAD [35] 16.7/72.3 21.6/64.4 36.9/82.9 12.5/63.3 05.8/58.2 03.5/60.9 27.6/71.6 10.9/44.0
RD++ [31] 25.2/78.1 17.6/64.5 42.8/84.8 33.6/80.8 08.5/61.4 04.4/62.7 29.4/72.3 11.8/46.4
DesTSeg [44] 28.4/72.6 20.3/63.9 29.0/77.0 25.6/71.9 04.5/52.8 04.1/53.6 21.2/61.0 08.5/40.7
4.2 Benchmark Results on Industrial, Medical, and General-purpose UAD Datasets
To thoroughly evaluate the effectiveness of different methods and their adaptability to various data domains, we conduct experiments on multiple datasets across three domains. Due to space constraints, we report the average metrics for the popular MVTec AD (see Tab. 3), VisA (see Tab. 4), and Real-IAD (see Tab. 5) datasets in the main paper. For the remaining datasets, we report the mAD and mIoU-max metrics (see Tab. 6). Full results for each category are provided in Appendix A to  V.
Quantitative results. InvAD [42] consistently shows excellent performance across all datasets. ViTAD [40] and MambaAD [15], specifically designed for multi-class settings, also achieve good results. In contrast, DiAD [16] and UniAD [35] require more epochs to converge and do not perform well under the 100/300 epoch standard we set. DeSTSeg [44] exhibits outstanding performance in pixel-level segmentation. Methods designed for single-class settings, such as RD [10], RD++ [31], CFLOW-AD [14], and RealNet [45], also perform well in multi-class settings. However, single-class methods like DRAEM [37], SimpleNet [27], CFA [21], and PyramidFlow [22] show significant performance gaps in multi-class anomaly detection and are not suitable for such tasks. Considering the training time, model parameters, and FLOPs shown in Fig. 2, InvAD, InvAD-lite, and ViTAD achieve a good balance of effectiveness and efficiency. RD, UniAD, and DeSTSeg also perform well in terms of both efficiency and effectiveness. On the other hand, methods like DiAD, PyramidFlow, CFLOW-AD, RD++, RealNet, MambaAD, and SimpleNet have significantly longer training times compared to other methods.
Convergence analysis. From Tab. 3 and Tab. 4, we analyze the convergence of different methods by comparing the results after training for 100 epochs and 300 epochs. The methods can be categorized into three groups: 1) Methods that show no significant improvement in performance after 300 epochs compared to 100 epochs, indicating rapid convergence within 100 epochs. These models include DRAEM [37], SimpleNet [27], CFA [21], PyramidFlow [22], ViTAD [40], InvAD-lite [42], RD++ [31], and DeSTSeg [44].2) Methods that show improvement with continued training on the VisA dataset but no improvement or a decline on the MVTec AD dataset, indicating slower convergence on larger datasets. These models include RealNet and RD. 3) Methods that show significant improvement after 300 epochs compared to 100 epochs, indicating slower convergence. These models include CFLOW-AD [14], DiAD [16], InvAD [42], MambaAD [15], and UniAD [35].
Stability analysis. For current anomaly detection algorithms, most authors select the best epoch’s results as the model’s performance. However, this method of epoch selection is unscientific and may indicate significant model instability. Therefore, we further analyze model stability using Tab. 3 and Tab. 4, comparing the results at 100 epochs and 300 epochs to identify any substantial differences. The results show that SimpleNet [27] and PyramidFlow [22] exhibit considerable differences, indicating poor model stability, while other methods do not show significant fluctuations.
Figure 4: A Pearson correlation coefficient analysis for different methods on several datasets.
Cross-domain dataset correlation. To analyze the adaptability of different methods across various datasets and the relationships and differences between different types of datasets, we employ Pearson correlation analysis to examine the correlations among these datasets. Specifically, we select four distinct datasets for analysis: MVTec AD [3], Uni-Medical [40], Real-IAD [32], and COCO-AD [42]. MVTec AD represents a fundamental industrial dataset, Uni-Medical consists of medical images from CT scans, Real-IAD is a large-scale multi-view industrial dataset from real-world scenarios, and COCO-AD is a large-scale panoptic segmentation dataset from real-life scenes. We evaluate four categories of methods using eight metrics: image and pixel mAU-ROC, mAP, mF1-max, region mAU-PRO, and segmentation mIoU-max. The results, as shown in Fig. 4, indicate that the COCO-AD and Uni-Medical datasets exhibit lower Pearson correlation coefficients due to significant differences in data distribution compared to general industrial datasets. Although the Pearson correlation coefficient between the Uni-Medical and Real-IAD datasets is relatively high, Tab. 5 and Tab. A11 in the Appendix reveal that this is because all methods perform poorly on these two datasets. Additionally, it is observed that the methods SimpleNet and DeSTSeg show considerable instability in their results across different datasets. This instability may be attributed to the inherent instability of the data augmentation algorithms they employ.
Training-free PatchCore PatchCore [29] does not require model training. It extracts all features from the training data, then selects a core subset and stores it in a Memory Bank. During testing, each test image is compared with the Memory Bank to compute an anomaly score. Because it stores the core subset of all normal features in the Memory Bank, PatchCore is only feasible for multi-class anomaly detection tasks on small-scale datasets. For large-scale datasets, it faces limitations due to insufficient GPU and memory resources. Although it achieves excellent results on the MVTec AD dataset, as shown in Tab. 1, its testing time is nearly a thousand times longer than other methods. In summary, PatchCore performs exceptionally well on small-scale datasets but is constrained by large-scale datasets and testing time.
Dataset Analysis The experimental results indicate that there is room for improvement in the VisA [47] and Real-IAD [32] datasets due to the very small defect areas, necessitating models with stronger capabilities for detecting minor defects. The MAD_Real and MAD_Sim [46] datasets, due to their small data volume and varying difficulty levels, result in similar performance across all models, particularly in the mF1-max metric. The Uni-Medical [40] dataset, consisting of images converted from CT scans, has a data distribution that significantly differs from other industrial datasets, suggesting the need for specialized detection networks tailored to medical datasets. COCO-AD [42], as a newly proposed large-scale dataset for general scenarios, presents high complexity. Current industrial AD networks are unable to achieve effective results on the COCO-AD dataset.
4.3 Challenges for Current VAD
Immature method. For challenging anomaly detection datasets such as MVTecLOCO, pose-agnostic MAD, and general-purpose COCO-AD, current methods perform poorly in a multi-class setting. Future research should focus on designing more robust methods to address this issue.
Efficiency. Most methods do not consider model complexity during design, resulting in high FLOPs. This issue becomes more pronounced when applied to real-world high-resolution scenarios. Incorporating lightweight characteristics in model design could be a potential solution.
Dataset scale. Mainstream datasets in the VAD field, such as MVTec AD and Real-IAD, are relatively small compared to those in detection and segmentation fields and are tailored to specific industrial scenarios. This limitation could hinder technological development. Collecting larger-scale, general-scene AD datasets is crucial for the advancement of the VAD field.
VAD-specific metric. Metrics like mAU-ROC and mAP are not uniquely designed for the BAD field. Developing more reliable evaluation methods to better meet practical application needs is essential.
Augmentation and tricks In fields such as classification, detection, and segmentation, data augmentation and tricks are extremely important for model training. However, few studies explore their role in the AD field, potentially limiting model performance.
Model interpretability In many applications, understanding why a model detects a particular anomaly is crucial. Providing effective visualization tools to display detection results and the model’s decision-making process remains a challenge.
model generalization UAD involves multiple domains, and further research is needed to enhance model generalization and cross-domain application capabilities.
privacy preserving Protecting data privacy is a significant issue when handling sensitive data. Developing privacy-preserving anomaly detection methods is necessary to ensure data security during transmission and storage, preventing data breaches and tampering.
5 Conclusion and Discussion
This paper addresses the urgent need for a comprehensive and fair benchmark in the field of visual anomaly detection. We introduce a modular and scalable ADer framework designed to fairly facilitate the evaluation of fifteen advanced VAD methods across multiple mainstream datasets. ensuring a thorough and unbiased assessment of each method’s performance. Our extensive experiments reveal the strengths and weaknesses of different methods, providing valuable insights into their efficiency and training resource consumption. We also develop and open-source a GPU-assisted ADEval package to reduce the evaluation time, enabling extensive assessments. Experimental results highlight the challenges of various VAD methods and offer valuable insights for future research directions.
Limitations. While the ADer framework integrates multiple datasets, it may not cover all possible real-world application scenarios. We plan to incorporate existing and newly released datasets continuously. Although the evaluation metrics are quite comprehensive, they may not capture all aspects of performance, especially in highly specialized applications. We will continue to explore more suitable metrics for the VAD field.
Broader Impacts. The open-sourcing ADer can accelerate the development of new VAD technology for the open-source community and become a valuable resource for practitioners in the field.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction have clearly included the motivations, important assumptions, and contributions made in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The authors have discussed the limitations of the work at the end of the Conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The authors have clearly provided the full set of assumptions and complete proofs in the Method section.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The authors have presented all the experimental details in the Implementation Details section. Also, the codes are submitted as Supplementary Material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The authors have submitted all the codes in Supplementary Material which provided sufficient instructions to faithfully reproduce the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The authors have presented all the training and test details in the Implementation Details section. Also, the codes are submitted as Supplementary Material.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: Due to the substantial workload and the high stability of most methods, with errors being nearly zero, this study does not use error bars, following the approach of previous RD [10].
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The authors have included the sufficient information on the computer resources in the Implementation Details and Ablation and Analysis sections.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The authors have explained the broader impacts of the work at the end of the Conclusion section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The authors have cited the original paper that produced the code package or dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The authors have submitted the details of the code/model which includes details about training, license, limitations, etc.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Overview
The supplementary material presents the following sections to strengthen the main manuscript:
  • Sec. A shows the average quantitative results on various datasets under 100 epochs.
  • Sec. B shows the average quantitative results on various datasets under 300 epochs.
  • Sec. C shows the specific quantitative results for each class under 100 epochs on MVTec AD [3] dataset.
  • Sec. D shows the specific quantitative results for each class under 100 epochs on MVTec AD 3D [6] dataset.
  • Sec. E shows the specific quantitative results for each class under 100 epochs on MVTec LOCO-AD [5] dataset.
  • Sec. F shows the specific quantitative results for each class under 100 epochs on VisA [47] dataset.
  • Sec. G shows the specific quantitative results for each class under 100 epochs on BTAD [28] dataset.
  • Sec. H shows the specific quantitative results for each class under 100 epochs on MPDD [20] dataset.
  • Sec. I shows the specific quantitative results for each class under 100 epochs on MAD_Real [46] dataset.
  • Sec. J shows the specific quantitative results for each class under 100 epochs on MAD_Sim [46] dataset.
  • Sec. K shows the specific quantitative results for each class under 100 epochs on Uni-Medical [40] dataset.
  • Sec. L shows the specific quantitative results for each class under 100 epochs on Real-IAD [32] dataset.
  • Sec. M shows the specific quantitative results for each class under 100 epochs on COCO-AD [42] dataset.
  • Sec. N shows the specific quantitative results for each class under 300 epochs on MVTec AD [3] dataset.
  • Sec. O shows the specific quantitative results for each class under 300 epochs on MVTec AD 3D [6] dataset.
  • Sec. P shows the specific quantitative results for each class under 300 epochs on MVTec LOCO-AD [5] dataset.
  • Sec. Q shows the specific quantitative results for each class under 300 epochs on VisA [47] dataset.
  • Sec. R shows the specific quantitative results for each class under 300 epochs on BTAD [28] dataset.
  • Sec. S shows the specific quantitative results for each class under 300 epochs on MPDD [20] dataset.
  • Sec. T shows the specific quantitative results for each class under 300 epochs on MAD_Real [46] dataset.
  • Sec. U shows the specific quantitative results for each class under 300 epochs on MAD_Sim [46] dataset.
  • Sec. V shows the specific quantitative results for each class under 300 epochs on Uni-Medical [40] dataset.
Appendix A Detailed Quantitative Results on All Dataset Under 100 epochs
Table A1: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 54.5 76.3 83.6 47.6 03.2 06.7 14.3 03.5 44.7
SimpleNet [27] 95.4 98.3 95.7 96.8 48.8 51.9 86.9 36.4 83.8
RealNet [45] 84.8 94.1 90.9 72.6 48.2 41.4 56.8 28.8 72.3
Emb. CFA [21] 57.6 78.3 84.7 54.8 11.9 14.7 25.3 08.9 50.1
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 88.6
CFLOW-AD [14] 91.6 96.7 93.4 95.7 45.9 48.6 88.3 33.2 81.8
PyramidalFlow [22] 70.2 85.5 85.5 80.0 22.3 22.0 47.5 12.8 61.7
Rec. RD [10] 93.6 97.2 95.6 95.8 48.2 53.6 91.2 37.0 83.8
DiAD [16] 88.9 95.8 93.5 89.3 27.0 32.5 63.9 21.1 70.1
ViTAD [40] 98.3 99.3 97.3 97.6 55.2 58.4 92.0 42.3 87.1
InvAD [42] 98.1 99.0 97.6 98.0 56.3 59.2 94.4 42.8 87.6
InvAD-lite [42] 97.9 99.2 96.8 97.3 54.4 57.8 93.3 41.4 86.8
MambaAD [15] 97.8 99.3 97.3 97.4 55.1 57.6 93.4 41.2 87.0
Hybrid UniAD [35] 92.5 97.3 95.4 95.8 42.7 48.0 89.3 32.5 82.0
RD++ [31] 97.9 98.8 96.4 97.3 54.7 58.0 93.2 41.5 86.8
DesTSeg [44] 96.4 98.6 96.2 92.0 71.1 68.2 83.4 52.8 87.9
Table A2: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.7 81.8 88.8 46.6 00.8 01.9 21.9 01.0 46.6
SimpleNet [27] 75.8 92.3 90.4 94.7 17.3 23.4 81.0 13.9 70.1
Emb. CFA [21] 67.1 89.3 89.6 71.6 09.7 16.1 43.2 09.3 58.6
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 78.4
CFLOW-AD [14] 73.1 91.0 90.2 96.8 21.6 26.6 89.0 15.8 71.6
PyramidalFlow [22] 58.7 85.6 88.6 90.8 07.2 12.0 74.9 06.4 61.9
Rec. RD [10] 75.8 91.2 91.5 98.4 29.6 35.5 93.9 22.2 75.3
DiAD [16] 75.4 91.0 90.9 91.7 05.3 10.2 74.9 05.4 62.8
ViTAD [40] 79.6 93.4 91.6 98.0 27.2 33.3 91.6 20.4 75.4
InvAD [42] 86.9 96.2 93.3 98.7 37.4 42.4 95.2 27.4 80.3
InvAD-lite [42] 84.7 95.6 92.6 98.6 38.3 41.9 94.2 26.9 79.6
MambaAD [15] 84.1 95.1 92.2 98.6 36.9 40.8 94.2 25.9 79.0
Hybrid UniAD [35] 75.3 92.1 90.3 96.6 21.2 27.9 88.9 16.7 72.3
RD++ [31] 81.8 93.9 92.8 98.5 34.4 39.6 94.1 25.2 78.1
DesTSeg [44] 79.1 93.8 90.7 81.1 41.3 42.7 64.0 28.4 72.6
Table A3: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 52.5 65.4 77.6 46.7 04.5 10.4 12.4 05.6 41.8
SimpleNet [27] 81.6 88.5 82.9 76.5 29.0 32.7 63.8 21.2 67.4
Emb. CFA [21] 66.5 78.6 77.9 58.5 14.2 16.3 31.1 09.1 52.2
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 67.7
CFLOW-AD [14] 75.1 84.2 79.8 70.7 25.0 27.9 69.5 17.3 64.0
PyramidalFlow [22] 53.4 69.3 77.5 57.7 15.0 14.5 25.7 08.0 47.5
Rec. RD [10] 74.6 85.3 78.9 70.7 21.0 26.1 67.7 15.8 63.0
DiAD [16] 71.3 79.5 80.9 71.8 19.0 24.8 48.4 14.9 56.5
ViTAD [40] 76.4 85.2 79.6 73.4 28.7 31.2 63.1 19.8 64.8
InvAD [42] 82.0 89.9 82.0 76.4 31.1 34.8 73.0 23.1 69.2
InvAD-lite [42] 79.4 87.8 80.9 75.8 28.5 32.2 69.4 20.6 67.1
MambaAD [15] 75.8 85.5 81.0 76.5 28.8 32.5 70.2 20.6 66.4
Hybrid UniAD [35] 73.2 82.1 79.6 73.1 30.7 33.4 64.7 21.6 64.4
RD++ [31] 76.4 86.8 79.2 71.5 24.3 28.3 68.8 17.6 64.5
DesTSeg [44] 77.9 86.7 81.3 64.1 35.2 30.5 53.4 20.3 63.9
Table A4: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.1 62.4 72.9 37.5 00.6 01.7 10.0 00.9 38.0
SimpleNet [27] 86.4 89.1 82.8 96.6 34.0 37.8 79.2 25.7 74.0
RealNet [45] 71.4 79.5 74.7 61.0 25.7 22.6 27.4 13.5 54.7
Emb. CFA [21] 66.3 74.3 74.2 81.3 22.1 26.2 50.8 17.0 58.4
CFLOW-AD [14] 86.5 88.8 84.9 97.7 33.9 37.2 86.8 24.9 75.3
PyramidalFlow [22] 58.2 66.3 74.4 77.0 07.2 09.6 42.8 05.6 50.2
Rec. RD [10] 90.6 90.9 89.3 98.0 35.4 42.5 91.9 27.9 78.6
DiAD [16] 84.8 88.5 86.9 82.5 17.9 23.2 44.5 14.9 61.2
ViTAD [40] 90.4 91.1 86.0 98.2 36.4 41.0 85.7 27.5 77.2
InvAD [42] 95.4 95.7 91.6 98.9 43.3 46.8 93.1 32.5 82.4
InvAD-lite [42] 94.9 95.2 90.7 98.6 40.2 44.0 93.1 29.8 81.3
MambaAD [15] 94.5 94.9 90.2 98.4 39.3 43.7 92.1 29.5 80.8
Hybrid UniAD [35] 89.0 91.0 85.8 98.3 34.5 39.6 86.5 26.4 76.7
RD++ [31] 93.9 94.7 90.2 98.4 42.3 46.3 91.9 31.2 81.3
DesTSeg [44] 89.9 91.4 86.7 86.7 46.6 47.2 61.1 32.7 74.9
Table A5: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 71.3 78.5 78.0 49.0 03.7 06.5 16.2 03.4 47.4
SimpleNet [27] 93.2 97.3 93.3 96.3 41.5 44.3 69.8 28.6 78.8
RealNet [45] 89.7 95.3 92.8 84.0 48.1 52.7 53.4 36.6 76.1
Emb. CFA [21] 92.7 97.5 93.5 96.3 47.4 50.2 69.5 33.6 80.2
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 83.3
CFLOW-AD [14] 91.2 94.8 88.3 96.8 45.6 50.1 72.7 33.8 78.9
PyramidalFlow [22] 87.0 83.1 81.0 90.9 29.6 26.9 64.1 18.3 68.3
Rec. RD [10] 94.4 96.6 94.0 98.1 59.6 59.2 80.7 42.1 84.7
DiAD [16] 90.1 88.4 92.6 91.7 19.6 26.7 70.4 15.7 68.5
ViTAD [40] 93.6 96.8 93.3 97.6 59.2 56.7 73.2 40.1 83.1
InvAD [42] 96.1 97.8 94.3 98.1 62.6 61.4 80.9 44.3 85.9
InvAD-lite [42] 93.1 97.4 94.6 97.9 59.2 59.7 78.6 42.6 84.4
MambaAD [15] 93.1 96.1 92.5 97.7 52.5 56.0 78.2 39.0 82.5
Hybrid UniAD [35] 94.8 98.3 94.2 97.2 50.3 53.8 78.8 36.9 82.9
RD++ [31] 94.6 97.8 94.1 98.0 59.6 59.8 79.0 42.8 84.8
DesTSeg [44] 92.8 95.9 92.3 92.2 34.8 44.3 70.0 29.0 77.0
Table A6: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 35.6 53.3 72.5 42.1 02.2 04.6 19.0 02.5 35.4
SimpleNet [27] 88.4 92.0 87.9 96.5 32.0 34.6 89.0 24.5 76.2
RealNet [45] 85.1 90.2 88.3 83.3 36.1 39.6 68.1 28.2 72.3
Emb. CFA [21] 81.6 87.7 85.7 84.9 19.6 22.9 53.5 16.6 65.1
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 83.5
CFLOW-AD [14] 75.7 80.1 81.7 96.8 26.3 28.0 89.5 20.1 69.7
PyramidalFlow [22] 73.6 77.0 79.4 94.1 21.1 17.8 77.2 10.4 64.6
Rec. RD [10] 91.3 93.6 91.8 98.3 40.4 41.8 95.5 31.4 80.6
DiAD [16] 68.3 77.9 80.1 90.4 10.9 13.1 66.1 08.2 58.1
ViTAD [40] 87.8 90.5 88.0 97.7 35.2 37.4 92.8 27.7 77.3
InvAD [42] 93.7 93.9 93.0 98.2 42.4 45.7 94.8 34.0 81.9
InvAD-lite [42] 90.9 92.9 89.5 98.0 39.7 42.6 94.0 30.9 79.8
MambaAD [15] 88.7 93.2 90.8 97.5 33.6 38.1 92.3 26.8 78.1
Hybrid UniAD [35] 70.5 76.0 78.0 93.9 13.7 19.9 79.7 12.5 63.3
RD++ [31] 90.2 93.3 90.5 98.5 43.0 44.1 95.5 33.6 80.8
DesTSeg [44] 91.3 90.8 90.2 82.0 32.6 34.6 63.3 25.6 71.9
Table A7: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 59.1 88.1 90.4 49.4 00.6 01.7 22.3 00.8 48.9
SimpleNet [27] 47.8 81.9 90.7 83.1 05.4 11.6 59.8 06.3 56.7
Emb. CFA [21] 49.3 81.7 91.4 83.3 10.5 15.2 64.2 08.7 58.7
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 68.2
CFLOW-AD [14] 59.8 86.0 91.2 90.9 10.7 15.4 74.4 08.8 63.4
PyramidalFlow [22] 52.1 83.6 90.3 88.4 05.0 09.4 68.2 05.1 59.0
Rec. RD [10] 55.5 85.8 90.6 87.0 08.2 12.8 67.8 07.2 60.6
DiAD [16] 49.9 85.1 90.8 87.7 02.9 06.8 67.4 03.6 55.8
ViTAD [40] 47.0 81.3 90.2 87.3 04.5 09.3 66.6 05.0 57.4
InvAD [42] 66.5 87.7 93.8 88.3 22.1 26.2 73.2 16.8 67.6
InvAD-lite [42] 60.4 84.3 92.2 88.2 09.0 14.3 72.9 08.1 62.5
MambaAD [15] 56.2 84.3 90.3 87.8 07.5 12.8 68.5 07.2 60.5
Hybrid UniAD [35] 47.6 81.3 90.4 88.2 05.8 10.6 68.6 05.8 58.2
RD++ [31] 54.7 83.9 92.3 88.5 09.4 14.9 70.4 08.5 61.4
DesTSeg [44] 45.4 80.3 90.7 72.7 05.0 08.2 48.3 04.5 52.8
Table A8: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 49.5 88.3 93.8 48.4 00.8 01.5 24.5 00.7 48.0
SimpleNet [27] 58.9 91.7 93.8 85.5 04.2 08.0 63.9 04.2 60.9
Emb. CFA [21] 56.4 91.1 93.8 76.6 04.2 08.7 54.7 04.6 58.2
CFLOW-AD [14] 55.8 90.6 93.8 86.9 03.4 05.2 63.9 02.7 60.0
PyramidalFlow [22] 50.8 88.9 93.8 82.0 02.3 04.8 58.8 02.5 57.4
Rec. RD [10] 58.6 91.7 93.8 88.3 04.7 08.5 74.2 04.5 62.6
DiAD [16] 65.7 93.3 94.1 87.2 03.9 07.9 60.1 04.2 58.9
ViTAD [40] 58.5 91.5 93.8 89.0 05.0 09.5 73.5 05.0 62.8
InvAD [42] 70.4 94.6 93.9 90.5 08.8 15.7 79.3 08.6 67.4
InvAD-lite [42] 64.7 93.3 93.8 89.4 06.4 11.2 76.1 06.0 64.9
MambaAD [15] 63.0 92.9 93.8 88.3 05.2 09.4 73.3 05.0 63.6
Hybrid UniAD [35] 54.2 90.0 93.8 88.1 03.5 06.8 71.8 03.5 60.9
RD++ [31] 59.3 91.7 93.8 88.3 04.6 08.5 73.8 04.4 62.7
DesTSeg [44] 57.9 91.8 93.8 54.5 06.2 07.7 35.8 04.1 53.6
Table A9: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 51.6 60.6 70.6 33.8 02.0 05.7 10.5 03.0 37.0
SimpleNet [27] 77.5 77.7 76.7 94.3 34.4 36.0 77.0 23.3 68.9
Emb. CFA [21] 71.0 73.3 73.1 78.3 19.5 24.9 44.7 14.7 57.2
CFLOW-AD [14] 79.2 79.6 77.2 95.4 40.4 26.9 81.8 17.7 69.9
PyramidalFlow [22] 53.9 57.3 70.5 78.8 08.4 17.1 34.2 09.4 47.6
Rec. RD [10] 76.1 75.7 78.2 96.5 38.8 39.8 86.8 26.9 71.1
DiAD [16] 78.8 77.2 77.7 95.8 34.2 35.5 84.3 23.2 69.1
ViTAD [40] 81.8 80.7 80.0 97.1 48.3 48.2 86.7 33.7 75.5
InvAD [42] 82.4 80.5 80.5 97.3 46.2 46.1 89.3 32.6 75.4
InvAD-lite [42] 77.9 77.3 79.1 96.2 39.0 39.3 85.8 26.5 71.6
MambaAD [15] 83.9 80.8 81.9 96.8 45.8 47.5 88.2 33.5 75.9
Hybrid UniAD [35] 79.0 76.1 77.1 96.6 39.3 41.1 86.0 27.6 71.6
RD++ [31] 77.2 76.5 79.2 96.7 41.5 42.2 87.8 29.4 72.3
DesTSeg [44] 78.5 77.0 78.2 65.7 41.7 34.0 35.3 21.2 61.0
Table A10: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 50.9 45.9 61.3 44.0 00.2 00.4 13.6 00.2 33.6
SimpleNet [27] 54.9 50.6 61.5 76.1 01.9 04.9 42.4 02.5 43.5
Emb. CFA [21] 55.7 50.5 61.9 81.3 01.6 03.8 48.8 02.0 45.0
CFLOW-AD [14] 77.0 75.8 69.9 94.8 17.6 21.7 80.4 12.4 63.9
PyramidalFlow [22] 54.4 48.0 62.0 71.1 01.2 01.1 34.9 00.5 40.9
Rec. RD [10] 82.7 79.3 74.1 97.2 25.2 32.8 90.0 20.0 70.0
DiAD [16] 75.6 66.4 69.9 88.0 02.9 07.1 58.1 03.7 52.6
ViTAD [40] 82.7 80.2 73.7 97.2 24.3 32.3 84.8 19.6 69.3
InvAD [42] 89.4 87.0 80.2 98.4 32.6 38.9 92.7 24.6 75.6
InvAD-lite [42] 87.2 85.2 77.8 98.0 31.7 37.9 92.0 23.8 74.2
MambaAD [15] 87.0 85.3 77.6 98.6 32.4 38.1 91.2 23.9 74.2
Hybrid UniAD [35] 83.1 81.2 74.5 97.4 23.3 30.9 87.1 18.6 69.6
RD++ [31] 83.6 80.6 74.8 97.7 25.9 33.6 90.7 20.5 70.8
DesTSeg [44] 79.3 76.7 70.7 80.3 36.9 40.3 56.1 26.2 64.5
Table A11: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 53.5 48.5 62.5 49.9 07.6 14.4 15.3 08.0 38.3
SimpleNet [27] 55.6 49.4 62.3 60.2 14.5 19.2 26.1 11.5 42.9
Emb. CFA [21] 56.7 51.2 62.6 56.2 10.3 16.0 17.9 08.9 41.0
CFLOW-AD [14] 67.7 57.9 64.5 76.0 20.3 26.4 47.7 16.0 53.0
PyramidalFlow [22] 51.6 47.4 61.6 50.0 13.4 14.4 15.0 08.0 38.4
Rec. RD [10] 57.6 49.9 62.0 66.5 13.9 20.0 39.8 11.5 45.8
DiAD [16] 59.0 53.0 63.2 68.1 20.5 14.2 30.8 11.6 44.1
ViTAD [40] 66.9 59.3 63.7 76.2 27.6 32.2 39.1 20.1 53.5
InvAD [42] 64.2 56.0 63.4 71.2 18.5 24.1 45.8 14.3 50.5
InvAD-lite [42] 63.3 56.0 62.9 69.9 18.2 23.1 39.7 13.7 49.2
MambaAD [15] 62.8 55.2 62.8 68.9 16.7 22.0 41.6 12.9 48.8
Hybrid UniAD [35] 55.2 49.3 61.7 64.6 12.8 19.0 34.3 10.9 44.0
RD++ [31] 57.5 49.3 62.0 68.2 15.5 20.1 42.2 11.8 46.4
DesTSeg [44] 54.4 49.4 61.8 54.5 10.6 15.4 24.4 08.5 40.7
Appendix B Detailed Quantitative Results on All Dataset Under 300 epochs
Table A12: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.2 77.0 83.9 48.7 03.1 06.3 15.8 03.3 45.3
SimpleNet [27] 79.2 90.8 87.6 82.4 24.0 29.0 62.0 17.8 67.6
RealNet [45] 82.9 93.3 90.9 69.8 50.0 40.4 51.2 28.5 70.9
Emb. CFA [21] 55.8 78.8 84.5 43.9 04.8 08.9 19.3 04.7 46.1
CFLOW-AD [14] 92.7 97.2 94.0 95.8 46.8 49.6 89.0 34.0 82.5
PyramidalFlow [22] 66.2 84.3 85.1 74.2 17.2 19.6 40.0 11.4 58.1
Rec. RD [10] 90.5 95.0 95.1 95.9 47.1 52.1 91.2 35.8 82.6
DiAD [16] 92.0 96.8 94.4 89.3 27.3 32.7 64.4 21.3 71.0
ViTAD [40] 98.4 99.4 97.5 97.5 55.2 58.1 91.7 42.0 87.0
InvAD [42] 98.9 99.6 98.2 98.1 57.1 59.6 94.4 43.1 88.1
InvAD-lite [42] 98.1 99.1 96.8 97.3 55.0 58.1 93.1 41.7 86.9
MambaAD [15] 98.5 99.5 97.7 97.6 56.1 58.7 93.6 42.3 87.5
Hybrid UniAD [35] 96.8 98.9 97.0 96.8 45.0 50.2 91.0 34.2 84.2
RD++ [31] 95.8 98.0 96.6 97.3 53.0 57.0 92.9 40.5 85.9
DesTSeg [44] 96.3 98.8 96.1 92.6 75.8 71.3 82.6 56.6 88.8
Table A13: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 49.2 79.4 89.0 50.7 00.7 02.2 23.4 01.1 45.9
SimpleNet [27] 74.9 91.8 90.2 94.0 14.4 21.4 79.2 12.4 68.9
Emb. CFA [21] 60.1 86.3 89.5 52.0 07.0 11.9 28.7 06.9 51.8
CFLOW-AD [14] 74.8 91.4 90.5 96.9 21.6 26.8 89.2 15.9 72.1
PyramidalFlow [22] 52.5 80.8 88.6 87.2 05.9 10.4 68.0 05.6 58.4
Rec. RD [10] 79.3 93.0 92.3 98.3 28.6 35.4 93.5 22.1 76.1
DiAD [16] 78.8 91.3 91.7 91.8 05.2 10.5 75.1 05.6 63.5
ViTAD [40] 78.7 93.3 91.5 98.0 27.0 33.0 91.3 20.2 75.1
InvAD [42] 86.3 95.9 93.4 98.7 36.2 41.1 94.9 26.4 79.8
InvAD-lite [42] 85.1 95.6 92.3 98.4 34.4 39.0 93.4 24.6 78.7
MambaAD [15] 85.8 95.5 92.6 98.6 37.1 40.8 94.1 26.0 79.5
Hybrid UniAD [35] 77.0 92.6 91.0 96.8 23.7 30.6 89.4 18.6 73.5
RD++ [31] 80.1 93.5 92.2 98.4 31.5 37.2 93.6 23.4 76.9
DesTSeg [44] 82.5 95.1 90.9 82.1 43.8 45.2 65.1 30.2 74.3
Table A14: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 51.4 66.9 77.5 48.0 05.2 10.5 13.7 05.6 42.3
SimpleNet [27] 81.8 89.1 82.5 70.9 28.3 32.1 61.2 20.9 66.3
Emb. CFA [21] 60.6 73.7 77.6 64.7 16.5 18.8 36.4 10.8 52.4
CFLOW-AD [14] 76.0 85.0 79.9 70.9 25.6 28.4 70.0 17.7 64.5
PyramidalFlow [22] 58.1 70.4 77.4 55.5 16.8 14.4 20.4 08.1 47.7
Rec. RD [10] 73.7 84.6 79.4 70.7 21.0 26.2 67.5 15.9 62.8
DiAD [16] 77.2 84.6 81.5 72.1 18.6 25.3 54.4 15.2 59.1
ViTAD [40] 76.2 85.0 80.1 73.2 27.9 30.6 62.0 19.3 64.4
InvAD [42] 82.3 89.9 82.4 77.7 30.9 34.1 72.8 22.4 69.4
InvAD-lite [42] 78.0 86.8 81.6 77.0 28.0 32.0 69.4 20.4 66.9
MambaAD [15] 78.2 87.1 81.8 77.5 28.0 32.9 68.6 20.8 67.1
Hybrid UniAD [35] 78.7 87.1 80.9 74.6 32.6 35.8 70.7 23.4 67.8
RD++ [31] 75.7 85.9 79.3 72.1 24.2 28.5 67.9 17.8 64.2
DesTSeg [44] 81.2 88.6 81.8 63.7 31.8 31.9 59.2 21.1 65.3
Table A15: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 56.2 64.6 74.9 45.0 00.7 01.8 16.0 00.9 40.6
SimpleNet [27] 80.7 83.8 79.3 94.4 29.2 33.1 74.2 22.1 69.5
RealNet [45] 79.2 84.8 78.3 65.4 29.2 27.9 33.9 17.4 59.9
Emb. CFA [21] 67.1 73.8 75.3 83.0 13.7 18.7 48.7 11.3 56.5
CFLOW-AD [14] 87.2 89.5 85.1 97.8 34.2 37.2 87.3 24.9 75.7
PyramidalFlow [22] 69.0 72.9 75.8 79.1 07.9 08.7 52.6 04.7 54.8
Rec. RD [10] 93.9 94.8 90.4 98.1 38.4 43.7 91.9 29.0 80.5
DiAD [16] 90.5 91.4 90.4 83.4 19.2 25.0 44.3 16.2 63.5
ViTAD [40] 90.3 91.2 86.4 98.2 36.4 40.9 85.8 27.5 77.3
InvAD [42] 95.6 96.0 92.3 99.0 43.7 46.9 93.0 32.6 82.6
InvAD-lite [42] 95.3 95.8 91.0 98.7 41.2 44.9 93.2 30.6 81.8
MambaAD [15] 93.6 93.9 89.8 98.2 34.0 39.3 90.5 25.9 79.0
Hybrid UniAD [35] 91.4 93.3 87.5 98.5 35.3 40.2 89.0 26.5 78.2
RD++ [31] 93.1 94.1 90.0 98.4 40.4 44.8 91.4 29.9 80.6
DesTSeg [44] 89.0 90.3 85.9 84.8 43.3 44.4 57.5 30.1 73.0
Table A16: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 69.3 76.8 79.5 47.7 03.2 06.1 16.1 03.2 46.7
SimpleNet [27] 94.0 97.9 93.9 96.2 41.0 43.7 69.6 28.1 78.9
RealNet [45] 93.1 96.3 92.5 87.2 48.0 55.5 57.9 38.7 78.1
Emb. CFA [21] 92.7 97.8 94.6 96.2 46.7 49.7 69.0 33.2 80.2
CFLOW-AD [14] 91.6 95.6 89.7 96.9 46.0 49.0 72.7 33.0 79.2
PyramidalFlow [22] 85.3 81.7 79.0 76.1 27.7 23.4 43.1 15.0 62.3
Rec. RD [10] 94.1 96.8 93.8 98.0 57.1 58.0 79.9 41.0 84.1
DiAD [16] 90.2 88.3 92.6 91.9 20.5 27.0 70.3 16.0 68.7
ViTAD [40] 94.0 97.0 93.7 97.6 58.3 56.5 72.8 39.9 83.1
InvAD [42] 95.3 97.3 93.7 97.9 58.7 58.8 78.8 41.7 84.5
InvAD-lite [42] 94.2 97.9 94.4 97.7 55.7 57.0 77.4 40.0 83.7
MambaAD [15] 92.9 96.2 93.0 97.6 51.2 55.1 77.3 38.2 82.2
Hybrid UniAD [35] 94.5 98.4 94.9 97.4 52.4 55.5 78.9 38.4 83.5
RD++ [31] 94.5 97.6 94.3 98.0 57.7 58.2 78.9 41.2 84.3
DesTSeg [44] 93.5 96.7 93.8 94.8 39.1 38.5 72.9 24.6 78.0
Table A17: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 60.2 67.5 75.1 50.4 05.4 04.7 21.8 02.6 44.1
SimpleNet [27] 90.6 94.1 89.7 97.1 33.6 35.7 90.0 25.6 77.8
RealNet [45] 86.0 90.0 87.3 74.7 39.2 39.7 52.3 28.0 69.6
Emb. CFA [21] 92.0 93.5 90.2 95.7 29.0 33.0 83.2 23.8 76.1
CFLOW-AD [14] 80.4 85.1 84.7 97.3 28.4 30.1 90.9 21.5 72.5
PyramidalFlow [22] 72.5 76.2 77.6 89.0 12.4 15.8 68.5 09.1 60.9
Rec. RD [10] 90.3 92.8 90.5 98.3 39.6 40.6 95.2 30.2 79.8
DiAD [16] 85.8 89.2 86.5 91.4 15.3 19.2 66.1 12.0 64.8
ViTAD [40] 87.4 90.8 87.0 97.8 34.6 37.8 92.9 28.0 77.1
InvAD [42] 96.0 97.1 94.5 98.5 44.1 46.4 95.3 34.4 83.5
InvAD-lite [42] 92.8 93.2 91.4 98.3 41.4 43.9 94.6 32.1 81.0
MambaAD [15] 89.2 93.1 90.3 97.7 33.5 38.6 92.8 27.2 78.3
Hybrid UniAD [35] 80.1 83.2 85.1 95.4 19.0 25.6 83.8 16.8 69.4
RD++ [31] 90.6 93.3 90.0 98.4 42.8 43.0 95.3 32.7 80.6
DesTSeg [44] 92.6 91.8 92.8 90.8 30.6 32.9 78.3 24.1 75.3
Table A18: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 46.3 82.5 90.2 42.5 00.5 01.2 14.9 00.6 43.9
SimpleNet [27] 36.1 77.0 89.6 77.2 02.6 06.3 48.9 03.3 50.7
Emb. CFA [21] 48.5 81.2 91.4 83.3 10.4 15.0 64.1 08.6 58.5
CFLOW-AD [14] 57.4 85.9 91.2 90.9 10.8 15.3 74.1 08.8 63.0
PyramidalFlow [22] 50.6 82.9 90.5 88.6 07.5 12.0 68.8 06.7 59.4
Rec. RD [10] 56.0 84.9 90.7 86.9 08.7 13.2 66.9 07.4 60.6
DiAD [16] 58.0 85.7 91.1 87.8 05.1 09.6 69.4 05.2 58.1
ViTAD [40] 44.5 81.3 90.2 86.9 04.5 08.9 65.0 04.8 56.7
InvAD [42] 63.5 86.5 92.7 91.6 18.9 24.2 77.2 14.9 66.9
InvAD-lite [42] 57.3 84.8 91.9 90.8 11.4 16.3 74.7 09.5 63.1
MambaAD [15] 54.7 83.1 92.2 90.8 13.1 19.2 74.7 11.4 63.1
Hybrid UniAD [35] 45.4 80.3 90.8 87.8 06.4 11.6 68.3 06.4 57.8
RD++ [31] 52.3 84.0 91.3 88.7 09.3 14.4 70.4 08.2 60.8
DesTSeg [44] 36.1 77.9 89.8 82.5 03.2 06.0 52.7 03.2 52.0
Table A19: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 48.8 88.1 93.8 47.9 00.7 01.5 20.0 00.8 47.2
SimpleNet [27] 51.5 88.6 93.8 86.1 02.3 04.5 63.9 02.3 58.6
Emb. CFA [21] 55.7 91.1 93.8 71.8 03.7 07.7 45.3 04.1 56.2
CFLOW-AD [14] 56.9 90.9 93.8 86.9 03.4 05.1 63.7 02.6 60.2
PyramidalFlow [22] 49.0 88.1 93.8 84.4 02.4 05.0 62.7 02.6 57.8
Rec. RD [10] 57.9 91.3 93.8 88.3 04.4 08.2 73.7 04.3 62.3
DiAD [16] 65.6 93.4 94.1 87.7 04.4 08.8 61.4 04.7 59.4
ViTAD [40] 59.0 91.4 93.8 89.0 04.8 09.3 73.6 04.9 62.8
InvAD [42] 70.4 94.5 94.0 90.7 08.9 15.3 79.1 08.4 67.4
InvAD-lite [42] 65.3 93.4 93.9 89.7 06.6 11.4 76.4 06.1 65.1
MambaAD [15] 61.4 92.4 93.8 88.3 04.8 09.0 73.0 04.8 63.2
Hybrid UniAD [35] 59.4 91.5 93.8 89.0 04.4 08.4 74.6 04.4 62.8
RD++ [31] 58.6 91.7 93.8 88.3 04.4 08.1 73.2 04.2 62.4
DesTSeg [44] 59.2 91.9 93.8 56.8 05.7 08.6 41.8 04.5 54.9
Table A20: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 58.5 63.2 71.7 32.5 02.0 06.1 09.1 03.2 38.4
SimpleNet [27] 72.5 75.3 76.7 94.9 35.9 37.0 79.2 24.2 68.3
Emb. CFA [21] 56.3 62.8 70.4 38.1 04.8 06.6 08.9 03.4 38.9
CFLOW-AD [14] 79.0 78.5 77.3 95.4 39.6 26.3 81.7 17.0 69.5
PyramidalFlow [22] 60.0 62.8 71.9 82.2 08.8 14.9 46.9 08.3 51.5
Rec. RD [10] 76.4 75.8 77.9 96.4 38.9 39.8 86.5 27.0 71.1
DiAD [16] 82.9 81.9 79.4 96.0 35.9 38.2 85.4 25.2 71.4
ViTAD [40] 81.5 80.6 80.1 97.0 46.8 46.9 86.5 32.7 75.0
InvAD [42] 81.3 80.0 79.9 97.2 45.4 45.5 89.0 32.3 74.8
InvAD-lite [42] 77.6 77.2 79.3 96.2 38.7 39.7 85.7 27.1 71.6
MambaAD [15] 82.3 78.9 81.3 96.7 43.0 46.1 87.9 32.4 74.6
Hybrid UniAD [35] 80.4 76.6 77.5 96.5 39.0 41.3 85.8 27.8 71.9
RD++ [31] 77.2 76.1 78.5 96.7 41.0 41.9 87.4 29.1 72.0
DesTSeg [44] 82.0 85.0 79.7 83.2 50.6 50.5 66.1 34.1 72.4
Appendix C Detailed Quantitative Results on MVTec AD Dataset Under 100 epochs
Table A21: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bottle Aug. DRAEM [37] 49.4 74.0 86.3 57.8 06.5 13.2 17.3 07.1 46.8
SimpleNet [27] 100. 100. 100. 97.4 57.3 65.9 89.1 49.1 88.7
RealNet [45] 95.6 98.8 95.1 69.8 53.9 46.6 60.9 30.4 77.2
Emb. CFA [21] 64.2 87.4 86.3 65.3 11.0 15.1 35.5 08.2 55.5
PatchCore [29] 100. 100. 100. 98.9 79.0 77.1 96.1 62.7 93.9
CFLOW-AD [14] 99.9 100. 99.2 97.4 61.8 63.6 91.9 46.6 89.2
PyramidalFlow [22] 77.1 93.1 86.3 77.7 16.9 23.9 40.9 13.6 62.7
Rec. RD [10] 40.6 77.5 86.3 93.4 29.6 48.5 80.4 32.0 65.6
DiAD [16] 97.7 99.1 99.2 93.5 45.7 49.5 76.0 32.9 80.1
ViTAD [40] 100. 100. 100. 98.8 79.9 75.8 94.9 61.0 93.7
InvAD [42] 100. 100. 100. 98.6 75.2 73.6 95.7 58.2 92.9
InvAD-lite [42] 100. 100. 100. 98.4 75.2 73.5 95.7 58.1 92.8
MambaAD [15] 99.8 100. 99.2 98.7 79.7 75.7 96.0 60.9 93.6
Hybrid UniAD [35] 99.4 99.8 98.4 98.1 73.5 70.7 95.3 54.7 91.8
RD++ [31] 99.8 99.9 98.4 98.2 72.5 71.6 95.0 55.8 91.8
DesTSeg [44] 100. 100. 100. 94.9 87.2 81.1 91.8 68.3 94.4
cable Aug. DRAEM [37] 50.2 61.7 76.0 55.6 03.2 05.7 16.6 02.9 41.5
SimpleNet [27] 97.1 98.3 93.9 96.6 48.5 55.0 86.0 37.9 84.0
RealNet [45] 70.0 83.0 76.0 61.5 23.5 25.9 33.3 14.9 56.2
Emb. CFA [21] 50.3 59.9 76.3 52.8 03.7 07.5 22.3 03.9 41.9
PatchCore [29] 99.8 99.9 99.5 97.7 51.6 55.2 93.6 38.1 87.1
CFLOW-AD [14] 89.6 93.7 86.0 89.8 27.3 33.3 78.0 20.0 73.4
PyramidalFlow [22] 58.2 73.7 76.0 82.8 19.0 16.4 41.6 08.9 54.6
Rec. RD [10] 81.9 88.7 81.4 84.0 26.6 34.4 77.3 20.8 69.8
DiAD [16] 88.5 94.3 88.9 90.7 27.0 33.3 63.9 20.0 69.5
ViTAD [40] 98.8 99.3 96.3 96.3 43.6 46.8 90.6 30.6 83.7
InvAD [42] 97.9 98.7 95.1 97.1 46.7 50.8 92.1 34.0 84.5
InvAD-lite [42] 97.0 98.2 94.0 90.9 39.4 47.5 87.7 31.2 81.4
MambaAD [15] 99.3 99.6 97.4 95.2 42.0 47.9 90.4 31.5 83.8
Hybrid UniAD [35] 93.0 95.7 89.3 96.1 45.0 50.7 84.9 34.0 80.9
RD++ [31] 93.7 96.3 91.3 93.2 37.7 45.1 86.8 29.1 79.7
DesTSeg [44] 97.5 98.7 94.6 95.2 67.2 64.0 85.1 47.1 87.4
capsule Aug. DRAEM [37] 58.0 86.3 90.5 48.3 00.8 01.8 14.8 00.9 47.3
SimpleNet [27] 85.4 96.7 92.7 98.1 36.9 46.8 87.1 30.6 79.4
RealNet [45] 64.4 90.3 90.5 54.6 23.7 12.1 23.4 06.4 55.1
Emb. CFA [21] 40.8 82.5 90.5 27.0 00.6 01.8 07.0 00.9 40.2
PatchCore [29] 99.2 99.8 98.6 99.0 49.0 52.9 95.7 35.9 86.7
CFLOW-AD [14] 86.1 96.7 93.4 98.5 41.0 44.2 92.8 28.4 80.6
PyramidalFlow [22] 55.6 87.0 90.5 90.3 13.5 19.6 57.3 10.9 61.4
Rec. RD [10] 98.4 99.7 96.9 98.8 45.4 51.8 96.1 34.9 85.7
DiAD [16] 88.6 97.7 91.8 90.9 16.0 23.5 54.5 13.3 66.1
ViTAD [40] 95.3 99.0 95.9 98.2 42.0 47.1 92.4 30.8 83.3
InvAD [42] 94.3 98.7 96.0 98.9 46.7 50.8 95.5 34.0 84.7
InvAD-lite [42] 93.9 98.6 94.8 98.4 42.2 47.4 94.4 31.1 83.2
MambaAD [15] 91.8 98.2 93.2 98.3 43.5 47.8 93.0 31.4 82.5
Hybrid UniAD [35] 69.8 90.4 91.6 97.5 32.6 36.9 86.9 22.6 73.7
RD++ [31] 98.3 99.7 97.3 99.0 47.1 52.1 96.3 35.3 86.0
DesTSeg [44] 90.3 97.6 94.6 89.3 56.4 58.1 60.6 41.0 80.1
carpet Aug. DRAEM [37] 66.7 87.4 87.0 51.9 01.7 03.2 16.6 01.6 49.4
SimpleNet [27] 97.1 99.1 94.6 97.7 42.8 46.9 90.0 30.6 83.1
RealNet [45] 96.5 99.1 96.0 89.2 69.0 64.3 84.0 47.4 86.9
Emb. CFA [21] 47.5 78.0 86.4 50.9 07.0 11.4 31.9 06.1 48.0
PatchCore [29] 98.4 99.5 97.8 99.1 66.4 64.6 95.2 47.7 89.9
CFLOW-AD [14] 99.2 99.8 97.2 98.8 55.9 59.4 94.3 42.3 87.9
PyramidalFlow [22] 44.6 73.5 86.4 79.2 08.5 15.1 52.3 08.2 53.5
Rec. RD [10] 99.0 99.7 98.3 99.1 59.8 61.1 95.8 44.0 89.0
DiAD [16] 64.5 87.8 86.4 87.9 17.0 26.5 67.0 15.3 62.4
ViTAD [40] 99.5 99.9 99.4 99.0 59.8 64.1 94.7 47.2 89.5
InvAD [42] 98.6 99.6 97.2 99.1 60.8 62.1 95.8 45.1 89.0
InvAD-lite [42] 99.8 99.9 99.4 99.1 58.0 61.4 96.4 44.3 89.2
MambaAD [15] 99.8 99.9 99.4 99.2 64.0 63.8 97.3 46.8 90.4
Hybrid UniAD [35] 99.8 99.9 99.4 98.5 53.6 55.6 95.9 38.5 87.8
RD++ [31] 97.6 99.3 96.0 98.9 57.2 58.5 94.8 41.3 87.5
DesTSeg [44] 95.9 98.7 95.0 95.7 68.7 64.9 93.4 48.0 88.6
Table A22: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
grid Aug. DRAEM [37] 47.3 72.2 84.4 46.4 00.6 01.4 13.5 00.7 41.7
SimpleNet [27] 98.2 99.5 98.2 96.9 26.4 33.0 88.3 19.7 79.9
RealNet [45] 97.2 99.2 96.4 82.6 41.2 45.7 77.7 29.7 79.7
Emb. CFA [21] 51.1 76.4 84.4 39.9 00.5 01.4 05.2 00.7 41.2
PatchCore [29] 99.4 99.8 98.2 98.9 40.8 44.3 94.0 28.4 84.3
CFLOW-AD [14] 86.7 95.5 88.0 92.9 18.2 25.1 81.0 14.4 72.2
PyramidalFlow [22] 75.5 89.9 85.1 85.7 09.8 16.6 66.9 09.0 64.1
Rec. RD [10] 98.7 99.6 97.4 99.2 47.6 48.8 97.5 32.3 85.9
DiAD [16] 99.0 99.7 98.2 85.7 04.0 09.5 56.7 05.0 64.7
ViTAD [40] 99.6 99.9 99.1 98.6 31.3 36.5 95.8 22.3 82.5
InvAD [42] 99.5 99.8 97.4 99.1 45.1 46.6 97.1 30.4 85.4
InvAD-lite [42] 100. 100. 100. 99.2 49.4 49.5 97.7 32.9 87.0
MambaAD [15] 99.8 99.9 99.1 98.9 48.4 48.5 96.9 32.0 86.4
Hybrid UniAD [35] 99.0 99.7 98.2 94.8 26.6 34.0 89.3 20.5 80.1
RD++ [31] 99.2 99.7 97.3 99.2 46.8 50.2 97.0 33.5 86.0
DesTSeg [44] 99.6 99.9 98.2 96.6 47.7 52.0 91.9 35.1 85.6
hazelnut Aug. DRAEM [37] 45.7 68.1 77.8 41.1 03.6 04.2 12.4 02.1 39.6
SimpleNet [27] 100. 100. 100. 98.1 49.0 52.4 93.9 35.5 86.7
RealNet [45] 100. 100. 100. 77.5 44.2 48.9 75.4 32.4 80.8
Emb. CFA [21] 51.8 67.7 78.0 45.7 04.5 08.3 17.3 04.3 42.4
PatchCore [29] 100. 100. 100. 99.0 61.2 63.6 96.2 46.6 90.0
CFLOW-AD [14] 99.9 100. 99.3 98.5 59.1 57.9 95.6 40.7 88.8
PyramidalFlow [22] 92.1 95.0 91.2 92.7 33.2 36.9 84.2 22.6 77.3
Rec. RD [10] 100. 100. 100. 98.7 58.6 59.6 96.5 42.5 89.2
DiAD [16] 96.9 98.6 96.3 95.3 27.2 36.9 81.3 22.6 76.1
ViTAD [40] 100. 100. 100. 99.0 64.8 64.3 95.6 47.4 90.5
InvAD [42] 100. 100. 100. 99.0 61.8 62.6 96.6 45.5 90.0
InvAD-lite [42] 99.7 99.8 99.3 98.9 58.4 63.0 95.7 46.0 89.3
MambaAD [15] 100. 100. 100. 99.1 67.0 66.1 95.3 49.4 90.9
Hybrid UniAD [35] 99.6 99.8 98.6 98.0 53.0 53.9 93.7 36.9 87.0
RD++ [31] 94.8 96.2 92.0 98.8 59.1 58.7 96.4 41.6 86.3
DesTSeg [44] 99.6 99.8 98.6 96.8 84.7 80.2 92.9 66.9 94.0
leather Aug. DRAEM [37] 53.2 81.4 85.2 45.4 01.1 02.9 16.3 01.5 44.8
SimpleNet [27] 100. 100. 100. 98.5 27.8 33.9 95.5 20.4 82.0
RealNet [45] 100. 100. 100. 97.9 70.4 68.0 98.0 51.5 91.8
Emb. CFA [21] 62.2 82.7 87.8 65.4 07.7 14.8 40.7 08.0 54.9
PatchCore [29] 100. 100. 100. 99.4 52.1 51.2 97.8 34.5 87.6
CFLOW-AD [14] 100. 100. 100. 99.2 45.0 46.1 98.1 30.0 86.0
PyramidalFlow [22] 67.2 77.3 89.9 87.7 06.4 15.2 74.0 08.2 62.0
Rec. RD [10] 100. 100. 100. 99.3 39.1 46.6 97.9 30.4 85.4
DiAD [16] 97.6 99.1 95.8 90.6 05.9 11.9 68.3 06.3 67.0
ViTAD [40] 100. 100. 100. 99.5 51.1 55.8 98.0 38.7 88.0
InvAD [42] 100. 100. 100. 99.5 54.3 54.3 98.1 37.3 88.3
InvAD-lite [42] 100. 100. 100. 99.3 45.6 47.4 98.5 31.0 86.3
MambaAD [15] 100. 100. 100. 99.3 50.6 50.4 98.7 33.7 87.4
Hybrid UniAD [35] 100. 100. 100. 99.1 41.1 43.5 98.3 27.8 85.2
RD++ [31] 100. 100. 100. 99.3 40.4 47.1 97.5 30.8 85.5
DesTSeg [44] 100. 100. 100. 99.3 79.6 71.9 98.4 56.2 93.7
metal_nut Aug. DRAEM [37] 58.1 82.5 89.4 38.8 08.9 21.3 11.0 11.9 48.3
SimpleNet [27] 99.0 99.8 97.8 97.7 82.2 79.2 87.6 65.6 92.8
RealNet [45] 78.6 94.7 89.4 52.5 32.3 21.0 39.6 11.7 62.0
Emb. CFA [21] 78.3 94.4 89.9 71.4 37.4 36.5 30.1 22.4 65.7
PatchCore [29] 100. 100. 100. 98.8 88.9 85.8 95.6 75.1 96.1
CFLOW-AD [14] 96.9 99.3 96.2 96.0 71.2 71.7 88.5 55.8 89.7
PyramidalFlow [22] 63.0 89.9 90.2 81.6 41.9 45.8 37.3 29.7 66.3
Rec. RD [10] 99.9 100. 99.5 93.9 62.8 64.7 92.0 47.9 89.1
DiAD [16] 88.1 97.4 89.9 94.5 74.0 72.7 53.5 57.2 81.4
ViTAD [40] 99.7 99.9 98.9 96.1 74.0 75.7 93.0 60.9 92.1
InvAD [42] 100. 100. 100. 97.4 79.0 81.2 93.1 68.4 93.8
InvAD-lite [42] 99.8 100. 98.9 97.2 78.3 79.5 93.8 66.0 93.4
MambaAD [15] 99.7 99.9 98.4 96.7 74.2 78.3 92.9 64.4 92.4
Hybrid UniAD [35] 98.0 99.5 98.4 93.4 55.0 67.2 80.8 50.6 86.4
RD++ [31] 100. 100. 100. 96.7 74.7 78.4 94.1 64.5 93.0
DesTSeg [44] 99.8 100. 98.9 95.8 88.2 82.5 94.8 70.2 94.9
Table A23: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pill Aug. DRAEM [37] 57.1 87.5 92.1 34.8 02.0 06.6 04.4 03.4 45.4
SimpleNet [27] 90.6 98.1 94.2 96.7 74.7 70.6 85.3 54.6 88.1
RealNet [45] 68.5 92.8 91.6 54.4 47.8 08.9 35.1 04.6 60.4
Emb. CFA [21] 71.3 93.4 92.2 88.7 66.4 63.7 42.1 46.8 75.4
PatchCore [29] 97.0 99.4 96.5 98.1 76.8 74.0 96.1 58.7 91.9
CFLOW-AD [14] 84.3 96.5 93.2 96.7 59.5 56.3 90.5 39.2 83.5
PyramidalFlow [22] 67.0 92.8 91.6 83.3 18.3 26.0 65.2 14.9 66.0
Rec. RD [10] 97.2 99.5 96.8 97.4 63.0 64.8 96.1 47.9 89.1
DiAD [16] 94.7 98.9 97.2 92.8 49.1 49.7 65.2 33.1 78.2
ViTAD [40] 96.9 99.5 96.1 98.7 77.7 75.0 95.5 59.9 92.1
InvAD [42] 98.6 99.8 97.5 97.9 67.7 68.7 96.6 52.3 90.7
InvAD-lite [42] 96.3 99.4 95.4 97.6 65.4 67.9 96.4 51.4 89.4
MambaAD [15] 94.7 99.0 96.5 96.2 55.0 58.9 95.2 41.7 86.5
Hybrid UniAD [35] 75.1 94.3 91.6 90.0 21.4 28.6 89.2 16.7 72.2
RD++ [31] 98.0 99.6 97.5 98.2 72.7 70.3 96.4 54.2 91.4
DesTSeg [44] 92.7 98.6 94.8 90.1 78.7 73.9 70.2 58.6 86.8
screw Aug. DRAEM [37] 48.4 71.5 85.3 41.4 00.2 00.5 09.3 00.3 40.6
SimpleNet [27] 77.5 90.6 86.7 95.8 15.8 23.8 83.1 13.5 69.8
RealNet [45] 71.3 86.6 87.0 51.8 15.4 04.5 18.5 02.3 52.1
Emb. CFA [21] 62.4 78.6 88.5 26.5 00.2 00.7 09.9 00.3 42.9
PatchCore [29] 98.3 99.4 97.1 99.5 41.5 44.6 97.2 28.7 84.5
CFLOW-AD [14] 63.4 85.1 85.3 96.5 13.6 18.5 87.7 10.2 66.0
PyramidalFlow [22] 63.7 78.7 87.0 71.4 00.9 02.1 21.5 01.0 50.2
Rec. RD [10] 97.8 99.3 96.3 99.4 42.3 47.3 97.1 31.0 84.7
DiAD [16] 63.2 82.8 86.9 86.3 03.3 09.2 57.4 04.8 55.6
ViTAD [40] 91.6 97.2 91.4 98.9 32.3 39.9 93.6 24.9 79.8
InvAD [42] 91.7 95.1 96.0 99.5 44.0 46.4 97.6 30.2 83.1
InvAD-lite [42] 93.0 97.4 93.6 99.4 47.2 48.9 97.5 32.4 84.0
MambaAD [15] 92.7 96.8 94.0 99.3 45.3 45.1 97.0 29.1 83.1
Hybrid UniAD [35] 88.4 94.4 92.3 98.2 19.2 28.4 90.9 16.5 75.4
RD++ [31] 97.5 99.2 95.8 99.5 45.5 48.9 97.7 32.3 85.2
DesTSeg [44] 86.1 95.2 88.6 76.2 52.1 50.0 53.3 33.3 73.9
tile Aug. DRAEM [37] 68.0 84.7 84.0 48.2 06.3 13.3 14.0 07.1 49.7
SimpleNet [27] 100. 100. 99.4 95.4 59.1 60.4 82.5 43.3 87.1
RealNet [45] 97.5 99.3 97.6 93.9 84.1 76.8 90.5 62.3 92.2
Emb. CFA [21] 46.0 74.5 83.6 40.3 08.5 13.2 16.6 07.0 43.8
PatchCore [29] 100. 100. 99.4 96.4 58.7 67.1 86.6 50.5 88.5
CFLOW-AD [14] 99.8 99.9 99.4 96.0 56.0 62.1 86.5 45.0 87.4
PyramidalFlow [22] 78.6 92.2 83.6 75.7 55.1 28.1 34.4 16.3 66.6
Rec. RD [10] 99.0 99.6 97.0 95.3 48.7 61.0 86.3 43.9 85.7
DiAD [16] 93.4 96.6 95.1 76.1 24.1 32.9 49.7 19.7 66.8
ViTAD [40] 100. 100. 100. 96.7 56.6 69.0 88.7 52.7 88.9
InvAD [42] 100. 100. 100. 95.8 49.8 62.4 88.3 45.4 87.0
InvAD-lite [42] 98.0 99.2 96.5 93.9 44.9 54.5 80.1 37.4 83.1
MambaAD [15] 96.8 98.8 93.8 93.0 43.9 52.6 79.5 35.7 81.9
Hybrid UniAD [35] 96.7 98.8 95.8 89.2 38.9 47.5 77.6 31.1 80.2
RD++ [31] 100. 100. 99.4 96.7 56.8 66.4 89.8 49.7 88.6
DesTSeg [44] 99.7 99.9 98.2 97.5 89.4 81.9 95.7 69.4 95.2
toothbrush Aug. DRAEM [37] 44.2 72.8 83.3 46.8 01.4 03.1 14.5 01.6 41.6
SimpleNet [27] 89.4 95.8 90.9 98.0 53.6 55.9 80.6 38.8 82.0
RealNet [45] 76.4 91.5 83.3 84.8 50.1 56.1 34.1 39.0 70.0
Emb. CFA [21] 61.4 72.0 87.9 69.6 03.3 08.5 43.8 04.4 52.5
PatchCore [29] 91.9 96.2 95.2 99.0 54.8 58.5 91.0 41.3 85.1
CFLOW-AD [14] 84.4 94.5 86.2 98.2 45.7 47.1 84.5 30.8 78.6
PyramidalFlow [22] 89.4 95.7 90.6 73.0 42.3 31.3 23.2 18.6 67.2
Rec. RD [10] 98.3 99.4 96.6 98.9 51.8 57.7 92.6 40.5 86.7
DiAD [16] 95.3 98.1 95.2 89.6 11.0 17.5 65.3 09.6 67.5
ViTAD [40] 98.6 99.5 96.8 99.1 51.8 62.2 91.3 45.1 87.2
InvAD [42] 93.9 97.4 95.2 99.0 50.6 59.7 91.6 42.6 85.4
InvAD-lite [42] 97.8 99.1 95.2 98.9 48.1 59.1 91.9 41.9 85.9
MambaAD [15] 96.4 98.5 96.8 98.9 47.5 59.7 92.0 42.5 85.9
Hybrid UniAD [35] 83.9 93.1 90.9 98.4 39.7 49.9 85.8 33.3 78.9
RD++ [31] 95.8 98.3 95.1 99.1 57.3 61.8 92.2 44.8 87.0
DesTSeg [44] 91.7 96.5 95.2 95.6 60.6 72.1 65.5 56.4 84.0
Table A24: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
transistor Aug. DRAEM [37] 52.8 44.1 57.1 51.7 04.6 09.3 17.8 04.9 36.1
SimpleNet [27] 97.8 96.6 91.6 95.4 60.4 57.6 82.5 40.4 84.7
RealNet [45] 79.7 83.0 73.3 60.9 40.2 28.3 44.6 16.5 61.1
Emb. CFA [21] 68.9 64.8 63.4 60.5 17.6 20.4 22.0 11.3 47.9
PatchCore [29] 99.8 99.7 97.5 96.0 62.3 59.2 90.8 42.0 88.0
CFLOW-AD [14] 90.5 91.4 84.2 84.8 38.3 39.0 73.0 24.2 73.7
PyramidalFlow [22] 50.8 53.8 57.1 75.9 13.2 19.5 26.1 10.8 43.8
Rec. RD [10] 94.2 94.8 89.5 85.1 41.8 44.7 74.5 28.8 77.2
DiAD [16] 99.0 98.1 97.5 97.8 71.8 67.6 85.7 51.1 88.2
ViTAD [40] 97.7 96.7 91.6 93.6 57.5 54.3 76.6 37.3 82.9
InvAD [42] 97.8 96.8 92.7 95.7 62.0 59.7 91.9 42.6 86.5
InvAD-lite [42] 98.2 97.4 92.3 95.0 61.0 58.9 88.3 41.8 85.9
MambaAD [15] 99.9 99.8 98.7 96.0 63.8 61.6 90.4 44.5 88.7
Hybrid UniAD [35] 99.3 98.9 97.6 97.3 66.8 66.4 92.2 49.7 89.6
RD++ [31] 95.2 95.0 89.5 89.3 47.7 49.0 77.9 32.4 79.6
DesTSeg [44] 96.9 96.2 89.5 73.9 54.1 50.5 76.6 33.8 79.0
wood Aug. DRAEM [37] 74.3 90.2 87.2 49.7 04.4 07.5 14.6 03.9 51.5
SimpleNet [27] 99.5 99.8 98.3 92.5 40.3 42.4 80.0 26.9 81.5
RealNet [45] 99.6 99.9 99.2 90.4 76.1 71.9 88.8 56.1 90.7
Emb. CFA [21] 72.5 90.1 87.0 54.4 06.8 09.3 31.6 04.9 54.4
PatchCore [29] 98.7 99.6 96.7 95.2 52.1 52.3 91.3 35.4 85.5
CFLOW-AD [14] 98.4 99.5 96.6 94.2 45.7 49.8 90.2 33.1 84.1
PyramidalFlow [22] 86.7 95.5 88.5 62.6 39.5 17.9 32.3 09.8 64.2
Rec. RD [10] 99.6 99.9 98.3 95.5 49.7 51.6 91.7 34.8 85.7
DiAD [16] 94.3 97.8 94.3 82.8 15.1 22.2 57.5 12.5 66.3
ViTAD [40] 98.8 99.6 96.7 96.3 61.3 59.2 89.3 42.0 87.4
InvAD [42] 99.2 99.8 98.3 95.3 51.3 52.1 91.7 35.2 85.8
InvAD-lite [42] 96.9 99.0 95.9 94.4 44.8 47.6 89.1 31.2 83.1
MambaAD [15] 98.5 99.5 96.0 94.0 46.9 48.4 92.0 32.0 84.2
Hybrid UniAD [35] 98.2 99.5 96.6 93.1 40.6 45.1 89.5 29.1 82.6
RD++ [31] 98.9 99.7 97.5 95.6 52.2 52.4 91.0 35.5 85.8
DesTSeg [44] 98.6 99.5 97.6 93.2 79.5 72.5 95.0 56.8 91.8
zipper Aug. DRAEM [37] 44.4 79.9 88.1 55.4 02.7 06.1 20.7 03.2 46.0
SimpleNet [27] 99.1 99.8 97.5 97.9 57.8 55.4 91.9 38.3 87.3
RealNet [45] 77.3 94.1 88.1 67.6 51.9 42.6 47.7 27.1 69.5
Emb. CFA [21] 34.8 72.3 88.1 64.4 03.2 07.5 23.4 03.9 44.8
PatchCore [29] 99.6 99.9 98.7 98.9 63.4 64.6 96.4 47.8 90.1
CFLOW-AD [14] 95.6 98.5 96.7 97.9 50.2 54.8 92.4 37.7 85.4
PyramidalFlow [22] 83.3 95.1 89.1 81.0 15.4 16.3 55.7 08.9 65.6
Rec. RD [10] 99.5 99.9 99.2 98.5 56.0 61.0 95.5 43.8 88.6
DiAD [16] 73.5 90.3 90.0 84.4 14.3 24.0 56.7 13.7 61.9
ViTAD [40] 97.5 99.2 97.0 95.9 44.1 50.1 89.5 33.4 83.9
InvAD [42] 99.4 99.8 98.3 98.2 50.2 57.7 94.5 40.5 87.2
InvAD-lite [42] 97.5 99.3 96.7 98.6 58.4 61.4 95.8 44.3 88.2
MambaAD [15] 98.2 99.5 96.7 98.1 55.0 58.9 94.5 41.7 87.4
Hybrid UniAD [35] 86.9 95.5 92.8 95.8 33.5 41.0 88.7 25.8 78.2
RD++ [31] 99.3 99.8 98.7 98.4 53.5 59.2 94.8 42.1 87.9
DesTSeg [44] 97.6 99.0 98.3 90.6 71.9 66.9 86.2 50.3 88.6
Avg Aug. DRAEM [37] 54.5 76.3 83.6 47.6 03.2 06.7 14.3 03.5 44.7
SimpleNet [27] 95.4 98.3 95.7 96.8 48.8 51.9 86.9 36.4 83.8
RealNet [45] 84.8 94.1 90.9 72.6 48.2 41.4 56.8 28.8 72.3
Emb. CFA [21] 57.6 78.3 84.7 54.8 11.9 14.7 25.3 08.9 50.1
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 88.6
CFLOW-AD [14] 91.6 96.7 93.4 95.7 45.9 48.6 88.3 33.2 81.8
PyramidalFlow [22] 70.2 85.5 85.5 80.0 22.3 22.0 47.5 12.8 61.7
Rec. RD [10] 93.6 97.2 95.6 95.8 48.2 53.6 91.2 37.0 83.8
DiAD [16] 88.9 95.8 93.5 89.3 27.0 32.5 63.9 21.1 70.1
ViTAD [40] 98.3 99.3 97.3 97.6 55.2 58.4 92.0 42.3 87.1
InvAD [42] 98.1 99.0 97.6 98.0 56.3 59.2 94.4 42.8 87.6
InvAD-lite [42] 97.9 99.2 96.8 97.3 54.4 57.8 93.3 41.4 86.8
MambaAD [15] 97.8 99.3 97.3 97.4 55.1 57.6 93.4 41.2 87.0
Hybrid UniAD [35] 92.5 97.3 95.4 95.8 42.7 48.0 89.3 32.5 82.0
RD++ [31] 97.9 98.8 96.4 97.3 54.7 58.0 93.2 41.5 86.8
DesTSeg [44] 96.4 98.6 96.2 92.0 71.1 68.2 83.4 52.8 87.9
Appendix D Detailed Quantitative Results on MVTec 3D Dataset Under 100 epochs
Table A25: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bagel Aug. DRAEM [37] 54.3 84.7 88.9 53.5 02.5 06.3 32.2 03.3 49.8
SimpleNet [27] 79.5 94.8 88.9 95.2 20.0 29.3 71.1 17.1 70.8
Emb. CFA [21] 82.6 95.6 89.3 84.3 19.1 30.6 48.8 18.1 67.4
PatchCore [29] 91.8 98.0 92.9 98.8 43.7 46.8 93.3 30.5 82.4
CFLOW-AD [14] 78.8 94.2 89.3 98.3 31.3 38.9 89.7 24.2 76.0
PyramidalFlow [22] 50.5 84.4 88.9 94.3 05.3 11.5 79.3 06.1 61.1
Rec. RD [10] 86.2 96.5 90.5 98.7 40.2 44.5 92.2 28.6 80.0
DiAD [16] 75.8 93.5 90.5 89.7 04.3 10.1 66.4 05.3 61.5
ViTAD [40] 92.1 98.0 93.4 98.7 40.3 44.2 92.4 28.4 81.7
InvAD [42] 95.3 98.8 94.6 99.1 45.4 49.6 95.0 33.0 84.3
InvAD-lite [42] 87.8 97.0 90.9 98.6 41.9 45.6 92.1 29.5 80.7
MambaAD [15] 93.4 98.3 93.3 98.7 41.0 45.3 94.6 29.3 82.5
Hybrid UniAD [35] 88.3 97.2 90.3 98.1 35.5 40.3 91.3 25.2 79.1
RD++ [31] 87.2 96.6 91.5 98.7 39.5 46.9 92.5 30.7 80.6
DesTSeg [44] 95.4 98.8 94.5 94.6 63.2 60.2 86.3 43.1 86.2
cable_gland Aug. DRAEM [37] 61.1 85.0 89.7 40.1 00.3 00.5 15.7 00.3 46.4
SimpleNet [27] 75.8 92.5 91.1 96.3 15.6 23.9 87.9 13.6 71.2
Emb. CFA [21] 61.2 87.5 89.2 71.0 01.8 04.6 38.7 02.4 54.2
PatchCore [29] 96.4 99.0 96.1 99.5 37.3 43.6 98.3 27.9 83.4
CFLOW-AD [14] 69.7 90.2 90.1 96.7 09.4 16.2 90.0 08.8 68.2
PyramidalFlow [22] 64.4 88.9 89.7 87.6 01.3 02.9 62.9 01.5 59.8
Rec. RD [10] 91.2 97.5 95.0 99.5 39.7 44.8 98.5 28.9 82.6
DiAD [16] 88.0 97.1 92.6 95.0 04.6 10.1 84.2 05.3 67.4
ViTAD [40] 84.5 94.2 92.1 98.7 22.6 31.4 95.7 18.6 76.2
InvAD [42] 96.6 99.1 96.6 99.6 40.6 47.1 98.8 30.8 84.5
InvAD-lite [42] 93.7 98.5 96.0 99.4 41.1 46.9 98.5 30.6 83.8
MambaAD [15] 91.1 97.8 93.6 99.3 38.8 44.6 98.2 28.7 82.2
Hybrid UniAD [35] 73.3 91.0 89.8 97.1 14.2 22.0 91.7 12.4 70.5
RD++ [31] 96.6 99.1 96.0 99.6 46.1 49.3 98.7 32.7 85.3
DesTSeg [44] 89.9 97.5 92.9 89.9 47.8 52.3 85.3 35.4 81.1
carrot Aug. DRAEM [37] 50.8 83.6 90.7 56.2 00.6 00.9 35.4 00.5 49.2
SimpleNet [27] 82.4 96.1 92.5 98.1 13.3 20.6 92.6 11.5 73.2
Emb. CFA [21] 58.6 88.6 91.0 83.1 03.0 09.4 54.0 05.0 58.4
PatchCore [29] 95.2 99.0 95.6 99.4 25.9 30.8 97.7 18.2 80.0
CFLOW-AD [14] 85.2 96.8 92.3 98.9 20.0 27.2 96.0 15.7 76.0
PyramidalFlow [22] 59.9 87.4 90.7 97.8 13.0 18.8 92.5 10.4 67.4
Rec. RD [10] 85.4 96.2 92.5 99.4 27.9 33.8 97.7 20.4 78.0
DiAD [16] 93.4 98.3 98.1 97.6 06.2 11.2 91.1 05.9 70.8
ViTAD [40] 77.4 93.0 92.2 98.8 19.3 25.9 95.9 14.8 73.8
InvAD [42] 90.5 97.6 94.5 99.3 24.9 30.1 97.6 17.7 78.6
InvAD-lite [42] 87.7 96.9 94.6 99.4 27.6 32.7 98.0 19.5 78.7
MambaAD [15] 90.1 97.7 93.9 99.3 29.0 32.7 97.9 19.6 79.3
Hybrid UniAD [35] 75.0 93.8 91.4 98.2 15.6 22.8 94.2 12.9 72.2
RD++ [31] 89.3 97.2 93.9 99.4 28.2 34.0 97.7 20.5 79.1
DesTSeg [44] 78.2 93.7 91.3 67.6 30.4 33.4 46.1 20.0 66.1
cookie Aug. DRAEM [37] 80.6 93.4 91.1 47.3 01.7 05.2 20.5 02.7 53.5
SimpleNet [27] 66.9 89.0 88.0 92.5 27.2 34.6 71.5 21.0 68.9
Emb. CFA [21] 61.2 86.8 88.4 72.4 19.4 27.6 38.3 16.0 59.1
PatchCore [29] 71.0 91.6 88.0 98.0 52.4 50.2 92.5 33.5 78.4
CFLOW-AD [14] 46.3 78.0 88.4 96.9 30.3 32.6 89.2 19.5 66.6
PyramidalFlow [22] 23.4 67.8 88.0 88.8 07.1 11.9 70.8 06.3 52.2
Rec. RD [10] 50.2 80.1 88.0 96.7 28.1 33.2 87.4 19.9 67.1
DiAD [16] 52.6 84.7 88.0 87.2 05.6 10.5 67.2 05.5 56.6
ViTAD [40] 70.7 90.9 88.2 97.7 47.6 48.8 87.6 32.3 76.8
InvAD [42] 63.5 88.5 88.0 98.4 48.8 49.3 92.5 32.7 76.1
InvAD-lite [42] 62.0 88.4 88.0 97.3 39.9 42.8 87.0 27.2 73.1
MambaAD [15] 58.3 85.6 88.4 96.8 38.8 42.0 84.3 26.6 71.5
Hybrid UniAD [35] 62.3 86.9 88.0 96.0 24.2 29.6 88.1 17.4 69.3
RD++ [31] 56.7 82.8 88.4 97.6 39.6 42.6 89.2 27.1 71.6
DesTSeg [44] 61.7 87.9 88.4 83.7 43.9 45.2 57.8 29.2 68.5
Table A26: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
dowel Aug. DRAEM [37] 38.5 75.0 88.9 42.1 00.7 02.2 24.6 01.1 42.4
SimpleNet [27] 90.3 97.5 92.5 97.8 16.2 21.5 90.2 12.0 74.9
Emb. CFA [21] 88.8 96.8 92.9 85.0 07.1 11.1 56.9 05.9 66.4
PatchCore [29] 98.2 99.6 97.5 99.3 44.5 46.7 96.3 30.5 85.1
CFLOW-AD [14] 92.1 98.1 92.9 98.8 28.3 32.7 94.4 19.5 79.0
PyramidalFlow [22] 66.5 89.3 89.7 93.9 06.3 10.5 74.9 05.5 64.1
Rec. RD [10] 97.3 99.3 96.7 99.7 51.1 50.6 98.8 33.9 86.4
DiAD [16] 65.3 89.1 91.9 79.7 04.6 09.2 47.9 04.8 55.4
ViTAD [40] 94.1 98.5 94.8 99.2 37.8 41.7 95.9 26.3 82.2
InvAD [42] 99.3 99.8 98.6 99.7 49.5 48.6 98.6 32.1 86.7
InvAD-lite [42] 96.8 99.3 96.0 99.6 51.4 51.5 97.7 34.7 86.2
MambaAD [15] 96.5 99.2 95.3 99.5 50.1 48.9 97.4 32.3 85.5
Hybrid UniAD [35] 87.1 96.9 90.1 98.9 34.0 39.0 95.3 24.2 79.1
RD++ [31] 99.6 99.9 98.6 99.7 53.9 51.3 98.8 34.5 87.6
DesTSeg [44] 86.0 96.5 90.2 82.8 49.4 50.1 60.8 33.4 75.8
foam Aug. DRAEM [37] 50.3 81.6 88.9 46.7 00.2 00.4 14.4 00.2 44.5
SimpleNet [27] 78.7 94.5 89.4 86.4 12.1 22.1 67.9 12.4 67.3
Emb. CFA [21] 66.1 90.6 89.4 45.1 08.0 21.0 32.6 11.8 54.4
PatchCore [29] 78.9 94.4 89.9 94.9 18.8 29.6 83.4 17.4 72.2
CFLOW-AD [14] 75.0 93.5 89.3 85.5 15.8 27.2 57.5 15.7 66.2
PyramidalFlow [22] 65.6 90.2 88.9 71.4 11.0 18.2 49.9 10.0 59.6
Rec. RD [10] 70.2 91.7 89.4 93.9 14.4 27.1 81.0 15.6 68.9
DiAD [16] 78.3 94.4 89.2 87.5 01.4 03.9 59.7 02.0 59.2
ViTAD [40] 82.9 95.4 90.8 94.0 14.3 26.0 81.3 14.9 71.8
InvAD [42] 83.5 95.8 91.2 93.8 18.8 31.0 81.1 18.3 73.2
InvAD-lite [42] 82.6 95.4 90.1 94.4 24.1 32.9 82.8 19.7 74.0
MambaAD [15] 79.9 94.7 89.9 94.9 23.8 33.4 83.8 20.1 73.6
Hybrid UniAD [35] 76.8 94.0 88.9 85.0 10.3 23.0 60.1 13.0 65.6
RD++ [31] 72.9 92.4 91.9 92.7 16.3 28.7 77.3 16.7 69.7
DesTSeg [44] 80.2 95.0 89.1 84.5 44.3 49.5 72.7 32.9 75.4
peach Aug. DRAEM [37] 61.6 86.9 89.5 46.7 00.5 01.0 22.5 00.5 48.5
SimpleNet [27] 70.0 90.5 90.1 92.2 05.5 11.0 73.3 05.8 64.5
Emb. CFA [21] 49.9 81.6 89.1 67.4 01.2 05.7 39.0 02.9 50.9
PatchCore [29] 75.2 92.7 91.2 98.2 26.4 32.4 93.3 19.3 74.5
CFLOW-AD [14] 72.2 91.3 90.4 97.9 17.1 18.9 92.1 10.4 70.6
PyramidalFlow [22] 62.0 86.7 89.8 96.4 06.6 11.0 87.3 05.8 64.9
Rec. RD [10] 64.0 85.0 91.6 98.4 15.6 21.7 93.7 12.2 68.8
DiAD [16] 92.4 93.5 99.1 95.5 04.4 08.0 84.9 04.2 68.3
ViTAD [40] 66.6 88.8 89.1 98.1 19.5 25.2 92.5 14.4 70.2
InvAD [42] 90.4 97.3 94.1 99.3 41.3 46.5 97.1 30.3 82.5
InvAD-lite [42] 85.4 95.3 94.1 99.3 42.8 44.0 97.2 28.2 81.2
MambaAD [15] 89.3 97.1 93.4 99.4 42.9 44.0 97.5 28.2 82.1
Hybrid UniAD [35] 73.7 91.8 90.4 98.0 19.6 25.5 92.5 14.6 72.1
RD++ [31] 74.2 89.8 92.0 98.8 26.2 31.7 94.8 18.8 74.1
DesTSeg [44] 70.6 91.5 89.4 76.7 46.7 45.5 56.1 29.5 70.0
potato Aug. DRAEM [37] 50.1 80.9 89.8 49.6 00.5 01.3 30.0 00.6 47.0
SimpleNet [27] 61.9 85.3 89.9 95.2 02.6 05.6 83.7 02.9 62.9
Emb. CFA [21] 39.1 77.4 89.3 64.9 00.3 01.0 30.7 00.5 46.4
PatchCore [29] 56.8 83.9 90.2 98.9 09.8 17.0 96.0 09.3 66.2
CFLOW-AD [14] 62.0 86.8 89.3 98.5 11.4 13.9 95.3 07.5 67.1
PyramidalFlow [22] 70.7 91.2 89.8 97.8 09.9 15.9 92.5 08.6 69.0
Rec. RD [10] 63.1 87.6 91.1 99.2 15.3 22.7 96.6 12.8 69.5
DiAD [16] 82.2 95.8 90.2 98.5 08.6 14.7 94.8 07.9 69.3
ViTAD [40] 58.2 86.7 89.8 98.7 11.6 17.9 95.1 09.8 67.0
InvAD [42] 66.9 90.4 90.2 99.1 17.8 24.8 96.4 14.1 71.0
InvAD-lite [42] 67.8 90.5 90.5 99.1 17.8 22.4 95.9 12.6 70.9
MambaAD [15] 59.1 86.0 90.2 99.0 17.7 22.8 95.4 12.8 68.6
Hybrid UniAD [35] 51.6 83.8 89.3 97.9 08.5 13.9 92.3 07.5 64.0
RD++ [31] 67.0 90.0 90.6 99.2 17.5 23.7 96.7 13.5 70.9
DesTSeg [44] 58.2 86.8 90.2 54.8 08.5 09.2 22.5 04.8 51.1
Table A27: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
rope Aug. DRAEM [37] 56.5 69.4 83.6 43.0 00.5 01.0 13.1 00.5 42.1
SimpleNet [27] 93.3 97.7 93.8 99.3 53.8 52.4 93.1 35.5 84.8
Emb. CFA [21] 92.5 96.8 90.1 90.4 35.7 44.1 74.5 28.3 77.2
PatchCore [29] 96.7 98.7 95.5 99.6 55.6 56.9 97.4 39.7 87.2
CFLOW-AD [14] 94.1 97.7 92.2 99.3 45.1 47.3 95.3 31.0 83.2
PyramidalFlow [22] 79.7 90.9 83.2 86.4 08.0 11.9 61.3 06.3 63.3
Rec. RD [10] 95.9 98.4 93.1 99.6 50.6 55.2 97.9 38.1 85.8
DiAD [16] 72.4 87.2 81.7 94.3 09.5 16.1 74.6 08.8 62.3
ViTAD [40] 98.5 99.4 96.4 98.5 41.7 43.3 89.5 27.6 83.2
InvAD [42] 97.0 98.8 94.9 99.6 57.8 58.9 97.1 41.7 87.6
InvAD-lite [42] 94.5 98.0 94.7 99.4 54.2 53.0 96.3 36.1 85.7
MambaAD [15] 96.1 98.4 93.2 99.3 45.1 48.1 96.1 31.7 84.0
Hybrid UniAD [35] 95.9 98.5 95.6 99.3 41.5 47.8 96.4 31.4 84.0
RD++ [31] 95.9 98.5 96.3 99.6 52.7 56.2 97.8 39.0 86.7
DesTSeg [44] 92.7 97.2 92.2 98.8 64.7 62.7 97.2 45.7 87.4
tire Aug. DRAEM [37] 52.7 76.9 87.4 40.6 00.1 00.3 10.2 00.2 42.6
SimpleNet [27] 59.5 85.7 87.4 93.9 06.9 13.1 78.3 07.0 62.8
Emb. CFA [21] 70.7 90.8 87.4 52.5 01.2 05.3 18.9 02.7 51.2
PatchCore [29] 80.1 93.6 87.9 99.1 22.6 31.1 95.8 18.4 74.7
CFLOW-AD [14] 56.1 83.9 87.4 97.5 07.0 11.2 90.7 05.9 63.7
PyramidalFlow [22] 44.2 79.2 87.4 93.7 03.7 07.3 78.0 03.8 58.0
Rec. RD [10] 54.2 79.2 87.4 99.0 13.5 21.5 95.5 12.1 65.5
DiAD [16] 54.0 76.1 87.4 92.5 03.4 08.1 78.1 04.2 57.1
ViTAD [40] 70.6 89.3 89.6 97.9 17.4 28.3 90.6 16.5 70.9
InvAD [42] 86.2 95.8 89.9 99.5 29.3 38.2 97.5 23.6 78.4
InvAD-lite [42] 89.0 96.7 91.3 99.3 42.3 47.4 96.4 31.1 81.8
MambaAD [15] 87.3 95.9 91.1 99.3 42.1 46.3 96.7 30.1 81.3
Hybrid UniAD [35] 68.8 87.5 88.7 97.0 08.4 15.4 87.5 08.4 66.9
RD++ [31] 78.4 92.6 88.7 99.4 23.6 31.8 97.0 18.9 74.8
DesTSeg [44] 78.1 93.0 88.4 77.9 13.7 18.7 55.7 10.3 64.0
Avg Aug. DRAEM [37] 55.7 81.8 88.8 46.6 00.8 01.9 21.9 01.0 46.6
SimpleNet [27] 75.8 92.3 90.4 94.7 17.3 23.4 81.0 13.9 70.1
Emb. CFA [21] 67.1 89.3 89.6 71.6 09.7 16.1 43.2 09.3 58.6
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 78.4
CFLOW-AD [14] 73.1 91.0 90.2 96.8 21.6 26.6 89.0 15.8 71.6
PyramidalFlow [22] 58.7 85.6 88.6 90.8 07.2 12.0 74.9 06.4 61.9
Rec. RD [10] 75.8 91.2 91.5 98.4 29.6 35.5 93.9 22.2 75.3
DiAD [16] 75.4 91.0 90.9 91.7 05.3 10.2 74.9 05.4 62.8
ViTAD [40] 79.6 93.4 91.6 98.0 27.2 33.3 91.6 20.4 75.4
InvAD [42] 86.9 96.2 93.3 98.7 37.4 42.4 95.2 27.4 80.3
InvAD-lite [42] 84.7 95.6 92.6 98.6 38.3 41.9 94.2 26.9 79.6
MambaAD [15] 84.1 95.1 92.2 98.6 36.9 40.8 94.2 25.9 79.0
Hybrid UniAD [35] 75.3 92.1 90.3 96.6 21.2 27.9 88.9 16.7 72.3
RD++ [31] 81.8 93.9 92.8 98.5 34.4 39.6 94.1 25.2 78.1
DesTSeg [44] 79.1 93.8 90.7 81.1 41.3 42.7 64.0 28.4 72.6
Appendix E Detailed Quantitative Results on MVTec LOCO Dataset Under 100 epochs
Table A28: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
breakfast_box Aug. DRAEM [37] 62.6 76.9 77.2 50.0 04.4 12.5 14.9 06.7 46.3
SimpleNet [27] 88.9 94.1 85.5 87.8 56.9 55.3 73.0 38.3 78.9
Emb. CFA [21] 64.3 79.1 77.6 68.0 29.7 28.5 33.9 16.6 56.8
PatchCore [29] 77.1 88.4 78.0 89.4 55.4 49.3 76.1 32.8 74.4
CFLOW-AD [14] 74.9 86.4 77.8 88.6 48.0 46.5 71.1 30.3 71.6
PyramidalFlow [22] 58.3 75.3 77.4 59.7 44.2 28.1 20.6 16.3 54.2
Rec. RD [10] 66.2 80.7 77.2 83.7 27.1 35.6 65.0 21.6 63.8
DiAD [16] 68.5 82.0 77.2 84.6 26.3 37.6 39.7 23.2 59.4
ViTAD [40] 77.7 88.0 78.8 89.1 56.1 53.1 70.3 36.1 74.3
InvAD [42] 82.7 91.1 80.5 90.2 63.7 60.8 79.7 43.6 79.2
InvAD-lite [42] 77.1 87.9 77.6 88.6 50.5 49.0 77.2 32.5 73.6
MambaAD [15] 69.7 83.6 77.5 89.1 54.5 50.1 74.5 33.5 72.0
Hybrid UniAD [35] 73.1 85.5 77.2 89.0 48.8 46.2 71.0 30.1 71.2
RD++ [31] 69.8 84.3 77.2 86.8 41.4 45.1 69.6 29.1 68.9
DesTSeg [44] 85.5 92.7 82.1 82.6 64.1 59.4 72.6 42.2 78.2
juice_bottle Aug. DRAEM [37] 52.1 74.3 83.4 31.3 03.9 10.1 07.3 05.3 41.5
SimpleNet [27] 96.7 98.8 94.8 91.6 48.9 54.6 86.1 37.5 83.5
Emb. CFA [21] 85.1 94.3 85.4 73.6 20.3 22.0 51.7 12.4 65.1
PatchCore [29] 95.2 98.3 94.1 89.9 56.2 57.8 87.1 40.6 84.3
CFLOW-AD [14] 95.9 98.4 93.6 91.7 44.1 47.3 84.1 31.0 81.4
PyramidalFlow [22] 48.6 74.9 83.5 54.2 09.7 10.1 24.9 05.3 46.9
Rec. RD [10] 90.1 96.6 89.3 89.2 46.9 47.7 84.4 31.4 79.5
DiAD [16] 97.3 98.8 96.1 83.2 40.0 42.4 74.9 26.9 76.1
ViTAD [40] 90.3 96.4 88.7 88.6 48.4 49.1 83.4 32.5 79.6
InvAD [42] 95.8 98.5 94.2 90.6 53.9 58.3 88.4 41.2 84.5
InvAD-lite [42] 93.0 97.3 92.7 90.6 53.4 57.2 87.5 40.0 83.3
MambaAD [15] 92.0 97.0 92.3 89.7 47.8 53.2 85.3 36.3 81.4
Hybrid UniAD [35] 92.2 97.0 90.5 90.6 52.5 54.6 87.3 37.6 82.2
RD++ [31] 90.6 96.5 88.8 90.1 49.2 50.9 85.2 34.2 80.4
DesTSeg [44] 98.5 99.5 97.2 78.8 57.0 57.2 79.7 40.1 83.3
pushpins Aug. DRAEM [37] 54.5 55.1 72.8 52.4 00.8 01.8 14.8 00.9 39.1
SimpleNet [27] 74.8 81.8 74.0 68.9 03.7 08.8 62.0 04.6 56.4
Emb. CFA [21] 53.7 63.2 71.4 43.6 00.8 01.7 22.5 00.9 40.0
PatchCore [29] 74.3 82.8 72.7 56.1 03.1 07.5 63.6 03.9 54.6
CFLOW-AD [14] 60.8 69.1 71.4 39.4 01.2 03.0 52.4 01.5 45.5
PyramidalFlow [22] 51.7 60.9 71.4 58.8 01.3 03.2 29.3 01.6 42.2
Rec. RD [10] 71.4 80.4 71.7 51.2 02.2 07.6 68.1 04.0 53.4
DiAD [16] 57.6 65.7 73.0 62.8 01.4 03.2 36.4 01.6 42.9
ViTAD [40] 66.8 73.6 72.2 46.1 02.2 06.0 57.1 03.1 49.4
InvAD [42] 81.0 87.4 76.7 67.2 03.8 09.1 75.1 04.8 60.2
InvAD-lite [42] 77.9 84.0 76.1 60.0 03.0 08.4 67.8 04.4 57.1
MambaAD [15] 79.8 85.9 76.3 58.8 02.6 09.4 72.7 05.0 58.3
Hybrid UniAD [35] 63.5 70.1 71.4 41.6 01.0 03.9 53.8 02.0 46.7
RD++ [31] 77.0 84.0 74.6 52.3 02.5 07.5 67.5 03.9 55.5
DesTSeg [44] 59.9 71.1 71.7 54.0 03.9 08.2 42.3 04.3 47.3
screw_bag Aug. DRAEM [37] 53.1 65.9 78.4 49.3 06.7 13.7 10.1 07.4 42.9
SimpleNet [27] 63.5 76.5 78.3 61.2 10.3 16.5 46.8 09.0 53.2
Emb. CFA [21] 63.2 77.2 78.4 50.2 07.5 13.6 30.2 07.3 49.2
PatchCore [29] 70.1 82.8 78.2 72.1 13.1 22.0 62.1 12.4 59.7
CFLOW-AD [14] 64.4 77.7 78.4 68.7 11.2 20.4 67.0 11.3 57.7
PyramidalFlow [22] 52.7 68.5 78.5 59.0 09.9 15.6 35.5 08.5 48.3
Rec. RD [10] 66.3 79.8 78.2 71.5 12.6 21.8 52.1 12.2 57.1
DiAD [16] 57.6 68.1 79.3 70.2 11.3 21.2 58.5 11.9 52.3
ViTAD [40] 65.2 77.7 79.0 72.5 13.2 21.9 42.2 12.3 55.7
InvAD [42] 68.5 82.1 78.7 71.9 13.2 21.9 53.0 12.3 58.2
InvAD-lite [42] 68.2 79.7 78.7 72.2 13.1 21.9 49.1 12.3 57.3
MambaAD [15] 55.2 70.5 78.2 73.1 13.2 22.7 53.6 12.8 54.3
Hybrid UniAD [35] 56.9 69.1 78.8 66.8 10.8 19.2 64.8 10.6 54.3
RD++ [31] 66.8 80.8 78.2 69.4 12.1 20.2 53.0 11.3 57.0
DesTSeg [44] 68.4 81.9 78.4 51.2 09.7 13.6 44.0 07.3 52.9
Table A29: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
splicing_connectors Aug. DRAEM [37] 40.1 54.7 76.4 50.4 06.5 14.1 15.0 07.6 39.3
SimpleNet [27] 83.8 91.4 81.9 73.2 25.1 28.5 51.1 16.6 65.1
Emb. CFA [21] 66.2 79.3 76.6 57.2 12.9 15.7 17.4 08.5 49.9
PatchCore [29] 85.6 92.9 84.4 68.1 21.5 22.3 60.8 12.6 65.4
CFLOW-AD [14] 79.3 89.3 77.8 65.2 20.3 22.3 73.1 12.5 63.7
PyramidalFlow [22] 55.9 67.0 76.4 56.7 09.8 15.7 18.2 08.5 45.8
Rec. RD [10] 79.2 89.1 78.0 58.0 16.3 17.9 68.8 09.9 61.2
DiAD [16] 75.7 83.0 78.8 58.1 15.9 19.6 32.4 10.9 51.9
ViTAD [40] 82.1 90.6 79.6 70.5 23.7 25.8 62.4 14.8 64.8
InvAD [42] 81.9 90.5 79.7 62.2 20.8 23.8 68.7 13.5 64.0
InvAD-lite [42] 80.8 89.9 79.3 67.9 22.6 24.5 65.5 14.0 64.2
MambaAD [15] 82.4 90.7 80.4 71.8 26.0 26.8 64.9 15.5 65.9
Hybrid UniAD [35] 80.4 88.7 80.1 77.7 40.5 43.2 46.5 27.5 67.5
RD++ [31] 77.7 88.5 77.3 59.2 16.2 17.5 68.9 09.6 60.8
DesTSeg [44] 77.5 88.1 77.0 53.8 41.3 14.0 28.4 07.5 57.6
Avg Aug. DRAEM [37] 52.5 65.4 77.6 46.7 04.5 10.4 12.4 05.6 41.8
SimpleNet [27] 81.6 88.5 82.9 76.5 29.0 32.7 63.8 21.2 67.4
Emb. CFA [21] 66.5 78.6 77.9 58.5 14.2 16.3 31.1 09.1 52.2
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 67.7
CFLOW-AD [14] 75.1 84.2 79.8 70.7 25.0 27.9 69.5 17.3 64.0
PyramidalFlow [22] 53.4 69.3 77.5 57.7 15.0 14.5 25.7 08.0 47.5
Rec. RD [10] 74.6 85.3 78.9 70.7 21.0 26.1 67.7 15.8 63.0
DiAD [16] 71.3 79.5 80.9 71.8 19.0 24.8 48.4 14.9 56.5
ViTAD [40] 76.4 85.2 79.6 73.4 28.7 31.2 63.1 19.8 64.8
InvAD [42] 82.0 89.9 82.0 76.4 31.1 34.8 73.0 23.1 69.2
InvAD-lite [42] 79.4 87.8 80.9 75.8 28.5 32.2 69.4 20.6 67.1
MambaAD [15] 75.8 85.5 81.0 76.5 28.8 32.5 70.2 20.6 66.4
Hybrid UniAD [35] 73.2 82.1 79.6 73.1 30.7 33.4 64.7 21.6 64.4
RD++ [31] 76.4 86.8 79.2 71.5 24.3 28.3 68.8 17.6 64.5
DesTSeg [44] 77.9 86.7 81.3 64.1 35.2 30.5 53.4 20.3 63.9
Appendix F Detailed Quantitative Results on VisA Dataset Under 100 epochs
Table A30: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
candle Aug. DRAEM [37] 36.0 45.5 66.7 32.0 00.1 00.3 05.5 00.2 29.4
SimpleNet [27] 93.6 94.5 84.9 96.4 09.9 19.9 88.4 11.1 72.3
RealNet [45] 52.7 57.8 66.7 51.8 09.8 05.2 28.5 02.7 41.4
Emb. CFA [21] 71.8 73.7 70.6 76.2 03.7 12.2 59.7 06.5 55.0
CFLOW-AD [14] 93.0 93.3 85.3 98.8 13.9 23.0 93.8 13.0 74.0
PyramidalFlow [22] 43.4 44.9 66.7 77.9 00.5 02.3 55.8 01.2 42.9
Rec. RD [10] 93.1 93.4 86.9 99.1 24.2 35.4 95.5 21.5 77.3
DiAD [16] 92.3 93.6 84.5 80.2 01.0 04.1 43.1 02.1 57.0
ViTAD [40] 90.0 91.3 81.6 96.1 16.2 27.9 86.6 16.2 72.2
InvAD [42] 95.3 94.8 90.6 99.2 23.3 34.1 95.8 20.5 78.3
InvAD-lite [42] 97.1 97.2 92.0 99.0 21.7 30.6 96.0 18.1 78.6
MambaAD [15] 95.8 96.4 89.6 99.0 21.3 31.5 95.8 18.7 77.9
Hybrid UniAD [35] 95.7 95.6 89.2 99.0 24.1 32.8 93.7 19.6 77.9
RD++ [31] 92.5 92.9 87.8 99.3 26.2 37.1 95.5 22.7 77.8
DesTSeg [44] 89.7 90.5 86.1 79.6 41.1 43.9 61.7 28.1 72.7
capsules Aug. DRAEM [37] 49.0 68.2 76.9 54.8 00.4 00.7 20.3 00.3 41.9
SimpleNet [27] 76.8 87.0 78.3 95.5 42.0 45.5 66.1 29.4 71.5
RealNet [45] 72.8 85.0 77.0 65.9 31.8 34.8 31.8 21.1 59.7
Emb. CFA [21] 56.5 71.2 77.2 55.2 00.4 00.9 05.9 00.4 41.9
CFLOW-AD [14] 54.8 72.8 76.9 94.2 26.4 30.4 64.3 17.9 61.0
PyramidalFlow [22] 55.8 69.4 76.9 86.9 03.3 08.7 62.4 04.5 53.8
Rec. RD [10] 81.8 90.5 81.3 99.3 60.2 60.7 93.0 43.5 81.4
DiAD [16] 51.5 71.2 78.1 90.6 02.9 06.5 49.7 03.4 50.1
ViTAD [40] 80.4 88.2 78.9 98.2 31.0 41.6 75.5 26.2 72.0
InvAD [42] 90.1 94.2 86.8 99.6 68.6 64.9 95.0 48.1 86.2
InvAD-lite [42] 91.7 94.7 90.1 99.3 60.5 58.5 95.2 41.3 85.3
MambaAD [15] 91.1 94.4 89.6 99.0 58.4 56.7 93.2 39.5 84.3
Hybrid UniAD [35] 71.2 83.9 78.4 97.5 42.2 45.5 73.6 29.4 71.3
RD++ [31] 84.3 91.4 83.5 99.4 64.1 63.6 93.2 46.6 83.2
DesTSeg [44] 85.4 91.8 83.3 95.9 39.0 43.6 82.6 27.9 76.1
cashew Aug. DRAEM [37] 84.4 92.9 84.0 32.3 00.7 02.2 08.7 01.1 49.0
SimpleNet [27] 93.2 96.8 90.2 98.6 67.8 65.0 82.0 48.1 85.9
RealNet [45] 79.8 89.6 84.5 51.0 22.4 03.1 19.7 01.6 54.3
Emb. CFA [21] 74.9 88.1 81.0 97.6 57.3 56.1 76.2 39.0 76.6
CFLOW-AD [14] 95.8 98.2 93.3 99.0 53.8 54.9 94.6 37.9 85.7
PyramidalFlow [22] 84.1 90.7 87.7 49.7 00.6 02.2 15.7 01.1 52.3
Rec. RD [10] 92.1 95.7 91.3 91.3 43.8 49.4 85.1 32.8 80.2
DiAD [16] 77.0 88.3 83.3 85.5 42.4 46.6 19.4 30.3 63.2
ViTAD [40] 89.1 94.9 87.5 98.2 62.5 61.5 79.9 44.4 83.0
InvAD [42] 96.1 98.3 93.2 96.5 58.4 61.1 90.6 44.0 86.3
InvAD-lite [42] 94.1 97.0 90.8 94.4 51.6 55.5 90.5 38.4 83.5
MambaAD [15] 93.8 97.0 91.7 94.0 46.7 51.7 88.2 34.9 82.2
Hybrid UniAD [35] 94.8 97.4 91.5 99.2 61.3 63.9 91.0 46.9 86.7
RD++ [31] 93.4 96.3 92.2 95.1 56.4 58.9 87.2 41.7 84.2
DesTSeg [44] 88.0 94.2 87.5 85.9 65.9 62.6 58.8 45.5 79.1
chewinggum Aug. DRAEM [37] 56.4 72.6 79.0 34.4 00.3 01.1 01.1 00.6 39.3
SimpleNet [27] 97.4 98.8 93.8 98.3 31.5 36.9 75.1 22.6 78.6
RealNet [45] 87.7 94.6 86.0 72.4 40.0 42.0 36.4 26.5 68.6
Emb. CFA [21] 76.3 88.7 80.7 95.6 33.1 36.5 70.6 22.3 70.4
CFLOW-AD [14] 97.3 98.9 95.3 99.2 60.3 60.3 87.9 43.2 87.0
PyramidalFlow [22] 36.3 63.6 80.0 68.5 01.3 01.1 39.2 00.6 43.7
Rec. RD [10] 95.8 98.0 92.6 98.7 60.4 62.6 78.7 45.5 85.3
DiAD [16] 82.5 90.9 83.5 64.0 14.7 29.3 15.7 17.2 54.4
ViTAD [40] 95.8 98.0 93.2 97.8 61.1 58.5 72.2 41.3 84.0
InvAD [42] 98.0 99.0 94.9 98.7 60.4 62.8 81.3 45.8 86.5
InvAD-lite [42] 97.5 98.9 94.8 98.3 56.0 57.6 80.6 40.4 85.1
MambaAD [15] 97.4 98.8 94.7 98.0 54.6 60.0 80.0 42.8 85.1
Hybrid UniAD [35] 97.4 98.9 96.3 99.1 60.3 58.5 85.9 41.4 86.7
RD++ [31] 97.1 98.7 94.2 98.7 58.8 61.8 79.8 44.7 85.7
DesTSeg [44] 96.3 98.6 95.8 98.0 79.2 71.9 68.9 56.1 88.2
Table A31: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
fryum Aug. DRAEM [37] 46.5 68.9 80.0 59.6 04.3 10.3 32.0 05.4 45.8
SimpleNet [27] 84.0 92.8 83.6 94.9 46.2 48.5 81.7 32.0 77.3
RealNet [45] 66.2 83.4 80.6 53.8 35.7 10.4 19.2 05.5 53.3
Emb. CFA [21] 68.7 84.0 81.0 79.4 26.1 32.5 49.8 19.4 62.4
CFLOW-AD [14] 90.1 95.5 88.3 97.5 53.6 53.3 87.4 36.3 82.1
PyramidalFlow [22] 68.7 83.5 81.0 81.7 12.7 19.0 62.7 10.5 60.9
Rec. RD [10] 95.4 97.9 92.5 96.9 47.5 51.6 93.2 34.8 83.8
DiAD [16] 83.7 92.6 84.3 97.1 55.7 58.1 69.0 40.9 77.2
ViTAD [40] 93.9 97.2 90.6 97.4 46.4 50.3 88.7 33.6 82.3
InvAD [42] 96.0 98.2 91.9 97.0 48.5 52.5 92.1 35.6 83.9
InvAD-lite [42] 95.6 98.0 90.9 96.8 47.2 51.6 92.8 34.7 83.5
MambaAD [15] 95.5 97.8 91.5 97.0 48.3 51.7 92.3 34.9 83.6
Hybrid UniAD [35] 86.0 93.4 85.7 97.2 46.3 53.1 81.9 36.2 79.0
RD++ [31] 94.8 97.8 91.9 96.9 47.3 52.7 92.5 35.8 83.6
DesTSeg [44] 92.2 96.6 88.8 68.9 48.9 43.8 49.2 28.0 72.6
macaroni1 Aug. DRAEM [37] 53.4 47.5 64.6 22.3 00.0 00.1 02.6 00.0 30.7
SimpleNet [27] 77.4 75.7 71.4 97.0 03.8 10.6 87.0 05.6 62.2
RealNet [45] 76.2 81.5 72.0 60.1 13.2 19.7 27.5 10.9 53.3
Emb. CFA [21] 57.9 61.6 68.1 73.5 00.4 02.7 58.8 01.4 48.2
CFLOW-AD [14] 80.6 77.5 75.3 99.1 07.6 13.3 94.9 07.1 65.8
PyramidalFlow [22] 41.9 43.4 66.0 81.2 00.1 00.2 39.9 00.1 40.4
Rec. RD [10] 94.8 92.8 88.9 99.8 23.8 31.1 97.4 18.4 77.6
DiAD [16] 97.9 97.2 92.1 54.1 00.0 00.1 03.3 00.0 49.3
ViTAD [40] 84.8 82.5 75.4 98.5 07.2 16.0 90.2 08.7 66.9
InvAD [42] 94.3 92.8 87.1 99.7 21.7 30.1 97.4 17.7 76.8
InvAD-lite [42] 92.9 90.4 86.0 99.7 18.6 25.6 97.2 14.7 75.0
MambaAD [15] 91.5 89.6 82.3 99.5 17.6 26.1 96.0 15.0 73.8
Hybrid UniAD [35] 84.2 81.4 77.0 99.0 08.0 16.4 94.0 08.9 67.6
RD++ [31] 93.4 91.6 86.0 99.7 24.1 34.2 96.9 20.7 77.0
DesTSeg [44] 95.2 93.3 92.7 79.7 15.7 24.4 46.2 13.9 67.6
macaroni2 Aug. DRAEM [37] 52.3 51.5 62.8 48.7 00.0 00.1 12.2 00.0 35.4
SimpleNet [27] 66.7 61.2 66.9 90.9 00.7 04.2 79.1 02.2 54.3
RealNet [45] 57.6 57.5 69.4 51.4 05.3 03.6 16.0 01.8 40.3
Emb. CFA [21] 50.9 51.1 67.6 79.0 00.2 01.6 52.5 00.8 44.9
CFLOW-AD [14] 64.3 62.7 66.2 97.2 01.5 05.5 88.2 02.8 56.2
PyramidalFlow [22] 67.8 71.0 69.2 63.9 00.0 00.1 08.3 00.0 43.7
Rec. RD [10] 87.4 83.0 82.8 99.6 12.8 22.4 97.2 12.6 71.2
DiAD [16] 50.6 42.5 63.4 56.2 00.0 00.1 12.3 00.0 32.1
ViTAD [40] 77.1 69.9 71.4 98.0 03.4 09.1 87.5 04.8 61.1
InvAD [42] 85.9 81.9 81.2 99.6 10.8 18.1 97.4 10.0 69.7
InvAD-lite [42] 86.5 83.0 80.4 99.6 10.9 17.9 97.9 09.8 69.9
MambaAD [15] 84.5 80.6 77.1 99.5 08.8 16.9 96.5 09.2 68.1
Hybrid UniAD [35] 76.5 76.6 69.4 97.3 05.0 12.4 88.0 06.6 62.4
RD++ [31] 83.7 78.6 77.8 99.5 10.3 18.7 96.3 10.3 68.1
DesTSeg [44] 70.6 70.0 71.7 69.2 11.3 19.1 44.7 10.6 53.4
pcb1 Aug. DRAEM [37] 56.1 52.2 68.4 33.6 00.3 00.9 11.4 00.5 35.2
SimpleNet [27] 92.8 93.9 87.9 98.7 80.0 74.9 78.9 59.8 87.3
RealNet [45] 76.6 79.6 71.2 70.5 36.9 41.9 25.1 26.5 59.7
Emb. CFA [21] 80.9 81.5 75.6 93.0 62.4 65.3 46.0 48.5 73.0
CFLOW-AD [14] 93.5 93.5 86.5 99.1 72.4 68.7 87.8 52.3 86.6
PyramidalFlow [22] 83.3 80.9 82.1 92.8 49.3 45.5 40.5 29.5 69.6
Rec. RD [10] 54.8 47.6 74.9 98.6 18.8 38.3 92.5 23.7 60.6
DiAD [16] 92.8 95.3 94.5 96.6 27.9 38.4 60.8 23.8 72.3
ViTAD [40] 95.3 93.7 90.7 99.4 63.2 61.1 89.0 43.9 85.7
InvAD [42] 96.5 96.1 93.1 99.8 81.3 75.9 96.1 61.2 91.8
InvAD-lite [42] 96.2 94.4 93.7 99.8 79.3 73.2 96.1 57.8 90.9
MambaAD [15] 95.7 94.5 93.3 99.8 79.0 72.9 94.3 57.4 90.5
Hybrid UniAD [35] 94.4 93.1 91.1 99.3 62.9 59.9 89.1 42.8 85.3
RD++ [31] 96.6 96.2 91.7 99.3 67.7 61.0 95.2 43.9 87.8
DesTSeg [44] 82.7 81.2 77.0 96.3 70.0 66.4 71.9 49.7 78.2
Table A32: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pcb2 Aug. DRAEM [37] 59.6 70.8 66.7 38.0 00.1 00.4 07.3 00.2 38.6
SimpleNet [27] 90.1 91.8 83.4 96.6 13.9 23.5 80.6 13.3 71.0
RealNet [45] 71.6 78.8 67.4 64.1 12.9 19.8 32.7 11.0 52.5
Emb. CFA [21] 74.3 75.9 71.5 74.7 06.5 14.9 46.5 08.1 54.8
CFLOW-AD [14] 92.4 93.0 85.8 96.6 12.5 18.7 82.3 10.3 71.5
PyramidalFlow [22] 72.8 76.3 70.8 93.9 08.8 15.9 71.9 08.7 60.5
Rec. RD [10] 97.3 96.9 94.1 97.8 22.1 30.0 91.6 17.6 78.2
DiAD [16] 97.7 98.0 92.8 83.8 03.8 09.9 50.9 05.2 62.4
ViTAD [40] 90.2 89.4 85.4 97.9 12.9 21.0 82.7 11.7 71.0
InvAD [42] 97.0 97.1 93.1 99.1 16.1 23.8 92.5 13.5 76.8
InvAD-lite [42] 94.7 94.7 89.9 98.9 12.9 23.0 91.4 13.0 74.8
MambaAD [15] 95.2 95.0 90.2 98.8 13.9 24.1 90.0 13.7 75.1
Hybrid UniAD [35] 90.8 91.5 85.4 97.9 09.0 17.3 83.6 09.4 70.6
RD++ [31] 98.3 98.5 94.9 97.9 28.6 34.6 91.1 20.9 80.1
DesTSeg [44] 90.1 89.1 83.0 93.6 20.9 29.6 66.3 17.4 70.0
pcb3 Aug. DRAEM [37] 36.1 46.1 66.4 38.5 00.2 00.5 07.9 00.3 30.7
SimpleNet [27] 86.4 88.3 78.0 97.6 19.7 27.7 80.8 16.1 70.3
RealNet [45] 78.8 84.0 72.5 71.5 29.8 36.6 35.7 22.4 60.9
Emb. CFA [21] 63.0 68.7 66.9 85.5 11.5 20.3 39.6 11.3 52.7
CFLOW-AD [14] 80.5 83.6 73.6 96.3 22.6 26.8 80.3 15.4 67.9
PyramidalFlow [22] 53.8 56.4 66.4 56.4 00.5 02.7 07.6 01.3 37.8
Rec. RD [10] 96.2 96.1 90.7 98.0 25.3 34.8 93.6 21.1 78.6
DiAD [16] 94.2 93.8 89.1 88.5 03.8 10.2 48.0 05.4 61.1
ViTAD [40] 90.4 90.6 84.6 98.2 23.5 29.6 87.9 17.4 74.2
InvAD [42] 96.6 97.0 92.1 99.2 17.3 28.7 93.1 16.8 77.4
InvAD-lite [42] 93.8 94.3 86.1 99.1 17.6 28.2 92.6 16.4 75.4
MambaAD [15] 95.0 95.2 88.8 99.1 21.7 27.6 92.0 16.0 76.5
Hybrid UniAD [35] 82.8 83.8 76.9 98.2 13.6 24.1 80.9 13.7 67.7
RD++ [31] 95.7 95.8 89.9 98.0 33.2 35.6 93.7 21.6 79.5
DesTSeg [44] 93.4 94.1 87.3 89.9 31.9 33.6 56.9 20.2 72.3
pcb4 Aug. DRAEM [37] 68.0 58.0 76.4 32.8 00.5 01.3 06.6 00.7 38.9
SimpleNet [27] 97.6 97.8 92.0 95.2 22.1 30.9 79.0 18.2 76.3
RealNet [45] 72.5 79.2 68.3 58.3 28.0 23.1 24.0 13.1 53.3
Emb. CFA [21] 72.5 75.7 69.8 75.2 10.0 19.1 44.2 10.6 54.9
CFLOW-AD [14] 98.5 98.4 96.1 96.7 21.9 30.9 85.4 18.3 78.2
PyramidalFlow [22] 48.9 52.7 66.4 89.5 05.0 08.9 66.3 04.7 49.2
Rec. RD [10] 100. 100. 99.5 97.8 30.2 36.0 89.2 22.0 81.6
DiAD [16] 99.7 99.7 99.0 95.4 09.9 18.2 83.5 10.0 72.2
ViTAD [40] 99.1 98.7 97.0 99.0 42.8 47.9 93.4 31.5 84.5
InvAD [42] 99.8 99.8 98.0 98.6 46.6 45.1 91.1 29.2 84.8
InvAD-lite [42] 99.8 99.8 98.0 98.7 50.6 49.4 90.8 32.8 85.8
MambaAD [15] 99.6 99.7 97.0 98.7 47.4 47.6 90.8 31.2 84.9
Hybrid UniAD [35] 99.2 99.2 95.5 97.1 29.1 33.7 83.1 20.3 79.4
RD++ [31] 100. 100. 99.0 97.4 32.8 38.7 86.8 24.0 81.8
DesTSeg [44] 98.8 98.7 95.1 94.3 53.4 52.2 74.3 35.3 83.0
pipe_fryum Aug. DRAEM [37] 63.5 74.5 83.1 22.9 00.7 02.5 04.9 01.2 40.7
SimpleNet [27] 80.7 90.9 83.2 98.8 70.8 66.2 71.3 49.5 80.9
RealNet [45] 64.0 83.5 80.5 61.3 43.2 31.4 32.3 18.6 59.0
Emb. CFA [21] 48.2 71.5 80.0 90.9 53.7 52.7 59.2 35.8 65.3
CFLOW-AD [14] 97.4 98.8 96.0 99.2 60.8 61.3 94.6 44.2 88.2
PyramidalFlow [22] 41.6 63.3 80.0 81.3 04.7 09.1 43.0 04.8 48.1
Rec. RD [10] 98.1 99.0 96.5 99.1 55.8 58.2 96.3 41.1 87.6
DiAD [16] 97.4 98.5 98.0 98.6 52.5 57.1 78.4 40.0 82.9
ViTAD [40] 98.1 99.1 95.5 99.5 67.0 67.2 95.0 50.6 89.9
InvAD [42] 99.2 99.6 97.5 99.4 66.3 64.7 94.8 47.9 90.0
InvAD-lite [42] 98.6 99.3 96.0 99.0 55.3 57.4 95.7 40.2 87.4
MambaAD [15] 99.0 99.5 97.0 99.0 53.9 57.5 95.6 40.3 87.5
Hybrid UniAD [35] 95.5 97.7 92.9 99.0 52.1 58.2 93.3 41.1 85.5
RD++ [31] 96.8 98.3 94.1 99.0 57.6 58.8 95.4 41.6 87.0
DesTSeg [44] 96.3 98.4 92.1 88.8 81.4 75.2 51.5 60.2 84.9
Table A33: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 55.1 62.4 72.9 37.5 00.6 01.7 10.0 00.9 38.0
SimpleNet [27] 86.4 89.1 82.8 96.6 34.0 37.8 79.2 25.7 74.0
RealNet [45] 71.4 79.5 74.7 61.0 25.7 22.6 27.4 13.5 54.7
Emb. CFA [21] 66.3 74.3 74.2 81.3 22.1 26.2 50.8 17.0 58.4
CFLOW-AD [14] 86.5 88.8 84.9 97.7 33.9 37.2 86.8 24.9 75.3
PyramidalFlow [22] 58.2 66.3 74.4 77.0 07.2 09.6 42.8 05.6 50.2
Rec. RD [10] 90.6 90.9 89.3 98.0 35.4 42.5 91.9 27.9 78.6
DiAD [16] 84.8 88.5 86.9 82.5 17.9 23.2 44.5 14.9 61.2
ViTAD [40] 90.4 91.1 86.0 98.2 36.4 41.0 85.7 27.5 77.2
InvAD [42] 95.4 95.7 91.6 98.9 43.3 46.8 93.1 32.5 82.4
InvAD-lite [42] 94.9 95.2 90.7 98.6 40.2 44.0 93.1 29.8 81.3
MambaAD [15] 94.5 94.9 90.2 98.4 39.3 43.7 92.1 29.5 80.8
Hybrid UniAD [35] 89.0 91.0 85.8 98.3 34.5 39.6 86.5 26.4 76.7
RD++ [31] 93.9 94.7 90.2 98.4 42.3 46.3 91.9 31.2 81.3
DesTSeg [44] 89.9 91.4 86.7 86.7 46.6 47.2 61.1 32.7 74.9
Appendix G Detailed Quantitative Results on BTAD Dataset Under 100 epochs
Table A34: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
01 Aug. DRAEM [37] 82.6 93.6 84.7 47.0 03.0 06.3 14.3 03.3 52.3
SimpleNet [27] 98.3 99.4 97.9 95.0 41.6 45.1 64.7 29.1 80.1
RealNet [45] 100. 100. 100. 95.7 69.5 67.6 81.3 51.1 89.3
Emb. CFA [21] 98.2 99.4 99.0 94.7 37.8 44.2 64.6 28.4 79.6
PatchCore [29] 99.6 99.8 99.0 97.1 55.5 56.6 75.8 39.5 85.4
CFLOW-AD [14] 94.8 98.1 93.5 94.7 36.6 43.4 66.2 27.7 77.8
PyramidalFlow [22] 94.5 96.7 98.0 89.2 18.9 06.3 66.6 03.3 70.8
Rec. RD [10] 99.8 99.9 99.0 97.8 57.5 58.5 84.4 41.4 87.1
DiAD [16] 99.4 99.8 99.0 90.5 24.3 31.4 73.7 18.6 74.0
ViTAD [40] 97.9 99.2 95.9 96.5 55.0 55.1 69.8 38.1 83.4
InvAD [42] 99.9 100. 99.0 98.0 63.7 62.9 84.1 45.9 88.4
InvAD-lite [42] 97.8 99.2 96.9 97.3 59.3 61.4 77.3 44.3 85.9
MambaAD [15] 97.4 99.0 94.7 97.1 55.3 58.1 76.5 41.0 84.4
Hybrid UniAD [35] 99.7 99.9 98.0 97.4 57.2 58.4 79.2 41.3 86.1
RD++ [31] 98.6 99.5 97.9 97.5 53.6 57.8 80.9 40.6 85.6
DesTSeg [44] 98.1 99.3 96.9 85.8 25.4 39.3 68.2 24.4 76.4
02 Aug. DRAEM [37] 70.6 94.7 92.8 51.7 07.9 11.5 18.2 06.1 54.2
SimpleNet [27] 81.7 96.8 93.6 94.5 46.3 49.4 47.4 32.8 75.1
RealNet [45] 72.9 95.3 92.8 83.9 38.8 50.3 38.6 33.6 70.0
Emb. CFA [21] 80.3 96.6 92.8 94.7 51.2 53.9 48.8 36.9 76.0
PatchCore [29] 83.8 97.3 92.8 96.0 59.3 59.8 54.1 42.7 79.3
CFLOW-AD [14] 79.6 96.7 92.8 96.4 61.4 61.0 54.5 43.9 79.0
PyramidalFlow [22] 73.7 95.6 92.8 86.8 64.1 62.5 38.0 45.4 75.1
Rec. RD [10] 84.0 97.3 93.4 96.7 64.5 64.3 58.8 47.4 81.3
DiAD [16] 72.6 95.4 92.8 86.9 27.3 35.5 44.8 21.6 65.0
ViTAD [40] 83.5 97.2 93.0 96.9 74.1 67.8 51.9 51.2 82.0
InvAD [42] 88.7 98.1 93.5 96.6 62.0 61.9 59.6 44.9 81.7
InvAD-lite [42] 81.8 97.0 93.2 96.6 61.8 62.4 60.2 45.4 80.5
MambaAD [15] 82.5 97.2 93.0 96.3 55.1 60.0 59.6 42.8 79.3
Hybrid UniAD [35] 85.0 97.6 92.8 94.8 41.8 51.8 59.4 35.0 76.9
RD++ [31] 85.5 97.6 93.2 96.7 65.2 64.2 57.2 47.2 81.5
DesTSeg [44] 81.2 96.9 92.8 94.6 45.9 58.4 50.9 41.2 76.4
03 Aug. DRAEM [37] 60.7 47.1 56.5 48.4 00.3 01.5 16.0 00.8 35.7
SimpleNet [27] 99.7 95.7 88.6 99.4 36.5 38.5 97.5 23.8 81.3
RealNet [45] 96.3 90.5 85.7 72.2 36.0 40.1 40.3 25.1 69.0
Emb. CFA [21] 99.7 96.6 88.9 99.4 53.3 52.5 95.2 35.6 85.1
PatchCore [29] 99.8 97.3 92.1 99.5 50.1 48.2 98.0 31.8 85.2
CFLOW-AD [14] 99.0 89.5 78.6 99.4 38.8 46.0 97.3 29.9 79.7
PyramidalFlow [22] 92.9 57.1 52.2 96.7 05.9 12.0 87.7 06.4 59.0
Rec. RD [10] 99.4 92.5 89.6 99.8 56.8 54.6 98.9 37.6 85.7
DiAD [16] 98.3 69.9 86.1 97.6 07.2 13.1 92.7 07.0 66.4
ViTAD [40] 99.6 94.2 90.9 99.6 48.5 47.2 98.0 30.9 84.1
InvAD [42] 99.7 95.2 90.3 99.8 62.0 59.4 98.9 42.2 87.5
InvAD-lite [42] 99.7 96.0 93.8 99.6 56.5 55.3 98.4 38.2 87.0
MambaAD [15] 99.5 92.2 89.9 99.6 47.1 49.9 98.4 33.2 83.8
Hybrid UniAD [35] 99.8 97.5 91.8 99.5 51.9 51.3 97.9 34.5 85.8
RD++ [31] 99.7 96.4 91.2 99.8 60.1 57.6 99.0 40.4 87.4
DesTSeg [44] 99.2 91.4 87.1 96.3 33.1 35.4 90.8 21.5 78.2
Avg Aug. DRAEM [37] 71.3 78.5 78.0 49.0 03.7 06.5 16.2 03.4 47.4
SimpleNet [27] 93.2 97.3 93.3 96.3 41.5 44.3 69.8 28.6 78.8
RealNet [45] 89.7 95.3 92.8 84.0 48.1 52.7 53.4 36.6 76.1
Emb. CFA [21] 92.7 97.5 93.5 96.3 47.4 50.2 69.5 33.6 80.2
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 83.3
CFLOW-AD [14] 91.2 94.8 88.3 96.8 45.6 50.1 72.7 33.8 78.9
PyramidalFlow [22] 87.0 83.1 81.0 90.9 29.6 26.9 64.1 18.3 68.3
Rec. RD [10] 94.4 96.6 94.0 98.1 59.6 59.2 80.7 42.1 84.7
DiAD [16] 90.1 88.4 92.6 91.7 19.6 26.7 70.4 15.7 68.5
ViTAD [40] 93.6 96.8 93.3 97.6 59.2 56.7 73.2 40.1 83.1
InvAD [42] 96.1 97.8 94.3 98.1 62.6 61.4 80.9 44.3 85.9
InvAD-lite [42] 93.1 97.4 94.6 97.9 59.2 59.7 78.6 42.6 84.4
MambaAD [15] 93.1 96.1 92.5 97.7 52.5 56.0 78.2 39.0 82.5
Hybrid UniAD [35] 94.8 98.3 94.2 97.2 50.3 53.8 78.8 36.9 82.9
RD++ [31] 94.6 97.8 94.1 98.0 59.6 59.8 79.0 42.8 84.8
DesTSeg [44] 92.8 95.9 92.3 92.2 34.8 44.3 70.0 29.0 77.0
Appendix H Detailed Quantitative Results on MPDD Dataset Under 100 epochs
Table A35: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bracket_black Aug. DRAEM [37] 39.0 51.7 74.6 45.1 00.1 00.4 24.0 00.2 36.2
SimpleNet [27] 74.0 83.3 78.5 92.8 03.1 09.7 86.1 05.1 63.3
RealNet [45] 63.0 72.7 75.2 65.9 02.3 06.8 50.7 03.5 50.9
Emb. CFA [21] 58.4 72.8 74.6 74.6 00.2 00.5 31.9 00.2 47.7
PatchCore [29] 88.4 92.7 83.9 98.5 11.0 21.0 96.7 11.8 72.6
CFLOW-AD [14] 60.8 70.3 75.8 95.5 01.7 04.1 85.2 02.1 57.8
PyramidalFlow [22] 65.8 72.7 77.0 95.3 01.8 05.0 81.4 02.6 58.9
Rec. RD [10] 84.2 88.3 83.1 97.7 06.6 12.7 94.1 06.8 69.0
DiAD [16] 70.9 81.7 79.0 89.8 00.4 01.0 65.3 00.5 55.5
ViTAD [40] 78.3 86.2 83.6 95.5 02.5 08.0 89.1 04.2 65.7
InvAD [42] 83.4 90.9 81.0 97.4 07.1 15.2 94.0 08.2 69.3
InvAD-lite [42] 75.4 84.6 78.6 95.2 07.9 16.8 89.6 09.1 66.0
MambaAD [15] 81.9 88.3 81.7 94.5 05.3 11.2 88.4 05.9 66.9
Hybrid UniAD [35] 78.1 84.9 79.2 93.0 00.6 01.6 79.2 00.8 62.2
RD++ [31] 82.7 86.6 82.9 98.0 04.2 10.1 94.1 05.3 67.8
DesTSeg [44] 83.4 87.3 85.7 72.4 05.7 13.5 59.0 07.3 61.6
bracket_brown Aug. DRAEM [37] 42.5 61.1 79.7 46.8 00.8 01.9 27.5 01.0 40.2
SimpleNet [27] 90.4 93.8 93.6 94.4 07.9 16.9 86.7 09.2 72.0
RealNet [45] 91.0 94.0 92.6 69.9 12.1 19.1 50.5 10.6 65.2
Emb. CFA [21] 88.7 94.9 88.9 85.1 03.2 06.9 52.6 03.6 63.9
PatchCore [29] 95.6 97.2 95.3 98.6 26.2 32.0 95.4 19.0 79.5
CFLOW-AD [14] 74.4 77.6 85.2 95.9 05.9 12.0 91.1 06.4 65.1
PyramidalFlow [22] 63.6 76.4 81.0 94.0 05.7 12.5 69.6 06.7 59.6
Rec. RD [10] 86.2 89.6 92.7 96.0 13.1 23.1 93.9 13.1 73.0
DiAD [16] 53.2 69.1 82.9 90.0 02.8 07.0 66.6 03.6 53.1
ViTAD [40] 86.5 86.0 92.6 98.0 16.3 24.9 95.0 14.2 73.5
InvAD [42] 93.7 95.5 95.2 95.5 15.2 22.8 90.8 12.8 75.4
InvAD-lite [42] 89.5 93.0 90.1 97.9 14.9 25.0 94.3 14.3 74.4
MambaAD [15] 95.7 97.5 95.3 97.5 26.3 30.8 92.8 18.2 79.0
Hybrid UniAD [35] 91.4 95.0 92.7 98.0 23.9 32.0 90.7 19.0 77.1
RD++ [31] 84.4 89.9 91.9 96.6 17.2 24.4 94.9 13.9 73.5
DesTSeg [44] 95.0 92.5 96.2 61.6 04.8 10.5 30.1 05.5 60.7
bracket_white Aug. DRAEM [37] 34.0 44.3 66.7 42.3 00.0 00.1 09.0 00.0 30.6
SimpleNet [27] 85.6 89.3 80.0 97.7 02.1 06.5 85.6 03.4 66.5
RealNet [45] 77.3 86.2 80.8 85.2 18.4 30.6 47.4 18.0 63.4
Emb. CFA [21] 80.2 86.0 77.8 89.8 03.9 14.9 47.3 08.0 60.2
PatchCore [29] 93.3 94.8 88.9 98.8 10.0 19.3 93.9 10.7 73.9
CFLOW-AD [14] 62.9 64.1 69.8 98.2 01.2 03.5 93.0 01.8 57.3
PyramidalFlow [22] 76.8 84.8 78.4 98.9 05.6 14.5 91.1 07.8 66.3
Rec. RD [10] 85.3 86.5 81.4 99.0 02.3 06.4 95.2 03.3 67.6
DiAD [16] 84.0 90.7 88.9 96.5 01.0 03.6 82.5 01.8 63.9
ViTAD [40] 78.9 84.0 78.7 95.7 00.9 02.9 86.7 01.5 63.5
InvAD [42] 92.1 94.7 90.0 98.9 09.8 21.9 94.1 12.3 74.2
InvAD-lite [42] 90.2 92.2 84.7 99.2 05.2 10.6 94.8 05.6 70.7
MambaAD [15] 97.1 97.5 90.9 99.3 12.2 20.4 95.0 11.3 75.9
Hybrid UniAD [35] 63.6 67.1 69.6 95.0 00.3 00.8 79.6 00.4 55.3
RD++ [31] 81.4 86.0 76.4 98.7 02.9 10.0 93.4 05.2 66.3
DesTSeg [44] 91.1 89.8 88.5 90.4 01.0 02.7 51.2 01.4 63.1
connector Aug. DRAEM [37] 21.7 22.2 48.3 35.1 00.6 02.1 16.8 01.1 22.2
SimpleNet [27] 96.2 92.0 90.3 98.3 46.4 47.0 94.5 30.7 82.2
RealNet [45] 99.0 97.9 96.6 85.8 38.9 44.1 75.8 28.3 79.5
Emb. CFA [21] 96.9 89.0 93.3 83.5 11.2 16.3 59.2 08.9 67.8
PatchCore [29] 100. 100. 100. 99.5 71.2 63.3 98.4 46.3 91.5
CFLOW-AD [14] 91.2 85.3 78.6 96.7 27.9 29.3 89.4 17.2 72.9
PyramidalFlow [22] 69.0 42.7 63.4 90.7 03.7 07.2 70.0 03.7 50.6
Rec. RD [10] 100. 100. 100. 99.2 60.6 55.5 97.5 38.4 89.1
DiAD [16] 57.1 48.0 56.3 90.1 02.9 06.5 68.4 03.4 47.1
ViTAD [40] 96.2 91.9 86.7 99.3 44.3 47.6 97.7 31.2 81.9
InvAD [42] 96.7 83.4 96.6 98.9 57.9 58.1 96.4 41.0 85.0
InvAD-lite [42] 97.1 90.8 93.3 99.2 62.2 58.1 97.3 41.0 86.5
MambaAD [15] 99.5 99.1 96.3 99.2 54.5 56.5 97.3 39.3 87.6
Hybrid UniAD [35] 69.8 49.9 58.5 95.2 07.6 14.9 83.9 08.1 54.9
RD++ [31] 100. 100. 100. 99.2 67.0 61.8 97.4 44.8 90.7
DesTSeg [44] 92.1 81.4 82.4 76.6 33.8 37.6 61.1 23.1 68.8
Table A36: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
metal_plate Aug. DRAEM [37] 41.9 76.3 84.5 36.3 10.8 21.2 13.3 11.9 44.0
SimpleNet [27] 100. 100. 100. 98.4 89.8 81.7 89.2 69.1 94.9
RealNet [45] 99.9 100. 99.3 97.7 90.5 84.1 93.6 72.6 95.6
Emb. CFA [21] 99.7 99.9 98.6 97.6 89.8 81.8 83.1 69.2 93.7
PatchCore [29] 100. 100. 100. 99.0 93.5 86.1 94.1 75.7 96.6
CFLOW-AD [14] 99.9 100. 99.3 98.5 89.0 83.3 92.4 71.4 95.3
PyramidalFlow [22] 96.7 98.7 95.2 88.7 61.3 21.1 60.8 11.8 77.4
Rec. RD [10] 100. 100. 100. 99.1 92.9 87.7 96.7 78.1 97.0
DiAD [16] 89.3 96.4 89.9 88.4 54.6 52.8 55.5 35.9 75.3
ViTAD [40] 100. 100. 100. 99.3 95.5 88.0 94.9 78.6 97.2
InvAD [42] 100. 100. 100. 99.0 92.2 87.2 96.0 77.2 96.8
InvAD-lite [42] 99.8 99.9 98.6 98.4 86.4 82.9 94.2 70.8 95.0
MambaAD [15] 99.8 99.9 98.6 98.2 83.8 82.2 94.3 69.8 94.5
Hybrid UniAD [35] 47.5 74.3 85.0 90.7 41.2 56.9 73.6 39.8 67.3
RD++ [31] 99.8 99.9 98.6 99.1 92.1 87.2 95.7 77.3 96.5
DesTSeg [44] 100. 100. 100. 95.6 91.4 87.0 88.8 77.0 95.3
tubes Aug. DRAEM [37] 34.8 64.0 81.2 47.0 01.1 02.0 23.2 01.0 39.2
SimpleNet [27] 84.0 93.9 85.0 97.6 42.8 45.8 91.7 29.7 78.6
RealNet [45] 80.1 90.1 85.2 95.3 54.3 53.0 90.4 36.1 79.2
Emb. CFA [21] 65.4 83.7 81.2 78.8 09.2 17.4 47.1 09.5 57.4
PatchCore [29] 90.4 96.6 92.4 98.8 64.5 63.8 95.5 46.9 86.9
CFLOW-AD [14] 65.0 83.3 81.7 96.2 32.0 35.6 86.0 21.7 69.6
PyramidalFlow [22] 69.8 86.6 81.2 97.1 48.2 46.4 90.3 30.2 74.8
Rec. RD [10] 91.9 97.1 93.3 99.0 67.1 65.6 96.0 48.8 88.0
DiAD [16] 55.4 81.1 83.6 87.7 03.9 07.6 58.2 04.0 53.9
ViTAD [40] 86.6 94.8 86.6 98.5 51.5 53.3 93.5 36.4 81.8
InvAD [42] 96.4 98.6 95.5 99.3 72.3 69.0 97.4 52.7 90.7
InvAD-lite [42] 93.0 96.9 91.3 98.3 61.4 62.0 94.0 44.9 86.3
MambaAD [15] 58.4 77.0 81.7 96.1 19.6 27.8 85.8 16.1 64.8
Hybrid UniAD [35] 72.5 84.8 83.1 91.4 08.6 13.5 71.1 07.2 63.1
RD++ [31] 93.0 97.5 93.0 99.3 74.6 70.8 97.2 54.8 90.0
DesTSeg [44] 86.4 93.8 88.4 95.6 59.2 56.1 89.4 39.0 82.3
Avg Aug. DRAEM [37] 35.6 53.3 72.5 42.1 02.2 04.6 19.0 02.5 35.4
SimpleNet [27] 88.4 92.0 87.9 96.5 32.0 34.6 89.0 24.5 76.2
RealNet [45] 85.1 90.2 88.3 83.3 36.1 39.6 68.1 28.2 72.3
Emb. CFA [21] 81.6 87.7 85.7 84.9 19.6 22.9 53.5 16.6 65.1
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 83.5
CFLOW-AD [14] 75.7 80.1 81.7 96.8 26.3 28.0 89.5 20.1 69.7
PyramidalFlow [22] 73.6 77.0 79.4 94.1 21.1 17.8 77.2 10.4 64.6
Rec. RD [10] 91.3 93.6 91.8 98.3 40.4 41.8 95.5 31.4 80.6
DiAD [16] 68.3 77.9 80.1 90.4 10.9 13.1 66.1 08.2 58.1
ViTAD [40] 87.8 90.5 88.0 97.7 35.2 37.4 92.8 27.7 77.3
InvAD [42] 93.7 93.9 93.0 98.2 42.4 45.7 94.8 34.0 81.9
InvAD-lite [42] 90.9 92.9 89.5 98.0 39.7 42.6 94.0 30.9 79.8
MambaAD [15] 88.7 93.2 90.8 97.5 33.6 38.1 92.3 26.8 78.1
Hybrid UniAD [35] 70.5 76.0 78.0 93.9 13.7 19.9 79.7 12.5 63.3
RD++ [31] 90.2 93.3 90.5 98.5 43.0 44.1 95.5 33.6 80.8
DesTSeg [44] 91.3 90.8 90.2 82.0 32.6 34.6 63.3 25.6 71.9
Appendix I Detailed Quantitative Results on MAD_Real Dataset Under 100 epochs
Table A37: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Bear Aug. DRAEM [37] 43.3 83.5 92.3 54.0 00.2 01.1 34.1 00.5 47.7
SimpleNet [27] 24.2 72.7 90.6 94.9 03.0 10.4 87.2 05.5 55.7
Emb. CFA [21] 15.0 69.1 90.6 99.1 04.2 10.3 96.5 05.5 55.4
PatchCore [29] 84.2 96.8 91.7 99.6 17.6 29.5 98.6 17.3 76.1
CFLOW-AD [14] 76.7 94.4 92.3 99.4 06.0 12.3 97.6 06.6 70.8
PyramidalFlow [22] 46.7 78.6 90.6 98.2 02.3 04.9 93.9 02.5 60.9
Rec. RD [10] 76.7 94.6 94.1 97.7 05.2 07.9 93.2 04.1 69.7
DiAD [16] 52.5 88.7 90.6 97.4 05.1 13.5 90.3 07.2 62.6
ViTAD [40] 35.0 76.8 90.6 97.3 01.9 06.9 91.0 03.6 58.4
InvAD [42] 95.8 99.1 98.0 99.7 39.9 44.1 98.9 28.3 84.1
InvAD-lite [42] 82.5 96.1 93.9 98.4 05.0 11.6 95.1 06.2 71.7
MambaAD [15] 32.5 75.1 90.6 93.3 01.0 04.0 80.3 02.0 55.4
Hybrid UniAD [35] 83.3 96.5 90.6 99.0 04.0 09.6 96.3 05.0 71.2
RD++ [31] 77.5 94.4 93.9 99.0 08.5 15.3 96.8 08.3 71.8
DesTSeg [44] 26.7 76.9 90.6 80.3 00.4 01.3 56.6 00.6 49.7
Bird Aug. DRAEM [37] 60.9 89.6 89.8 57.7 00.7 01.0 22.1 00.5 50.2
SimpleNet [27] 58.2 87.8 91.7 82.7 02.4 06.2 53.7 03.2 57.7
Emb. CFA [21] 55.5 83.1 89.8 86.6 04.6 10.9 56.9 05.7 57.9
PatchCore [29] 71.8 92.6 93.6 96.4 32.2 39.7 84.2 24.8 74.6
CFLOW-AD [14] 35.5 74.4 89.8 89.3 02.7 06.3 64.4 03.3 53.6
PyramidalFlow [22] 17.3 69.0 89.8 91.2 02.9 06.4 66.7 03.3 50.2
Rec. RD [10] 51.8 85.9 89.8 87.1 07.7 16.6 57.2 09.1 59.0
DiAD [16] 09.1 70.5 89.8 93.1 04.3 10.4 71.1 05.5 49.8
ViTAD [40] 67.3 89.7 93.6 89.6 09.8 17.6 62.6 09.7 64.2
InvAD [42] 84.5 96.4 93.3 89.0 10.1 20.0 66.1 11.1 68.8
InvAD-lite [42] 53.6 78.6 91.7 92.8 06.2 13.4 73.8 07.2 60.6
MambaAD [15] 86.4 97.1 91.3 95.1 17.8 28.5 77.9 16.6 73.2
Hybrid UniAD [35] 38.2 74.3 89.8 87.1 02.4 06.5 63.3 03.4 53.6
RD++ [31] 45.5 82.2 89.8 89.6 06.8 15.4 63.2 08.4 58.1
DesTSeg [44] 60.0 82.7 93.6 66.2 05.4 13.8 35.4 07.4 54.5
Elephant Aug. DRAEM [37] 77.8 93.9 90.0 50.1 00.4 01.2 18.7 00.6 52.4
SimpleNet [27] 20.0 68.9 87.8 76.4 06.1 14.5 42.9 07.8 46.9
Emb. CFA [21] 35.6 71.5 87.8 86.8 05.7 10.6 58.4 05.6 52.7
PatchCore [29] 27.8 69.5 87.8 93.7 11.1 18.3 78.8 10.1 56.1
CFLOW-AD [14] 23.3 71.6 87.8 90.6 08.6 11.5 66.4 06.1 52.6
PyramidalFlow [22] 51.1 84.2 87.8 93.0 02.8 06.0 76.5 03.1 59.5
Rec. RD [10] 43.3 82.7 87.8 88.2 02.1 05.4 65.6 02.8 55.8
DiAD [16] 36.7 76.9 90.0 91.8 02.0 04.3 74.1 02.2 53.7
ViTAD [40] 46.7 77.4 87.8 96.1 09.1 18.0 85.6 09.9 61.4
InvAD [42] 32.2 72.6 87.8 91.3 06.5 13.2 71.4 07.0 54.9
InvAD-lite [42] 36.7 72.7 87.8 91.5 04.7 10.1 75.2 05.3 55.6
MambaAD [15] 47.8 79.2 87.8 90.7 04.7 09.2 72.5 04.8 57.9
Hybrid UniAD [35] 11.1 63.1 87.8 79.1 01.3 04.0 38.7 02.0 42.4
RD++ [31] 41.1 80.3 87.8 85.2 01.7 04.1 59.6 02.1 53.7
DesTSeg [44] 58.9 83.6 90.0 66.6 01.2 05.0 32.6 02.6 51.9
Parrot Aug. DRAEM [37] 64.3 87.0 92.0 49.2 00.6 01.3 14.6 00.6 48.8
SimpleNet [27] 16.5 71.6 90.2 61.9 00.9 01.8 32.7 00.9 41.9
Emb. CFA [21] 04.3 65.9 90.2 42.3 00.5 01.3 17.0 00.7 34.4
PatchCore [29] 05.2 66.2 90.2 81.7 02.1 05.1 64.0 02.6 46.0
CFLOW-AD [14] 16.5 68.0 90.2 66.7 01.0 02.5 41.9 01.3 43.1
PyramidalFlow [22] 25.2 77.3 90.2 64.2 00.9 02.2 31.8 01.1 44.5
Rec. RD [10] 01.7 64.6 90.2 64.3 00.9 02.0 34.3 01.0 38.8
DiAD [16] 60.9 91.6 90.2 63.6 00.9 02.0 41.7 01.0 50.1
ViTAD [40] 16.5 69.8 90.2 69.8 01.1 02.6 40.8 01.3 43.7
InvAD [42] 11.3 68.6 90.2 52.1 00.7 01.4 32.7 00.7 39.2
InvAD-lite [42] 04.3 65.4 90.2 60.4 00.8 01.8 38.4 00.9 39.3
MambaAD [15] 06.1 66.6 90.2 72.4 01.1 02.6 45.6 01.3 42.4
Hybrid UniAD [35] 33.9 76.4 90.2 61.2 00.8 01.8 37.7 00.9 46.1
RD++ [31] 07.0 66.2 90.2 63.1 00.9 02.0 33.5 01.0 39.7
DesTSeg [44] 03.5 65.1 90.2 48.7 00.6 01.3 18.5 00.6 35.1
Table A38: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Pig Aug. DRAEM [37] 67.1 89.5 87.2 52.2 00.5 01.2 23.8 00.6 50.3
SimpleNet [27] 49.4 76.5 87.2 93.9 10.1 21.9 72.9 12.3 60.4
Emb. CFA [21] 55.3 77.6 89.5 93.0 05.0 11.2 73.0 05.9 59.8
PatchCore [29] 56.5 78.8 89.5 98.5 19.9 28.7 89.7 16.7 67.1
CFLOW-AD [14] 57.6 78.8 87.2 96.5 05.8 12.0 85.2 06.4 62.2
PyramidalFlow [22] 57.6 84.2 87.2 95.7 05.2 10.0 81.9 05.3 62.3
Rec. RD [10] 60.0 80.2 89.5 98.4 21.1 31.0 92.5 18.3 68.7
DiAD [16] 56.5 87.9 87.2 92.9 02.7 05.9 71.7 03.0 57.8
ViTAD [40] 37.6 71.6 87.2 97.7 08.6 17.2 89.8 09.4 59.4
InvAD [42] 58.8 77.1 89.5 98.1 11.7 22.1 91.1 12.4 65.4
InvAD-lite [42] 47.1 73.7 89.5 98.4 13.0 24.1 92.9 13.7 63.6
MambaAD [15] 61.2 80.3 87.2 97.7 12.2 22.8 87.1 12.8 65.6
Hybrid UniAD [35] 50.6 83.9 87.2 88.8 01.6 04.2 62.2 02.1 56.5
RD++ [31] 61.2 80.3 91.4 98.0 14.3 23.0 91.0 13.0 67.1
DesTSeg [44] 48.2 74.8 89.5 71.1 02.3 06.0 43.6 03.1 50.8
Puppy Aug. DRAEM [37] 53.0 82.9 88.9 46.3 00.4 01.9 23.9 01.0 46.5
SimpleNet [27] 68.0 90.6 90.9 94.5 04.1 10.3 72.4 05.4 64.2
Emb. CFA [21] 83.0 95.3 93.0 98.6 07.1 14.5 92.3 07.8 71.8
PatchCore [29] 100. 100. 100. 99.2 13.4 25.1 92.1 14.3 78.7
CFLOW-AD [14] 79.0 94.3 93.0 99.2 27.9 33.3 94.2 20.0 76.2
PyramidalFlow [22] 56.0 85.5 88.9 95.1 07.6 16.9 75.0 09.3 62.7
Rec. RD [10] 78.0 94.4 90.0 95.8 04.4 09.1 87.2 04.7 68.3
DiAD [16] 76.0 94.7 88.9 94.6 01.8 05.0 78.3 02.6 62.8
ViTAD [40] 58.0 87.6 88.9 92.5 01.2 02.9 75.8 01.5 60.6
InvAD [42] 100. 100. 100. 99.8 45.6 44.9 98.1 28.9 86.0
InvAD-lite [42] 96.0 99.0 97.6 96.9 05.9 11.9 90.3 06.3 74.4
MambaAD [15] 82.0 94.7 92.3 88.1 01.8 05.3 71.1 02.7 65.6
Hybrid UniAD [35] 74.0 92.9 90.9 98.6 14.1 21.9 91.0 12.3 71.2
RD++ [31] 66.0 87.6 93.0 98.1 04.8 11.2 92.4 05.9 66.9
DesTSeg [44] 49.0 81.8 88.9 97.6 11.2 11.6 87.7 06.2 62.6
Scorpion Aug. DRAEM [37] 33.0 80.7 90.2 46.8 01.4 02.9 14.6 01.5 42.2
SimpleNet [27] 43.5 84.0 90.2 70.9 06.0 13.0 33.4 06.9 51.7
Emb. CFA [21] 30.4 81.6 90.2 60.9 05.6 13.4 26.4 07.2 47.0
PatchCore [29] 13.0 70.7 90.2 88.2 08.8 17.0 61.6 09.3 50.9
CFLOW-AD [14] 64.3 90.1 90.2 79.1 04.0 07.8 35.9 04.1 56.6
PyramidalFlow [22] 47.8 85.3 90.2 68.0 02.7 05.9 22.1 03.0 49.6
Rec. RD [10] 64.3 91.8 90.2 64.7 02.0 04.8 20.3 02.5 52.5
DiAD [16] 56.5 79.2 93.9 73.2 02.9 07.0 30.8 03.6 49.1
ViTAD [40] 46.1 84.4 90.2 62.9 01.9 04.2 17.3 02.2 47.6
InvAD [42] 45.2 86.6 90.2 68.4 02.3 05.0 22.0 02.6 49.2
InvAD-lite [42] 50.4 83.8 90.2 67.4 02.2 04.9 18.6 02.5 49.0
MambaAD [15] 53.0 84.0 90.2 66.6 02.1 04.8 17.0 02.4 49.2
Hybrid UniAD [35] 32.2 75.9 90.2 81.2 04.8 10.1 44.9 05.3 50.7
RD++ [31] 22.6 73.5 90.2 70.9 02.5 05.8 26.5 03.0 44.3
DesTSeg [44] 31.3 81.2 90.2 48.6 01.3 02.9 11.1 01.5 41.8
Turtle Aug. DRAEM [37] 73.3 93.3 89.4 47.9 01.3 04.0 31.0 02.0 53.2
SimpleNet [27] 95.2 99.0 95.2 98.2 15.2 24.7 93.1 14.1 77.1
Emb. CFA [21] 100. 100. 100. 99.3 41.9 43.0 96.8 27.4 85.1
PatchCore [29] 100. 100. 100. 99.5 43.1 48.4 97.9 31.9 86.1
CFLOW-AD [14] 99.0 99.8 97.7 99.1 30.0 34.8 96.7 21.1 82.0
PyramidalFlow [22] 81.9 95.6 90.5 97.7 15.1 24.5 92.8 13.9 73.4
Rec. RD [10] 90.5 98.0 92.3 98.4 22.8 29.0 94.4 17.0 77.4
DiAD [16] 96.2 99.1 95.5 93.2 07.0 13.8 82.6 07.4 69.6
ViTAD [40] 83.8 96.0 91.3 95.6 06.3 12.8 85.6 06.9 70.2
InvAD [42] 100. 100. 100. 99.7 53.8 55.4 98.7 38.3 88.5
InvAD-lite [42] 96.2 99.2 95.2 97.9 17.7 25.3 94.8 14.5 77.9
MambaAD [15] 87.6 97.1 91.3 96.3 13.7 20.1 90.2 11.2 73.5
Hybrid UniAD [35] 94.3 98.7 95.5 97.8 18.7 28.5 89.6 16.6 77.4
RD++ [31] 100. 100. 100. 99.2 24.3 38.2 96.7 23.6 82.3
DesTSeg [44] 67.6 90.7 89.4 95.1 20.6 26.3 86.2 15.2 69.8
Table A39: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Unicorn Aug. DRAEM [37] 72.4 91.3 91.3 47.5 00.2 00.5 14.5 00.2 50.3
SimpleNet [27] 17.1 70.1 89.4 61.4 00.4 01.5 23.8 00.7 40.3
Emb. CFA [21] 34.3 76.8 89.4 68.2 00.8 03.5 33.8 01.8 46.7
PatchCore [29] 43.8 74.1 93.3 97.3 06.8 14.1 90.3 07.6 61.3
CFLOW-AD [14] 72.4 93.2 91.3 90.3 01.4 03.0 66.6 01.5 63.0
PyramidalFlow [22] 46.7 77.8 89.4 83.6 00.9 02.3 52.2 01.2 53.0
Rec. RD [10] 40.0 76.4 89.4 77.0 00.6 01.4 39.3 00.7 49.1
DiAD [16] 42.9 85.6 89.4 83.4 00.8 02.0 56.4 01.0 51.5
ViTAD [40] 20.0 68.6 89.4 74.7 00.5 01.1 32.1 00.6 43.2
InvAD [42] 37.1 76.5 89.4 85.3 01.0 02.5 54.8 01.3 51.8
InvAD-lite [42] 40.0 75.4 89.4 79.0 00.9 02.3 51.5 01.2 50.8
MambaAD [15] 18.1 70.9 89.4 79.1 00.9 03.1 47.0 01.6 46.0
Hybrid UniAD [35] 21.9 68.2 89.4 91.8 01.7 04.1 72.6 02.1 51.2
RD++ [31] 37.1 77.2 89.4 83.1 00.8 01.8 48.4 00.9 50.7
DesTSeg [44] 33.3 74.0 89.4 59.0 00.4 01.3 29.3 00.7 44.0
Whale Aug. DRAEM [37] 46.2 89.5 92.8 42.6 00.3 01.5 25.6 00.7 46.8
SimpleNet [27] 86.2 97.4 94.1 96.4 05.6 11.7 86.4 06.2 71.3
Emb. CFA [21] 80.0 96.2 93.5 98.6 29.5 32.8 90.5 19.6 76.4
PatchCore [29] 95.0 99.1 98.5 99.7 46.8 47.4 97.8 31.0 85.2
CFLOW-AD [14] 73.1 95.3 92.8 99.1 19.6 30.1 95.0 17.7 74.0
PyramidalFlow [22] 90.6 98.1 98.5 97.4 09.3 15.4 88.7 08.3 74.2
Rec. RD [10] 48.8 89.5 92.8 98.7 15.6 21.0 93.7 11.7 67.1
DiAD [16] 11.9 76.2 92.8 94.0 01.7 04.4 76.4 02.2 51.0
ViTAD [40] 58.8 91.4 92.8 96.8 04.1 09.8 85.2 05.2 65.0
InvAD [42] 100. 100. 100. 99.8 49.4 53.9 98.6 36.9 87.7
InvAD-lite [42] 96.9 99.5 97.0 99.6 34.2 37.8 98.5 23.3 82.7
MambaAD [15] 86.9 97.9 92.8 99.1 19.3 28.1 96.5 16.4 76.6
Hybrid UniAD [35] 36.2 83.4 92.8 97.8 08.2 14.9 89.7 08.0 61.7
RD++ [31] 88.8 97.8 97.0 99.3 29.9 32.3 96.0 19.2 79.5
DesTSeg [44] 75.6 92.1 95.4 94.1 06.3 12.9 81.9 06.9 68.2
Avg Aug. DRAEM [37] 59.1 88.1 90.4 49.4 00.6 01.7 22.3 00.8 48.9
SimpleNet [27] 47.8 81.9 90.7 83.1 05.4 11.6 59.8 06.3 56.7
Emb. CFA [21] 49.3 81.7 91.4 83.3 10.5 15.2 64.2 08.7 58.7
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 68.2
CFLOW-AD [14] 59.8 86.0 91.2 90.9 10.7 15.4 74.4 08.8 63.4
PyramidalFlow [22] 52.1 83.6 90.3 88.4 05.0 09.4 68.2 05.1 59.0
Rec. RD [10] 55.5 85.8 90.6 87.0 08.2 12.8 67.8 07.2 60.6
DiAD [16] 49.9 85.1 90.8 87.7 02.9 06.8 67.4 03.6 55.8
ViTAD [40] 47.0 81.3 90.2 87.3 04.5 09.3 66.6 05.0 57.4
InvAD [42] 66.5 87.7 93.8 88.3 22.1 26.2 73.2 16.8 67.6
InvAD-lite [42] 60.4 84.3 92.2 88.2 09.0 14.3 72.9 08.1 62.5
MambaAD [15] 56.2 84.3 90.3 87.8 07.5 12.8 68.5 07.2 60.5
Hybrid UniAD [35] 47.6 81.3 90.4 88.2 05.8 10.6 68.6 05.8 58.2
RD++ [31] 54.7 83.9 92.3 88.5 09.4 14.9 70.4 08.5 61.4
DesTSeg [44] 45.4 80.3 90.7 72.7 05.0 08.2 48.3 04.5 52.8
Appendix J Detailed Quantitative Results on MAD_Sim Dataset Under 100 epochs
Table A40: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
01Gorilla Aug. DRAEM [37] 62.6 94.3 94.9 50.1 00.7 01.8 20.9 00.9 51.2
SimpleNet [27] 62.7 94.7 94.9 91.1 08.0 15.0 69.5 08.1 65.0
Emb. CFA [21] 56.4 93.3 94.9 80.3 09.9 19.5 63.7 10.8 62.4
CFLOW-AD [14] 51.4 92.1 95.1 90.8 04.6 05.2 65.1 02.7 60.5
PyramidalFlow [22] 39.2 89.3 94.9 87.3 03.3 05.0 70.3 02.6 58.0
Rec. RD [10] 57.6 92.0 94.9 93.3 04.9 09.6 77.2 05.0 63.9
DiAD [16] 78.2 97.3 94.9 91.3 03.4 08.1 63.2 04.2 62.3
ViTAD [40] 62.4 92.9 95.1 93.5 05.3 10.0 76.1 05.3 64.8
InvAD [42] 66.9 95.3 94.9 95.1 15.8 25.2 82.8 14.4 70.2
InvAD-lite [42] 68.9 95.0 95.1 93.6 06.7 12.9 78.5 06.9 67.1
MambaAD [15] 72.0 95.5 94.9 92.8 05.5 10.4 75.6 05.5 66.8
Hybrid UniAD [35] 49.6 91.1 94.9 91.9 04.1 07.9 76.1 04.1 61.8
RD++ [31] 54.9 91.1 94.9 92.9 04.2 08.5 75.8 04.5 62.8
DesTSeg [44] 55.4 92.5 94.9 54.4 04.2 06.3 32.1 03.2 52.6
02Unicorn Aug. DRAEM [37] 51.6 91.4 95.2 57.1 00.7 00.9 23.6 00.5 50.0
SimpleNet [27] 76.0 96.7 95.2 83.7 02.0 05.0 59.4 02.6 63.4
Emb. CFA [21] 66.3 95.7 95.5 72.7 01.3 02.9 43.6 01.5 58.0
CFLOW-AD [14] 53.7 93.0 95.2 84.5 02.6 03.7 55.2 01.9 58.6
PyramidalFlow [22] 42.9 89.5 95.5 73.4 00.8 01.8 46.3 00.9 53.3
Rec. RD [10] 60.1 93.9 95.2 85.3 02.0 04.2 71.4 02.1 61.9
DiAD [16] 71.0 96.6 95.2 78.9 01.4 03.9 45.8 02.0 56.1
ViTAD [40] 63.8 95.2 95.2 86.4 02.4 05.6 67.6 02.9 62.6
InvAD [42] 79.4 97.5 95.4 87.5 04.1 11.1 75.6 05.9 67.7
InvAD-lite [42] 72.4 96.6 95.2 86.2 02.5 05.4 73.2 02.8 64.9
MambaAD [15] 66.8 95.7 95.2 84.6 02.7 05.4 69.6 02.8 63.2
Hybrid UniAD [35] 64.7 95.1 95.2 86.2 02.1 05.3 61.9 02.7 61.9
RD++ [31] 64.4 94.5 95.2 85.3 02.2 04.8 69.5 02.5 62.6
DesTSeg [44] 68.3 95.5 95.5 53.2 04.7 06.4 29.3 03.3 54.9
03Mallard Aug. DRAEM [37] 38.8 88.1 95.4 49.8 01.3 01.9 18.2 01.0 46.0
SimpleNet [27] 36.2 87.6 95.4 84.9 04.8 12.1 55.4 06.4 56.2
Emb. CFA [21] 32.2 87.5 95.4 75.4 04.6 12.8 46.7 06.9 53.3
CFLOW-AD [14] 41.3 91.4 95.4 84.9 03.7 05.1 53.1 02.6 56.4
PyramidalFlow [22] 46.0 88.7 95.4 82.1 02.4 04.9 55.0 02.5 56.4
Rec. RD [10] 47.9 91.2 95.4 84.3 04.5 09.1 57.9 04.8 58.6
DiAD [16] 74.2 96.4 95.4 82.4 02.3 04.5 47.5 02.3 57.5
ViTAD [40] 44.9 91.3 95.4 85.5 05.8 11.9 60.4 06.3 59.0
InvAD [42] 44.7 90.0 95.4 85.9 07.2 14.5 61.6 07.8 59.5
InvAD-lite [42] 39.9 89.3 95.4 85.1 05.2 11.7 59.6 06.2 57.6
MambaAD [15] 57.4 93.6 95.4 84.1 05.1 10.5 57.6 05.5 60.7
Hybrid UniAD [35] 29.6 86.4 95.4 85.3 03.4 08.6 59.5 04.5 54.8
RD++ [31] 41.2 88.7 95.4 84.2 03.5 07.8 57.7 04.1 56.7
DesTSeg [44] 32.1 88.0 95.4 56.5 03.7 07.8 32.2 04.0 48.4
04Turtle Aug. DRAEM [37] 40.6 87.6 94.9 53.6 01.0 01.3 27.2 00.7 47.6
SimpleNet [27] 41.9 90.6 94.9 82.8 02.9 06.0 51.7 03.1 55.8
Emb. CFA [21] 27.8 87.2 94.9 71.7 02.6 03.9 45.7 02.0 50.5
CFLOW-AD [14] 64.5 94.5 94.9 86.5 04.2 06.4 58.0 03.3 61.7
PyramidalFlow [22] 33.7 86.7 94.9 80.1 01.6 04.0 48.3 02.0 52.6
Rec. RD [10] 62.7 93.7 94.9 91.3 09.4 16.9 72.5 09.2 65.6
DiAD [16] 86.2 98.4 95.8 91.2 06.1 12.6 60.9 06.7 64.4
ViTAD [40] 59.0 93.6 94.9 90.8 07.8 15.6 70.0 08.5 64.3
InvAD [42] 75.8 96.0 95.1 93.1 13.5 24.6 78.5 14.0 70.7
InvAD-lite [42] 61.8 93.7 94.9 91.8 10.8 19.7 74.2 10.9 66.3
MambaAD [15] 68.9 95.5 94.9 91.0 09.3 18.0 71.2 09.9 66.9
Hybrid UniAD [35] 45.1 88.6 94.9 87.2 03.2 05.8 66.0 03.0 58.4
RD++ [31] 61.9 94.4 94.9 90.5 06.9 12.8 71.2 06.8 64.5
DesTSeg [44] 35.2 89.3 94.9 52.3 02.6 04.0 36.6 02.0 48.5
Table A41: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
05Whale Aug. DRAEM [37] 54.6 91.4 94.4 47.3 00.7 01.4 25.9 00.7 49.5
SimpleNet [27] 63.1 93.9 94.4 82.6 02.3 04.9 60.5 02.5 60.7
Emb. CFA [21] 65.2 94.4 94.4 73.4 02.3 03.8 54.3 01.9 59.1
CFLOW-AD [14] 54.5 91.0 94.4 85.4 02.7 04.4 63.3 02.2 59.5
PyramidalFlow [22] 37.9 84.7 94.4 73.6 01.0 02.4 49.4 01.2 52.0
Rec. RD [10] 56.5 92.6 94.5 87.0 03.6 07.1 74.8 03.7 62.2
DiAD [16] 48.7 91.0 94.4 80.3 02.3 06.6 50.3 03.4 53.4
ViTAD [40] 58.0 92.5 94.4 88.3 04.0 08.4 76.0 04.4 62.9
InvAD [42] 75.7 96.6 94.4 89.7 08.6 17.6 81.5 09.6 69.1
InvAD-lite [42] 64.9 94.5 94.4 89.0 06.1 11.7 78.5 06.2 65.5
MambaAD [15] 60.2 93.5 94.4 87.5 04.9 09.8 74.4 05.1 63.4
Hybrid UniAD [35] 72.4 95.7 94.5 89.5 03.8 08.2 76.9 04.3 66.1
RD++ [31] 52.4 91.7 94.4 86.7 03.7 08.0 74.4 04.1 61.4
DesTSeg [44] 52.0 91.5 94.4 58.1 04.2 09.0 45.7 04.7 54.3
06Bird Aug. DRAEM [37] 44.4 88.1 94.4 44.1 01.0 02.0 23.7 01.0 46.7
SimpleNet [27] 63.6 93.5 94.4 90.3 05.5 09.1 71.1 04.8 63.9
Emb. CFA [21] 65.7 92.3 94.7 83.8 04.6 10.8 62.3 05.7 62.3
CFLOW-AD [14] 63.8 91.5 94.4 90.4 03.9 07.1 70.9 03.7 63.2
PyramidalFlow [22] 75.7 96.1 94.4 92.6 07.0 13.5 76.6 07.2 68.1
Rec. RD [10] 58.5 91.8 94.4 90.6 04.7 08.3 75.4 04.3 63.2
DiAD [16] 79.5 97.3 95.4 93.3 06.0 12.4 77.0 06.6 65.9
ViTAD [40] 64.4 93.3 94.5 91.2 04.8 09.8 75.9 05.1 64.7
InvAD [42] 83.3 97.6 94.7 91.9 06.7 12.5 78.5 06.7 69.6
InvAD-lite [42] 74.8 95.6 94.7 91.5 05.6 09.9 77.2 05.2 67.2
MambaAD [15] 64.5 93.7 94.5 91.2 05.0 09.1 76.4 04.7 64.8
Hybrid UniAD [35] 51.3 89.6 94.4 90.2 03.9 07.5 73.0 03.9 61.0
RD++ [31] 62.9 92.1 94.5 90.9 04.7 08.3 75.7 04.4 64.0
DesTSeg [44] 57.1 91.6 94.4 52.2 04.0 04.6 25.6 02.4 51.3
07Owl Aug. DRAEM [37] 59.2 88.7 92.8 47.6 00.7 01.8 27.1 00.9 49.8
SimpleNet [27] 67.0 92.0 93.4 88.8 03.5 07.5 66.8 03.9 62.9
Emb. CFA [21] 58.3 90.8 93.0 73.4 03.2 07.4 53.8 03.8 57.6
CFLOW-AD [14] 64.1 91.9 92.8 90.0 05.7 05.0 66.4 02.6 62.4
PyramidalFlow [22] 44.7 85.0 93.0 85.3 02.2 05.2 64.2 02.7 56.7
Rec. RD [10] 63.4 92.2 93.0 91.9 05.0 09.2 77.9 04.8 64.4
DiAD [16] 62.6 92.5 92.8 93.8 04.9 10.2 71.6 05.3 61.2
ViTAD [40] 66.8 92.3 92.8 92.5 05.0 10.2 77.6 05.4 65.1
InvAD [42] 71.0 93.8 93.4 94.0 10.3 19.4 82.5 10.8 68.8
InvAD-lite [42] 67.4 93.7 92.8 92.4 06.5 11.9 78.6 06.3 66.0
MambaAD [15] 67.8 93.6 92.8 91.2 04.6 08.2 76.0 04.3 64.9
Hybrid UniAD [35] 61.1 91.5 93.0 91.5 04.1 07.9 74.3 04.1 63.2
RD++ [31] 69.4 93.4 92.8 92.0 05.1 09.2 78.2 04.8 65.7
DesTSeg [44] 65.2 92.6 93.2 53.2 10.7 07.9 35.5 04.1 55.2
08Sabertooth Aug. DRAEM [37] 60.4 90.6 93.2 45.8 00.5 00.9 24.3 00.4 49.6
SimpleNet [27] 54.4 90.2 93.0 83.8 01.7 02.8 57.6 01.4 57.8
Emb. CFA [21] 55.6 89.9 93.0 76.3 02.5 05.1 54.6 02.6 57.1
CFLOW-AD [14] 61.5 90.9 93.2 86.7 01.8 03.4 58.7 01.7 59.8
PyramidalFlow [22] 65.2 93.3 93.0 77.6 01.1 02.4 49.2 01.2 58.2
Rec. RD [10] 67.4 93.2 93.2 87.9 02.1 03.6 67.2 01.8 62.4
DiAD [16] 89.2 98.4 93.8 85.1 01.5 03.0 48.5 01.5 59.9
ViTAD [40] 61.8 91.9 93.2 88.8 02.1 03.8 67.5 02.0 61.4
InvAD [42] 73.7 94.9 93.0 90.5 03.5 05.8 74.0 03.0 65.3
InvAD-lite [42] 69.3 93.7 93.2 88.3 02.8 04.0 68.0 02.1 63.1
MambaAD [15] 63.3 92.5 93.2 86.9 02.5 03.7 63.9 01.9 61.1
Hybrid UniAD [35] 55.5 89.1 93.0 88.2 02.1 03.7 71.2 01.9 60.2
RD++ [31] 64.0 92.2 93.0 87.7 02.2 03.7 66.6 01.9 61.6
DesTSeg [44] 61.1 92.1 93.0 51.6 04.5 04.5 31.0 02.3 52.5
Table A42: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
09Swan Aug. DRAEM [37] 41.6 83.8 93.1 44.7 00.6 01.0 23.4 00.5 45.1
SimpleNet [27] 52.6 89.0 93.1 83.7 02.0 04.8 58.1 02.4 57.7
Emb. CFA [21] 56.6 89.3 93.3 75.1 02.5 07.4 48.4 03.8 56.5
CFLOW-AD [14] 47.4 86.7 93.1 85.8 02.2 04.0 60.6 02.0 56.9
PyramidalFlow [22] 43.9 85.2 93.1 75.7 01.1 02.6 42.6 01.3 52.3
Rec. RD [10] 55.6 89.6 93.1 87.7 02.5 04.5 71.1 02.3 60.4
DiAD [16] 59.6 90.5 93.8 86.8 05.9 11.8 54.7 06.3 57.6
ViTAD [40] 52.8 88.1 93.2 90.0 03.0 06.0 73.1 03.1 60.5
InvAD [42] 68.9 93.7 93.1 91.6 05.9 11.5 80.1 06.1 66.3
InvAD-lite [42] 59.9 90.7 93.1 89.8 04.0 07.8 76.1 04.1 62.8
MambaAD [15] 54.7 88.7 93.1 87.2 02.4 04.3 69.0 02.2 59.8
Hybrid UniAD [35] 52.5 87.5 93.1 88.4 02.7 05.0 71.3 02.5 59.8
RD++ [31] 53.9 88.9 93.1 87.6 02.4 04.5 70.6 02.3 60.0
DesTSeg [44] 54.7 89.3 93.1 50.9 02.5 02.3 29.0 01.2 50.1
10Sheep Aug. DRAEM [37] 44.9 86.4 92.7 44.1 00.8 01.7 27.5 00.9 46.6
SimpleNet [27] 60.4 91.3 92.7 90.5 03.8 06.8 76.8 03.5 63.0
Emb. CFA [21] 65.9 92.2 93.1 85.1 03.9 08.6 67.7 04.5 62.5
CFLOW-AD [14] 61.4 89.8 92.9 92.0 03.8 06.1 79.3 03.1 63.3
PyramidalFlow [22] 54.4 89.6 92.7 90.6 02.3 05.0 71.8 02.6 60.7
Rec. RD [10] 64.4 92.6 92.7 92.6 05.3 09.2 84.5 04.8 65.6
DiAD [16] 85.0 96.8 94.1 93.3 04.1 08.5 76.0 04.4 65.4
ViTAD [40] 64.8 92.0 92.7 93.4 04.8 09.3 84.8 04.9 65.6
InvAD [42] 84.6 97.1 93.6 94.0 08.2 13.6 87.1 07.3 71.2
InvAD-lite [42] 78.7 95.9 93.4 93.5 06.7 10.1 85.2 05.3 69.1
MambaAD [15] 70.1 93.8 92.9 92.8 05.5 09.4 84.0 04.9 66.8
Hybrid UniAD [35] 57.4 90.3 92.7 93.3 04.8 08.2 83.3 04.3 63.8
RD++ [31] 71.4 94.2 92.9 92.7 06.2 10.6 84.3 05.6 67.3
DesTSeg [44] 67.5 93.7 92.7 53.1 08.5 08.1 39.0 04.2 55.9
11Pig Aug. DRAEM [37] 51.2 90.2 94.0 54.8 01.0 01.6 28.9 00.8 50.0
SimpleNet [27] 54.8 90.9 94.0 94.0 06.4 12.8 75.7 06.8 63.6
Emb. CFA [21] 54.7 90.6 94.0 87.4 07.5 14.9 61.9 08.1 61.3
CFLOW-AD [14] 51.4 89.3 94.0 92.3 03.5 05.5 74.5 02.8 61.1
PyramidalFlow [22] 56.4 91.0 94.0 89.8 02.7 05.5 66.9 02.8 60.8
Rec. RD [10] 58.2 91.8 94.0 94.3 06.5 11.8 83.5 06.3 65.2
DiAD [16] 44.0 90.1 94.0 93.7 09.4 16.8 70.9 09.2 59.9
ViTAD [40] 58.4 91.1 94.0 95.9 08.8 17.0 83.3 09.3 66.2
InvAD [42] 72.3 95.1 94.4 96.9 18.9 27.6 87.9 16.0 72.5
InvAD-lite [42] 71.8 95.1 94.2 96.3 16.4 24.4 86.4 13.9 71.5
MambaAD [15] 67.4 94.2 94.2 95.2 10.6 18.5 84.5 10.2 68.7
Hybrid UniAD [35] 46.7 86.8 94.0 93.6 03.8 07.3 78.8 03.8 60.9
RD++ [31] 58.7 92.1 94.0 94.7 07.2 14.1 83.9 07.6 65.8
DesTSeg [44] 63.0 92.9 94.0 57.5 07.1 11.7 37.9 06.2 55.9
12Zalika Aug. DRAEM [37] 43.6 84.1 92.9 43.4 01.1 02.3 25.0 01.2 45.7
SimpleNet [27] 57.1 88.2 93.1 84.2 04.0 08.4 59.5 04.4 59.2
Emb. CFA [21] 54.8 90.0 92.9 73.9 04.2 09.2 53.5 04.8 57.2
CFLOW-AD [14] 51.4 88.0 92.9 86.0 04.4 06.9 55.3 03.6 57.8
PyramidalFlow [22] 43.8 85.3 92.9 79.1 02.2 04.1 57.5 02.1 54.9
Rec. RD [10] 59.3 92.0 92.9 87.2 04.9 08.6 69.9 04.5 62.0
DiAD [16] 46.6 86.6 92.9 85.2 04.4 06.9 42.0 03.6 52.1
ViTAD [40] 54.9 91.0 92.9 87.3 04.7 08.7 67.5 04.6 60.8
InvAD [42] 64.0 92.3 92.9 90.2 10.0 18.1 76.5 10.0 65.9
InvAD-lite [42] 60.9 92.1 92.9 88.6 07.0 11.5 72.3 06.1 63.4
MambaAD [15] 58.3 91.6 92.9 87.2 05.2 08.7 68.9 04.5 61.7
Hybrid UniAD [35] 52.1 88.9 92.9 85.9 03.4 06.0 64.8 03.1 59.0
RD++ [31] 55.5 90.5 92.9 87.0 05.1 08.2 69.3 04.3 61.0
DesTSeg [44] 52.3 88.9 92.9 52.2 04.0 05.7 38.5 02.9 51.6
Table A43: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
13Pheonix Aug. DRAEM [37] 51.9 89.6 94.3 45.4 00.6 01.0 22.8 00.5 48.0
SimpleNet [27] 55.7 90.2 94.4 80.7 01.5 03.3 56.9 01.7 57.9
Emb. CFA [21] 54.1 90.4 94.3 77.4 02.0 05.5 58.0 02.8 57.7
CFLOW-AD [14] 52.1 89.7 94.3 82.4 01.4 02.7 57.2 01.4 57.3
PyramidalFlow [22] 48.0 87.6 94.3 72.8 00.9 02.1 54.1 01.0 54.6
Rec. RD [10] 53.5 90.5 94.3 83.6 02.5 03.5 70.0 01.8 59.7
DiAD [16] 41.5 86.8 94.3 83.2 01.4 03.0 49.1 01.5 51.3
ViTAD [40] 52.6 89.9 94.3 84.8 01.8 03.5 68.4 01.8 59.3
InvAD [42] 62.7 92.9 94.6 87.0 03.6 07.2 77.5 03.7 63.6
InvAD-lite [42] 56.0 91.6 94.3 85.3 03.2 05.3 74.0 02.7 61.3
MambaAD [15] 55.7 91.7 94.4 83.7 02.8 03.7 70.7 01.9 60.4
Hybrid UniAD [35] 55.1 90.2 94.3 83.8 01.5 03.1 64.5 01.6 59.0
RD++ [31] 54.7 90.6 94.4 84.1 01.9 03.6 71.0 01.8 60.0
DesTSeg [44] 53.3 91.0 94.3 52.5 02.2 04.2 35.5 02.1 51.6
14Elephant Aug. DRAEM [37] 50.4 89.5 93.9 52.4 01.1 01.9 25.9 00.9 49.1
SimpleNet [27] 58.7 91.9 93.9 64.9 02.2 04.9 54.8 02.5 56.6
Emb. CFA [21] 54.2 90.9 93.9 61.8 01.9 03.8 45.8 01.9 54.0
CFLOW-AD [14] 55.7 91.4 93.9 68.2 02.2 03.8 49.7 01.9 55.7
PyramidalFlow [22] 44.5 87.3 93.9 63.7 01.4 03.1 45.2 01.6 51.8
Rec. RD [10] 58.3 91.7 93.9 69.9 02.9 06.6 64.6 03.4 58.6
DiAD [16] 68.7 95.1 93.9 74.2 01.9 03.9 44.7 02.0 54.6
ViTAD [40] 57.4 91.9 93.9 70.8 03.1 07.4 64.0 03.8 58.7
InvAD [42] 63.1 93.3 93.9 72.3 04.4 10.8 68.4 05.7 61.2
InvAD-lite [42] 66.2 94.2 93.9 70.5 03.6 08.2 66.0 04.3 60.9
MambaAD [15] 66.2 94.4 93.9 69.6 03.1 07.2 63.5 03.7 60.3
Hybrid UniAD [35] 55.7 90.5 93.9 68.5 01.9 04.1 59.2 02.1 56.7
RD++ [31] 60.7 92.5 93.9 70.2 03.0 06.9 63.5 03.6 59.1
DesTSeg [44] 60.0 91.9 93.9 53.5 05.3 08.1 46.9 04.2 55.2
15Parrot Aug. DRAEM [37] 52.8 86.4 91.8 48.6 00.8 01.4 24.1 00.7 47.9
SimpleNet [27] 62.2 90.3 91.8 80.3 02.4 04.8 53.2 02.4 58.3
Emb. CFA [21] 55.6 88.6 91.8 69.4 02.2 05.0 44.5 02.5 54.5
CFLOW-AD [14] 52.9 86.3 92.0 83.5 02.9 05.2 60.5 02.7 57.5
PyramidalFlow [22] 62.3 90.4 92.0 81.5 03.0 06.6 66.1 03.4 60.4
Rec. RD [10] 54.6 87.9 92.0 83.9 03.2 06.7 67.2 03.5 59.2
DiAD [16] 52.4 88.4 91.8 83.6 02.5 05.6 63.4 02.9 55.4
ViTAD [40] 50.5 87.3 91.8 83.7 03.2 07.1 66.6 03.7 58.3
InvAD [42] 60.8 90.9 92.0 86.3 06.5 13.4 73.9 07.2 63.1
InvAD-lite [42] 57.0 89.6 91.8 85.1 04.2 08.4 69.1 04.4 60.6
MambaAD [15] 55.1 88.9 91.8 84.1 03.5 06.5 65.9 03.4 59.3
Hybrid UniAD [35] 42.9 83.7 91.8 84.4 02.9 05.4 65.8 02.8 56.2
RD++ [31] 49.2 86.2 92.0 84.1 03.2 06.6 66.7 03.4 58.0
DesTSeg [44] 52.4 88.7 91.8 52.0 09.9 05.6 24.8 02.9 50.4
16Cat Aug. DRAEM [37] 50.5 89.9 94.3 43.0 00.6 01.5 24.4 00.7 47.8
SimpleNet [27] 58.3 91.8 94.3 92.1 02.5 05.7 79.3 02.9 63.2
Emb. CFA [21] 54.6 90.9 94.4 82.6 02.7 08.1 59.8 04.2 59.1
CFLOW-AD [14] 52.0 90.1 94.3 93.1 02.3 04.8 80.5 02.5 62.0
PyramidalFlow [22] 49.2 88.6 94.3 91.1 01.7 03.8 70.2 01.9 59.5
Rec. RD [10] 52.4 90.1 94.3 93.6 03.1 06.3 84.8 03.3 62.9
DiAD [16] 45.8 88.0 94.3 93.7 03.0 05.3 81.8 02.7 58.8
ViTAD [40] 57.2 92.0 94.3 93.8 03.2 06.3 84.6 03.3 64.1
InvAD [42] 58.9 92.1 94.3 94.5 05.3 11.0 86.9 05.8 65.6
InvAD-lite [42] 58.5 92.5 94.3 94.2 03.7 07.4 85.6 03.9 64.7
MambaAD [15] 54.5 91.0 94.4 94.0 03.5 07.0 84.9 03.6 63.7
Hybrid UniAD [35] 51.7 89.5 94.3 93.8 02.7 05.6 82.4 02.9 62.3
RD++ [31] 53.9 90.8 94.3 93.4 03.4 06.3 84.5 03.3 63.3
DesTSeg [44] 56.2 90.5 94.3 52.8 01.4 03.1 30.6 01.6 51.2
Table A44: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
17Scorpion Aug. DRAEM [37] 48.4 86.4 92.9 51.5 00.7 01.4 25.6 00.7 47.9
SimpleNet [27] 72.7 94.5 93.1 88.4 04.0 06.5 67.9 03.4 64.2
Emb. CFA [21] 68.3 92.4 93.0 78.9 04.1 07.0 55.9 03.6 60.5
CFLOW-AD [14] 59.7 91.0 92.9 89.6 04.6 04.9 68.1 02.5 61.5
PyramidalFlow [22] 57.6 87.3 93.1 83.4 01.9 05.1 53.1 02.6 57.6
Rec. RD [10] 60.1 91.2 92.9 91.3 04.8 07.4 82.3 03.8 63.9
DiAD [16] 84.8 96.9 93.8 87.0 01.9 03.5 56.5 01.8 60.6
ViTAD [40] 60.1 90.5 93.1 91.2 03.3 07.7 77.6 04.0 63.1
InvAD [42] 81.5 96.7 93.3 93.6 07.2 10.5 86.1 05.6 69.9
InvAD-lite [42] 73.5 94.7 92.9 93.3 06.1 08.6 84.8 04.5 67.6
MambaAD [15] 73.0 94.6 93.0 93.0 05.5 08.2 83.6 04.3 67.2
Hybrid UniAD [35] 63.5 90.9 92.9 91.7 04.4 07.2 79.1 03.8 64.0
RD++ [31] 64.1 92.3 92.9 91.7 05.4 07.9 82.4 04.1 65.0
DesTSeg [44] 77.8 95.9 93.1 60.6 05.4 09.7 44.3 05.1 59.5
18Obesobeso Aug. DRAEM [37] 43.9 87.1 94.1 47.5 01.0 01.8 26.9 00.9 47.2
SimpleNet [27] 63.3 93.6 94.1 92.4 15.8 21.7 81.6 12.2 68.3
Emb. CFA [21] 62.9 93.2 94.1 81.8 12.6 20.0 64.6 11.1 64.1
CFLOW-AD [14] 59.6 91.6 94.1 92.5 05.8 08.8 79.4 04.6 64.2
PyramidalFlow [22] 61.7 93.2 94.1 91.9 04.5 08.8 74.7 04.6 64.0
Rec. RD [10] 64.6 94.0 94.1 93.5 11.6 20.7 87.8 11.5 68.8
DiAD [16] 87.5 98.3 94.4 95.1 10.2 17.7 82.9 09.7 69.4
ViTAD [40] 67.5 94.4 94.1 94.3 18.2 24.7 88.3 14.1 70.9
InvAD [42] 83.5 97.6 94.2 94.8 16.6 26.3 90.8 15.1 74.4
InvAD-lite [42] 70.3 95.3 94.2 93.7 12.9 21.2 88.1 11.8 70.3
MambaAD [15] 69.6 95.0 94.2 93.0 10.9 20.2 86.4 11.3 69.4
Hybrid UniAD [35] 64.9 93.4 94.1 92.2 06.9 12.9 82.4 06.9 66.4
RD++ [31] 67.7 94.0 94.4 93.2 09.1 17.4 86.5 09.5 68.5
DesTSeg [44] 63.1 93.1 94.2 61.1 16.2 22.4 45.0 12.6 59.8
19Bear Aug. DRAEM [37] 54.1 89.3 93.3 49.3 00.8 01.3 23.0 00.7 48.7
SimpleNet [27] 63.0 92.8 93.5 88.4 07.0 12.8 68.6 06.8 63.7
Emb. CFA [21] 63.0 92.4 93.5 80.7 05.5 12.1 57.5 06.4 61.0
CFLOW-AD [14] 60.9 91.3 93.3 88.9 03.7 06.4 64.6 03.3 61.4
PyramidalFlow [22] 52.6 89.6 93.6 87.2 03.6 06.3 61.0 03.3 59.1
Rec. RD [10] 62.9 92.2 93.6 90.8 06.2 10.8 76.8 05.7 64.5
DiAD [16] 59.1 92.9 93.3 89.4 04.2 10.5 62.6 05.5 58.8
ViTAD [40] 60.9 91.2 93.5 91.1 05.5 11.5 75.1 06.1 63.8
InvAD [42] 73.9 95.4 93.3 92.5 13.8 21.7 82.0 12.2 70.0
InvAD-lite [42] 67.7 93.2 93.3 92.2 11.3 18.1 78.9 10.0 67.4
MambaAD [15] 64.2 92.9 93.5 91.2 07.9 13.4 76.1 07.2 65.3
Hybrid UniAD [35] 56.8 90.6 93.3 90.0 04.2 07.9 77.2 04.1 62.5
RD++ [31] 65.3 93.0 93.5 90.8 07.7 11.8 76.9 06.3 65.4
DesTSeg [44] 66.8 93.5 93.7 58.5 09.4 13.7 43.7 07.3 58.0
20Puppy Aug. DRAEM [37] 44.7 84.1 92.7 47.9 00.5 00.9 20.7 00.4 45.7
SimpleNet [27] 54.6 89.8 92.9 83.1 02.6 04.9 54.4 02.5 57.7
Emb. CFA [21] 56.3 90.2 92.9 71.2 03.0 06.9 51.5 03.6 56.5
CFLOW-AD [14] 57.5 90.1 92.7 85.2 01.9 04.5 57.1 02.3 58.6
PyramidalFlow [22] 56.0 89.4 92.9 81.1 01.5 03.2 53.1 01.6 57.1
Rec. RD [10] 54.0 89.6 93.0 86.8 03.6 06.3 66.5 03.3 59.8
DiAD [16] 48.7 88.8 92.7 83.4 01.4 02.7 51.9 01.4 52.8
ViTAD [40] 52.0 87.4 93.0 87.6 02.4 04.9 66.3 02.5 58.9
InvAD [42] 62.9 92.5 92.9 89.3 05.9 11.0 73.3 05.8 63.8
InvAD-lite [42] 54.0 90.0 93.0 87.5 03.5 06.2 68.2 03.2 60.2
MambaAD [15] 49.2 87.7 92.7 86.6 02.6 05.3 64.8 02.7 58.2
Hybrid UniAD [35] 55.6 89.8 92.7 86.5 03.4 07.9 68.4 04.1 60.5
RD++ [31] 59.0 91.5 92.9 87.1 04.1 08.1 66.8 04.2 61.3
DesTSeg [44] 64.6 93.0 92.9 53.7 13.3 09.5 32.9 05.0 55.4
Avg Aug. DRAEM [37] 49.5 88.3 93.8 48.4 00.8 01.5 24.5 00.7 48.0
SimpleNet [27] 58.9 91.7 93.8 85.5 04.2 08.0 63.9 04.2 60.9
Emb. CFA [21] 56.4 91.1 93.8 76.6 04.2 08.7 54.7 04.6 58.2
CFLOW-AD [14] 55.8 90.6 93.8 86.9 03.4 05.2 63.9 02.7 60.0
PyramidalFlow [22] 50.8 88.9 93.8 82.0 02.3 04.8 58.8 02.5 57.4
Rec. RD [10] 58.6 91.7 93.8 88.3 04.7 08.5 74.2 04.5 62.6
DiAD [16] 65.7 93.3 94.1 87.2 03.9 07.9 60.1 04.2 58.9
ViTAD [40] 58.5 91.5 93.8 89.0 05.0 09.5 73.5 05.0 62.8
InvAD [42] 70.4 94.6 93.9 90.5 08.8 15.7 79.3 08.6 67.4
InvAD-lite [42] 64.7 93.3 93.8 89.4 06.4 11.2 76.1 06.0 64.9
MambaAD [15] 63.0 92.9 93.8 88.3 05.2 09.4 73.3 05.0 63.6
Hybrid UniAD [35] 54.2 90.0 93.8 88.1 03.5 06.8 71.8 03.5 60.9
RD++ [31] 59.3 91.7 93.8 88.3 04.6 08.5 73.8 04.4 62.7
DesTSeg [44] 57.9 91.8 93.8 54.5 06.2 07.7 35.8 04.1 53.6
Appendix K Detailed Quantitative Results on Uni-Medical Dataset Under 100 epochs
Table A45: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
brain Aug. DRAEM [37] 47.1 82.8 90.6 33.8 02.2 06.1 07.9 03.1 43.0
SimpleNet [27] 84.8 96.3 91.0 94.5 43.4 43.9 71.4 28.1 77.0
Emb. CFA [21] 82.5 95.9 90.6 75.6 33.0 39.0 36.5 24.2 67.8
CFLOW-AD [14] 85.7 96.2 91.9 94.5 56.6 17.2 73.8 09.4 75.9
PyramidalFlow [22] 62.4 86.1 90.9 85.7 10.4 18.0 50.5 09.9 60.5
Rec. RD [10] 83.9 94.6 91.9 96.5 45.6 48.9 82.8 32.4 79.3
DiAD [16] 93.2 97.7 95.2 94.9 35.2 39.8 79.5 24.8 76.5
ViTAD [40] 89.4 97.3 93.0 97.7 61.8 58.8 84.0 41.7 84.4
InvAD [42] 90.8 97.7 93.2 97.7 59.1 58.3 86.2 41.2 84.6
InvAD-lite [42] 84.9 95.9 91.8 96.5 46.2 49.2 82.5 32.6 79.7
MambaAD [15] 93.7 98.5 94.2 97.9 60.8 61.5 87.5 44.4 86.2
Hybrid UniAD [35] 90.2 97.7 92.6 97.4 56.2 55.6 82.9 38.5 83.3
RD++ [31] 84.8 95.5 91.8 96.9 48.5 51.5 84.2 34.6 80.5
DesTSeg [44] 87.1 95.9 92.2 70.2 52.2 42.7 33.9 27.2 70.7
liver Aug. DRAEM [37] 53.1 49.7 61.6 20.0 00.4 01.5 10.2 00.7 31.4
SimpleNet [27] 60.5 50.6 62.2 94.2 06.0 11.6 81.5 06.2 53.0
Emb. CFA [21] 52.9 48.1 61.3 88.0 04.7 11.1 60.5 05.9 47.6
CFLOW-AD [14] 63.9 56.1 63.0 95.9 02.9 05.6 87.4 02.9 54.5
PyramidalFlow [22] 47.2 43.2 60.9 75.0 03.3 11.5 09.0 06.1 37.6
Rec. RD [10] 55.5 45.8 64.4 96.6 05.5 10.4 91.1 05.5 53.1
DiAD [16] 56.5 51.7 62.0 96.4 05.5 10.6 89.6 05.6 53.2
ViTAD [40] 63.9 54.9 65.1 97.9 13.2 21.4 92.6 12.0 58.8
InvAD [42] 64.2 52.8 65.9 97.0 07.8 14.5 92.9 07.8 57.0
InvAD-lite [42] 57.0 46.6 63.8 96.4 05.3 09.8 90.6 05.1 53.2
MambaAD [15] 64.2 54.6 64.5 97.0 10.0 17.4 92.6 09.5 57.7
Hybrid UniAD [35] 60.3 48.3 63.1 97.1 07.9 14.0 92.7 07.5 55.1
RD++ [31] 55.6 44.8 64.4 96.6 05.5 10.3 91.5 05.4 53.0
DesTSeg [44] 58.5 47.8 61.4 57.6 13.8 15.6 31.4 08.5 42.8
retinal Aug. DRAEM [37] 54.7 49.4 59.7 47.8 03.5 09.6 13.3 05.1 36.6
SimpleNet [27] 87.0 86.2 76.8 94.2 53.9 52.5 78.2 35.6 76.5
Emb. CFA [21] 77.7 75.8 67.4 71.4 20.7 24.5 37.3 14.0 56.1
CFLOW-AD [14] 88.0 86.5 76.8 95.8 61.6 57.9 84.1 40.7 79.3
PyramidalFlow [22] 52.3 42.7 59.7 75.9 11.4 21.8 43.2 12.2 44.8
Rec. RD [10] 89.0 86.8 78.2 96.4 65.1 60.1 86.5 43.0 80.8
DiAD [16] 86.6 82.0 76.0 96.1 61.8 56.2 83.8 39.0 77.5
ViTAD [40] 92.1 90.0 81.9 95.6 69.8 64.2 83.5 47.3 83.1
InvAD [42] 92.2 91.0 82.4 97.1 71.9 65.5 88.7 48.7 84.7
InvAD-lite [42] 91.7 89.2 81.8 95.7 65.3 59.0 84.4 41.9 81.8
MambaAD [15] 93.8 89.1 86.9 95.5 66.4 63.5 84.6 46.6 83.7
Hybrid UniAD [35] 86.3 82.4 75.8 95.3 53.8 53.7 82.5 36.7 76.4
RD++ [31] 91.2 89.2 81.3 96.7 70.4 64.9 87.6 48.1 83.6
DesTSeg [44] 90.0 87.4 81.0 69.2 59.0 43.7 40.4 27.9 69.6
Avg Aug. DRAEM [37] 51.6 60.6 70.6 33.8 02.0 05.7 10.5 03.0 37.0
SimpleNet [27] 77.5 77.7 76.7 94.3 34.4 36.0 77.0 23.3 68.9
Emb. CFA [21] 71.0 73.3 73.1 78.3 19.5 24.9 44.7 14.7 57.2
CFLOW-AD [14] 79.2 79.6 77.2 95.4 40.4 26.9 81.8 17.7 69.9
PyramidalFlow [22] 53.9 57.3 70.5 78.8 08.4 17.1 34.2 09.4 47.6
Rec. RD [10] 76.1 75.7 78.2 96.5 38.8 39.8 86.8 26.9 71.1
DiAD [16] 78.8 77.2 77.7 95.8 34.2 35.5 84.3 23.2 69.1
ViTAD [40] 81.8 80.7 80.0 97.1 48.3 48.2 86.7 33.7 75.5
InvAD [42] 82.4 80.5 80.5 97.3 46.2 46.1 89.3 32.6 75.4
InvAD-lite [42] 77.9 77.3 79.1 96.2 39.0 39.3 85.8 26.5 71.6
MambaAD [15] 83.9 80.8 81.9 96.8 45.8 47.5 88.2 33.5 75.9
Hybrid UniAD [35] 79.0 76.1 77.1 96.6 39.3 41.1 86.0 27.6 71.6
RD++ [31] 77.2 76.5 79.2 96.7 41.5 42.2 87.8 29.4 72.3
DesTSeg [44] 78.5 77.0 78.2 65.7 41.7 34.0 35.3 21.2 61.0
Appendix L Detailed Quantitative Results on Real-IAD Dataset Under 100 epochs
Table A46: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
audiojack Aug. DRAEM [37] 53.0 36.9 51.4 42.8 00.1 00.3 13.7 00.2 30.7
SimpleNet [27] 51.4 36.3 50.9 66.0 00.2 00.9 37.6 00.5 36.2
Emb. CFA [21] 57.3 41.9 52.1 86.5 00.5 01.6 57.1 00.8 43.4
CFLOW-AD [14] 77.4 70.8 60.9 95.4 16.4 15.4 73.2 08.3 59.9
PyramidalFlow [22] 51.7 34.9 50.7 81.7 00.2 00.4 40.7 00.2 38.3
Rec. RD [10] 76.0 62.8 60.6 96.7 13.1 22.1 80.6 12.4 59.8
DiAD [16] 76.5 54.3 65.7 91.6 01.0 03.9 63.3 02.0 50.9
ViTAD [40] 80.6 72.8 63.5 96.8 21.3 31.5 79.5 18.7 64.8
InvAD [42] 85.2 80.7 69.3 98.5 33.9 43.8 89.6 28.0 72.4
InvAD-lite [42] 84.3 78.7 67.8 98.1 32.4 40.8 87.0 25.6 70.8
MambaAD [15] 84.9 77.7 68.2 98.0 22.2 30.4 86.2 17.9 68.1
Hybrid UniAD [35] 81.7 76.2 65.4 97.6 20.7 31.1 84.1 18.4 66.4
RD++ [31] 76.0 60.0 61.5 96.6 10.6 20.1 80.4 11.2 58.9
DesTSeg [44] 81.6 74.7 66.5 88.3 44.9 50.7 64.5 34.0 68.2
bottle_cap Aug. DRAEM [37] 46.9 40.9 60.2 35.2 00.0 00.1 09.7 00.0 30.3
SimpleNet [27] 47.4 41.0 60.2 80.1 00.5 03.7 35.8 01.9 39.8
Emb. CFA [21] 55.7 49.7 60.2 85.2 00.5 02.4 57.0 01.2 45.7
CFLOW-AD [14] 80.0 78.3 71.7 98.8 13.6 22.6 91.9 12.7 66.7
PyramidalFlow [22] 55.4 46.7 60.3 93.2 00.7 02.6 72.0 01.3 48.1
Rec. RD [10] 89.6 86.3 81.2 99.4 18.6 28.6 96.3 16.7 73.2
DiAD [16] 91.6 94.0 87.9 94.6 04.9 11.4 73.0 06.0 65.3
ViTAD [40] 82.2 78.8 74.8 98.7 16.7 27.5 91.5 16.0 68.6
InvAD [42] 95.5 94.9 85.9 99.6 24.9 33.9 97.2 20.4 78.0
InvAD-lite [42] 92.6 91.2 81.9 99.6 21.2 28.5 97.5 16.6 75.1
MambaAD [15] 93.2 92.6 82.7 99.7 29.9 33.8 97.4 20.3 77.3
Hybrid UniAD [35] 91.6 90.6 80.4 99.4 21.0 30.5 95.8 18.0 74.6
RD++ [31] 89.7 86.7 81.5 99.3 18.9 30.4 95.6 17.9 73.5
DesTSeg [44] 75.5 74.5 65.9 71.1 27.3 31.7 43.2 18.9 57.6
button_battery Aug. DRAEM [37] 34.4 46.9 72.4 51.8 00.5 00.9 17.8 00.4 34.5
SimpleNet [27] 60.4 68.6 72.4 78.0 09.4 15.4 38.3 08.4 51.2
Emb. CFA [21] 53.2 57.9 72.5 84.3 01.6 03.2 54.8 01.6 48.6
CFLOW-AD [14] 66.0 75.1 72.7 95.6 29.2 25.8 74.8 14.8 63.8
PyramidalFlow [22] 52.5 56.4 72.5 50.2 17.5 00.8 15.0 00.4 40.7
Rec. RD [10] 74.7 79.7 75.9 97.7 34.5 39.4 86.4 24.5 70.6
DiAD [16] 80.5 71.3 70.6 84.1 01.4 05.3 66.9 02.7 54.3
ViTAD [40] 77.0 82.2 77.6 98.3 49.1 49.6 82.7 32.9 74.4
InvAD [42] 90.4 92.5 84.6 99.1 57.2 56.1 92.6 39.0 82.7
InvAD-lite [42] 82.0 86.9 78.4 98.5 51.1 52.4 89.0 35.5 77.6
MambaAD [15] 82.8 87.4 79.2 98.3 48.6 49.6 88.1 33.0 77.1
Hybrid UniAD [35] 72.8 79.8 74.6 96.5 29.5 33.7 78.7 20.3 67.7
RD++ [31] 76.5 82.2 75.8 97.9 38.4 42.7 87.1 27.2 72.3
DesTSeg [44] 84.6 87.8 81.2 92.2 55.7 53.7 60.4 36.7 75.0
end_cap Aug. DRAEM [37] 48.1 57.5 72.9 54.9 00.1 00.3 21.0 00.1 39.3
SimpleNet [27] 52.4 59.6 72.9 65.5 00.3 01.8 39.0 00.9 44.1
Emb. CFA [21] 55.3 62.5 73.3 72.8 00.4 02.7 40.0 01.4 46.3
CFLOW-AD [14] 63.7 72.7 72.9 87.6 03.5 05.9 60.3 03.0 54.5
PyramidalFlow [22] 55.7 63.8 73.0 77.2 00.4 01.9 33.2 01.0 46.2
Rec. RD [10] 80.0 84.0 78.2 96.5 12.3 21.9 89.4 12.3 67.9
DiAD [16] 85.1 83.4 84.8 81.3 02.0 06.9 38.2 03.6 54.5
ViTAD [40] 72.6 77.1 75.2 93.8 05.7 12.2 75.9 06.5 60.9
InvAD [42] 84.1 87.7 80.3 98.0 14.7 22.7 93.5 12.8 70.6
InvAD-lite [42] 79.8 84.2 78.3 97.4 13.5 21.0 91.5 11.8 68.3
MambaAD [15] 78.6 83.1 77.4 97.2 12.2 19.5 90.3 10.8 67.2
Hybrid UniAD [35] 80.9 86.3 77.8 95.5 09.5 17.7 86.4 09.7 67.0
RD++ [31] 79.7 84.3 77.8 96.8 12.6 22.6 90.4 12.7 68.1
DesTSeg [44] 72.7 76.8 74.9 75.3 10.3 18.7 53.2 10.3 57.1
Table A47: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
eraser Aug. DRAEM [37] 53.8 42.8 55.9 42.6 00.1 00.2 11.2 00.1 32.2
SimpleNet [27] 53.4 46.3 55.8 87.6 03.8 08.7 59.8 04.6 45.9
Emb. CFA [21] 51.2 43.0 55.9 78.5 01.8 05.8 57.6 03.0 43.0
CFLOW-AD [14] 88.6 87.5 77.6 98.9 23.2 26.9 93.4 15.6 72.6
PyramidalFlow [22] 57.8 42.5 57.3 88.8 00.7 02.6 60.3 01.3 45.3
Rec. RD [10] 90.0 88.5 79.4 99.5 30.6 35.9 96.3 21.9 75.8
DiAD [16] 80.0 80.0 77.3 91.1 07.7 15.4 67.5 08.4 59.9
ViTAD [40] 86.0 85.1 74.5 98.5 24.8 32.4 89.3 19.3 71.6
InvAD [42] 92.1 90.3 81.7 99.4 30.4 36.1 95.1 22.0 76.6
InvAD-lite [42] 88.7 87.2 77.4 99.3 28.9 33.6 94.8 20.2 74.3
MambaAD [15] 88.4 86.9 76.7 99.2 27.4 35.7 93.8 21.7 74.0
Hybrid UniAD [35] 90.2 88.8 80.2 99.3 25.4 32.0 94.4 19.1 74.6
RD++ [31] 90.6 89.4 79.7 99.5 29.2 35.9 96.8 21.9 76.0
DesTSeg [44] 80.5 79.5 68.4 83.3 53.8 53.0 56.0 36.1 68.8
fire_hood Aug. DRAEM [37] 56.3 39.5 53.4 42.7 00.1 00.2 09.9 00.1 31.5
SimpleNet [27] 51.3 39.9 52.9 76.8 01.1 04.8 39.9 02.5 39.3
Emb. CFA [21] 56.2 40.6 52.9 79.6 00.4 01.4 42.0 00.7 40.4
CFLOW-AD [14] 80.3 73.8 68.2 98.1 21.6 22.7 86.0 12.8 65.6
PyramidalFlow [22] 56.5 39.1 54.1 50.0 00.2 00.1 15.4 00.1 33.2
Rec. RD [10] 78.4 70.8 64.5 98.9 28.3 34.6 88.4 20.9 66.9
DiAD [16] 83.3 81.7 80.5 91.8 03.2 09.2 66.7 04.8 59.5
ViTAD [40] 75.6 67.8 60.6 98.0 18.0 27.8 82.5 16.1 62.3
InvAD [42] 82.1 75.1 68.3 99.0 28.2 35.3 89.4 21.5 69.1
InvAD-lite [42] 80.0 73.6 65.6 98.8 24.4 31.6 87.5 18.7 66.8
MambaAD [15] 79.9 73.5 65.4 98.7 25.9 33.0 86.8 19.7 67.0
Hybrid UniAD [35] 80.3 74.5 66.8 98.6 25.6 33.5 84.8 20.1 67.2
RD++ [31] 80.5 73.7 66.7 98.9 29.2 36.6 89.4 22.4 68.6
DesTSeg [44] 75.6 65.6 61.5 78.4 26.7 34.1 45.9 20.6 56.9
mint Aug. DRAEM [37] 51.8 46.9 64.0 45.3 00.1 00.1 13.9 00.1 34.5
SimpleNet [27] 55.7 51.9 63.8 73.1 00.5 03.0 26.8 01.5 41.5
Emb. CFA [21] 51.2 49.0 63.8 77.1 00.1 00.3 39.2 00.2 41.9
CFLOW-AD [14] 63.6 64.1 63.7 92.8 08.7 12.8 61.4 06.8 53.9
PyramidalFlow [22] 57.0 51.2 63.8 50.0 00.0 00.1 15.0 00.1 36.8
Rec. RD [10] 66.3 64.6 65.1 94.9 12.8 23.8 73.5 13.5 58.3
DiAD [16] 76.7 76.7 76.0 91.1 05.7 11.6 64.2 06.2 57.4
ViTAD [40] 69.2 70.1 64.9 94.7 13.9 26.6 70.0 15.3 59.7
InvAD [42] 76.8 77.1 68.9 98.2 18.5 29.1 82.8 17.0 65.7
InvAD-lite [42] 72.0 73.1 66.2 97.3 17.9 28.6 80.0 16.7 63.2
MambaAD [15] 72.6 73.8 66.2 97.1 16.7 27.2 76.9 15.8 62.7
Hybrid UniAD [35] 67.6 68.6 64.4 94.6 09.3 22.2 63.2 12.5 57.1
RD++ [31] 67.6 66.6 65.2 96.2 13.7 25.2 76.9 14.4 59.7
DesTSeg [44] 62.8 60.6 64.2 66.5 21.1 27.5 27.5 15.9 49.1
mounts Aug. DRAEM [37] 49.8 34.4 52.1 37.5 00.1 00.2 10.3 00.1 28.7
SimpleNet [27] 49.9 39.2 52.1 87.2 01.4 04.2 63.0 02.2 43.0
Emb. CFA [21] 62.8 49.5 53.9 89.9 03.3 09.3 68.6 04.9 49.1
CFLOW-AD [14] 82.9 74.5 70.3 96.1 16.9 25.5 86.4 14.6 66.1
PyramidalFlow [22] 56.1 41.1 52.1 82.7 01.0 04.0 54.3 02.0 42.6
Rec. RD [10] 88.6 79.8 75.0 99.3 30.7 37.5 95.1 23.0 73.4
DiAD [16] 75.3 74.5 82.5 84.3 00.4 01.1 48.8 00.5 52.4
ViTAD [40] 84.9 74.8 71.4 98.8 25.1 32.2 91.7 19.2 69.5
InvAD [42] 88.9 79.3 77.2 99.4 31.2 35.7 94.2 21.7 73.5
InvAD-lite [42] 87.4 76.6 75.7 99.3 27.4 32.7 94.2 19.5 71.6
MambaAD [15] 87.4 78.3 74.2 99.1 30.9 35.2 92.6 21.3 72.2
Hybrid UniAD [35] 87.2 75.6 76.2 99.4 28.1 33.6 95.3 20.2 71.9
RD++ [31] 88.5 80.6 74.5 99.3 28.8 36.2 95.3 22.1 73.1
DesTSeg [44] 72.8 53.9 62.6 79.6 27.5 37.6 58.0 23.1 56.9
Table A48: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pcb Aug. DRAEM [37] 56.8 64.0 76.2 45.3 00.1 00.3 13.9 00.2 40.3
SimpleNet [27] 54.2 63.4 75.5 76.8 00.4 01.0 45.8 00.5 47.7
Emb. CFA [21] 51.5 62.4 75.5 88.6 01.2 03.5 58.6 01.8 50.6
CFLOW-AD [14] 74.3 83.2 76.6 95.1 15.0 23.8 75.6 13.5 65.2
PyramidalFlow [22] 54.8 64.5 75.7 81.6 00.3 00.7 45.0 00.4 48.5
Rec. RD [10] 79.2 85.4 79.7 97.5 16.1 24.3 88.6 13.8 69.0
DiAD [16] 86.0 85.1 85.4 92.0 03.7 07.4 66.5 03.8 60.9
ViTAD [40] 83.3 89.2 81.2 98.2 25.7 33.6 88.2 20.2 73.0
InvAD [42] 91.9 95.2 86.9 99.4 50.7 53.2 94.9 36.2 82.9
InvAD-lite [42] 90.3 94.3 85.4 99.3 49.1 51.9 94.5 35.1 81.8
MambaAD [15] 90.3 94.3 85.1 99.2 48.4 51.7 93.8 34.9 81.6
Hybrid UniAD [35] 81.0 88.2 78.9 97.2 22.0 32.1 81.7 19.1 70.5
RD++ [31] 82.2 88.3 80.6 97.9 22.6 31.0 89.0 18.3 71.9
DesTSeg [44] 82.0 89.2 79.0 90.0 50.6 52.6 66.1 35.7 74.1
phone_battery Aug. DRAEM [37] 51.7 46.1 58.0 45.8 00.1 00.2 14.6 00.1 33.6
SimpleNet [27] 55.2 48.7 58.1 75.4 01.5 06.0 46.8 03.1 43.2
Emb. CFA [21] 49.9 43.5 58.0 64.0 00.9 03.1 40.6 01.6 38.8
CFLOW-AD [14] 74.9 73.4 64.2 75.2 13.6 22.7 83.9 12.8 59.8
PyramidalFlow [22] 45.5 36.1 58.4 71.6 00.4 01.2 27.2 00.6 35.9
Rec. RD [10] 87.5 82.5 77.4 77.3 22.3 32.1 94.6 19.1 69.5
DiAD [16] 82.3 77.7 75.9 96.8 05.3 11.4 85.4 06.0 62.1
ViTAD [40] 90.7 89.3 79.4 98.8 24.1 32.3 91.0 19.3 74.0
InvAD [42] 91.2 87.7 80.8 86.2 22.2 30.4 95.1 17.9 72.5
InvAD-lite [42] 91.2 89.6 81.8 80.3 25.5 33.8 95.7 20.3 73.2
MambaAD [15] 90.2 88.6 80.6 99.4 35.2 39.9 95.5 24.9 77.0
Hybrid UniAD [35] 84.9 81.2 73.5 89.1 10.5 18.4 88.7 10.1 65.8
RD++ [31] 88.6 85.1 78.4 83.4 22.5 31.9 95.1 19.0 71.1
DesTSeg [44] 83.4 78.5 71.9 75.6 10.5 20.6 62.3 11.5 60.1
plastic_nut Aug. DRAEM [37] 53.7 34.0 51.7 51.4 00.1 00.1 16.7 00.0 31.8
SimpleNet [27] 61.9 43.0 52.6 74.0 00.5 02.3 40.9 01.2 41.0
Emb. CFA [21] 60.2 39.8 52.9 75.3 00.2 00.8 39.1 00.4 39.9
CFLOW-AD [14] 69.7 60.5 53.6 95.6 13.4 14.9 79.1 08.0 56.0
PyramidalFlow [22] 45.0 29.3 49.7 76.3 00.1 00.4 29.3 00.2 33.9
Rec. RD [10] 81.3 70.0 64.5 98.8 22.4 30.4 92.6 17.9 66.5
DiAD [16] 71.9 58.2 65.6 81.1 00.4 03.4 38.6 01.7 45.6
ViTAD [40] 82.3 73.8 64.8 97.6 20.8 28.3 87.1 16.5 66.0
InvAD [42] 89.7 82.7 74.6 99.5 29.0 32.7 96.9 19.6 73.4
InvAD-lite [42] 88.7 81.6 73.3 99.5 30.3 34.6 97.0 20.9 73.3
MambaAD [15] 87.8 82.0 71.9 99.5 34.3 38.0 96.8 23.4 73.9
Hybrid UniAD [35] 80.1 70.5 63.0 98.4 23.0 28.4 89.9 16.5 65.6
RD++ [31] 79.8 69.1 63.7 98.9 21.9 28.8 92.5 16.8 65.7
DesTSeg [44] 84.8 78.4 68.9 85.2 49.2 49.5 62.7 32.9 69.5
plastic_plug Aug. DRAEM [37] 64.1 52.9 54.8 41.6 00.0 00.1 09.4 00.0 35.0
SimpleNet [27] 51.3 40.0 54.6 75.9 00.2 00.8 38.6 00.4 38.8
Emb. CFA [21] 42.0 33.4 54.6 85.6 00.2 00.9 53.7 00.5 39.2
CFLOW-AD [14] 78.8 75.2 64.9 97.2 11.8 20.7 85.7 11.5 63.4
PyramidalFlow [22] 46.2 35.2 54.6 79.7 00.1 00.3 40.5 00.2 37.7
Rec. RD [10] 82.6 75.5 68.8 99.1 21.3 30.2 95.5 17.8 68.6
DiAD [16] 88.7 89.2 90.9 92.9 08.7 15.0 66.1 08.1 64.5
ViTAD [40] 80.8 74.7 66.7 97.5 13.3 22.6 84.6 12.7 64.3
InvAD [42] 89.0 82.9 76.2 99.0 22.9 30.7 94.3 18.1 72.2
InvAD-lite [42] 88.0 83.9 74.6 99.1 26.2 33.7 94.3 20.2 72.7
MambaAD [15] 86.5 83.4 72.8 99.0 25.2 32.3 93.0 19.3 71.6
Hybrid UniAD [35] 80.9 76.6 66.8 98.6 18.4 27.1 89.9 15.7 66.6
RD++ [31] 85.5 77.9 73.2 99.2 21.3 28.8 95.8 16.8 70.1
DesTSeg [44] 73.6 64.7 61.6 71.9 04.4 08.7 41.7 04.6 49.2
Table A49: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
porcelain_doll Aug. DRAEM [37] 45.1 29.1 49.8 46.6 00.1 00.1 14.6 00.1 28.3
SimpleNet [27] 63.5 50.7 51.2 81.4 02.4 07.4 48.9 03.8 45.1
Emb. CFA [21] 69.4 55.1 55.2 80.1 00.1 00.3 41.7 00.2 45.2
CFLOW-AD [14] 83.6 78.1 68.2 97.1 14.0 20.9 86.6 11.7 65.6
PyramidalFlow [22] 48.7 34.0 49.8 84.5 00.2 00.3 48.7 00.2 38.8
Rec. RD [10] 87.3 78.2 72.3 99.2 25.1 33.8 96.1 20.3 71.4
DiAD [16] 72.6 66.8 65.2 93.1 01.4 04.8 70.4 02.5 53.5
ViTAD [40] 86.4 77.5 70.8 99.0 21.6 29.7 94.7 17.4 69.7
InvAD [42] 87.1 78.2 70.5 98.5 23.3 33.0 92.7 19.7 70.2
InvAD-lite [42] 88.6 82.8 74.4 99.3 32.0 36.6 96.4 22.4 74.0
MambaAD [15] 88.4 82.7 74.4 99.2 32.2 36.9 95.9 22.6 73.9
Hybrid UniAD [35] 84.9 76.0 70.3 98.6 14.9 24.8 93.3 14.1 67.5
RD++ [31] 88.4 80.7 72.6 99.3 26.3 34.7 96.7 21.0 72.4
DesTSeg [44] 75.0 61.4 59.4 76.3 35.4 38.7 44.5 24.0 57.0
regulator Aug. DRAEM [37] 46.1 26.1 43.8 38.8 00.0 00.1 11.4 00.0 25.6
SimpleNet [27] 47.5 26.2 43.8 79.9 00.1 00.7 40.8 00.4 34.8
Emb. CFA [21] 55.7 32.2 45.1 81.4 00.4 01.4 41.5 00.7 37.8
CFLOW-AD [14] 50.5 29.5 43.9 88.2 01.2 02.1 58.0 01.0 39.3
PyramidalFlow [22] 55.1 31.3 44.9 76.9 00.1 00.3 42.9 00.1 36.9
Rec. RD [10] 67.9 49.9 48.9 98.1 09.0 17.7 89.6 09.7 54.6
DiAD [16] 72.1 71.4 78.2 84.2 00.4 01.5 44.4 00.7 50.3
ViTAD [40] 61.3 45.6 44.7 95.6 07.6 17.8 78.3 09.8 50.2
InvAD [42] 84.5 78.6 68.8 99.5 25.9 33.7 96.6 20.3 70.6
InvAD-lite [42] 75.7 67.1 57.3 98.7 18.1 29.2 92.6 17.1 63.2
MambaAD [15] 72.1 62.7 53.4 97.9 21.7 30.1 88.5 17.7 61.1
Hybrid UniAD [35] 51.9 34.4 44.3 94.2 07.3 14.9 72.1 08.1 45.3
RD++ [31] 70.3 54.4 51.2 98.3 11.1 20.1 90.8 11.2 56.9
DesTSeg [44] 68.9 56.3 48.6 67.9 29.1 32.6 35.5 19.5 49.6
rolled_strip_base Aug. DRAEM [37] 53.3 68.5 79.8 37.1 00.1 00.3 09.4 00.1 39.5
SimpleNet [27] 66.7 80.1 79.9 77.7 01.5 04.9 50.3 02.5 54.6
Emb. CFA [21] 54.9 68.8 80.1 85.0 00.7 01.6 51.5 00.8 51.3
CFLOW-AD [14] 92.6 96.6 88.9 97.8 10.6 15.0 93.2 08.1 73.4
PyramidalFlow [22] 59.9 74.9 79.8 88.1 01.3 03.5 69.1 01.8 56.0
Rec. RD [10] 97.8 98.8 95.1 99.7 30.5 39.3 98.7 24.4 82.1
DiAD [16] 68.4 55.9 56.8 87.7 00.6 03.2 63.4 01.6 48.0
ViTAD [40] 98.3 99.1 95.5 99.5 27.1 35.5 97.8 21.6 81.3
InvAD [42] 99.3 99.6 97.7 99.8 32.8 42.5 99.0 27.0 83.7
InvAD-lite [42] 98.0 99.0 95.3 99.8 31.6 41.1 99.0 25.9 82.7
MambaAD [15] 98.3 99.2 95.6 99.6 27.4 32.3 98.2 19.2 81.0
Hybrid UniAD [35] 98.7 99.3 96.1 99.5 21.6 32.9 98.2 19.7 80.5
RD++ [31] 96.3 97.9 93.5 99.7 26.8 36.4 98.5 22.2 80.6
DesTSeg [44] 96.5 98.4 93.2 91.2 55.9 55.5 81.4 38.5 83.5
sim_card_set Aug. DRAEM [37] 56.7 62.2 70.7 47.5 00.2 00.3 18.6 00.1 39.9
SimpleNet [27] 50.3 57.9 70.7 73.9 02.8 07.1 30.3 03.7 44.1
Emb. CFA [21] 50.0 57.1 70.7 79.5 00.9 03.0 39.3 01.5 45.0
CFLOW-AD [14] 91.5 92.9 85.3 98.2 30.3 35.2 87.9 21.3 76.4
PyramidalFlow [22] 77.9 75.3 77.1 66.0 01.0 00.6 31.0 00.3 50.7
Rec. RD [10] 91.7 92.0 85.2 98.4 39.3 43.6 89.0 27.9 78.6
DiAD [16] 72.6 53.7 61.5 89.9 01.7 05.8 60.4 03.0 49.4
ViTAD [40] 92.9 92.4 89.5 96.2 23.3 31.1 75.4 18.4 74.1
InvAD [42] 94.1 94.2 87.6 98.0 37.7 43.0 84.7 27.4 78.9
InvAD-lite [42] 94.5 95.0 88.2 98.6 46.7 46.9 88.9 30.6 81.4
MambaAD [15] 94.7 95.4 87.9 98.7 51.0 50.4 89.1 33.7 82.5
Hybrid UniAD [35] 91.4 91.9 84.3 97.8 34.7 41.2 85.8 25.9 77.0
RD++ [31] 93.7 94.1 87.2 98.6 40.4 43.2 89.9 27.6 79.8
DesTSeg [44] 92.9 94.4 86.3 95.7 69.4 65.7 81.0 49.0 84.6
Table A50: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
switch Aug. DRAEM [37] 46.5 52.1 68.6 44.5 00.3 00.6 12.9 00.3 35.2
SimpleNet [27] 47.8 53.1 68.6 69.9 01.2 03.1 49.3 01.6 43.7
Emb. CFA [21] 52.2 56.1 68.6 76.0 01.1 02.9 44.1 01.5 45.0
CFLOW-AD [14] 75.3 79.9 72.5 90.9 13.9 17.7 78.8 09.7 63.1
PyramidalFlow [22] 60.8 62.0 69.9 50.0 00.3 00.6 15.0 00.3 40.4
Rec. RD [10] 84.6 87.5 78.2 94.4 19.4 26.6 91.0 15.3 70.6
DiAD [16] 73.4 49.4 61.2 90.5 01.4 05.3 64.2 02.7 49.3
ViTAD [40] 87.8 90.0 80.6 96.4 28.8 39.2 88.5 24.4 74.7
InvAD [42] 95.8 96.8 91.1 98.9 43.9 51.5 96.0 34.7 83.6
InvAD-lite [42] 92.8 94.5 86.7 99.0 44.3 48.6 95.1 32.1 81.5
MambaAD [15] 92.4 94.3 86.3 98.4 34.9 42.2 93.8 26.7 79.2
Hybrid UniAD [35] 85.0 88.4 77.8 98.8 51.4 55.4 91.1 38.3 79.0
RD++ [31] 82.4 86.1 75.8 95.9 22.0 29.7 90.9 17.5 70.5
DesTSeg [44] 89.4 92.4 82.1 87.9 61.3 57.9 68.3 40.8 78.4
tape Aug. DRAEM [37] 50.0 35.4 54.6 38.0 00.1 00.2 11.4 00.1 29.5
SimpleNet [27] 46.3 37.4 54.6 82.8 01.3 04.0 42.8 02.0 39.4
Emb. CFA [21] 54.5 42.1 55.1 80.2 00.5 01.9 51.1 01.0 42.0
CFLOW-AD [14] 93.5 92.6 84.3 99.2 24.6 24.8 95.6 14.2 75.6
PyramidalFlow [22] 60.4 43.3 58.3 66.9 00.2 00.4 10.6 00.2 36.8
Rec. RD [10] 96.2 95.4 88.4 99.7 42.0 47.4 98.4 31.1 82.6
DiAD [16] 73.9 57.8 66.1 81.7 00.4 02.7 47.3 01.3 47.1
ViTAD [40] 93.6 92.6 84.6 99.4 30.7 38.7 95.9 24.0 78.2
InvAD [42] 97.8 96.9 91.9 99.7 38.9 44.0 98.3 28.2 82.9
InvAD-lite [42] 97.3 96.4 90.1 99.7 41.1 45.8 98.4 29.7 82.9
MambaAD [15] 97.1 96.2 89.6 99.8 45.9 48.4 98.4 31.9 83.7
Hybrid UniAD [35] 96.9 95.9 88.7 99.6 31.6 38.4 97.5 23.8 80.3
RD++ [31] 96.8 95.8 89.8 99.8 40.7 46.1 98.5 30.0 82.7
DesTSeg [44] 90.1 88.7 79.2 87.1 58.3 55.9 67.7 38.8 76.6
terminalblock Aug. DRAEM [37] 44.5 49.0 68.7 46.6 00.1 00.1 13.6 00.1 34.6
SimpleNet [27] 53.6 56.1 68.7 85.0 00.7 02.0 57.7 01.0 47.9
Emb. CFA [21] 55.9 59.9 68.7 89.9 00.7 02.8 64.0 01.4 50.4
CFLOW-AD [14] 81.1 84.3 76.0 97.0 12.1 17.7 86.5 09.7 66.9
PyramidalFlow [22] 57.8 57.5 70.0 91.0 00.5 01.7 67.3 00.8 50.9
Rec. RD [10] 89.5 90.0 83.2 99.5 27.9 35.6 98.0 21.7 76.4
DiAD [16] 62.1 36.4 47.8 75.5 00.1 01.1 38.5 00.5 37.4
ViTAD [40] 89.6 89.3 85.1 98.9 20.1 30.1 94.8 17.7 74.5
InvAD [42] 97.4 98.0 92.5 99.8 35.5 39.1 98.8 24.3 82.1
InvAD-lite [42] 96.9 97.6 91.4 99.8 34.6 38.2 98.8 23.6 81.6
MambaAD [15] 95.3 95.5 89.8 99.6 26.8 32.8 97.6 19.6 78.9
Hybrid UniAD [35] 87.0 88.6 80.6 99.1 22.7 30.8 94.6 18.2 73.6
RD++ [31] 89.3 90.2 82.3 99.5 27.7 35.0 97.8 21.2 76.1
DesTSeg [44] 68.4 73.3 69.4 69.3 22.2 27.2 40.3 15.7 55.1
toothbrush Aug. DRAEM [37] 47.1 50.4 69.8 39.4 00.6 01.6 14.2 00.8 34.9
SimpleNet [27] 61.7 63.5 70.2 79.1 03.0 07.3 46.3 03.8 49.5
Emb. CFA [21] 58.4 60.3 70.4 84.6 03.8 06.7 48.3 03.5 49.4
CFLOW-AD [14] 70.3 74.9 71.7 94.7 18.8 14.0 78.0 07.6 61.8
PyramidalFlow [22] 48.1 50.9 70.1 50.8 08.4 02.6 15.6 01.3 37.9
Rec. RD [10] 81.8 83.9 77.4 96.9 26.3 33.6 89.0 20.2 71.2
DiAD [16] 91.2 93.7 90.9 82.0 01.9 06.6 54.5 03.4 60.1
ViTAD [40] 81.8 84.8 76.5 96.7 25.3 34.9 87.1 21.1 71.0
InvAD [42] 88.2 90.0 81.6 97.6 29.5 37.8 91.0 23.3 75.3
InvAD-lite [42] 87.0 88.5 80.9 97.4 29.3 37.0 90.8 22.7 74.5
MambaAD [15] 86.2 87.5 80.7 97.6 30.1 37.9 91.9 23.4 74.6
Hybrid UniAD [35] 81.9 83.9 77.9 96.1 20.5 30.2 85.9 17.8 69.7
RD++ [31] 83.6 85.5 78.7 97.2 27.5 35.8 90.0 21.8 72.6
DesTSeg [44] 72.4 76.2 71.5 67.3 21.8 27.9 42.0 16.2 56.6
Table A51: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
toy Aug. DRAEM [37] 55.3 60.6 73.6 41.6 00.1 00.2 10.6 00.1 38.1
SimpleNet [27] 53.7 61.8 73.4 75.2 00.2 00.9 36.9 00.5 45.6
Emb. CFA [21] 47.8 56.6 73.4 78.3 00.3 01.0 40.6 00.5 44.7
CFLOW-AD [14] 60.6 68.8 73.7 87.5 02.1 08.2 56.9 04.3 53.2
PyramidalFlow [22] 56.2 65.2 73.4 79.8 00.4 01.6 47.4 00.8 48.6
Rec. RD [10] 70.4 74.8 75.7 95.1 05.1 13.0 82.6 07.0 61.3
DiAD [16] 66.2 57.3 59.8 82.1 01.1 04.2 50.3 02.1 45.9
ViTAD [40] 76.5 82.9 76.0 92.8 09.2 19.8 79.2 11.0 64.4
InvAD [42] 86.8 89.8 83.5 98.0 16.6 25.7 91.8 14.8 72.4
InvAD-lite [42] 84.4 88.9 80.3 96.6 18.1 29.0 89.2 17.0 71.4
MambaAD [15] 83.7 88.4 79.7 96.2 16.8 26.1 88.0 15.0 70.4
Hybrid UniAD [35] 68.5 75.1 74.4 93.6 04.8 11.8 73.0 06.3 59.2
RD++ [31] 73.4 78.1 76.4 96.0 07.8 17.7 85.6 09.7 63.9
DesTSeg [44] 71.5 77.5 75.8 67.5 12.9 19.6 47.5 10.8 55.9
toy_brick Aug. DRAEM [37] 51.5 42.9 57.9 48.5 00.1 00.3 11.3 00.1 32.9
SimpleNet [27] 57.4 48.8 57.9 85.9 04.5 10.6 47.0 05.6 45.8
Emb. CFA [21] 56.7 45.9 58.3 78.7 01.3 03.9 40.6 02.0 42.4
CFLOW-AD [14] 74.2 69.7 64.0 96.1 24.2 26.7 79.2 15.4 62.9
PyramidalFlow [22] 54.6 43.4 58.2 50.1 00.1 00.3 15.3 00.2 34.3
Rec. RD [10] 63.4 56.2 58.9 96.2 16.6 24.4 75.8 13.9 56.4
DiAD [16] 68.4 45.3 55.9 93.5 03.1 08.1 66.4 04.2 48.7
ViTAD [40] 69.1 64.3 60.2 94.6 20.6 29.4 72.9 17.2 59.5
InvAD [42] 74.1 67.4 64.4 97.6 25.4 31.4 82.4 18.6 63.9
InvAD-lite [42] 69.1 62.4 61.1 96.4 17.5 24.2 78.0 13.8 59.1
MambaAD [15] 70.6 64.0 61.8 96.6 18.9 26.8 75.9 15.5 60.0
Hybrid UniAD [35] 78.9 74.6 68.1 97.7 21.8 30.4 82.7 17.9 66.0
RD++ [31] 66.7 58.3 60.6 96.7 17.1 25.3 77.3 14.5 58.0
DesTSeg [44] 72.7 68.5 62.8 82.5 30.1 35.7 55.5 21.8 59.5
transistor1 Aug. DRAEM [37] 57.2 59.6 72.2 43.3 00.2 00.6 13.0 00.3 38.6
SimpleNet [27] 61.7 66.8 72.6 84.5 05.4 10.1 58.5 05.3 53.3
Emb. CFA [21] 53.9 63.0 72.1 87.7 03.5 08.7 56.8 04.5 51.1
CFLOW-AD [14] 91.8 94.7 86.2 98.1 26.2 28.3 90.7 16.5 75.9
PyramidalFlow [22] 49.1 56.2 72.4 77.1 00.4 00.9 35.3 00.4 43.8
Rec. RD [10] 90.9 93.7 85.2 99.1 30.5 36.5 95.5 22.3 77.7
DiAD [16] 73.1 63.1 62.7 88.6 07.2 15.3 58.1 08.3 52.6
ViTAD [40] 91.7 93.9 86.3 98.8 32.0 36.4 93.5 22.2 77.9
InvAD [42] 97.6 98.2 94.2 99.5 38.7 39.9 97.5 24.9 82.8
InvAD-lite [42] 96.4 97.5 92.1 99.5 40.5 42.4 97.4 26.9 82.6
MambaAD [15] 94.9 96.4 89.2 99.4 38.1 39.4 96.8 24.5 81.0
Hybrid UniAD [35] 93.7 95.9 88.8 98.9 27.3 33.0 94.2 19.8 78.1
RD++ [31] 93.4 95.5 89.0 99.3 34.6 38.5 96.2 23.8 79.9
DesTSeg [44] 88.0 91.6 82.5 87.9 44.0 44.1 67.0 28.3 74.1
u_block Aug. DRAEM [37] 55.2 35.2 49.4 43.5 00.1 00.2 10.0 00.1 30.0
SimpleNet [27] 58.4 40.0 51.2 72.8 00.6 02.5 40.2 01.3 39.4
Emb. CFA [21] 54.9 38.6 48.9 74.7 00.3 01.8 36.4 00.9 37.9
CFLOW-AD [14] 80.3 73.0 63.8 98.4 19.9 24.7 89.3 14.1 65.2
PyramidalFlow [22] 44.9 28.6 48.8 81.9 00.5 01.2 39.5 00.6 35.8
Rec. RD [10] 90.1 85.5 75.4 99.6 40.3 45.8 97.2 29.7 77.2
DiAD [16] 75.2 68.4 67.9 88.8 01.6 05.4 54.2 02.8 51.6
ViTAD [40] 84.3 78.2 67.8 99.0 30.3 38.4 90.8 23.8 70.7
InvAD [42] 92.9 89.5 80.5 99.6 33.4 41.3 96.7 26.0 77.7
InvAD-lite [42] 90.4 86.4 76.2 99.6 34.5 43.0 96.8 27.4 76.4
MambaAD [15] 90.0 85.8 74.8 99.5 33.2 42.8 96.1 27.2 75.7
Hybrid UniAD [35] 88.2 83.7 74.4 99.3 24.2 32.7 94.9 19.6 72.4
RD++ [31] 90.2 85.9 75.7 99.6 34.1 41.2 97.3 26.0 76.0
DesTSeg [44] 76.7 70.4 61.0 83.6 57.6 55.0 53.3 37.9 65.9
Table A52: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
usb Aug. DRAEM [37] 45.8 43.8 63.0 48.2 00.1 00.5 19.9 00.3 34.0
SimpleNet [27] 52.2 49.0 63.0 80.4 00.8 02.8 50.8 01.4 44.2
Emb. CFA [21] 56.6 48.6 63.2 85.9 00.7 02.4 51.6 01.2 45.6
CFLOW-AD [14] 68.0 69.8 63.4 94.8 13.0 15.8 75.0 08.6 58.4
PyramidalFlow [22] 48.3 44.5 63.0 79.2 00.2 00.2 39.8 00.1 40.9
Rec. RD [10] 85.2 84.6 75.6 98.1 26.8 35.2 91.2 21.3 72.3
DiAD [16] 58.9 37.4 45.7 78.0 01.0 03.1 28.0 01.6 36.0
ViTAD [40] 86.9 87.3 77.7 98.3 26.1 34.8 90.6 21.1 73.2
InvAD [42] 94.1 93.3 86.8 99.5 38.5 44.0 97.2 28.2 80.6
InvAD-lite [42] 93.1 92.7 85.8 99.5 37.0 43.7 96.8 27.9 79.9
MambaAD [15] 92.7 92.7 85.3 99.3 39.7 44.8 96.0 28.9 80.1
Hybrid UniAD [35] 80.8 81.8 70.9 98.0 21.0 30.9 85.8 18.3 68.4
RD++ [31] 85.5 85.8 75.6 98.0 27.2 36.3 90.8 22.2 72.7
DesTSeg [44] 87.6 88.3 78.5 91.5 45.0 50.9 70.5 34.2 74.6
usb_adaptor Aug. DRAEM [37] 54.0 45.6 56.5 49.5 00.1 00.1 20.0 00.0 34.7
SimpleNet [27] 49.9 40.6 56.7 52.3 00.1 00.4 19.2 00.2 33.5
Emb. CFA [21] 52.1 41.8 56.6 78.4 00.1 00.8 39.5 00.4 39.9
CFLOW-AD [14] 69.9 64.1 59.5 96.2 08.1 16.6 80.8 09.1 57.5
PyramidalFlow [22] 53.4 40.5 56.8 84.2 00.2 00.6 48.4 00.3 41.8
Rec. RD [10] 70.9 61.0 61.9 94.4 09.7 18.7 74.4 10.3 57.0
DiAD [16] 76.9 60.2 67.2 94.0 02.3 06.6 75.5 03.4 54.7
ViTAD [40] 76.1 66.1 66.2 90.8 05.9 13.9 56.2 07.5 55.6
InvAD [42] 80.0 74.4 66.8 96.5 15.8 24.3 79.3 13.8 63.9
InvAD-lite [42] 79.8 74.2 67.3 97.0 16.6 24.8 82.8 14.2 64.5
MambaAD [15] 79.1 75.7 66.0 97.0 15.8 24.9 81.6 14.2 64.2
Hybrid UniAD [35] 76.9 71.5 64.3 96.5 10.9 19.8 80.1 11.0 61.4
RD++ [31] 69.9 62.7 61.1 94.7 10.2 20.1 76.2 11.2 57.4
DesTSeg [44] 78.0 73.3 66.1 80.8 27.6 36.7 51.2 22.4 60.8
vcpill Aug. DRAEM [37] 47.3 37.7 56.2 31.2 00.3 00.8 07.2 00.4 28.5
SimpleNet [27] 62.8 52.2 58.6 82.0 06.5 12.0 47.6 06.4 47.4
Emb. CFA [21] 58.4 47.1 57.3 85.5 04.3 09.3 54.5 04.9 46.3
CFLOW-AD [14] 81.4 79.1 67.5 97.3 35.6 41.3 84.6 26.0 70.3
PyramidalFlow [22] 55.7 43.9 57.8 72.3 00.6 00.8 28.4 00.4 39.0
Rec. RD [10] 85.1 80.3 72.5 98.3 43.7 48.2 89.1 31.8 74.6
DiAD [16] 64.1 40.4 56.2 90.2 01.3 05.2 60.8 02.7 45.5
ViTAD [40] 82.2 80.5 70.4 97.6 44.4 50.0 81.3 33.3 73.0
InvAD [42] 90.1 88.0 78.2 98.9 50.4 55.2 91.1 38.1 79.7
InvAD-lite [42] 84.6 82.2 71.3 98.4 41.4 48.1 87.9 31.7 74.2
MambaAD [15] 88.5 87.6 77.6 98.7 47.9 52.0 89.4 35.1 78.3
Hybrid UniAD [35] 88.6 86.3 75.8 99.2 47.3 47.7 92.7 31.3 77.6
RD++ [31] 85.9 82.6 73.0 98.5 46.0 50.7 90.2 34.0 75.9
DesTSeg [44] 69.5 65.1 58.6 75.3 44.1 45.1 43.5 29.1 58.2
wooden_beads Aug. DRAEM [37] 52.8 44.9 60.2 43.9 00.1 00.3 12.5 00.1 33.4
SimpleNet [27] 56.2 52.2 60.2 75.5 01.2 04.4 36.2 02.2 42.8
Emb. CFA [21] 55.3 46.2 60.2 83.6 00.8 03.1 48.4 01.6 43.9
CFLOW-AD [14] 79.7 79.3 68.6 96.5 20.2 27.3 79.5 15.8 65.9
PyramidalFlow [22] 66.4 56.3 64.8 50.1 00.6 00.3 15.1 00.1 39.5
Rec. RD [10] 81.1 79.1 70.1 97.9 27.2 35.5 85.8 21.6 69.2
DiAD [16] 62.1 56.4 65.9 85.0 01.1 04.7 45.6 02.4 45.8
ViTAD [40] 82.8 82.3 72.9 97.1 31.0 37.4 79.4 23.0 70.3
InvAD [42] 86.6 85.6 76.2 98.4 35.5 40.3 88.0 25.3 74.2
InvAD-lite [42] 85.6 84.6 75.1 98.2 30.5 37.3 87.9 22.9 72.6
MambaAD [15] 83.3 82.3 72.7 98.1 32.4 39.7 85.3 24.8 71.7
Hybrid UniAD [35] 80.9 80.1 70.3 97.7 20.3 27.9 85.0 16.2 67.4
RD++ [31] 83.4 81.4 72.5 98.2 28.9 36.5 87.8 22.3 71.0
DesTSeg [44] 78.2 73.6 69.8 77.9 30.8 34.9 50.5 21.1 61.2
Table A53: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
woodstick Aug. DRAEM [37] 52.9 28.3 45.2 57.3 00.1 00.2 22.9 00.1 31.1
SimpleNet [27] 55.7 35.7 44.6 74.0 02.8 09.1 32.1 04.8 37.4
Emb. CFA [21] 57.8 33.2 46.5 70.2 00.9 04.1 26.5 02.1 35.6
CFLOW-AD [14] 71.3 60.5 52.8 92.6 32.1 38.3 67.9 23.7 59.6
PyramidalFlow [22] 50.1 28.1 44.4 50.0 00.0 00.2 15.0 00.1 28.6
Rec. RD [10] 77.5 61.9 58.5 97.8 31.2 38.8 85.0 24.0 64.6
DiAD [16] 74.1 66.0 62.1 90.9 02.6 08.0 60.7 04.2 52.1
ViTAD [40] 75.6 64.0 56.3 96.3 38.4 43.7 77.8 28.0 64.7
InvAD [42] 80.8 67.6 62.7 98.0 39.0 45.1 85.9 29.1 68.7
InvAD-lite [42] 78.7 65.9 59.3 97.6 36.3 41.0 83.8 25.8 66.3
MambaAD [15] 81.8 71.0 64.4 97.9 42.8 47.0 85.5 30.7 70.3
Hybrid UniAD [35] 81.1 73.7 64.6 94.8 41.5 46.0 79.2 29.9 69.3
RD++ [31] 77.2 61.2 58.2 97.9 30.6 38.0 85.3 23.4 64.2
DesTSeg [44] 80.2 70.6 60.5 91.9 61.5 60.5 69.9 43.4 70.7
zipper Aug. DRAEM [37] 46.5 61.6 76.7 36.4 00.7 01.5 11.3 00.8 37.0
SimpleNet [27] 56.8 67.7 76.7 52.7 01.2 03.8 24.3 01.9 43.8
Emb. CFA [21] 80.6 88.0 79.9 92.5 15.6 22.6 78.3 12.7 67.5
CFLOW-AD [14] 94.3 97.1 90.3 98.2 34.9 35.6 91.0 21.6 79.4
PyramidalFlow [22] 49.2 61.7 77.5 49.9 00.7 01.4 15.6 00.7 39.9
Rec. RD [10] 95.4 97.2 91.6 99.0 43.7 49.9 96.5 33.2 83.5
DiAD [16] 86.0 87.0 84.0 90.2 12.5 18.8 53.5 10.4 61.7
ViTAD [40] 98.8 99.3 95.9 99.1 47.8 50.3 96.4 33.6 85.7
InvAD [42] 98.8 99.2 96.0 99.2 54.3 56.8 97.0 39.7 87.4
InvAD-lite [42] 98.9 99.4 95.6 99.2 54.0 57.4 97.0 40.3 87.4
MambaAD [15] 99.4 99.6 97.1 99.3 58.4 60.9 97.6 43.8 88.9
Hybrid UniAD [35] 98.2 98.9 95.3 98.4 33.1 36.9 95.1 22.6 81.7
RD++ [31] 97.0 98.3 93.0 99.1 47.1 51.6 96.4 34.7 84.8
DesTSeg [44] 93.4 96.0 89.6 70.5 18.9 25.2 71.2 14.4 69.7
Avg Aug. DRAEM [37] 50.9 45.9 61.3 44.0 00.2 00.4 13.6 00.2 33.6
SimpleNet [27] 54.9 50.6 61.5 76.1 01.9 04.9 42.4 02.5 43.5
Emb. CFA [21] 55.7 50.5 61.9 81.3 01.6 03.8 48.8 02.0 45.0
CFLOW-AD [14] 77.0 75.8 69.9 94.8 17.6 21.7 80.4 12.4 63.9
PyramidalFlow [22] 54.4 48.0 62.0 71.1 01.2 01.1 34.9 00.5 40.9
Rec. RD [10] 82.7 79.3 74.1 97.2 25.2 32.8 90.0 20.0 70.0
DiAD [16] 75.6 66.4 69.9 88.0 02.9 07.1 58.1 03.7 52.6
ViTAD [40] 82.7 80.2 73.7 97.2 24.3 32.3 84.8 19.6 69.3
InvAD [42] 89.4 87.0 80.2 98.4 32.6 38.9 92.7 24.6 75.6
InvAD-lite [42] 87.2 85.2 77.8 98.0 31.7 37.9 92.0 23.8 74.2
MambaAD [15] 87.0 85.3 77.6 98.6 32.4 38.1 91.2 23.9 74.2
Hybrid UniAD [35] 83.1 81.2 74.5 97.4 23.3 30.9 87.1 18.6 69.6
RD++ [31] 83.6 80.6 74.8 97.7 25.9 33.6 90.7 20.5 70.8
DesTSeg [44] 79.3 76.7 70.7 80.3 36.9 40.3 56.1 26.2 64.5
Appendix M Detailed Quantitative Results on COCO-AD Dataset Under 100 epochs
Table A54: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
0 Aug. DRAEM [37] 48.8 74.5 85.0 51.9 16.6 27.7 18.1 16.1 49.0
SimpleNet [27] 63.4 82.1 85.0 78.2 39.1 44.6 44.4 28.7 64.2
Emb. CFA [21] 52.1 76.1 85.0 59.3 22.2 29.6 19.4 17.4 51.8
CFLOW-AD [14] 80.2 91.2 87.1 81.9 46.4 48.9 51.8 32.4 71.7
PyramidalFlow [22] 48.5 73.3 85.0 50.0 08.1 27.7 15.0 16.1 47.1
Rec. RD [10] 65.7 82.0 85.0 72.0 30.8 38.3 47.9 23.7 62.4
DiAD [16] 57.5 77.5 85.3 67.0 33.4 26.2 28.8 20.1 53.7
ViTAD [40] 84.8 93.9 88.0 83.1 55.3 53.2 51.0 36.2 74.8
InvAD [42] 74.3 88.1 85.1 78.0 41.6 44.6 52.3 28.7 68.3
InvAD-lite [42] 75.7 89.7 85.3 76.6 40.8 43.0 47.8 27.4 67.8
MambaAD [15] 74.0 88.9 85.2 75.4 38.2 41.5 48.9 26.2 66.9
Hybrid UniAD [35] 66.2 84.5 85.1 70.4 29.3 36.7 39.8 22.4 61.3
RD++ [31] 70.3 85.4 85.1 74.9 34.9 41.1 50.3 25.9 65.3
DesTSeg [44] 57.1 77.4 85.0 54.2 21.2 27.7 28.2 16.1 53.0
1 Aug. DRAEM [37] 47.4 42.7 60.9 48.1 02.4 05.8 13.6 03.0 33.9
SimpleNet [27] 45.0 39.8 60.9 53.3 03.5 06.5 19.3 03.4 34.6
Emb. CFA [21] 52.4 44.8 61.0 56.6 03.4 06.9 13.1 03.6 36.4
CFLOW-AD [14] 59.5 50.0 61.7 76.1 08.2 14.2 46.4 07.6 46.6
PyramidalFlow [22] 52.4 45.7 60.9 50.1 19.5 05.8 15.0 03.0 37.8
Rec. RD [10] 54.6 46.3 61.0 70.1 05.9 10.7 40.9 05.6 42.9
DiAD [16] 54.4 49.8 62.2 71.3 11.8 07.8 28.8 06.2 40.9
ViTAD [40] 58.5 51.2 61.8 78.9 16.3 23.4 38.1 13.3 48.2
InvAD [42] 55.7 47.9 61.0 73.4 07.9 13.4 43.0 07.2 44.6
InvAD-lite [42] 54.0 47.0 61.1 71.9 07.3 12.2 36.6 06.5 43.0
MambaAD [15] 54.9 47.7 61.0 70.8 06.2 11.0 38.4 05.8 43.1
Hybrid UniAD [35] 55.9 47.8 61.0 69.1 05.9 10.4 37.3 05.5 42.8
RD++ [31] 55.8 46.6 61.1 69.2 06.2 08.8 41.1 04.6 42.9
DesTSeg [44] 54.2 47.0 61.1 59.1 05.3 09.6 29.5 05.0 40.0
2 Aug. DRAEM [37] 58.1 37.0 51.0 50.4 05.6 11.9 15.2 06.3 34.7
SimpleNet [27] 55.3 36.8 49.9 55.7 07.8 12.9 22.7 06.9 36.1
Emb. CFA [21] 62.2 42.3 51.6 60.4 09.2 15.1 26.8 08.2 40.0
CFLOW-AD [14] 65.5 44.8 53.7 75.8 14.0 22.8 46.1 12.9 47.2
PyramidalFlow [22] 50.2 32.4 49.4 50.0 09.4 11.9 15.0 06.3 32.8
Rec. RD [10] 58.6 37.3 50.9 67.3 11.3 17.3 40.7 09.5 41.5
DiAD [16] 63.8 43.4 52.5 68.0 19.2 12.2 33.2 10.6 41.8
ViTAD [40] 67.5 53.9 52.6 80.5 28.9 35.9 40.2 21.9 52.2
InvAD [42] 66.5 45.8 54.4 73.5 15.5 23.5 49.5 13.3 48.0
InvAD-lite [42] 67.4 48.9 53.7 72.1 16.2 23.0 43.1 13.0 47.6
MambaAD [15] 65.2 45.3 53.3 70.9 13.9 21.4 45.0 12.0 46.2
Hybrid UniAD [35] 48.8 31.6 49.5 60.0 07.6 14.6 27.9 07.9 35.4
RD++ [31] 51.5 31.0 50.3 70.3 12.7 16.9 43.5 09.2 40.1
DesTSeg [44] 56.6 37.8 50.0 53.8 08.8 11.9 22.3 06.3 36.2
3 Aug. DRAEM [37] 59.6 39.8 53.3 49.1 05.8 12.4 14.3 06.6 35.6
SimpleNet [27] 58.6 38.7 53.4 53.7 07.5 12.9 18.0 06.9 36.6
Emb. CFA [21] 60.0 41.5 52.9 48.4 06.4 12.4 12.2 06.6 35.7
CFLOW-AD [14] 65.5 45.7 55.5 70.1 12.6 19.7 46.3 10.9 46.4
PyramidalFlow [22] 55.2 38.1 51.2 50.0 16.7 12.4 15.0 06.6 35.8
Rec. RD [10] 51.4 34.1 51.3 56.7 07.8 13.7 29.7 07.4 36.3
DiAD [16] 60.1 41.4 52.9 65.9 17.5 10.6 32.3 09.6 40.1
ViTAD [40] 56.6 38.2 52.3 62.3 10.0 16.4 27.1 09.0 39.0
InvAD [42] 60.2 42.0 53.1 59.7 09.0 14.9 38.5 08.0 41.1
InvAD-lite [42] 56.2 38.6 51.6 59.1 08.6 14.3 31.3 07.7 38.6
MambaAD [15] 57.1 38.9 51.5 58.6 08.5 14.2 34.0 07.6 39.0
Hybrid UniAD [35] 49.9 33.2 51.3 59.0 08.2 14.4 32.1 07.8 36.6
RD++ [31] 52.5 34.3 51.6 58.5 08.3 13.5 33.7 07.3 37.3
DesTSeg [44] 49.8 35.3 51.1 50.8 07.2 12.4 17.6 06.6 33.7
Table A55: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 53.5 48.5 62.5 49.9 07.6 14.4 15.3 08.0 38.3
SimpleNet [27] 55.6 49.4 62.3 60.2 14.5 19.2 26.1 11.5 42.9
Emb. CFA [21] 56.7 51.2 62.6 56.2 10.3 16.0 17.9 08.9 41.0
CFLOW-AD [14] 67.7 57.9 64.5 76.0 20.3 26.4 47.7 16.0 53.0
PyramidalFlow [22] 51.6 47.4 61.6 50.0 13.4 14.4 15.0 08.0 38.4
Rec. RD [10] 57.6 49.9 62.0 66.5 13.9 20.0 39.8 11.5 45.8
DiAD [16] 59.0 53.0 63.2 68.1 20.5 14.2 30.8 11.6 44.1
ViTAD [40] 66.9 59.3 63.7 76.2 27.6 32.2 39.1 20.1 53.5
InvAD [42] 64.2 56.0 63.4 71.2 18.5 24.1 45.8 14.3 50.5
InvAD-lite [42] 63.3 56.0 62.9 69.9 18.2 23.1 39.7 13.7 49.2
MambaAD [15] 62.8 55.2 62.8 68.9 16.7 22.0 41.6 12.9 48.8
Hybrid UniAD [35] 55.2 49.3 61.7 64.6 12.8 19.0 34.3 10.9 44.0
RD++ [31] 57.5 49.3 62.0 68.2 15.5 20.1 42.2 11.8 46.4
DesTSeg [44] 54.4 49.4 61.8 54.5 10.6 15.4 24.4 08.5 40.7
Appendix N Detailed Quantitative Results on MVTec AD Dataset Under 300 epochs
Table A56: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bottle Aug. DRAEM [37] 42.8 73.8 86.3 55.0 06.2 11.3 15.8 06.0 44.9
SimpleNet [27] 85.4 95.6 87.6 80.8 21.9 26.6 50.7 15.3 67.3
RealNet [45] 98.7 99.6 96.9 72.5 60.5 53.4 61.5 36.4 80.2
Emb. CFA [21] 36.7 71.5 86.3 23.6 03.5 10.9 02.4 05.8 37.5
PatchCore [29] 100. 100. 100. 98.9 79.0 77.1 96.1 62.7 93.9
CFLOW-AD [14] 99.9 100. 99.2 97.3 62.3 63.9 92.1 47.0 89.3
PyramidalFlow [22] 75.1 91.3 86.9 54.8 07.4 11.8 17.9 06.3 53.7
Rec. RD [10] 99.6 99.9 98.4 97.8 68.0 67.9 94.6 51.5 90.7
DiAD [16] 97.8 99.1 99.2 93.5 45.9 49.4 75.9 32.8 80.1
ViTAD [40] 100. 100. 100. 98.8 80.0 75.3 94.7 60.4 93.6
InvAD [42] 100. 100. 100. 98.4 74.8 72.2 95.2 56.6 92.6
InvAD-lite [42] 100. 100. 100. 98.4 74.8 73.2 95.5 57.7 92.7
MambaAD [15] 100. 100. 100. 98.9 80.3 77.1 96.2 62.7 94.1
Hybrid UniAD [35] 99.8 100. 99.2 98.0 68.9 69.8 94.5 53.6 91.2
RD++ [31] 99.8 100. 99.2 98.2 71.0 71.3 94.8 55.4 91.7
DesTSeg [44] 100. 100. 100. 95.7 88.0 80.8 92.6 67.8 94.6
cable Aug. DRAEM [37] 53.6 66.2 76.0 52.4 02.7 05.6 14.7 02.9 42.1
SimpleNet [27] 71.3 81.6 77.1 76.8 13.0 18.2 51.5 10.0 58.3
RealNet [45] 63.3 80.1 76.0 59.8 29.4 26.6 29.1 15.3 54.7
Emb. CFA [21] 56.4 71.8 76.0 57.1 07.4 08.1 16.3 04.2 45.1
PatchCore [29] 99.8 99.9 99.5 97.7 51.6 55.2 93.6 38.1 87.1
CFLOW-AD [14] 90.3 94.1 88.2 89.9 27.7 33.8 79.4 20.3 74.3
PyramidalFlow [22] 59.6 75.3 76.3 74.7 17.9 17.8 37.5 09.8 53.7
Rec. RD [10] 81.9 88.4 80.6 84.9 25.9 33.0 76.8 19.8 69.4
DiAD [16] 85.9 92.4 85.7 90.8 27.2 33.0 64.4 19.8 68.5
ViTAD [40] 98.8 99.2 96.3 95.4 41.0 45.0 88.9 29.1 82.8
InvAD [42] 99.2 99.5 96.8 97.3 47.0 51.1 92.1 34.3 85.2
InvAD-lite [42] 96.3 97.7 92.6 90.2 36.9 45.7 86.5 29.6 80.2
MambaAD [15] 98.9 99.3 95.6 95.9 42.6 48.0 90.9 31.6 83.6
Hybrid UniAD [35] 96.6 98.0 91.8 97.0 49.3 55.2 86.6 38.2 83.7
RD++ [31] 94.5 96.3 91.6 93.7 37.2 44.5 86.4 28.6 79.8
DesTSeg [44] 93.0 96.2 88.3 92.2 62.8 62.3 70.9 45.3 82.3
capsule Aug. DRAEM [37] 41.4 81.7 90.5 47.9 01.1 01.8 17.6 00.9 44.2
SimpleNet [27] 73.6 92.1 91.1 93.4 19.8 26.6 77.1 15.3 69.9
RealNet [45] 51.0 85.4 90.5 51.8 23.3 05.7 18.8 02.9 50.3
Emb. CFA [21] 40.2 79.1 90.5 81.8 04.0 08.4 56.0 04.4 53.7
PatchCore [29] 99.2 99.8 98.6 99.0 49.0 52.9 95.7 35.9 86.7
CFLOW-AD [14] 87.8 97.2 93.4 98.5 39.9 43.7 93.3 28.0 80.8
PyramidalFlow [22] 54.6 87.9 90.5 89.5 09.1 15.1 55.7 08.2 60.0
Rec. RD [10] 98.2 99.6 96.9 98.8 44.2 50.4 96.0 33.7 85.3
DiAD [16] 97.9 99.6 97.3 90.3 16.2 23.1 54.2 13.0 68.4
ViTAD [40] 95.8 99.1 95.4 98.1 41.6 47.5 92.2 31.2 83.3
InvAD [42] 98.2 99.6 96.9 99.0 47.5 51.1 95.8 34.3 85.8
InvAD-lite [42] 96.5 99.2 96.8 98.6 43.9 49.0 95.1 32.5 84.6
MambaAD [15] 94.8 98.9 95.4 98.4 43.8 47.1 93.9 30.8 83.6
Hybrid UniAD [35] 87.5 96.9 93.5 98.7 45.6 48.5 92.5 32.0 82.0
RD++ [31] 97.6 99.5 96.8 98.8 45.4 51.2 95.7 34.4 85.4
DesTSeg [44] 91.3 98.2 93.4 86.0 57.5 58.4 59.0 41.3 79.8
carpet Aug. DRAEM [37] 34.4 72.8 86.4 50.2 01.6 03.2 15.2 01.6 41.0
SimpleNet [27] 88.8 96.7 89.6 95.4 37.3 44.0 80.9 28.2 78.1
RealNet [45] 97.3 99.2 96.1 87.8 73.0 68.3 82.3 51.8 87.7
Emb. CFA [21] 54.3 81.4 86.4 41.0 03.1 08.0 25.7 04.2 46.7
PatchCore [29] 98.4 99.5 97.8 99.1 66.4 64.6 95.2 47.7 89.9
CFLOW-AD [14] 99.2 99.8 98.3 98.9 59.0 61.4 94.7 44.3 88.8
PyramidalFlow [22] 56.7 80.3 86.4 81.9 07.3 14.9 49.0 08.0 56.4
Rec. RD [10] 98.4 99.5 97.2 99.0 57.7 60.1 95.6 42.9 88.2
DiAD [16] 91.3 97.6 90.5 87.7 16.9 26.6 65.5 15.4 68.0
ViTAD [40] 99.4 99.8 99.4 98.9 61.1 64.2 94.3 47.3 89.6
InvAD [42] 98.8 99.7 97.2 99.1 60.9 62.2 95.6 45.2 89.0
InvAD-lite [42] 99.5 99.8 98.9 99.0 55.6 60.1 96.0 42.9 88.5
MambaAD [15] 99.9 100. 99.4 99.2 62.7 63.3 97.2 46.3 90.2
Hybrid UniAD [35] 99.8 99.9 99.4 98.5 52.9 52.6 95.2 35.7 87.2
RD++ [31] 97.4 99.3 96.6 98.8 52.1 56.3 94.5 39.2 86.6
DesTSeg [44] 93.6 97.9 94.4 96.1 72.3 67.5 94.2 50.9 88.9
Table A57: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
grid Aug. DRAEM [37] 64.2 84.1 85.5 49.3 00.6 01.4 14.3 00.7 47.2
SimpleNet [27] 40.5 69.4 85.7 38.2 00.5 01.4 03.6 00.7 38.1
RealNet [45] 97.4 99.3 98.2 81.0 47.8 49.2 70.3 32.6 80.2
Emb. CFA [21] 51.2 79.2 85.1 27.5 00.4 01.4 01.7 00.7 39.8
PatchCore [29] 99.4 99.8 98.2 98.9 40.8 44.3 94.0 28.4 84.3
CFLOW-AD [14] 89.4 96.5 88.9 93.6 21.3 26.9 82.3 15.5 73.8
PyramidalFlow [22] 78.4 90.2 88.7 90.0 13.9 22.4 75.5 12.6 68.1
Rec. RD [10] 97.8 99.3 96.6 99.2 45.9 47.6 97.3 31.2 85.2
DiAD [16] 99.7 99.9 98.3 85.7 03.9 09.7 57.4 05.1 65.0
ViTAD [40] 99.6 99.9 99.1 98.6 30.6 36.6 95.8 22.4 82.5
InvAD [42] 100. 100. 100. 99.2 45.5 47.3 97.3 31.0 86.2
InvAD-lite [42] 100. 100. 100. 99.3 50.0 50.2 97.7 33.5 87.2
MambaAD [15] 100. 100. 100. 99.2 48.6 48.3 97.6 31.8 86.7
Hybrid UniAD [35] 98.8 99.6 96.6 97.0 23.8 30.0 92.1 17.6 79.5
RD++ [31] 99.1 99.7 98.2 99.1 43.7 48.1 96.7 31.6 85.5
DesTSeg [44] 99.8 99.9 99.1 99.3 65.9 64.9 97.8 48.0 90.8
hazelnut Aug. DRAEM [37] 51.0 69.7 77.8 32.8 02.2 04.2 11.2 02.1 39.4
SimpleNet [27] 93.0 96.1 91.5 92.0 24.7 33.0 81.2 19.8 75.6
RealNet [45] 99.4 99.6 98.6 70.2 47.1 44.9 64.3 29.0 77.9
Emb. CFA [21] 66.1 79.8 78.9 38.0 04.7 08.1 26.7 04.2 47.2
PatchCore [29] 100. 100. 100. 99.0 61.2 63.6 96.2 46.6 90.0
CFLOW-AD [14] 99.9 99.9 99.3 98.5 59.9 56.9 95.7 39.8 88.7
PyramidalFlow [22] 88.5 93.6 87.7 88.1 31.8 39.3 75.5 24.5 74.3
Rec. RD [10] 100. 100. 100. 98.6 58.3 59.3 96.4 42.1 89.1
DiAD [16] 96.6 98.5 96.3 95.3 27.9 37.2 81.0 22.9 76.1
ViTAD [40] 99.8 99.9 98.6 98.9 64.1 62.6 95.2 45.5 89.8
InvAD [42] 100. 100. 100. 98.9 58.7 60.3 96.6 43.2 89.3
InvAD-lite [42] 100. 100. 100. 98.8 58.1 60.9 95.6 43.8 89.2
MambaAD [15] 100. 100. 100. 99.0 64.8 65.6 95.8 48.8 90.7
Hybrid UniAD [35] 99.9 100. 99.3 98.1 53.5 55.2 94.6 38.1 87.5
RD++ [31] 93.4 94.0 94.6 98.7 57.1 58.0 96.4 40.9 85.8
DesTSeg [44] 99.7 99.8 99.3 97.9 90.2 84.2 90.5 72.7 95.2
leather Aug. DRAEM [37] 80.3 91.5 87.8 48.8 00.8 01.3 16.7 00.7 51.7
SimpleNet [27] 87.3 95.2 90.1 82.2 05.3 11.7 74.6 06.2 67.2
RealNet [45] 100. 100. 100. 95.8 73.8 70.4 93.9 54.3 91.7
Emb. CFA [21] 58.1 80.3 86.7 23.7 00.4 01.3 01.7 00.6 40.9
PatchCore [29] 100. 100. 100. 99.4 52.1 51.2 97.8 34.5 87.6
CFLOW-AD [14] 100. 100. 100. 99.3 48.8 49.0 98.3 32.4 86.9
PyramidalFlow [22] 39.1 65.1 87.2 47.9 00.3 01.3 11.6 00.6 39.5
Rec. RD [10] 100. 100. 100. 99.3 38.2 45.4 97.8 29.4 85.1
DiAD [16] 97.4 99.0 96.8 91.5 06.4 13.3 77.9 07.1 68.9
ViTAD [40] 100. 100. 100. 99.5 51.3 55.4 98.0 38.3 88.0
InvAD [42] 100. 100. 100. 99.4 49.9 52.1 97.9 35.2 87.4
InvAD-lite [42] 100. 100. 100. 99.2 41.2 46.4 98.4 30.2 85.7
MambaAD [15] 100. 100. 100. 99.3 47.7 50.1 98.6 33.4 87.0
Hybrid UniAD [35] 100. 100. 100. 98.9 36.3 38.1 97.6 23.5 83.9
RD++ [31] 100. 100. 100. 99.2 39.7 44.4 97.4 28.5 85.1
DesTSeg [44] 100. 100. 99.5 99.5 80.6 73.4 98.8 58.0 94.0
metal_nut Aug. DRAEM [37] 61.2 86.4 89.4 39.8 08.4 21.2 10.8 11.9 49.5
SimpleNet [27] 74.3 92.3 90.1 79.9 33.0 36.8 52.7 22.5 68.1
RealNet [45] 63.4 90.4 89.4 51.6 37.6 21.0 31.2 11.7 58.2
Emb. CFA [21] 55.7 87.0 89.4 58.4 16.2 24.9 10.0 14.2 52.4
PatchCore [29] 100. 100. 100. 98.8 88.9 85.8 95.6 75.1 96.1
CFLOW-AD [14] 97.8 99.5 97.3 96.0 71.4 68.3 88.8 51.9 89.7
PyramidalFlow [22] 68.0 91.1 89.9 80.1 32.8 43.5 31.7 27.8 65.0
Rec. RD [10] 59.9 83.6 93.9 92.1 50.0 52.2 87.5 35.3 74.8
DiAD [16] 94.6 98.8 94.2 94.4 73.7 72.5 53.9 56.9 83.2
ViTAD [40] 99.9 100. 99.5 96.1 74.2 75.9 92.9 61.2 92.3
InvAD [42] 100. 100. 100. 97.7 81.3 82.0 93.7 69.5 94.3
InvAD-lite [42] 99.7 99.9 98.9 97.4 79.6 80.4 93.8 67.2 93.7
MambaAD [15] 99.8 100. 98.9 97.1 77.3 80.2 94.3 66.9 93.4
Hybrid UniAD [35] 98.3 99.5 98.9 92.6 48.8 63.0 79.4 46.0 84.9
RD++ [31] 100. 100. 99.5 96.6 74.0 77.4 93.4 63.1 92.6
DesTSeg [44] 100. 100. 100. 96.9 91.2 86.1 93.2 75.6 95.9
Table A58: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pill Aug. DRAEM [37] 68.0 91.5 92.5 46.0 03.3 06.6 18.4 03.4 51.3
SimpleNet [27] 71.4 93.5 91.6 94.1 53.0 50.7 77.7 34.0 77.2
RealNet [45] 63.3 91.6 91.6 51.1 41.1 06.5 28.3 03.4 57.0
Emb. CFA [21] 62.4 89.5 91.6 48.5 04.8 09.5 27.3 05.0 51.8
PatchCore [29] 97.0 99.4 96.5 98.1 76.8 74.0 96.1 58.7 91.9
CFLOW-AD [14] 85.0 96.8 92.3 96.5 56.2 55.8 90.9 38.7 83.1
PyramidalFlow [22] 68.7 93.0 91.8 77.6 13.9 24.2 61.2 13.8 64.4
Rec. RD [10] 97.0 99.5 97.2 97.5 63.1 64.5 96.0 47.6 89.1
DiAD [16] 94.7 99.0 97.2 93.0 50.4 50.6 65.7 33.9 78.6
ViTAD [40] 96.2 99.4 96.4 98.7 77.5 75.4 95.5 60.6 92.1
InvAD [42] 97.6 99.6 97.2 98.0 69.9 68.3 96.4 51.9 90.6
InvAD-lite [42] 96.5 99.4 96.1 98.0 71.7 70.6 96.6 54.6 90.8
MambaAD [15] 95.5 99.2 96.5 97.1 61.3 64.1 95.6 47.2 88.3
Hybrid UniAD [35] 94.1 98.9 95.1 95.1 40.7 45.9 94.7 29.8 82.6
RD++ [31] 97.6 99.5 97.9 98.2 73.4 70.9 96.3 54.9 91.5
DesTSeg [44] 88.9 97.8 93.0 89.5 83.5 77.4 48.9 63.1 84.0
screw Aug. DRAEM [37] 53.0 75.7 85.6 51.6 00.4 00.5 17.5 00.3 44.5
SimpleNet [27] 63.0 84.6 85.3 91.3 01.7 03.8 70.5 01.9 59.7
RealNet [45] 72.7 89.0 85.6 50.6 09.2 01.9 17.0 01.0 51.1
Emb. CFA [21] 64.2 82.6 85.9 32.2 00.2 00.8 13.1 00.4 44.6
PatchCore [29] 98.3 99.4 97.1 99.5 41.5 44.6 97.2 28.7 84.5
CFLOW-AD [14] 71.4 88.4 85.8 97.0 17.2 24.2 89.3 13.7 69.4
PyramidalFlow [22] 46.7 72.0 85.3 58.5 00.3 01.3 10.7 00.6 42.9
Rec. RD [10] 98.3 99.5 96.1 99.4 41.3 44.9 97.1 29.0 84.3
DiAD [16] 72.4 87.1 87.5 86.4 03.1 08.2 55.4 04.3 57.1
ViTAD [40] 91.2 97.0 92.0 98.9 33.2 39.8 93.6 24.9 79.9
InvAD [42] 97.2 99.0 95.9 99.6 49.5 51.6 97.5 34.8 86.0
InvAD-lite [42] 93.9 97.8 93.5 99.5 46.9 50.4 97.7 33.7 84.3
MambaAD [15] 93.9 97.6 94.0 99.4 50.1 50.2 97.4 33.5 84.7
Hybrid UniAD [35] 92.4 96.6 93.3 98.9 31.5 37.8 94.2 23.3 79.8
RD++ [31] 98.7 99.6 96.7 99.5 41.7 47.5 97.7 31.1 85.0
DesTSeg [44] 86.4 95.3 89.8 74.4 49.2 46.7 50.1 30.5 72.8
tile Aug. DRAEM [37] 63.4 83.5 83.8 48.8 06.4 13.2 14.2 07.0 48.8
SimpleNet [27] 86.5 94.9 88.1 75.0 27.3 35.7 49.0 21.7 68.3
RealNet [45] 98.7 99.6 98.2 91.9 84.8 78.4 87.3 64.5 92.2
Emb. CFA [21] 52.5 76.5 83.6 22.5 04.2 13.2 05.3 07.0 41.1
PatchCore [29] 100. 100. 99.4 96.4 58.7 67.1 86.6 50.5 88.5
CFLOW-AD [14] 99.7 99.9 99.4 95.9 55.9 62.1 86.7 45.0 87.4
PyramidalFlow [22] 82.9 93.8 85.4 79.5 46.6 41.2 49.2 25.9 70.7
Rec. RD [10] 97.9 99.2 95.2 95.2 48.4 60.4 85.7 43.2 84.9
DiAD [16] 93.0 96.6 96.3 76.2 25.0 33.8 50.5 20.3 67.4
ViTAD [40] 99.9 100. 98.8 96.5 56.5 68.7 88.0 52.3 88.5
InvAD [42] 100. 100. 100. 95.5 50.0 61.3 87.6 44.2 86.8
InvAD-lite [42] 98.2 99.3 96.5 94.4 47.9 55.8 82.4 38.7 84.1
MambaAD [15] 98.2 99.3 96.0 93.7 44.7 54.0 80.8 37.0 83.1
Hybrid UniAD [35] 100. 100. 100. 92.6 43.2 51.4 81.5 34.6 83.6
RD++ [31] 99.9 100. 99.4 96.3 52.8 64.3 88.7 47.4 87.6
DesTSeg [44] 100. 100. 100. 98.9 95.9 89.5 98.0 81.0 97.8
toothbrush Aug. DRAEM [37] 43.6 68.4 83.3 49.9 01.7 03.8 16.1 02.0 41.5
SimpleNet [27] 91.4 96.5 90.9 94.4 36.7 40.9 72.1 25.7 77.0
RealNet [45] 85.0 94.6 86.6 72.3 41.5 43.2 25.6 27.6 67.2
Emb. CFA [21] 43.6 72.2 83.3 52.5 05.6 09.5 39.9 05.0 46.6
PatchCore [29] 91.9 96.2 95.2 99.0 54.8 58.5 91.0 41.3 85.1
CFLOW-AD [14] 83.3 94.3 84.6 98.2 46.0 52.3 85.7 35.5 79.0
PyramidalFlow [22] 60.3 82.9 83.3 75.0 17.5 18.1 16.5 10.0 53.6
Rec. RD [10] 35.8 62.6 85.7 97.3 21.3 40.0 86.1 25.0 61.3
DiAD [16] 95.0 98.0 95.2 89.7 11.1 18.4 64.9 10.1 67.5
ViTAD [40] 100. 100. 100. 99.1 54.7 62.5 91.0 45.4 88.4
InvAD [42] 94.2 97.6 95.2 99.0 52.4 59.7 91.5 42.5 85.7
InvAD-lite [42] 98.1 99.2 96.8 99.0 50.3 59.7 91.2 42.6 86.5
MambaAD [15] 98.9 99.6 96.8 99.0 47.8 60.3 92.5 43.2 86.7
Hybrid UniAD [35] 90.3 95.6 95.2 98.5 39.1 49.9 87.9 33.2 81.3
RD++ [31] 96.7 98.7 95.1 99.1 57.0 61.5 92.2 44.4 87.1
DesTSeg [44] 93.9 97.7 92.3 97.5 74.5 70.7 81.1 54.6 87.8
Table A59: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
transistor Aug. DRAEM [37] 42.0 36.8 57.6 50.4 04.2 09.2 14.9 04.8 32.6
SimpleNet [27] 81.8 81.9 70.1 75.8 31.9 38.0 57.0 23.5 64.3
RealNet [45] 73.2 76.8 68.8 57.5 48.6 22.7 31.6 12.8 56.5
Emb. CFA [21] 70.0 59.8 66.0 53.8 08.0 12.0 24.0 06.4 44.9
PatchCore [29] 99.8 99.7 97.5 96.0 62.3 59.2 90.8 42.0 88.0
CFLOW-AD [14] 92.4 93.4 88.9 84.6 38.8 40.3 74.1 25.3 75.5
PyramidalFlow [22] 66.9 64.9 59.8 73.4 17.9 20.4 25.2 11.3 49.0
Rec. RD [10] 93.7 94.4 90.9 85.6 41.7 44.7 74.8 28.8 77.3
DiAD [16] 99.0 98.7 97.5 98.1 72.8 68.5 85.3 52.1 88.6
ViTAD [40] 98.5 97.9 92.1 93.7 57.6 54.5 77.0 37.5 83.4
InvAD [42] 99.7 99.6 96.4 97.1 69.4 66.3 93.1 49.6 90.0
InvAD-lite [42] 96.6 95.6 87.8 95.0 62.4 59.7 85.7 42.5 84.5
MambaAD [15] 100. 99.9 98.8 95.7 64.4 63.5 86.4 46.5 88.5
Hybrid UniAD [35] 100. 100. 100. 97.6 65.2 65.8 94.4 49.0 90.4
RD++ [31] 95.4 95.0 92.3 89.5 48.8 50.2 78.3 33.5 80.5
DesTSeg [44] 99.5 99.3 96.3 76.6 60.4 56.6 78.2 39.5 83.2
wood Aug. DRAEM [37] 52.4 82.0 86.3 47.8 03.7 07.5 13.7 03.9 45.9
SimpleNet [27] 84.4 93.4 91.5 74.6 15.1 23.3 53.6 13.2 65.7
RealNet [45] 99.6 99.9 99.2 87.4 77.0 73.7 85.4 58.3 90.2
Emb. CFA [21] 81.7 92.4 90.3 64.8 07.7 12.9 31.0 06.9 58.6
PatchCore [29] 98.7 99.6 96.7 95.2 52.1 52.3 91.3 35.4 85.5
CFLOW-AD [14] 98.9 99.7 97.5 94.4 47.0 50.4 91.0 33.7 84.7
PyramidalFlow [22] 86.8 95.2 89.9 71.2 35.9 15.2 50.3 08.2 66.9
Rec. RD [10] 99.4 99.8 98.3 95.5 49.4 51.7 91.6 34.9 85.6
DiAD [16] 93.5 97.5 94.4 83.1 15.0 22.8 58.0 12.8 66.3
ViTAD [40] 98.9 99.7 96.8 96.1 60.8 57.9 88.6 40.8 87.2
InvAD [42] 98.9 99.6 98.3 94.8 48.3 50.0 89.9 33.3 84.8
InvAD-lite [42] 97.3 99.1 95.9 94.9 47.3 48.4 88.7 31.9 83.6
MambaAD [15] 98.7 99.6 96.8 94.2 46.7 47.7 92.1 31.3 84.3
Hybrid UniAD [35] 98.6 99.6 96.7 93.3 38.4 43.5 88.0 27.8 82.0
RD++ [31] 99.3 99.8 98.3 95.3 50.2 51.1 90.3 34.3 85.4
DesTSeg [44] 99.2 99.7 98.4 94.1 83.6 76.8 94.8 62.3 93.2
zipper Aug. DRAEM [37] 77.0 90.6 89.8 60.1 03.7 04.1 26.4 02.1 54.7
SimpleNet [27] 95.0 98.7 93.5 92.6 39.2 44.6 78.3 28.7 79.7
RealNet [45] 80.6 94.8 88.1 65.5 55.0 40.5 42.0 25.4 69.3
Emb. CFA [21] 43.4 78.6 88.1 32.7 01.6 04.2 09.1 02.1 41.0
PatchCore [29] 99.6 99.9 98.7 98.9 63.4 64.6 96.4 47.8 90.1
CFLOW-AD [14] 96.0 98.7 97.1 98.0 50.1 55.7 92.8 38.6 85.7
PyramidalFlow [22] 60.9 87.5 88.1 70.0 06.0 08.1 32.7 04.2 54.0
Rec. RD [10] 99.5 99.9 99.2 98.4 52.4 59.9 95.2 42.8 88.0
DiAD [16] 71.8 89.5 90.2 84.0 13.5 23.8 55.5 13.5 61.2
ViTAD [40] 97.7 99.3 97.5 95.9 43.6 50.2 89.4 33.6 84.0
InvAD [42] 99.6 99.9 99.2 98.4 50.8 58.0 95.0 40.9 87.6
InvAD-lite [42] 99.3 99.8 97.5 98.5 57.7 61.0 95.5 43.9 88.5
MambaAD [15] 99.1 99.8 97.4 98.2 58.2 60.5 95.0 43.3 88.4
Hybrid UniAD [35] 95.3 98.5 96.2 97.0 37.7 46.6 91.4 30.4 82.4
RD++ [31] 68.2 89.3 92.2 98.3 50.8 57.9 94.5 40.7 79.3
DesTSeg [44] 99.4 99.8 98.3 94.7 80.8 74.1 90.9 58.8 92.1
Avg Aug. DRAEM [37] 55.2 77.0 83.9 48.7 03.1 06.3 15.8 03.3 45.3
SimpleNet [27] 79.2 90.8 87.6 82.4 24.0 29.0 62.0 17.8 67.6
RealNet [45] 82.9 93.3 90.9 69.8 50.0 40.4 51.2 28.5 70.9
Emb. CFA [21] 55.8 78.8 84.5 43.9 04.8 08.9 19.3 04.7 46.1
PatchCore [29] 98.8 99.5 98.4 98.3 59.9 61.0 94.2 44.9 88.6
CFLOW-AD [14] 92.7 97.2 94.0 95.8 46.8 49.6 89.0 34.0 82.5
PyramidalFlow [22] 66.2 84.3 85.1 74.2 17.2 19.6 40.0 11.4 58.1
Rec. RD [10] 90.5 95.0 95.1 95.9 47.1 52.1 91.2 35.8 82.6
DiAD [16] 92.0 96.8 94.4 89.3 27.3 32.7 64.4 21.3 71.0
ViTAD [40] 98.4 99.4 97.5 97.5 55.2 58.1 91.7 42.0 87.0
InvAD [42] 98.9 99.6 98.2 98.1 57.1 59.6 94.4 43.1 88.1
InvAD-lite [42] 98.1 99.1 96.8 97.3 55.0 58.1 93.1 41.7 86.9
MambaAD [15] 98.5 99.5 97.7 97.6 56.1 58.7 93.6 42.3 87.5
Hybrid UniAD [35] 96.8 98.9 97.0 96.8 45.0 50.2 91.0 34.2 84.2
RD++ [31] 95.8 98.0 96.6 97.3 53.0 57.0 92.9 40.5 85.9
DesTSeg [44] 96.3 98.8 96.1 92.6 75.8 71.3 82.6 56.6 88.8
Appendix O Detailed Quantitative Results on MVTec 3D Dataset Under 300 epochs
Table A60: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bagel Aug. DRAEM [37] 59.8 86.3 88.9 53.2 01.5 05.3 28.4 02.7 50.2
SimpleNet [27] 80.8 94.3 91.3 95.7 19.2 28.0 78.1 16.3 72.0
Emb. CFA [21] 68.2 90.2 88.9 71.5 12.8 23.4 34.8 13.3 59.0
PatchCore [29] 91.8 98.0 92.9 98.8 43.7 46.8 93.3 30.5 82.4
CFLOW-AD [14] 79.6 94.4 89.8 98.3 31.0 39.1 89.6 24.3 76.2
PyramidalFlow [22] 59.9 87.0 88.9 93.9 11.9 17.7 76.8 09.7 64.3
Rec. RD [10] 91.8 97.9 94.3 98.6 39.2 44.8 91.2 28.8 81.6
DiAD [16] 89.6 97.6 91.5 89.6 04.5 10.3 66.2 05.4 64.2
ViTAD [40] 89.3 97.1 93.4 98.6 38.6 43.6 91.8 27.9 80.7
InvAD [42] 91.3 97.8 92.5 98.9 44.7 48.5 93.8 32.0 82.7
InvAD-lite [42] 81.1 94.8 91.1 98.3 40.1 44.5 89.6 28.7 78.6
MambaAD [15] 91.9 98.0 92.0 98.6 38.3 43.6 93.1 27.9 81.2
Hybrid UniAD [35] 84.9 96.1 89.9 98.0 32.9 37.6 89.2 23.2 77.4
RD++ [31] 85.6 96.2 91.8 98.7 39.3 45.9 91.8 29.8 80.1
DesTSeg [44] 96.8 99.3 95.2 96.6 66.6 64.1 89.2 47.1 88.1
cable_gland Aug. DRAEM [37] 26.1 68.5 89.2 56.9 00.4 00.9 26.6 00.5 41.2
SimpleNet [27] 87.1 96.6 92.9 96.9 13.8 20.3 90.3 11.3 73.8
Emb. CFA [21] 42.9 80.5 89.2 43.1 01.3 05.5 23.3 02.8 44.6
PatchCore [29] 96.4 99.0 96.1 99.5 37.3 43.6 98.3 27.9 83.4
CFLOW-AD [14] 75.4 92.1 89.7 97.2 11.4 16.9 91.5 09.2 70.0
PyramidalFlow [22] 55.4 83.9 89.2 75.4 00.8 02.2 47.2 01.1 53.8
Rec. RD [10] 91.5 97.7 93.6 99.4 36.8 41.9 98.2 26.5 81.7
DiAD [16] 88.1 96.5 94.6 94.5 04.1 10.0 81.3 05.3 67.0
ViTAD [40] 85.9 95.0 92.7 98.8 24.2 32.4 96.2 19.4 77.1
InvAD [42] 97.3 99.3 97.7 99.5 40.6 46.9 98.7 30.6 84.8
InvAD-lite [42] 95.1 98.8 94.8 99.4 40.9 47.7 98.4 31.3 83.9
MambaAD [15] 94.6 98.7 94.9 99.4 40.5 46.1 98.4 29.9 83.6
Hybrid UniAD [35] 78.8 91.7 91.6 98.5 23.5 32.4 95.3 19.3 74.9
RD++ [31] 94.6 98.6 95.6 99.5 41.0 45.3 98.4 29.3 83.7
DesTSeg [44] 88.2 97.1 91.3 86.8 45.8 52.1 80.2 35.2 79.2
carrot Aug. DRAEM [37] 32.5 76.1 90.7 57.7 00.7 01.4 31.5 00.7 44.6
SimpleNet [27] 68.2 91.5 91.0 96.5 08.8 15.7 86.3 08.5 67.7
Emb. CFA [21] 48.3 85.8 90.7 56.3 02.6 05.8 32.4 03.0 49.6
PatchCore [29] 95.2 99.0 95.6 99.4 25.9 30.8 97.7 18.2 80.0
CFLOW-AD [14] 86.8 97.2 92.6 98.9 18.6 25.6 96.0 14.7 76.0
PyramidalFlow [22] 64.8 88.4 91.3 96.0 11.6 18.0 86.4 09.9 67.2
Rec. RD [10] 85.4 96.3 93.1 99.3 25.8 32.3 97.6 19.2 77.7
DiAD [16] 92.7 98.1 98.1 97.6 06.3 12.0 91.0 06.4 70.8
ViTAD [40] 79.4 94.0 92.5 98.8 19.5 26.2 95.9 15.1 74.4
InvAD [42] 89.9 97.5 94.9 99.2 24.8 29.6 97.4 17.4 78.4
InvAD-lite [42] 89.0 97.4 93.5 99.3 27.1 32.5 98.0 19.4 78.8
MambaAD [15] 90.7 97.6 95.2 99.4 27.1 32.0 98.0 19.0 79.3
Hybrid UniAD [35] 73.1 93.0 92.0 98.2 13.6 20.5 94.0 11.4 71.3
RD++ [31] 88.8 97.3 93.5 99.3 24.4 29.8 97.2 17.5 77.9
DesTSeg [44] 77.4 94.0 90.7 66.8 34.7 33.5 42.8 20.1 65.9
cookie Aug. DRAEM [37] 51.8 80.7 89.2 44.9 01.2 03.5 14.5 01.8 45.0
SimpleNet [27] 62.0 85.7 88.0 93.1 21.9 29.6 73.1 17.4 66.5
Emb. CFA [21] 48.2 81.2 88.0 49.8 05.2 11.2 22.9 05.9 47.4
PatchCore [29] 71.0 91.6 88.0 98.0 52.4 50.2 92.5 33.5 78.4
CFLOW-AD [14] 46.9 78.7 88.4 96.9 30.8 31.6 89.1 18.8 66.7
PyramidalFlow [22] 24.9 69.8 88.0 85.7 06.3 12.3 62.7 06.6 51.3
Rec. RD [10] 49.9 81.2 88.0 96.3 24.4 31.4 86.4 18.6 66.3
DiAD [16] 53.6 79.3 88.0 87.4 06.1 10.8 67.5 05.7 56.1
ViTAD [40] 71.9 91.7 88.5 97.5 46.3 47.5 87.1 31.2 76.8
InvAD [42] 62.8 87.8 88.0 98.0 46.3 46.6 91.5 30.4 75.1
InvAD-lite [42] 67.1 89.2 88.3 96.2 29.1 34.2 83.9 20.7 71.2
MambaAD [15] 67.2 89.2 88.0 96.5 36.7 39.2 83.4 24.4 72.7
Hybrid UniAD [35] 72.1 90.8 88.5 97.4 42.5 46.8 89.8 30.5 76.5
RD++ [31] 56.1 83.8 88.0 97.3 36.7 39.6 88.5 24.7 70.7
DesTSeg [44] 70.7 90.5 88.6 81.0 50.2 50.4 51.7 33.7 70.8
Table A61: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
dowel Aug. DRAEM [37] 41.9 77.5 88.9 34.5 00.5 02.5 17.0 01.2 41.5
SimpleNet [27] 86.2 96.5 90.0 96.6 13.7 21.9 85.2 12.3 72.6
Emb. CFA [21] 60.1 82.8 90.4 58.5 00.4 00.9 21.1 00.4 49.0
PatchCore [29] 98.2 99.6 97.5 99.3 44.5 46.7 96.3 30.5 85.1
CFLOW-AD [14] 94.9 98.8 95.2 98.9 29.9 34.1 94.8 20.6 80.4
PyramidalFlow [22] 61.3 85.5 89.3 93.5 05.6 10.7 74.2 05.7 62.4
Rec. RD [10] 99.4 99.9 98.6 99.7 51.6 51.4 98.8 34.5 87.3
DiAD [16] 69.9 90.8 92.4 78.6 04.5 09.7 48.2 05.1 56.3
ViTAD [40] 94.3 98.6 93.8 99.2 37.0 41.1 95.9 25.9 81.9
InvAD [42] 99.2 99.8 99.0 99.7 50.5 50.6 98.3 33.8 87.1
InvAD-lite [42] 96.5 99.2 96.2 99.6 50.0 50.0 97.4 33.3 85.8
MambaAD [15] 97.6 99.4 96.1 99.5 51.6 50.4 97.4 33.7 86.2
Hybrid UniAD [35] 94.0 98.6 93.0 99.1 35.6 41.2 96.1 25.9 81.6
RD++ [31] 99.6 99.9 99.0 99.7 52.2 50.6 98.7 33.9 87.4
DesTSeg [44] 82.3 95.5 89.8 84.4 39.8 45.8 63.0 29.7 73.7
foam Aug. DRAEM [37] 50.1 83.0 88.9 57.0 00.3 00.4 19.2 00.2 46.6
SimpleNet [27] 74.3 93.3 88.9 86.4 10.8 21.9 64.2 12.3 65.7
Emb. CFA [21] 61.7 88.7 88.9 44.4 06.4 16.1 32.5 08.8 52.3
PatchCore [29] 78.9 94.4 89.9 94.9 18.8 29.6 83.4 17.4 72.2
CFLOW-AD [14] 73.8 93.3 89.3 85.3 16.2 26.4 57.3 15.2 65.9
PyramidalFlow [22] 55.5 81.3 89.9 61.2 00.4 00.5 32.7 00.3 49.6
Rec. RD [10] 74.7 92.9 89.9 93.5 14.1 27.3 79.6 15.8 69.7
DiAD [16] 79.9 95.3 88.9 88.3 02.0 06.3 62.1 03.2 60.4
ViTAD [40] 79.4 94.4 90.8 94.0 14.2 25.6 80.7 14.7 70.9
InvAD [42] 87.4 96.8 91.6 93.8 18.9 30.1 81.3 17.7 74.0
InvAD-lite [42] 86.4 96.5 90.5 94.9 24.2 32.9 83.6 19.7 75.0
MambaAD [15] 84.3 95.9 89.9 95.3 25.2 33.6 84.8 20.2 74.9
Hybrid UniAD [35] 75.0 93.2 88.9 83.2 09.8 22.4 57.1 12.6 64.4
RD++ [31] 69.2 91.5 89.0 92.4 14.7 26.6 75.8 15.4 67.8
DesTSeg [44] 85.0 96.4 89.8 85.5 44.7 49.6 74.4 33.0 77.0
peach Aug. DRAEM [37] 48.5 82.0 89.1 47.4 00.6 02.3 19.3 01.2 45.3
SimpleNet [27] 65.4 87.9 89.1 94.3 05.8 12.1 80.3 06.5 64.5
Emb. CFA [21] 61.2 87.5 89.4 43.3 01.7 07.6 24.4 04.0 49.3
PatchCore [29] 75.2 92.7 91.2 98.2 26.4 32.4 93.3 19.3 74.5
CFLOW-AD [14] 74.4 91.8 90.5 97.9 16.2 18.4 92.3 10.2 70.9
PyramidalFlow [22] 68.7 90.0 89.6 95.2 06.1 11.7 83.8 06.2 66.0
Rec. RD [10] 64.3 86.3 91.7 98.3 12.9 20.5 93.2 11.4 68.5
DiAD [16] 92.2 94.1 97.7 95.6 04.3 08.7 84.7 04.5 68.2
ViTAD [40] 61.2 87.2 89.1 98.0 18.5 24.8 92.0 14.1 68.7
InvAD [42] 82.5 94.2 93.3 99.2 36.9 39.5 96.8 24.6 79.0
InvAD-lite [42] 85.6 96.1 90.6 99.0 33.2 36.9 95.6 22.6 78.5
MambaAD [15] 89.5 96.9 94.2 99.4 44.0 44.3 97.2 28.5 82.4
Hybrid UniAD [35] 70.0 89.8 91.2 98.0 18.5 23.5 92.6 13.3 70.9
RD++ [31] 73.0 89.2 92.6 98.8 22.4 29.9 94.7 17.6 73.2
DesTSeg [44] 80.3 94.9 89.8 84.9 46.6 46.6 70.6 30.4 75.3
potato Aug. DRAEM [37] 49.0 82.9 89.3 58.8 00.9 03.3 36.1 01.7 49.3
SimpleNet [27] 61.1 86.8 89.3 95.4 03.2 08.0 85.0 04.2 63.5
Emb. CFA [21] 41.6 75.9 89.3 14.3 00.5 01.4 06.9 00.7 37.4
PatchCore [29] 56.8 83.9 90.2 98.9 09.8 17.0 96.0 09.3 66.2
CFLOW-AD [14] 63.9 86.8 89.3 98.5 10.3 13.8 95.0 07.4 67.2
PyramidalFlow [22] 62.3 87.3 90.5 97.0 08.2 15.1 90.5 08.2 66.4
Rec. RD [10] 62.9 87.4 91.1 99.1 14.6 21.7 96.4 12.2 69.2
DiAD [16] 87.9 97.0 93.0 98.5 08.4 14.1 94.7 07.6 70.5
ViTAD [40] 57.9 86.4 89.8 98.8 13.0 18.2 95.6 10.0 67.2
InvAD [42] 67.0 90.2 90.6 99.1 15.7 23.6 96.4 13.4 70.7
InvAD-lite [42] 63.3 87.8 90.4 98.9 11.6 17.5 95.2 09.6 68.2
MambaAD [15] 58.6 85.3 90.2 99.0 16.4 22.8 95.5 12.9 68.2
Hybrid UniAD [35] 51.6 82.7 89.3 97.9 07.2 11.9 92.6 06.3 63.5
RD++ [31] 62.3 88.0 90.1 99.1 15.2 22.1 96.5 12.4 69.2
DesTSeg [44] 64.7 89.2 90.2 54.3 12.8 10.9 21.7 05.7 53.1
Table A62: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
rope Aug. DRAEM [37] 54.8 67.5 82.4 41.6 00.8 01.0 18.6 00.5 41.9
SimpleNet [27] 95.5 98.2 92.9 98.9 41.7 46.1 90.4 30.0 82.4
Emb. CFA [21] 93.0 97.2 92.4 89.8 38.0 45.6 70.3 29.6 77.6
PatchCore [29] 96.7 98.7 95.5 99.6 55.6 56.9 97.4 39.7 87.2
CFLOW-AD [14] 93.8 97.6 92.2 99.3 43.6 47.2 95.4 30.9 83.0
PyramidalFlow [22] 15.9 51.0 81.2 79.2 01.1 01.9 43.2 01.0 40.4
Rec. RD [10] 96.0 98.4 93.1 99.6 48.2 53.8 97.7 36.8 85.3
DiAD [16] 76.9 86.1 85.1 94.5 06.8 13.8 74.5 07.4 62.5
ViTAD [40] 97.0 98.9 95.6 98.5 41.3 42.7 89.0 27.1 82.5
InvAD [42] 96.0 98.5 95.5 99.6 52.4 56.0 96.5 38.9 86.4
InvAD-lite [42] 95.4 98.2 94.7 99.2 45.1 46.8 95.4 30.5 83.9
MambaAD [15] 95.4 98.0 93.1 99.3 49.2 48.9 95.8 32.4 84.4
Hybrid UniAD [35] 97.4 99.0 96.2 99.3 39.4 45.5 96.5 29.5 83.9
RD++ [31] 95.4 98.2 94.2 99.6 49.2 54.1 97.6 37.1 85.5
DesTSeg [44] 94.1 97.9 93.3 96.9 55.6 56.8 92.3 39.6 85.2
tire Aug. DRAEM [37] 77.4 89.3 93.5 55.1 00.3 00.9 22.9 00.5 53.3
SimpleNet [27] 68.4 87.4 89.0 86.0 05.0 10.0 59.0 05.3 60.8
Emb. CFA [21] 75.9 92.9 87.4 48.8 00.7 01.7 17.8 00.8 51.3
PatchCore [29] 80.1 93.6 87.9 99.1 22.6 31.1 95.8 18.4 74.7
CFLOW-AD [14] 58.0 83.3 87.9 97.6 07.7 14.8 91.0 08.0 64.6
PyramidalFlow [22] 56.6 83.6 88.3 94.7 07.2 13.9 82.5 07.4 62.9
Rec. RD [10] 76.8 91.6 89.1 99.1 18.9 28.6 96.0 16.7 73.2
DiAD [16] 57.7 78.0 87.4 93.2 04.9 09.6 80.6 05.0 58.7
ViTAD [40] 71.0 89.2 88.7 97.6 17.0 28.2 89.1 16.4 70.5
InvAD [42] 89.7 96.8 91.0 99.5 31.3 39.6 97.7 24.7 79.8
InvAD-lite [42] 91.5 97.6 93.2 99.4 42.9 46.5 97.0 30.3 82.8
MambaAD [15] 88.2 96.3 92.3 99.5 42.3 47.2 97.1 30.9 81.9
Hybrid UniAD [35] 72.8 91.3 88.8 98.1 14.0 24.6 90.9 14.0 70.6
RD++ [31] 76.5 92.1 88.0 99.2 20.3 27.7 96.4 16.1 73.2
DesTSeg [44] 85.6 95.9 90.1 84.3 40.9 42.5 65.1 27.0 74.4
Avg Aug. DRAEM [37] 49.2 79.4 89.0 50.7 00.7 02.2 23.4 01.1 45.9
SimpleNet [27] 74.9 91.8 90.2 94.0 14.4 21.4 79.2 12.4 68.9
Emb. CFA [21] 60.1 86.3 89.5 52.0 07.0 11.9 28.7 06.9 51.8
PatchCore [29] 84.1 95.1 92.5 98.6 33.7 38.5 94.4 24.5 78.4
CFLOW-AD [14] 74.8 91.4 90.5 96.9 21.6 26.8 89.2 15.9 72.1
PyramidalFlow [22] 52.5 80.8 88.6 87.2 05.9 10.4 68.0 05.6 58.4
Rec. RD [10] 79.3 93.0 92.3 98.3 28.6 35.4 93.5 22.1 76.1
DiAD [16] 78.8 91.3 91.7 91.8 05.2 10.5 75.1 05.6 63.5
ViTAD [40] 78.7 93.3 91.5 98.0 27.0 33.0 91.3 20.2 75.1
InvAD [42] 86.3 95.9 93.4 98.7 36.2 41.1 94.9 26.4 79.8
InvAD-lite [42] 85.1 95.6 92.3 98.4 34.4 39.0 93.4 24.6 78.7
MambaAD [15] 85.8 95.5 92.6 98.6 37.1 40.8 94.1 26.0 79.5
Hybrid UniAD [35] 77.0 92.6 91.0 96.8 23.7 30.6 89.4 18.6 73.5
RD++ [31] 80.1 93.5 92.2 98.4 31.5 37.2 93.6 23.4 76.9
DesTSeg [44] 82.5 95.1 90.9 82.1 43.8 45.2 65.1 30.2 74.3
Appendix P Detailed Quantitative Results on MVTec LOCO Dataset Under 300 epochs
Table A63: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
breakfast_box Aug. DRAEM [37] 48.8 68.9 77.2 50.1 06.5 12.5 14.3 06.7 42.9
SimpleNet [27] 89.4 94.5 85.9 89.6 58.6 56.1 75.4 39.0 79.9
Emb. CFA [21] 72.0 83.2 78.7 75.6 34.3 32.6 39.1 19.5 61.7
PatchCore [29] 77.1 88.4 78.0 89.4 55.4 49.3 76.1 32.8 74.4
CFLOW-AD [14] 75.8 86.7 77.8 88.8 50.1 48.1 71.7 31.7 72.4
PyramidalFlow [22] 61.2 78.0 77.2 70.6 34.9 32.5 34.4 19.4 57.6
Rec. RD [10] 63.4 79.2 77.2 83.7 26.6 35.6 63.9 21.6 62.9
DiAD [16] 56.8 74.3 77.2 81.6 22.9 34.2 35.5 20.6 54.6
ViTAD [40] 75.0 86.2 78.7 88.1 52.4 50.8 65.7 34.1 72.1
InvAD [42] 78.8 88.5 79.2 89.7 58.9 56.2 75.7 39.1 76.1
InvAD-lite [42] 71.7 83.9 78.0 88.3 47.6 47.7 73.6 31.3 71.1
MambaAD [15] 69.3 82.9 77.2 87.4 37.0 43.6 70.7 27.9 68.1
Hybrid UniAD [35] 77.5 88.4 77.4 88.0 51.8 50.1 72.8 33.4 73.4
RD++ [31] 68.2 81.7 77.2 86.6 39.7 43.9 67.0 28.2 67.5
DesTSeg [44] 84.0 91.1 82.2 77.5 64.0 57.9 68.0 40.8 76.3
juice_bottle Aug. DRAEM [37] 52.9 75.6 83.4 40.0 05.3 10.2 14.7 05.4 44.1
SimpleNet [27] 96.0 98.6 93.6 86.4 47.5 55.1 84.8 38.0 82.3
Emb. CFA [21] 73.0 88.2 83.4 85.1 26.9 29.1 57.4 17.1 65.6
PatchCore [29] 95.2 98.3 94.1 89.9 56.2 57.8 87.1 40.6 84.3
CFLOW-AD [14] 96.2 98.6 93.9 91.7 44.7 47.8 85.4 31.4 81.8
PyramidalFlow [22] 77.3 91.1 83.7 51.9 36.9 10.1 17.6 05.3 56.6
Rec. RD [10] 89.5 96.4 90.3 89.4 47.3 48.2 84.2 31.8 79.7
DiAD [16] 96.0 98.3 94.1 81.1 41.0 45.2 76.0 29.2 76.0
ViTAD [40] 90.5 96.4 88.6 88.6 48.1 49.2 83.4 32.6 79.6
InvAD [42] 96.1 98.6 94.8 90.9 56.1 58.7 88.2 41.5 85.0
InvAD-lite [42] 94.7 98.1 93.7 90.8 52.9 56.8 87.6 39.7 83.8
MambaAD [15] 93.2 97.6 91.7 89.8 53.9 55.8 86.5 38.7 82.8
Hybrid UniAD [35] 95.4 98.3 93.2 90.7 55.5 58.3 89.3 41.2 84.5
RD++ [31] 90.7 96.7 89.6 90.1 49.7 52.4 84.7 35.5 80.8
DesTSeg [44] 99.5 99.8 97.9 79.5 59.9 59.1 79.6 41.9 84.3
pushpins Aug. DRAEM [37] 62.2 66.8 72.2 53.9 00.8 01.9 11.5 01.0 42.0
SimpleNet [27] 76.5 83.6 74.9 55.8 02.7 07.4 59.6 03.8 54.9
Emb. CFA [21] 47.2 57.8 71.4 51.1 00.9 01.7 22.7 00.9 39.0
PatchCore [29] 74.3 82.8 72.7 56.1 03.1 07.5 63.6 03.9 54.6
CFLOW-AD [14] 63.3 71.4 71.4 39.5 01.2 03.4 52.8 01.7 46.5
PyramidalFlow [22] 45.7 52.1 71.5 55.2 00.8 01.7 18.8 00.9 37.8
Rec. RD [10] 71.5 79.6 72.3 51.1 02.2 07.5 68.2 03.9 53.4
DiAD [16] 69.0 73.8 72.9 61.4 01.6 04.6 49.7 02.4 47.6
ViTAD [40] 68.3 74.6 73.6 46.5 02.3 05.7 57.4 02.9 50.1
InvAD [42] 82.6 88.3 77.9 72.2 04.2 09.3 77.4 04.9 61.9
InvAD-lite [42] 78.2 84.1 76.4 64.0 03.2 08.5 70.3 04.5 58.0
MambaAD [15] 78.8 85.1 76.9 61.1 02.8 08.7 70.5 04.5 58.0
Hybrid UniAD [35] 71.3 79.9 72.2 46.3 02.3 08.1 62.0 04.2 52.1
RD++ [31] 75.7 82.7 73.4 54.1 02.7 07.7 67.4 04.0 55.1
DesTSeg [44] 72.2 81.3 71.4 55.1 04.3 10.1 64.0 05.3 54.2
screw_bag Aug. DRAEM [37] 49.1 66.4 78.4 45.6 06.0 13.8 10.6 07.4 41.8
SimpleNet [27] 65.3 78.1 78.3 55.5 09.2 14.8 48.8 08.0 53.0
Emb. CFA [21] 48.9 65.2 78.2 51.1 08.0 13.6 26.4 07.3 44.4
PatchCore [29] 70.1 82.8 78.2 72.1 13.1 22.0 62.1 12.4 59.7
CFLOW-AD [14] 64.8 78.7 78.4 68.9 11.4 20.6 66.8 11.5 57.9
PyramidalFlow [22] 54.2 66.8 78.2 49.9 05.8 13.6 15.9 07.3 43.8
Rec. RD [10] 63.7 78.1 78.2 71.6 12.6 21.8 52.0 12.3 56.4
DiAD [16] 81.7 91.1 80.6 71.4 12.0 22.2 58.9 12.5 59.7
ViTAD [40] 64.1 76.9 79.6 72.4 13.1 21.6 40.9 12.1 55.3
InvAD [42] 70.3 82.4 78.2 73.4 13.7 22.8 52.8 12.9 58.8
InvAD-lite [42] 63.7 77.4 78.5 72.7 13.2 22.4 48.1 12.6 56.2
MambaAD [15] 62.6 76.4 78.2 73.2 13.4 22.7 49.0 12.8 56.0
Hybrid UniAD [35] 65.0 77.1 78.5 67.5 11.3 19.3 62.8 10.7 56.9
RD++ [31] 64.9 79.3 78.2 69.8 12.2 20.5 51.3 11.4 56.3
DesTSeg [44] 70.0 81.1 79.0 50.3 08.6 13.6 46.4 07.3 53.2
Table A64: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
splicing_connectors Aug. DRAEM [37] 44.0 56.7 76.4 50.5 07.4 14.1 17.5 07.6 40.7
SimpleNet [27] 81.7 90.9 80.0 67.0 23.5 27.2 37.6 15.8 61.5
Emb. CFA [21] 61.6 74.3 76.5 60.3 12.4 16.6 36.1 09.1 51.1
PatchCore [29] 85.6 92.9 84.4 68.1 21.5 22.3 60.8 12.6 65.4
CFLOW-AD [14] 80.0 89.7 78.2 65.4 20.4 21.9 73.5 12.3 64.0
PyramidalFlow [22] 52.3 64.1 76.4 49.8 05.9 14.0 15.1 07.5 42.7
Rec. RD [10] 80.5 89.6 78.9 58.0 16.4 18.0 69.0 09.9 61.7
DiAD [16] 82.3 85.5 82.7 65.2 15.6 20.2 51.7 11.2 57.6
ViTAD [40] 82.8 91.0 79.8 70.6 23.6 25.4 62.5 14.6 65.0
InvAD [42] 83.8 91.6 81.7 62.4 21.4 23.7 70.1 13.4 65.0
InvAD-lite [42] 81.5 90.3 81.4 69.0 23.1 24.4 67.5 13.9 65.2
MambaAD [15] 87.3 93.7 84.8 76.0 32.7 33.7 66.1 20.3 70.4
Hybrid UniAD [35] 84.2 91.6 83.0 80.5 42.2 43.2 66.6 27.5 72.2
RD++ [31] 78.7 89.0 78.1 60.0 16.6 18.0 69.3 09.9 61.5
DesTSeg [44] 80.1 89.8 78.8 56.3 22.1 18.5 37.9 10.2 58.3
Avg Aug. DRAEM [37] 51.4 66.9 77.5 48.0 05.2 10.5 13.7 05.6 42.3
SimpleNet [27] 81.8 89.1 82.5 70.9 28.3 32.1 61.2 20.9 66.3
Emb. CFA [21] 60.6 73.7 77.6 64.7 16.5 18.8 36.4 10.8 52.4
PatchCore [29] 80.5 89.0 81.5 75.1 29.9 31.8 69.9 20.4 67.7
CFLOW-AD [14] 76.0 85.0 79.9 70.9 25.6 28.4 70.0 17.7 64.5
PyramidalFlow [22] 58.1 70.4 77.4 55.5 16.8 14.4 20.4 08.1 47.7
Rec. RD [10] 73.7 84.6 79.4 70.7 21.0 26.2 67.5 15.9 62.8
DiAD [16] 77.2 84.6 81.5 72.1 18.6 25.3 54.4 15.2 59.1
ViTAD [40] 76.2 85.0 80.1 73.2 27.9 30.6 62.0 19.3 64.4
InvAD [42] 82.3 89.9 82.4 77.7 30.9 34.1 72.8 22.4 69.4
InvAD-lite [42] 78.0 86.8 81.6 77.0 28.0 32.0 69.4 20.4 66.9
MambaAD [15] 78.2 87.1 81.8 77.5 28.0 32.9 68.6 20.8 67.1
Hybrid UniAD [35] 78.7 87.1 80.9 74.6 32.6 35.8 70.7 23.4 67.8
RD++ [31] 75.7 85.9 79.3 72.1 24.2 28.5 67.9 17.8 64.2
DesTSeg [44] 81.2 88.6 81.8 63.7 31.8 31.9 59.2 21.1 65.3
Appendix Q Detailed Quantitative Results on VisA Dataset Under 300 epochs
Table A65: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
candle Aug. DRAEM [37] 58.1 59.6 67.6 45.7 00.1 00.6 16.7 00.3 38.8
SimpleNet [27] 83.3 85.8 77.7 91.3 07.6 17.6 80.4 09.7 65.7
RealNet [45] 53.6 63.1 66.7 53.5 08.8 09.0 53.6 04.7 46.2
Emb. CFA [21] 86.6 85.0 80.0 79.4 02.6 09.0 58.0 04.7 60.6
CFLOW-AD [14] 92.4 93.2 85.4 98.8 14.7 24.0 93.7 13.6 74.1
PyramidalFlow [22] 80.6 75.9 77.0 84.5 02.2 07.7 79.7 04.0 60.7
Rec. RD [10] 91.8 92.7 83.9 99.0 19.1 30.5 95.2 18.0 75.2
DiAD [16] 99.3 99.3 96.0 80.1 01.0 02.6 41.4 01.3 60.0
ViTAD [40] 90.0 91.5 82.4 96.1 17.1 27.6 86.5 16.0 72.4
InvAD [42] 95.2 94.9 92.2 99.3 23.3 33.5 95.7 20.1 78.5
InvAD-lite [42] 97.2 97.5 92.9 99.1 20.5 30.9 96.2 18.3 78.8
MambaAD [15] 95.5 95.8 89.6 99.1 22.4 32.4 96.1 19.3 78.1
Hybrid UniAD [35] 97.3 97.4 91.6 99.3 24.4 33.9 94.8 20.4 79.3
RD++ [31] 92.7 92.8 87.6 99.3 25.5 36.3 95.4 22.2 77.6
DesTSeg [44] 89.8 89.8 83.0 74.5 39.5 41.4 49.9 26.1 69.4
capsules Aug. DRAEM [37] 71.9 80.7 77.5 39.1 00.3 01.0 10.7 00.5 44.7
SimpleNet [27] 66.9 79.4 77.2 90.0 21.2 28.3 57.5 16.5 61.9
RealNet [45] 81.6 90.5 79.8 74.6 34.2 41.1 39.6 25.9 65.7
Emb. CFA [21] 59.6 70.8 78.7 60.6 00.5 01.0 24.3 00.5 45.7
CFLOW-AD [14] 57.5 73.8 77.5 94.5 27.9 27.8 65.9 16.1 61.8
PyramidalFlow [22] 60.5 72.8 77.8 81.9 02.3 06.5 61.3 03.3 54.2
Rec. RD [10] 81.0 89.7 81.6 99.3 57.3 59.4 91.9 42.3 80.5
DiAD [16] 66.7 80.5 79.7 94.3 08.0 16.4 55.3 08.9 57.3
ViTAD [40] 80.2 88.2 80.2 98.3 31.1 41.3 76.0 26.0 72.3
InvAD [42] 87.2 92.8 84.5 99.6 67.1 61.9 94.9 44.8 84.5
InvAD-lite [42] 88.8 92.5 87.4 99.3 60.7 59.1 94.9 42.0 84.0
MambaAD [15] 83.4 88.4 83.8 98.8 32.8 41.6 87.6 26.2 75.2
Hybrid UniAD [35] 73.5 85.6 79.1 98.4 52.1 53.6 80.8 36.6 75.3
RD++ [31] 79.5 88.6 79.8 99.1 56.8 59.7 91.4 42.6 79.7
DesTSeg [44] 86.5 92.9 85.6 95.7 38.7 42.8 77.8 27.2 76.0
cashew Aug. DRAEM [37] 19.8 54.8 80.0 53.0 01.5 03.4 40.8 01.7 38.1
SimpleNet [27] 87.1 93.3 87.2 98.5 68.0 65.2 85.0 48.3 84.2
RealNet [45] 89.5 95.4 86.8 53.5 25.4 10.1 29.0 05.3 60.0
Emb. CFA [21] 78.3 87.5 83.8 82.3 23.2 29.5 28.8 17.3 62.1
CFLOW-AD [14] 95.6 98.1 92.7 99.0 52.5 54.4 94.6 37.4 85.3
PyramidalFlow [22] 89.0 93.8 90.6 49.8 00.6 02.2 15.2 01.1 54.0
Rec. RD [10] 93.3 96.5 91.5 92.3 44.9 50.1 84.5 33.4 80.9
DiAD [16] 99.8 99.9 99.0 89.1 45.6 50.1 21.5 33.4 72.1
ViTAD [40] 87.6 94.0 87.5 98.3 62.6 61.4 79.3 44.3 82.6
InvAD [42] 96.0 98.3 93.9 97.6 65.5 65.5 89.7 48.7 87.8
InvAD-lite [42] 92.7 96.4 88.8 94.6 55.4 57.9 88.9 40.7 83.4
MambaAD [15] 92.5 96.3 90.3 94.7 50.8 55.0 86.2 37.9 82.4
Hybrid UniAD [35] 90.4 95.2 87.7 97.9 44.5 48.9 89.0 32.4 80.6
RD++ [31] 93.5 96.4 92.7 95.9 56.3 56.7 85.1 39.6 83.8
DesTSeg [44] 86.6 93.1 86.5 84.4 48.2 52.0 55.8 35.2 74.4
chewinggum Aug. DRAEM [37] 33.5 58.0 79.0 32.3 00.3 01.1 04.2 00.6 33.2
SimpleNet [27] 94.6 97.6 91.3 97.9 39.7 45.3 74.7 29.3 79.5
RealNet [45] 89.0 95.3 87.0 82.2 63.8 61.5 42.3 44.5 76.4
Emb. CFA [21] 75.9 87.8 82.7 84.4 16.6 26.5 43.1 15.3 62.4
CFLOW-AD [14] 97.2 98.8 95.2 99.2 61.9 60.3 87.8 43.1 87.2
PyramidalFlow [22] 74.5 84.3 83.3 50.1 00.7 01.1 15.1 00.6 48.7
Rec. RD [10] 95.5 98.0 93.8 98.7 58.4 61.0 78.1 43.9 84.9
DiAD [16] 97.6 98.4 96.6 68.5 15.6 30.6 18.6 18.0 60.8
ViTAD [40] 95.2 97.8 92.3 97.7 60.7 58.5 71.4 41.3 83.6
InvAD [42] 97.7 99.0 95.8 98.6 57.2 59.0 80.0 41.9 85.6
InvAD-lite [42] 97.6 99.0 95.3 98.3 56.7 56.4 78.8 39.3 84.9
MambaAD [15] 97.3 98.8 94.7 97.3 27.4 38.0 78.2 23.4 78.6
Hybrid UniAD [35] 99.1 99.5 96.8 99.3 62.4 59.6 88.2 42.5 87.9
RD++ [31] 96.2 98.3 93.0 98.8 58.8 61.8 80.1 44.7 85.4
DesTSeg [44] 96.8 98.7 95.3 97.5 78.3 70.2 68.4 54.1 87.8
Table A66: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
fryum Aug. DRAEM [37] 46.2 65.7 80.0 51.2 03.5 08.7 25.5 04.6 43.1
SimpleNet [27] 82.0 91.5 84.0 93.4 40.6 41.7 78.5 26.4 74.7
RealNet [45] 80.1 90.6 81.0 57.0 32.4 17.8 22.8 09.8 58.2
Emb. CFA [21] 65.4 81.3 80.6 76.7 16.2 22.1 41.3 12.4 57.4
CFLOW-AD [14] 91.2 96.0 90.3 97.6 53.8 55.8 87.4 38.7 83.1
PyramidalFlow [22] 80.2 90.0 82.2 89.3 19.2 18.3 74.5 10.1 67.2
Rec. RD [10] 95.8 98.1 92.4 97.0 48.1 51.8 93.1 35.0 84.0
DiAD [16] 97.0 98.3 95.6 97.5 59.0 61.8 68.9 44.7 82.6
ViTAD [40] 94.3 97.5 91.1 97.5 47.2 52.0 88.9 35.1 82.8
InvAD [42] 97.0 98.6 93.1 97.5 52.5 56.2 92.1 39.1 85.4
InvAD-lite [42] 96.0 98.1 92.8 97.3 50.6 53.2 93.2 36.3 84.6
MambaAD [15] 96.3 98.2 93.1 97.1 49.7 52.9 92.4 36.0 84.4
Hybrid UniAD [35] 91.0 96.0 87.9 97.3 45.3 52.4 85.5 35.5 80.9
RD++ [31] 95.7 98.1 92.2 97.2 49.9 53.6 92.0 36.6 84.3
DesTSeg [44] 92.3 96.6 88.5 67.0 47.6 41.3 35.4 26.0 70.1
macaroni1 Aug. DRAEM [37] 89.4 86.6 81.5 38.7 00.0 00.1 06.2 00.0 48.5
SimpleNet [27] 65.8 60.9 65.6 91.4 01.5 04.6 66.5 02.4 52.5
RealNet [45] 69.9 75.2 66.7 54.7 14.6 12.4 19.7 06.6 48.0
Emb. CFA [21] 56.0 57.4 66.9 87.7 00.4 02.4 56.0 01.2 48.4
CFLOW-AD [14] 81.8 78.7 75.7 99.1 07.9 13.2 95.2 07.1 66.3
PyramidalFlow [22] 57.6 53.5 70.0 91.0 00.5 02.0 60.0 01.0 49.4
Rec. RD [10] 95.2 93.7 88.8 99.8 23.6 31.1 97.2 18.4 77.7
DiAD [16] 96.3 94.9 92.4 55.2 00.0 00.1 03.6 00.0 48.9
ViTAD [40] 85.3 83.5 76.0 98.6 07.1 15.4 90.7 08.3 67.3
InvAD [42] 95.2 93.4 88.7 99.7 21.8 30.5 97.3 18.0 77.4
InvAD-lite [42] 94.2 92.5 86.2 99.8 19.3 27.7 97.7 16.1 76.0
MambaAD [15] 91.9 89.4 83.2 99.3 13.4 21.9 95.4 12.3 72.8
Hybrid UniAD [35] 88.1 86.2 80.0 99.4 10.8 20.1 95.6 11.2 70.6
RD++ [31] 92.8 91.2 86.5 99.7 24.0 32.4 96.7 19.3 76.7
DesTSeg [44] 93.5 89.9 90.6 73.3 15.2 25.7 36.6 14.7 64.5
macaroni2 Aug. DRAEM [37] 57.7 53.4 63.4 58.9 00.0 00.1 25.1 00.1 39.6
SimpleNet [27] 57.9 53.2 63.3 89.8 00.4 01.7 66.4 00.9 48.9
RealNet [45] 59.0 60.2 67.3 53.5 06.8 08.3 20.1 04.3 42.2
Emb. CFA [21] 58.6 56.7 66.9 82.1 00.1 00.8 48.5 00.4 46.8
CFLOW-AD [14] 67.1 65.5 65.6 97.4 01.9 06.3 89.9 03.3 57.5
PyramidalFlow [22] 62.5 59.1 71.6 83.2 00.1 00.2 35.0 00.1 47.0
Rec. RD [10] 85.6 79.3 81.0 99.6 12.3 20.2 97.0 11.3 69.6
DiAD [16] 56.0 46.8 68.2 56.8 00.0 00.1 15.8 00.0 34.8
ViTAD [40] 77.2 70.3 74.2 98.1 03.5 09.5 88.1 05.0 61.8
InvAD [42] 88.3 83.9 83.2 99.6 11.4 19.1 97.8 10.6 71.1
InvAD-lite [42] 88.1 84.6 80.4 99.7 11.7 19.8 98.2 11.0 70.9
MambaAD [15] 80.5 74.3 75.7 96.5 02.0 07.3 87.1 03.8 62.5
Hybrid UniAD [35] 81.3 81.1 71.9 98.3 05.6 13.2 92.8 07.0 65.3
RD++ [31] 82.8 76.9 80.2 99.5 09.6 18.3 96.2 10.1 67.9
DesTSeg [44] 72.2 70.4 73.1 62.8 10.3 17.2 36.0 09.4 51.7
pcb1 Aug. DRAEM [37] 80.6 78.0 78.2 53.2 00.6 00.9 21.4 00.5 49.0
SimpleNet [27] 85.2 87.0 78.1 98.6 73.1 69.7 79.0 53.5 81.8
RealNet [45] 82.7 83.6 77.5 73.6 40.1 43.6 27.3 27.9 63.7
Emb. CFA [21] 58.0 63.9 66.7 88.4 19.7 29.9 39.1 17.5 53.6
CFLOW-AD [14] 94.9 95.2 88.2 99.1 73.0 68.8 88.3 52.4 87.5
PyramidalFlow [22] 67.7 64.6 71.9 53.4 34.1 11.3 15.3 06.0 48.3
Rec. RD [10] 96.4 96.4 91.8 99.5 66.8 62.2 96.1 45.1 88.0
DiAD [16] 91.7 94.1 89.5 95.7 27.7 39.5 56.5 24.6 70.7
ViTAD [40] 95.6 94.1 91.3 99.4 65.6 61.9 89.2 44.8 86.3
InvAD [42] 96.8 96.2 92.9 99.8 80.3 75.4 96.7 60.5 91.7
InvAD-lite [42] 97.7 97.2 94.2 99.8 79.3 73.5 97.1 58.1 91.9
MambaAD [15] 96.0 94.7 94.1 99.8 79.8 72.8 94.9 57.2 90.9
Hybrid UniAD [35] 96.3 95.5 92.8 99.4 68.1 63.3 90.5 46.3 87.6
RD++ [31] 96.3 96.0 90.9 99.4 64.9 59.1 95.5 42.0 87.1
DesTSeg [44] 78.6 77.3 74.5 96.0 66.5 64.3 68.6 47.3 75.3
Table A67: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
pcb2 Aug. DRAEM [37] 35.0 44.1 66.7 44.9 00.2 00.4 10.0 00.2 31.2
SimpleNet [27] 82.5 83.8 75.6 94.9 07.2 15.1 76.0 08.2 64.5
RealNet [45] 87.0 89.3 79.2 72.2 13.9 23.0 40.3 13.0 61.3
Emb. CFA [21] 57.2 65.9 66.7 80.0 02.7 07.5 47.7 03.9 48.9
CFLOW-AD [14] 92.1 93.0 85.6 96.5 12.6 18.9 82.5 10.5 71.4
PyramidalFlow [22] 66.8 71.1 68.7 92.7 07.4 08.6 74.9 04.5 57.4
Rec. RD [10] 97.3 96.9 94.2 97.7 21.1 30.1 91.7 17.7 78.1
DiAD [16] 89.9 91.7 82.9 82.2 04.4 12.2 48.1 06.5 58.8
ViTAD [40] 90.0 89.0 84.2 98.0 12.5 20.4 82.6 11.4 70.6
InvAD [42] 97.0 97.3 93.9 99.1 15.7 23.4 92.8 13.2 76.9
InvAD-lite [42] 96.2 96.4 90.1 99.0 11.7 23.4 92.5 13.3 75.4
MambaAD [15] 96.1 96.3 91.0 99.0 14.2 24.1 91.2 13.7 75.8
Hybrid UniAD [35] 93.5 94.1 87.2 98.3 13.0 19.5 86.0 10.8 72.9
RD++ [31] 97.9 98.3 96.0 98.0 25.5 33.4 91.7 20.0 79.8
DesTSeg [44] 82.8 83.2 78.1 94.4 19.1 24.0 69.8 13.6 66.6
pcb3 Aug. DRAEM [37] 45.6 50.3 66.7 41.2 00.2 00.5 11.0 00.3 33.7
SimpleNet [27] 84.2 86.2 76.6 97.3 12.9 22.4 82.7 12.6 68.1
RealNet [45] 87.0 90.8 82.9 76.9 24.6 32.2 40.9 19.2 65.3
Emb. CFA [21] 56.3 60.8 67.1 85.6 13.8 22.5 49.6 12.7 52.1
CFLOW-AD [14] 80.8 84.1 74.2 96.4 20.6 26.0 81.1 14.9 67.9
PyramidalFlow [22] 66.3 68.6 67.6 90.7 05.0 12.3 58.0 06.5 54.5
Rec. RD [10] 96.8 96.7 90.6 98.0 23.8 34.6 93.6 20.9 78.6
DiAD [16] 94.0 94.4 88.1 87.2 04.0 10.7 41.8 05.7 60.0
ViTAD [40] 91.2 91.1 86.1 98.3 22.6 28.7 88.0 16.8 74.4
InvAD [42] 97.7 97.9 93.3 99.2 15.8 28.0 93.2 16.2 77.7
InvAD-lite [42] 96.4 96.5 90.6 99.2 17.8 28.0 93.6 16.3 77.1
MambaAD [15] 95.2 95.5 88.7 99.1 18.9 27.4 92.5 15.8 76.3
Hybrid UniAD [35] 89.8 90.6 83.3 98.5 13.4 23.9 86.1 13.5 71.7
RD++ [31] 95.7 95.6 90.8 98.0 30.9 34.1 93.3 20.5 79.1
DesTSeg [44] 95.8 96.3 89.8 90.6 27.5 31.6 61.2 18.7 73.3
pcb4 Aug. DRAEM [37] 73.1 66.8 77.3 36.1 00.4 01.3 06.1 00.7 41.7
SimpleNet [27] 95.4 95.1 91.1 92.2 13.9 23.4 67.1 13.2 71.5
RealNet [45] 92.1 92.1 84.4 68.6 37.2 38.7 31.4 24.0 66.8
Emb. CFA [21] 89.5 87.4 83.7 91.4 07.6 14.1 70.0 07.6 66.3
CFLOW-AD [14] 98.9 98.8 95.7 96.8 22.2 29.7 85.6 17.4 78.2
PyramidalFlow [22] 62.6 64.8 69.5 92.7 06.9 12.4 78.4 06.6 56.6
Rec. RD [10] 99.9 99.9 98.5 97.7 30.0 35.4 88.4 21.5 81.2
DiAD [16] 99.8 99.7 99.0 95.4 09.9 18.4 82.4 10.1 72.1
ViTAD [40] 98.9 98.5 97.0 99.0 42.2 47.7 93.5 31.3 84.4
InvAD [42] 99.9 99.9 98.5 98.5 44.1 43.4 90.6 27.7 84.3
InvAD-lite [42] 99.7 99.7 97.5 98.8 49.7 48.7 92.0 32.2 85.6
MambaAD [15] 99.7 99.7 96.4 98.6 44.0 44.9 90.6 28.9 84.1
Hybrid UniAD [35] 99.4 99.4 96.2 97.6 34.3 39.0 85.2 24.2 81.2
RD++ [31] 99.9 99.9 99.0 97.1 27.4 35.9 85.8 21.9 80.6
DesTSeg [44] 98.0 98.0 93.0 94.4 53.8 52.1 74.9 35.3 82.6
pipe_fryum Aug. DRAEM [37] 63.2 76.9 80.8 46.0 01.3 03.1 14.3 01.6 44.9
SimpleNet [27] 83.7 92.1 84.2 97.9 64.1 61.6 76.6 44.5 80.9
RealNet [45] 79.1 91.2 80.3 64.1 48.5 37.7 39.6 23.2 65.5
Emb. CFA [21] 63.7 81.4 80.0 96.8 61.0 59.4 77.8 42.3 74.4
CFLOW-AD [14] 97.6 98.8 95.0 99.2 61.6 61.4 94.9 44.3 88.2
PyramidalFlow [22] 59.7 76.7 80.0 89.8 15.7 21.5 64.2 12.1 60.0
Rec. RD [10] 98.2 99.1 96.6 99.1 55.1 58.1 95.9 41.0 87.5
DiAD [16] 97.7 98.7 98.0 98.6 55.1 58.1 77.9 41.0 83.5
ViTAD [40] 97.6 98.9 94.6 99.5 65.1 66.3 94.9 49.6 89.2
InvAD [42] 99.5 99.7 98.0 99.5 69.7 67.1 95.1 50.5 91.0
InvAD-lite [42] 98.3 99.2 95.6 99.2 61.5 60.5 95.1 43.4 88.4
MambaAD [15] 98.8 99.3 97.6 98.8 52.4 53.0 94.2 36.0 86.6
Hybrid UniAD [35] 97.0 98.5 95.1 98.9 50.1 55.0 93.9 37.9 85.7
RD++ [31] 94.1 96.9 91.9 98.9 54.9 57.0 94.2 39.9 85.3
DesTSeg [44] 95.0 97.5 93.1 86.8 74.3 69.7 55.9 53.5 83.4
Table A68: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Avg Aug. DRAEM [37] 56.2 64.6 74.9 45.0 00.7 01.8 16.0 00.9 40.6
SimpleNet [27] 80.7 83.8 79.3 94.4 29.2 33.1 74.2 22.1 69.5
RealNet [45] 79.2 84.8 78.3 65.4 29.2 27.9 33.9 17.4 59.9
Emb. CFA [21] 67.1 73.8 75.3 83.0 13.7 18.7 48.7 11.3 56.5
CFLOW-AD [14] 87.2 89.5 85.1 97.8 34.2 37.2 87.3 24.9 75.7
PyramidalFlow [22] 69.0 72.9 75.8 79.1 07.9 08.7 52.6 04.7 54.8
Rec. RD [10] 93.9 94.8 90.4 98.1 38.4 43.7 91.9 29.0 80.5
DiAD [16] 90.5 91.4 90.4 83.4 19.2 25.0 44.3 16.2 63.5
ViTAD [40] 90.3 91.2 86.4 98.2 36.4 40.9 85.8 27.5 77.3
InvAD [42] 95.6 96.0 92.3 99.0 43.7 46.9 93.0 32.6 82.6
InvAD-lite [42] 95.3 95.8 91.0 98.7 41.2 44.9 93.2 30.6 81.8
MambaAD [15] 93.6 93.9 89.8 98.2 34.0 39.3 90.5 25.9 79.0
Hybrid UniAD [35] 91.4 93.3 87.5 98.5 35.3 40.2 89.0 26.5 78.2
RD++ [31] 93.1 94.1 90.0 98.4 40.4 44.8 91.4 29.9 80.6
DesTSeg [44] 89.0 90.3 85.9 84.8 43.3 44.4 57.5 30.1 73.0
Appendix R Detailed Quantitative Results on BTAD Dataset Under 300 epochs
Table A69: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
01 Aug. DRAEM [37] 75.0 88.7 86.7 44.8 02.8 06.3 13.0 03.3 50.1
SimpleNet [27] 99.2 99.7 98.0 94.6 38.0 42.6 64.2 27.1 79.4
RealNet [45] 100. 100. 100. 94.2 63.8 63.6 84.3 46.6 88.2
Emb. CFA [21] 98.2 99.4 99.0 94.6 37.3 43.0 64.5 27.4 79.4
PatchCore [29] 99.6 99.8 99.0 97.1 55.5 56.6 75.8 39.5 85.4
CFLOW-AD [14] 95.5 98.4 93.6 94.8 36.2 45.6 66.4 29.5 78.3
PyramidalFlow [22] 90.3 94.1 94.2 51.8 14.7 06.3 17.1 03.3 57.7
Rec. RD [10] 99.9 100. 99.0 97.6 55.3 57.2 82.6 40.1 86.4
DiAD [16] 99.5 99.8 99.0 90.1 24.3 30.8 72.6 18.2 73.7
ViTAD [40] 98.3 99.3 96.9 96.5 55.0 55.2 69.2 38.2 83.6
InvAD [42] 99.9 100. 99.0 97.5 57.4 58.8 80.7 41.7 86.6
InvAD-lite [42] 98.2 99.4 96.9 97.1 56.7 59.5 75.4 42.3 85.2
MambaAD [15] 97.7 99.2 96.0 97.1 55.3 59.0 75.9 41.8 84.7
Hybrid UniAD [35] 99.8 99.9 99.0 97.4 55.7 57.7 80.1 40.6 86.1
RD++ [31] 99.4 99.8 97.9 97.5 53.5 56.9 80.6 39.8 85.6
DesTSeg [44] 98.3 99.4 96.0 91.9 29.9 43.9 69.9 28.1 78.4
02 Aug. DRAEM [37] 63.6 92.6 92.8 47.9 06.6 10.4 16.2 05.5 51.6
SimpleNet [27] 83.2 97.1 93.2 94.6 49.2 50.3 47.2 33.6 75.7
RealNet [45] 80.0 96.7 93.0 90.4 31.8 51.8 44.0 35.0 72.2
Emb. CFA [21] 80.3 96.5 93.2 94.4 47.5 51.2 47.6 34.5 75.1
PatchCore [29] 83.8 97.3 92.8 96.0 59.3 59.8 54.1 42.7 79.3
CFLOW-AD [14] 79.9 96.7 92.8 96.3 59.3 62.8 54.3 45.7 79.0
PyramidalFlow [22] 73.7 95.2 92.8 81.9 63.2 52.9 32.5 35.9 72.4
Rec. RD [10] 82.9 97.2 92.8 96.7 63.3 64.2 58.3 47.2 80.8
DiAD [16] 72.9 95.2 92.8 87.8 29.4 37.1 45.1 22.8 65.8
ViTAD [40] 84.0 97.3 93.2 96.7 72.9 67.2 51.2 50.6 81.8
InvAD [42] 86.4 97.7 93.8 96.4 60.2 61.1 57.1 44.0 80.7
InvAD-lite [42] 84.6 97.5 92.8 96.4 59.8 60.7 58.5 43.5 80.2
MambaAD [15] 81.6 97.0 92.8 96.1 54.2 59.2 57.6 42.0 78.6
Hybrid UniAD [35] 84.0 97.3 92.8 95.1 43.5 53.6 57.9 36.6 76.9
RD++ [31] 84.3 97.4 93.2 96.6 64.1 63.4 57.3 46.4 81.0
DesTSeg [44] 82.7 97.0 92.8 93.6 58.6 51.0 52.7 34.2 77.4
03 Aug. DRAEM [37] 69.2 49.0 59.1 50.4 00.3 01.7 19.2 00.9 38.5
SimpleNet [27] 99.7 96.9 90.6 99.4 35.6 38.1 97.2 23.5 81.7
RealNet [45] 99.3 92.2 84.5 76.9 48.6 51.1 45.3 34.4 73.7
Emb. CFA [21] 99.8 97.5 91.5 99.5 55.4 54.9 95.0 37.8 86.2
PatchCore [29] 99.8 97.3 92.1 99.5 50.1 48.2 98.0 31.8 85.2
CFLOW-AD [14] 99.3 91.8 82.7 99.4 42.5 38.5 97.5 23.9 80.4
PyramidalFlow [22] 92.0 55.9 50.0 94.8 05.2 11.2 79.7 05.9 56.8
Rec. RD [10] 99.6 93.3 89.6 99.7 52.7 52.6 98.9 35.7 85.1
DiAD [16] 98.3 69.9 86.1 97.9 07.7 13.1 93.3 07.0 66.6
ViTAD [40] 99.6 94.3 90.9 99.6 47.0 47.1 98.1 30.8 83.9
InvAD [42] 99.6 94.1 88.2 99.7 58.6 56.4 98.7 39.3 86.2
InvAD-lite [42] 99.8 96.8 93.5 99.6 50.7 51.0 98.2 34.2 85.8
MambaAD [15] 99.5 92.3 90.3 99.6 44.2 47.1 98.3 30.8 83.2
Hybrid UniAD [35] 99.8 97.8 93.1 99.7 57.8 55.1 98.7 38.0 87.4
RD++ [31] 99.7 95.6 91.8 99.8 55.4 54.3 98.8 37.3 86.4
DesTSeg [44] 99.6 93.9 92.5 98.9 28.8 20.6 96.1 11.5 78.2
Avg Aug. DRAEM [37] 69.3 76.8 79.5 47.7 03.2 06.1 16.1 03.2 46.7
SimpleNet [27] 94.0 97.9 93.9 96.2 41.0 43.7 69.6 28.1 78.9
RealNet [45] 93.1 96.3 92.5 87.2 48.0 55.5 57.9 38.7 78.1
Emb. CFA [21] 92.7 97.8 94.6 96.2 46.7 49.7 69.0 33.2 80.2
PatchCore [29] 94.4 98.2 94.6 97.5 55.0 54.9 76.0 38.0 83.3
CFLOW-AD [14] 91.6 95.6 89.7 96.9 46.0 49.0 72.7 33.0 79.2
PyramidalFlow [22] 85.3 81.7 79.0 76.1 27.7 23.4 43.1 15.0 62.3
Rec. RD [10] 94.1 96.8 93.8 98.0 57.1 58.0 79.9 41.0 84.1
DiAD [16] 90.2 88.3 92.6 91.9 20.5 27.0 70.3 16.0 68.7
ViTAD [40] 94.0 97.0 93.7 97.6 58.3 56.5 72.8 39.9 83.1
InvAD [42] 95.3 97.3 93.7 97.9 58.7 58.8 78.8 41.7 84.5
InvAD-lite [42] 94.2 97.9 94.4 97.7 55.7 57.0 77.4 40.0 83.7
MambaAD [15] 92.9 96.2 93.0 97.6 51.2 55.1 77.3 38.2 82.2
Hybrid UniAD [35] 94.5 98.4 94.9 97.4 52.4 55.5 78.9 38.4 83.5
RD++ [31] 94.5 97.6 94.3 98.0 57.7 58.2 78.9 41.2 84.3
DesTSeg [44] 93.5 96.7 93.8 94.8 39.1 38.5 72.9 24.6 78.0
Appendix S Detailed Quantitative Results on MPDD Dataset Under 300 epochs
Table A70: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
bracket_black Aug. DRAEM [37] 45.3 54.3 77.0 52.8 00.2 00.5 24.9 00.3 39.2
SimpleNet [27] 78.7 85.3 80.4 94.9 03.6 09.1 89.3 04.7 65.3
RealNet [45] 72.0 84.5 76.4 64.4 09.5 16.7 45.9 09.1 55.9
Emb. CFA [21] 87.9 92.7 84.1 93.8 02.9 08.8 83.8 04.6 67.8
PatchCore [29] 88.4 92.7 83.9 98.5 11.0 21.0 96.7 11.8 72.6
CFLOW-AD [14] 64.8 74.6 76.7 95.8 02.5 05.4 86.5 02.8 59.8
PyramidalFlow [22] 75.3 81.8 79.3 89.5 00.4 01.0 66.2 00.5 59.0
Rec. RD [10] 83.7 87.0 85.7 97.6 05.9 12.3 93.3 06.5 68.9
DiAD [16] 100. 100. 100. 90.8 00.5 01.4 63.8 00.7 65.2
ViTAD [40] 78.7 86.1 81.5 96.0 02.7 07.8 89.8 04.1 65.6
InvAD [42] 88.2 93.2 83.8 98.1 08.2 15.0 95.5 08.1 71.3
InvAD-lite [42] 81.2 87.5 81.7 95.8 07.4 16.2 90.6 08.8 68.0
MambaAD [15] 82.8 88.6 80.7 94.8 04.6 13.1 89.1 07.0 67.2
Hybrid UniAD [35] 96.3 96.6 96.9 94.8 00.9 02.2 84.5 01.1 71.1
RD++ [31] 82.3 86.2 81.7 97.6 04.4 09.6 93.8 05.0 67.4
DesTSeg [44] 92.9 95.0 91.3 91.9 06.0 14.2 87.6 07.6 71.5
bracket_brown Aug. DRAEM [37] 53.9 69.1 80.3 54.4 00.7 01.3 28.8 00.6 44.5
SimpleNet [27] 92.4 95.3 93.5 94.9 08.8 17.0 87.4 09.3 72.9
RealNet [45] 87.4 90.4 92.6 51.4 10.1 03.5 17.8 01.8 55.4
Emb. CFA [21] 93.5 95.8 96.2 91.5 07.4 17.7 69.6 09.7 70.9
PatchCore [29] 95.6 97.2 95.3 98.6 26.2 32.0 95.4 19.0 79.5
CFLOW-AD [14] 77.8 81.6 87.7 96.4 07.2 14.0 92.5 07.5 67.4
PyramidalFlow [22] 70.0 78.6 84.4 93.9 05.5 12.6 75.4 06.8 62.3
Rec. RD [10] 83.1 87.0 91.7 96.0 14.2 23.2 94.0 13.1 72.1
DiAD [16] 88.6 94.2 89.5 94.8 22.5 29.9 69.8 17.6 69.9
ViTAD [40] 86.0 87.8 91.9 98.3 16.2 26.2 95.0 15.1 73.7
InvAD [42] 96.8 98.0 98.1 96.3 18.7 29.4 91.4 17.2 78.3
InvAD-lite [42] 93.4 96.2 92.7 98.5 21.4 31.6 95.1 18.8 77.9
MambaAD [15] 97.2 98.5 97.1 97.8 19.6 27.3 92.8 15.8 78.5
Hybrid UniAD [35] 93.1 95.8 93.5 98.6 39.3 47.0 90.7 30.8 81.5
RD++ [31] 85.7 90.5 92.7 96.8 19.8 27.7 95.1 16.1 74.7
DesTSeg [44] 91.4 92.4 96.2 78.9 05.1 08.8 66.9 04.6 66.6
bracket_white Aug. DRAEM [37] 69.4 70.9 74.0 66.1 00.1 00.2 23.6 00.1 47.0
SimpleNet [27] 88.4 91.4 83.0 97.8 02.2 05.6 86.4 02.9 67.8
RealNet [45] 72.0 82.3 78.4 81.4 30.1 39.6 40.1 24.7 62.7
Emb. CFA [21] 89.2 91.7 81.5 94.9 00.8 02.9 72.4 01.4 65.1
PatchCore [29] 93.3 94.8 88.9 98.8 10.0 19.3 93.9 10.7 73.9
CFLOW-AD [14] 75.3 76.4 73.0 98.6 02.1 04.6 93.8 02.4 62.3
PyramidalFlow [22] 76.3 78.5 76.9 99.0 06.3 14.2 90.1 07.7 64.8
Rec. RD [10] 85.0 86.9 80.6 98.9 01.7 05.0 94.4 02.5 67.1
DiAD [16] 81.0 82.5 75.4 94.2 00.2 00.6 80.0 00.3 59.1
ViTAD [40] 79.4 85.4 77.2 95.4 00.8 02.6 86.6 01.3 63.5
InvAD [42] 95.8 96.9 93.1 99.3 10.7 20.9 94.7 11.7 75.8
InvAD-lite [42] 93.6 95.1 87.3 99.3 06.8 10.7 94.9 05.6 72.5
MambaAD [15] 93.6 95.1 89.7 99.3 08.9 18.9 95.0 10.5 74.2
Hybrid UniAD [35] 82.4 80.5 79.3 95.9 00.8 02.8 82.0 01.4 63.1
RD++ [31] 83.4 86.4 76.2 98.7 01.6 04.3 93.4 02.2 65.8
DesTSeg [44] 96.3 93.5 94.9 91.6 00.9 02.2 52.6 01.1 65.9
connector Aug. DRAEM [37] 63.3 47.7 53.1 30.3 00.5 01.5 12.2 00.8 32.9
SimpleNet [27] 99.3 98.5 96.6 98.9 56.0 55.2 96.5 38.2 87.4
RealNet [45] 93.3 86.8 85.7 70.7 47.9 42.9 48.0 27.3 70.5
Emb. CFA [21] 95.7 86.4 93.3 97.1 31.9 39.9 90.8 25.0 78.4
PatchCore [29] 100. 100. 100. 99.5 71.2 63.3 98.4 46.3 91.5
CFLOW-AD [14] 96.7 92.9 89.7 97.4 33.1 33.2 91.5 19.9 78.5
PyramidalFlow [22] 58.3 38.3 54.5 91.3 04.4 08.5 73.7 04.4 47.4
Rec. RD [10] 99.8 99.5 96.6 99.1 58.3 54.5 97.2 37.5 88.0
DiAD [16] 88.3 81.5 74.3 96.2 11.6 22.0 87.7 12.4 66.0
ViTAD [40] 96.0 91.7 86.7 99.3 39.8 46.9 97.6 30.6 81.2
InvAD [42] 98.3 95.7 96.6 99.2 62.3 57.6 97.3 40.5 88.0
InvAD-lite [42] 96.7 83.4 96.6 99.4 61.5 57.4 97.9 40.2 85.6
MambaAD [15] 98.8 97.9 92.9 99.2 55.4 53.4 97.5 36.4 86.5
Hybrid UniAD [35] 82.9 65.0 72.7 97.6 14.0 24.0 92.1 13.7 65.2
RD++ [31] 100. 100. 100. 99.3 66.3 59.8 97.7 42.7 90.4
DesTSeg [44] 90.7 77.3 87.5 89.1 31.4 38.3 81.2 23.7 72.6
Table A71: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
metal_plate Aug. DRAEM [37] 71.7 89.7 84.5 57.5 30.0 21.2 25.7 11.8 57.8
SimpleNet [27] 100. 100. 100. 98.2 88.9 81.3 88.1 68.5 94.6
RealNet [45] 100. 100. 100. 90.2 84.6 80.4 80.7 67.2 92.0
Emb. CFA [21] 100. 100. 100. 98.8 92.6 85.3 90.8 74.4 95.9
PatchCore [29] 100. 100. 100. 99.0 93.5 86.1 94.1 75.7 96.6
CFLOW-AD [14] 100. 100. 100. 98.6 89.7 83.7 92.9 72.0 95.6
PyramidalFlow [22] 78.3 92.4 87.5 68.0 49.7 40.2 29.1 25.1 66.4
Rec. RD [10] 100. 100. 100. 99.1 92.6 87.4 96.6 77.6 97.0
DiAD [16] 98.3 99.5 98.6 87.2 54.3 55.1 45.0 38.0 76.9
ViTAD [40] 100. 100. 100. 99.3 95.5 88.2 94.8 78.9 97.2
InvAD [42] 100. 100. 100. 99.0 91.6 86.6 95.7 76.3 96.6
InvAD-lite [42] 100. 100. 100. 98.6 88.9 84.5 94.5 73.2 95.8
MambaAD [15] 99.8 99.9 99.3 98.5 86.2 84.3 94.7 72.8 95.3
Hybrid UniAD [35] 58.8 79.6 85.0 93.3 50.4 62.8 80.8 45.8 73.1
RD++ [31] 99.9 100. 99.3 99.0 91.1 87.0 95.5 77.0 96.4
DesTSeg [44] 100. 100. 100. 96.9 93.8 87.2 91.4 77.2 96.2
tubes Aug. DRAEM [37] 57.4 73.4 81.7 40.9 01.0 03.5 15.6 01.8 43.0
SimpleNet [27] 84.6 94.1 84.6 97.9 42.4 46.1 92.6 29.9 78.8
RealNet [45] 91.3 95.9 90.6 90.3 53.0 55.2 81.2 38.1 81.3
Emb. CFA [21] 85.6 94.4 85.9 97.8 38.1 43.7 91.7 27.9 78.2
PatchCore [29] 90.4 96.6 92.4 98.8 64.5 63.8 95.5 46.9 86.9
CFLOW-AD [14] 68.0 85.0 81.2 96.9 35.7 39.6 88.2 24.7 71.6
PyramidalFlow [22] 76.8 87.4 83.2 92.5 08.4 18.0 76.8 09.9 65.7
Rec. RD [10] 90.4 96.2 88.5 98.9 64.7 61.2 95.7 44.1 85.9
DiAD [16] 58.7 77.3 81.2 85.6 02.8 05.9 50.3 03.0 51.7
ViTAD [40] 84.3 93.6 84.5 98.5 52.5 55.0 93.6 38.0 81.2
InvAD [42] 96.6 98.7 95.5 99.3 73.1 69.0 97.4 52.7 90.8
InvAD-lite [42] 91.6 96.8 90.1 98.4 62.6 63.0 94.3 45.9 86.2
MambaAD [15] 62.9 78.6 81.9 96.5 26.7 34.5 87.4 20.8 67.9
Hybrid UniAD [35] 67.3 81.6 83.3 92.1 08.3 14.7 72.7 07.9 62.2
RD++ [31] 92.1 96.9 90.0 99.2 73.6 69.7 96.6 53.5 88.9
DesTSeg [44] 84.0 92.3 87.2 96.5 46.3 46.7 89.9 30.4 78.8
Avg Aug. DRAEM [37] 60.2 67.5 75.1 50.4 05.4 04.7 21.8 02.6 44.1
SimpleNet [27] 90.6 94.1 89.7 97.1 33.6 35.7 90.0 25.6 77.8
RealNet [45] 86.0 90.0 87.3 74.7 39.2 39.7 52.3 28.0 69.6
Emb. CFA [21] 92.0 93.5 90.2 95.7 29.0 33.0 83.2 23.8 76.1
PatchCore [29] 94.6 96.9 93.4 98.9 46.1 47.6 95.7 35.0 83.5
CFLOW-AD [14] 80.4 85.1 84.7 97.3 28.4 30.1 90.9 21.5 72.5
PyramidalFlow [22] 72.5 76.2 77.6 89.0 12.4 15.8 68.5 09.1 60.9
Rec. RD [10] 90.3 92.8 90.5 98.3 39.6 40.6 95.2 30.2 79.8
DiAD [16] 85.8 89.2 86.5 91.4 15.3 19.2 66.1 12.0 64.8
ViTAD [40] 87.4 90.8 87.0 97.8 34.6 37.8 92.9 28.0 77.1
InvAD [42] 96.0 97.1 94.5 98.5 44.1 46.4 95.3 34.4 83.5
InvAD-lite [42] 92.8 93.2 91.4 98.3 41.4 43.9 94.6 32.1 81.0
MambaAD [15] 89.2 93.1 90.3 97.7 33.5 38.6 92.8 27.2 78.3
Hybrid UniAD [35] 80.1 83.2 85.1 95.4 19.0 25.6 83.8 16.8 69.4
RD++ [31] 90.6 93.3 90.0 98.4 42.8 43.0 95.3 32.7 80.6
DesTSeg [44] 92.6 91.8 92.8 90.8 30.6 32.9 78.3 24.1 75.3
Appendix T Detailed Quantitative Results on MAD_Real Dataset Under 300 epochs
Table A72: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Bear Aug. DRAEM [37] 62.5 91.6 90.6 47.4 00.3 01.8 25.3 00.9 50.1
SimpleNet [27] 17.5 69.8 90.6 74.6 01.2 04.9 56.3 02.5 46.8
Emb. CFA [21] 13.3 68.8 90.6 99.0 03.9 09.6 96.2 05.1 54.9
PatchCore [29] 84.2 96.8 91.7 99.6 17.6 29.5 98.6 17.3 76.1
CFLOW-AD [14] 77.5 94.6 92.3 99.4 06.3 12.2 97.7 06.5 71.0
PyramidalFlow [22] 07.5 67.7 90.6 98.5 02.8 05.3 94.9 02.7 52.8
Rec. RD [10] 86.7 96.4 96.0 97.7 06.9 10.5 92.9 05.5 72.5
DiAD [16] 85.0 97.0 90.6 97.5 01.8 05.4 92.0 02.8 67.0
ViTAD [40] 42.5 79.5 90.6 96.3 03.1 04.4 88.0 02.2 59.4
InvAD [42] 80.8 96.0 92.0 99.6 15.9 25.0 98.5 14.3 74.7
InvAD-lite [42] 47.5 82.3 90.6 97.8 04.2 05.7 93.3 02.9 61.9
MambaAD [15] 58.3 87.2 90.6 99.2 11.2 18.8 97.1 10.4 67.6
Hybrid UniAD [35] 78.3 95.5 92.3 99.0 05.3 12.8 96.4 06.8 71.0
RD++ [31] 74.2 93.9 92.3 99.0 08.1 15.0 96.6 08.1 70.7
DesTSeg [44] 22.5 75.9 90.6 91.4 00.5 01.2 71.4 00.6 52.1
Bird Aug. DRAEM [37] 23.6 74.8 89.8 58.7 00.7 01.4 20.0 00.7 41.5
SimpleNet [27] 60.0 86.9 89.8 58.3 00.6 01.4 20.8 00.7 49.6
Emb. CFA [21] 54.5 83.0 89.8 86.7 04.7 10.9 57.1 05.8 57.8
PatchCore [29] 71.8 92.6 93.6 96.4 32.2 39.7 84.2 24.8 74.6
CFLOW-AD [14] 38.2 80.2 89.8 89.5 02.9 06.3 64.2 03.2 55.1
PyramidalFlow [22] 15.5 68.6 89.8 90.7 02.7 06.4 66.5 03.3 49.8
Rec. RD [10] 23.6 72.4 89.8 87.0 06.8 16.3 56.7 08.9 51.8
DiAD [16] 90.9 98.3 95.2 94.5 19.5 28.0 77.3 16.3 72.0
ViTAD [40] 62.7 87.5 91.7 89.9 08.3 16.9 61.9 09.2 62.4
InvAD [42] 89.1 97.6 91.7 93.4 25.0 37.3 75.4 22.9 75.3
InvAD-lite [42] 84.5 96.4 91.7 96.2 23.1 29.3 83.5 17.2 74.4
MambaAD [15] 77.3 93.1 95.7 95.0 11.3 21.4 78.7 12.0 70.1
Hybrid UniAD [35] 41.8 78.2 89.8 86.9 02.6 06.3 65.7 03.2 55.2
RD++ [31] 50.9 86.6 89.8 90.0 08.2 17.8 64.4 09.8 60.4
DesTSeg [44] 72.7 93.3 91.7 86.2 10.5 15.2 59.3 08.2 64.3
Elephant Aug. DRAEM [37] 62.2 87.9 87.8 47.6 00.3 00.7 13.5 00.4 47.4
SimpleNet [27] 05.6 61.1 87.8 64.1 00.8 02.8 24.7 01.4 37.3
Emb. CFA [21] 33.3 70.9 87.8 87.5 06.0 11.0 59.6 05.8 52.5
PatchCore [29] 27.8 69.5 87.8 93.7 11.1 18.3 78.8 10.1 56.1
CFLOW-AD [14] 30.0 74.7 87.8 90.9 07.0 11.4 67.3 06.1 54.2
PyramidalFlow [22] 60.0 86.0 87.8 92.5 02.7 05.4 75.1 02.8 60.9
Rec. RD [10] 48.9 84.7 87.8 88.4 02.1 05.2 63.2 02.7 56.8
DiAD [16] 20.0 64.3 90.0 91.8 02.0 04.4 73.1 02.2 49.4
ViTAD [40] 43.3 79.9 87.8 95.5 09.8 15.9 83.6 08.7 60.8
InvAD [42] 35.6 72.2 87.8 92.9 05.0 11.3 76.6 06.0 55.8
InvAD-lite [42] 35.6 78.4 87.8 91.6 04.9 11.3 73.7 06.0 56.3
MambaAD [15] 24.4 67.4 87.8 89.2 02.3 05.6 70.6 02.9 50.9
Hybrid UniAD [35] 13.3 63.1 87.8 78.4 01.6 05.6 38.8 02.9 42.9
RD++ [31] 40.0 77.1 87.8 87.5 02.1 05.0 63.9 02.6 54.0
DesTSeg [44] 12.2 64.6 87.8 68.7 00.8 02.3 36.5 01.2 41.0
Parrot Aug. DRAEM [37] 30.4 76.4 92.0 49.3 00.6 01.3 13.5 00.6 41.2
SimpleNet [27] 24.3 73.9 90.2 58.3 00.8 01.8 35.1 00.9 43.4
Emb. CFA [21] 05.2 66.2 90.2 42.0 00.5 01.3 16.9 00.7 34.5
PatchCore [29] 05.2 66.2 90.2 81.7 02.1 05.1 64.0 02.6 46.0
CFLOW-AD [14] 15.7 67.9 90.2 66.7 01.0 02.5 41.7 01.2 43.0
PyramidalFlow [22] 09.6 69.4 90.2 67.7 01.0 02.2 32.4 01.1 41.1
Rec. RD [10] 05.2 65.5 90.2 64.5 00.9 02.1 37.0 01.1 39.9
DiAD [16] 00.0 64.3 90.2 56.3 00.7 01.7 37.9 00.9 35.9
ViTAD [40] 10.4 67.8 90.2 70.3 01.1 02.6 40.7 01.3 42.4
InvAD [42] 05.2 65.7 90.2 68.7 01.1 02.4 44.6 01.2 41.5
InvAD-lite [42] 03.5 65.1 90.2 71.4 01.1 02.5 42.2 01.3 41.1
MambaAD [15] 04.3 65.4 90.2 72.5 01.1 02.5 46.1 01.3 41.9
Hybrid UniAD [35] 10.4 67.2 90.2 61.5 00.8 01.9 38.0 00.9 40.7
RD++ [31] 02.6 64.8 90.2 64.2 00.9 02.0 35.7 01.0 39.1
DesTSeg [44] 07.8 67.0 90.2 66.4 01.1 02.5 30.5 01.3 40.1
Table A73: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Pig Aug. DRAEM [37] 43.5 72.6 89.5 61.9 00.6 00.8 26.9 00.4 45.5
SimpleNet [27] 14.1 63.4 87.2 76.6 01.7 06.1 37.0 03.2 42.6
Emb. CFA [21] 56.5 78.2 89.5 93.0 05.0 11.3 72.6 06.0 60.1
PatchCore [29] 56.5 78.8 89.5 98.5 19.9 28.7 89.7 16.7 67.1
CFLOW-AD [14] 57.6 78.4 87.2 96.8 07.1 12.7 86.1 06.8 62.5
PyramidalFlow [22] 50.6 80.7 89.5 96.4 06.8 13.4 83.4 07.2 61.8
Rec. RD [10] 60.0 80.1 87.2 98.0 20.0 29.6 90.2 17.4 67.6
DiAD [16] 57.6 79.8 87.2 96.4 09.0 18.9 81.6 10.4 61.5
ViTAD [40] 44.7 75.3 87.2 97.6 09.2 19.4 88.0 10.7 61.3
InvAD [42] 64.7 81.2 88.9 98.0 11.2 24.2 88.1 13.8 66.8
InvAD-lite [42] 57.6 78.1 88.2 98.6 17.5 31.3 93.2 18.5 67.4
MambaAD [15] 42.4 74.5 87.2 97.3 08.5 18.6 87.3 10.2 60.5
Hybrid UniAD [35] 48.2 76.9 89.5 89.5 02.1 04.9 61.6 02.5 55.5
RD++ [31] 56.5 79.2 89.5 98.0 13.8 23.6 91.2 13.3 65.9
DesTSeg [44] 45.9 75.2 87.2 84.8 01.4 03.3 50.5 01.7 52.2
Puppy Aug. DRAEM [37] 57.0 87.0 88.9 31.2 00.2 00.7 13.7 00.3 44.5
SimpleNet [27] 52.0 82.2 88.9 90.8 01.7 04.4 65.1 02.3 57.4
Emb. CFA [21] 82.0 94.8 93.0 98.5 06.7 14.0 92.0 07.5 71.4
PatchCore [29] 100. 100. 100. 99.2 13.4 25.1 92.1 14.3 78.7
CFLOW-AD [14] 70.0 91.3 93.0 99.2 31.8 34.5 93.1 20.9 74.7
PyramidalFlow [22] 68.0 91.2 88.9 95.9 11.7 20.7 77.6 11.5 67.1
Rec. RD [10] 82.0 95.6 88.9 95.7 05.5 08.5 83.8 04.4 68.6
DiAD [16] 91.0 97.7 95.2 95.8 03.4 09.1 80.8 04.8 67.6
ViTAD [40] 65.0 90.2 88.9 90.5 01.0 02.6 69.6 01.3 61.1
InvAD [42] 100. 100. 100. 99.4 28.6 31.3 93.7 18.5 81.6
InvAD-lite [42] 97.0 99.3 97.4 95.9 03.4 09.2 82.8 04.8 72.9
MambaAD [15] 92.0 97.8 95.2 98.7 26.1 36.1 94.2 22.0 79.4
Hybrid UniAD [35] 78.0 94.7 88.9 98.7 12.0 20.2 92.8 11.2 71.6
RD++ [31] 65.0 90.3 90.9 97.3 04.0 07.5 90.5 03.9 65.9
DesTSeg [44] 54.0 82.7 88.9 93.1 01.5 03.9 66.6 02.0 58.2
Scorpion Aug. DRAEM [37] 46.1 83.9 90.2 41.4 01.2 02.9 10.8 01.5 43.7
SimpleNet [27] 47.0 84.1 90.2 83.5 09.0 15.8 50.3 08.6 56.7
Emb. CFA [21] 27.8 80.3 90.2 61.0 05.4 13.1 26.3 07.0 46.3
PatchCore [29] 13.0 70.7 90.2 88.2 08.8 17.0 61.6 09.3 50.9
CFLOW-AD [14] 54.8 87.2 90.2 77.9 04.2 07.5 34.0 03.9 54.1
PyramidalFlow [22] 47.0 84.3 90.2 65.5 02.6 05.8 20.3 03.0 48.7
Rec. RD [10] 55.7 87.9 90.2 66.3 02.1 04.9 21.8 02.5 50.9
DiAD [16] 11.3 71.8 90.2 71.9 03.1 06.8 30.7 03.5 40.8
ViTAD [40] 29.6 79.5 90.2 62.7 01.9 04.2 16.8 02.1 43.9
InvAD [42] 31.3 80.4 90.2 74.4 03.1 06.5 30.8 03.4 48.0
InvAD-lite [42] 32.2 79.2 90.2 74.6 03.0 06.0 28.1 03.1 47.6
MambaAD [15] 42.6 80.8 90.2 75.0 02.9 06.3 27.1 03.2 49.5
Hybrid UniAD [35] 27.0 73.9 90.2 78.6 04.2 09.1 41.2 04.8 48.5
RD++ [31] 25.2 75.8 90.2 70.7 02.5 05.8 24.6 03.0 44.8
DesTSeg [44] 27.8 77.9 90.2 68.1 03.2 08.0 25.3 04.2 45.7
Turtle Aug. DRAEM [37] 58.1 85.3 91.3 24.4 00.3 01.5 09.2 00.8 43.5
SimpleNet [27] 63.8 87.9 89.4 94.6 07.0 15.1 81.8 08.1 65.0
Emb. CFA [21] 100. 100. 100. 99.3 41.5 42.7 96.8 27.2 85.0
PatchCore [29] 100. 100. 100. 99.5 43.1 48.4 97.9 31.9 86.1
CFLOW-AD [14] 99.0 99.8 97.7 99.1 28.4 38.0 96.5 23.5 82.2
PyramidalFlow [22] 93.3 98.5 93.3 98.4 23.9 33.3 95.3 20.0 78.9
Rec. RD [10] 95.2 99.0 95.2 98.5 23.2 30.4 94.9 18.0 79.1
DiAD [16] 74.3 94.2 89.4 93.2 07.5 13.5 83.5 07.2 65.1
ViTAD [40] 83.8 96.1 93.3 95.6 06.2 13.2 85.6 07.0 70.6
InvAD [42] 100. 100. 100. 99.7 51.9 50.0 98.4 33.3 87.5
InvAD-lite [42] 99.0 99.8 97.7 98.0 15.5 21.6 94.5 12.1 78.1
MambaAD [15] 100. 100. 100. 99.3 32.8 39.2 97.2 24.4 83.6
Hybrid UniAD [35] 98.1 99.6 97.6 98.2 22.8 31.5 91.0 18.7 79.7
RD++ [31] 100. 100. 100. 99.2 25.1 35.8 97.0 21.8 82.1
DesTSeg [44] 54.3 86.7 89.4 95.2 09.8 16.1 77.0 08.8 63.2
Table A74: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
Unicorn Aug. DRAEM [37] 36.2 79.8 89.4 43.0 00.2 00.5 09.9 00.2 40.9
SimpleNet [27] 19.0 68.8 89.4 81.8 00.9 02.2 49.5 01.1 46.3
Emb. CFA [21] 31.4 73.8 89.4 67.8 00.8 03.6 33.2 01.9 45.6
PatchCore [29] 43.8 74.1 93.3 97.3 06.8 14.1 90.3 07.6 61.3
CFLOW-AD [14] 63.8 90.9 91.3 90.1 01.5 03.3 65.7 01.7 61.1
PyramidalFlow [22] 65.7 84.4 89.4 82.2 00.9 02.2 51.8 01.1 57.1
Rec. RD [10] 21.0 71.0 89.4 73.5 00.5 01.3 33.8 00.6 43.9
DiAD [16] 62.9 91.4 89.4 86.0 01.0 02.4 60.4 01.2 56.2
ViTAD [40] 20.0 72.2 89.4 73.9 00.5 01.1 31.7 00.6 43.7
InvAD [42] 31.4 72.8 89.4 89.9 01.6 04.3 67.3 02.2 52.7
InvAD-lite [42] 32.4 72.3 89.4 84.4 01.4 04.3 57.0 02.2 50.7
MambaAD [15] 18.1 67.6 89.4 81.7 00.8 02.1 50.0 01.1 46.0
Hybrid UniAD [35] 16.2 66.6 89.4 89.3 01.4 03.7 66.1 01.9 48.8
RD++ [31] 29.5 75.7 89.4 81.4 00.7 01.5 43.7 00.8 48.3
DesTSeg [44] 21.0 69.4 89.4 76.4 00.5 01.3 31.0 00.6 43.6
Whale Aug. DRAEM [37] 43.1 85.6 92.8 20.2 00.1 00.4 06.4 00.2 40.3
SimpleNet [27] 58.1 91.5 92.8 89.1 02.5 08.2 68.5 04.3 61.4
Emb. CFA [21] 80.6 96.2 93.5 98.6 29.3 32.2 90.5 19.2 76.4
PatchCore [29] 95.0 99.1 98.5 99.7 46.8 47.4 97.8 31.0 85.2
CFLOW-AD [14] 67.5 94.4 92.8 99.1 17.5 24.7 95.1 14.1 72.0
PyramidalFlow [22] 89.4 98.2 95.5 98.0 20.2 25.6 90.6 14.7 76.5
Rec. RD [10] 81.2 96.7 92.8 98.9 19.2 23.5 94.9 13.3 74.7
DiAD [16] 86.9 98.0 94.1 94.6 02.9 05.4 77.2 02.8 65.6
ViTAD [40] 42.5 85.4 92.8 96.3 03.6 08.5 83.6 04.5 60.8
InvAD [42] 96.9 99.5 97.0 99.8 45.3 50.1 98.7 33.4 85.6
InvAD-lite [42] 83.8 96.6 95.5 99.5 39.4 41.9 98.2 26.5 80.9
MambaAD [15] 87.5 97.6 95.5 99.7 33.8 41.7 98.6 26.3 81.0
Hybrid UniAD [35] 42.5 87.3 92.8 98.1 11.1 19.6 91.1 10.9 64.6
RD++ [31] 79.4 95.9 92.8 99.2 27.9 29.8 96.1 17.5 76.3
DesTSeg [44] 42.5 86.4 92.8 95.1 02.6 06.0 79.3 03.1 59.8
Avg Aug. DRAEM [37] 46.3 82.5 90.2 42.5 00.5 01.2 14.9 00.6 43.9
SimpleNet [27] 36.1 77.0 89.6 77.2 02.6 06.3 48.9 03.3 50.7
Emb. CFA [21] 48.5 81.2 91.4 83.3 10.4 15.0 64.1 08.6 58.5
PatchCore [29] 59.7 84.8 93.5 95.4 20.2 27.3 85.5 16.6 68.2
CFLOW-AD [14] 57.4 85.9 91.2 90.9 10.8 15.3 74.1 08.8 63.0
PyramidalFlow [22] 50.6 82.9 90.5 88.6 07.5 12.0 68.8 06.7 59.4
Rec. RD [10] 56.0 84.9 90.7 86.9 08.7 13.2 66.9 07.4 60.6
DiAD [16] 58.0 85.7 91.1 87.8 05.1 09.6 69.4 05.2 58.1
ViTAD [40] 44.5 81.3 90.2 86.9 04.5 08.9 65.0 04.8 56.7
InvAD [42] 63.5 86.5 92.7 91.6 18.9 24.2 77.2 14.9 66.9
InvAD-lite [42] 57.3 84.8 91.9 90.8 11.4 16.3 74.7 09.5 63.1
MambaAD [15] 54.7 83.1 92.2 90.8 13.1 19.2 74.7 11.4 63.1
Hybrid UniAD [35] 45.4 80.3 90.8 87.8 06.4 11.6 68.3 06.4 57.8
RD++ [31] 52.3 84.0 91.3 88.7 09.3 14.4 70.4 08.2 60.8
DesTSeg [44] 36.1 77.9 89.8 82.5 03.2 06.0 52.7 03.2 52.0
Appendix U Detailed Quantitative Results on MAD_Sim Dataset Under 300 epochs
Table A75: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
01Gorilla Aug. DRAEM [37] 54.4 91.6 95.1 45.4 00.7 02.0 19.4 01.0 48.6
SimpleNet [27] 53.0 91.5 94.9 90.9 02.7 05.7 65.5 02.9 60.5
Emb. CFA [21] 55.9 93.6 94.9 79.0 07.2 15.4 55.8 08.3 60.4
CFLOW-AD [14] 54.4 92.8 94.9 90.8 04.8 05.1 65.4 02.6 61.1
PyramidalFlow [22] 46.5 90.4 94.9 91.5 03.7 06.0 75.1 03.1 60.7
Rec. RD [10] 62.7 93.3 95.1 93.1 04.5 09.0 76.4 04.7 64.7
DiAD [16] 64.4 94.6 94.9 91.6 03.6 08.3 64.6 04.3 60.3
ViTAD [40] 66.1 93.9 95.1 93.7 05.9 10.6 76.3 05.6 65.8
InvAD [42] 67.4 95.7 94.9 95.2 17.1 24.1 82.0 13.7 70.3
InvAD-lite [42] 70.6 95.5 94.9 94.1 08.0 15.1 78.3 08.2 67.9
MambaAD [15] 65.2 93.8 94.9 92.8 05.0 10.7 75.6 05.7 65.3
Hybrid UniAD [35] 52.6 92.3 94.9 92.7 06.3 11.0 79.7 05.8 63.7
RD++ [31] 53.6 91.1 94.9 92.7 04.0 08.3 74.6 04.3 62.4
DesTSeg [44] 46.5 90.3 94.9 58.2 05.0 09.0 34.4 04.7 51.9
02Unicorn Aug. DRAEM [37] 41.9 90.6 95.2 47.3 00.4 00.8 14.1 00.4 45.8
SimpleNet [27] 55.0 91.6 95.2 83.2 01.3 02.6 61.3 01.3 58.9
Emb. CFA [21] 63.8 94.8 95.5 67.8 01.4 03.9 32.1 02.0 55.5
CFLOW-AD [14] 59.0 93.5 95.2 84.4 02.3 03.4 55.4 01.7 59.5
PyramidalFlow [22] 37.0 87.7 95.5 80.3 01.1 02.3 52.2 01.2 53.7
Rec. RD [10] 61.4 94.1 95.2 85.3 02.0 04.1 70.7 02.1 62.0
DiAD [16] 79.3 97.2 95.4 80.0 01.6 04.5 48.0 02.3 58.0
ViTAD [40] 63.9 94.7 95.2 86.5 02.4 05.8 67.8 03.0 62.6
InvAD [42] 79.7 97.6 95.4 87.3 03.8 09.8 74.6 05.2 67.4
InvAD-lite [42] 71.3 96.0 95.2 86.6 02.9 06.4 73.2 03.3 64.9
MambaAD [15] 65.7 95.2 95.4 84.3 02.1 04.7 68.7 02.4 62.7
Hybrid UniAD [35] 71.4 96.2 95.4 87.1 02.5 06.3 64.5 03.3 63.9
RD++ [31] 60.6 93.8 95.2 85.1 02.3 04.6 69.2 02.3 61.8
DesTSeg [44] 72.0 96.1 95.5 55.9 06.4 10.5 40.0 05.6 58.0
03Mallard Aug. DRAEM [37] 36.6 88.0 95.4 46.6 01.3 01.5 15.3 00.8 44.8
SimpleNet [27] 55.2 91.6 95.4 85.5 02.6 05.6 58.6 02.9 59.4
Emb. CFA [21] 31.3 87.1 95.4 63.8 02.9 10.0 36.7 05.3 49.8
CFLOW-AD [14] 40.1 90.5 95.4 84.8 03.8 05.1 53.1 02.6 56.0
PyramidalFlow [22] 44.2 89.9 95.4 83.9 02.5 05.0 56.3 02.6 56.7
Rec. RD [10] 44.4 89.1 95.4 83.5 03.4 07.9 56.7 04.1 57.1
DiAD [16] 73.9 96.0 95.4 83.0 02.4 04.6 48.7 02.4 57.7
ViTAD [40] 46.6 91.4 95.4 85.2 05.3 11.2 60.4 05.9 59.2
InvAD [42] 46.5 91.3 95.4 85.9 07.8 13.4 61.5 07.2 59.9
InvAD-lite [42] 40.9 90.2 95.4 85.2 06.0 11.1 60.2 05.9 58.1
MambaAD [15] 55.2 93.6 95.4 84.2 04.9 10.5 57.9 05.6 60.4
Hybrid UniAD [35] 33.9 87.0 95.4 86.1 03.8 08.9 61.8 04.7 56.1
RD++ [31] 37.1 87.9 95.4 84.0 03.4 07.8 57.1 04.1 55.8
DesTSeg [44] 33.2 88.1 95.4 58.7 03.4 08.0 39.0 04.1 49.8
04Turtle Aug. DRAEM [37] 45.6 87.8 94.9 47.2 00.6 01.3 12.0 00.7 45.7
SimpleNet [27] 48.7 90.2 94.9 85.7 02.3 04.7 60.9 02.4 58.2
Emb. CFA [21] 40.8 90.8 94.9 71.1 02.4 03.6 38.6 01.8 52.2
CFLOW-AD [14] 63.1 94.7 94.9 86.3 04.1 06.3 57.1 03.2 61.3
PyramidalFlow [22] 16.8 81.8 94.9 76.6 01.2 03.1 39.7 01.6 47.3
Rec. RD [10] 64.1 94.1 94.9 91.1 09.3 16.8 72.6 09.2 65.9
DiAD [16] 85.5 98.3 96.3 91.8 07.1 14.3 63.1 07.7 65.2
ViTAD [40] 59.0 93.5 94.9 91.0 07.9 15.4 70.5 08.3 64.3
InvAD [42] 83.0 97.7 95.1 93.4 15.5 25.9 78.8 14.9 72.7
InvAD-lite [42] 76.8 96.9 95.6 92.3 12.4 21.3 75.4 11.9 70.1
MambaAD [15] 66.6 95.2 94.9 91.0 09.3 17.7 70.8 09.7 66.4
Hybrid UniAD [35] 55.7 92.0 94.9 88.9 04.3 07.6 70.9 04.0 61.9
RD++ [31] 60.1 94.2 94.9 90.2 07.0 13.2 69.8 07.1 64.1
DesTSeg [44] 34.1 88.5 94.9 55.0 01.6 03.6 38.8 01.8 48.6
Table A76: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
05Whale Aug. DRAEM [37] 50.5 90.3 94.4 47.6 00.6 01.5 20.6 00.8 48.0
SimpleNet [27] 50.4 89.2 94.4 85.1 01.8 03.7 62.4 01.9 58.1
Emb. CFA [21] 51.6 91.5 94.4 71.3 02.2 04.1 41.0 02.1 54.4
CFLOW-AD [14] 54.0 91.1 94.4 85.3 02.8 04.4 63.2 02.2 59.4
PyramidalFlow [22] 27.7 82.7 94.4 78.9 01.2 02.8 52.5 01.4 51.1
Rec. RD [10] 52.4 91.0 94.4 86.8 03.3 07.1 74.0 03.7 61.0
DiAD [16] 51.0 90.8 94.4 81.1 02.6 07.5 51.9 03.9 54.2
ViTAD [40] 58.5 92.7 94.4 88.3 04.1 08.3 76.5 04.3 63.1
InvAD [42] 74.0 96.3 94.4 89.9 08.8 16.9 81.1 09.2 68.7
InvAD-lite [42] 68.1 95.2 94.4 89.3 06.2 11.2 78.5 05.9 66.1
MambaAD [15] 60.4 93.5 94.4 87.3 04.4 08.9 74.1 04.6 63.2
Hybrid UniAD [35] 79.4 97.1 94.4 90.8 05.2 11.6 80.1 06.2 68.6
RD++ [31] 54.1 91.8 94.4 86.5 03.5 07.0 73.0 03.6 61.3
DesTSeg [44] 57.1 92.1 94.4 61.4 03.2 08.3 55.5 04.3 56.6
06Bird Aug. DRAEM [37] 44.1 87.6 94.5 48.6 01.0 02.3 27.2 01.1 47.6
SimpleNet [27] 52.9 91.4 94.4 91.8 05.1 09.4 75.9 04.9 62.6
Emb. CFA [21] 65.1 92.8 94.5 73.3 04.0 09.7 47.2 05.1 58.8
CFLOW-AD [14] 63.8 91.8 94.4 90.4 03.8 07.1 70.8 03.7 63.2
PyramidalFlow [22] 77.2 96.3 94.4 92.7 05.9 11.1 77.8 05.9 68.1
Rec. RD [10] 60.4 92.0 94.4 90.6 04.9 08.4 75.2 04.4 63.5
DiAD [16] 77.3 97.0 94.4 93.6 06.5 12.9 77.3 06.9 65.6
ViTAD [40] 65.7 93.5 94.5 91.2 04.9 09.8 76.1 05.2 65.0
InvAD [42] 82.8 97.4 94.9 91.9 06.3 12.8 78.0 06.8 69.5
InvAD-lite [42] 72.0 95.5 94.5 91.6 05.8 10.0 77.0 05.3 66.7
MambaAD [15] 66.9 94.7 94.7 91.1 04.7 08.6 76.1 04.5 65.3
Hybrid UniAD [35] 67.0 93.2 94.4 90.8 04.8 09.1 74.8 04.8 64.9
RD++ [31] 60.0 91.4 94.5 90.8 04.8 08.3 75.4 04.3 63.4
DesTSeg [44] 68.2 93.5 94.5 60.3 04.1 08.8 40.5 04.6 56.9
07Owl Aug. DRAEM [37] 56.3 87.8 92.8 45.4 00.5 01.2 16.4 00.6 47.4
SimpleNet [27] 55.0 87.5 92.8 87.7 01.9 04.1 64.6 02.1 59.0
Emb. CFA [21] 60.8 91.7 93.0 77.1 03.2 07.0 47.4 03.6 57.8
CFLOW-AD [14] 66.9 92.8 92.8 89.9 03.8 05.0 66.2 02.6 62.7
PyramidalFlow [22] 48.4 85.1 93.0 90.0 03.5 07.8 71.1 04.1 59.3
Rec. RD [10] 62.2 91.8 93.0 91.9 05.3 09.4 77.7 04.9 64.2
DiAD [16] 62.0 91.4 92.8 94.4 06.1 13.1 72.8 07.0 61.8
ViTAD [40] 65.6 91.9 92.8 92.7 05.2 10.9 77.8 05.7 65.0
InvAD [42] 72.6 94.3 93.0 94.6 12.5 22.7 83.6 12.8 70.0
InvAD-lite [42] 66.6 93.1 93.2 93.3 07.6 13.7 80.1 07.3 66.5
MambaAD [15] 67.5 93.5 92.8 91.5 04.8 09.3 76.0 04.9 65.0
Hybrid UniAD [35] 63.7 92.1 93.0 92.5 04.9 09.8 77.6 05.2 64.6
RD++ [31] 66.1 92.5 92.8 92.2 05.2 09.0 78.4 04.7 65.0
DesTSeg [44] 64.8 92.4 93.2 54.3 11.1 10.0 39.1 05.3 56.0
08Sabertooth Aug. DRAEM [37] 58.0 88.3 93.0 47.1 00.5 01.1 18.6 00.6 48.3
SimpleNet [27] 55.1 88.3 93.2 87.3 01.9 03.9 65.6 02.0 59.3
Emb. CFA [21] 53.8 90.6 93.0 72.0 02.2 03.8 36.8 01.9 53.9
CFLOW-AD [14] 60.8 90.8 93.2 86.6 01.7 03.4 58.3 01.7 59.6
PyramidalFlow [22] 65.2 93.2 93.0 82.5 01.2 02.6 58.6 01.3 60.0
Rec. RD [10] 66.0 93.2 93.4 88.0 02.0 03.6 67.2 01.9 62.2
DiAD [16] 69.2 94.0 94.5 86.1 01.9 03.4 51.3 01.7 57.2
ViTAD [40] 61.6 91.3 93.2 89.0 02.2 04.0 68.0 02.0 61.4
InvAD [42] 78.0 95.8 93.4 91.3 03.9 06.2 74.7 03.2 66.5
InvAD-lite [42] 73.7 94.6 93.4 89.4 03.1 04.6 69.4 02.4 64.4
MambaAD [15] 64.4 92.6 93.4 87.1 02.2 03.5 63.9 01.8 61.3
Hybrid UniAD [35] 62.5 91.0 93.2 89.5 02.5 04.2 74.5 02.1 62.5
RD++ [31] 63.2 91.9 93.0 88.1 02.3 03.7 66.6 01.9 61.4
DesTSeg [44] 70.0 93.6 93.2 52.0 02.3 04.0 33.9 02.0 54.3
Table A77: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
09Swan Aug. DRAEM [37] 46.6 86.6 93.1 47.7 00.6 01.4 19.9 00.7 46.4
SimpleNet [27] 49.8 85.9 93.1 86.1 01.9 03.9 62.4 02.0 57.4
Emb. CFA [21] 54.9 88.2 93.1 72.0 02.0 06.7 37.8 03.5 54.2
CFLOW-AD [14] 47.0 86.9 93.1 85.7 02.2 03.9 60.1 02.0 56.8
PyramidalFlow [22] 43.7 85.4 93.1 80.2 01.3 02.9 53.9 01.5 54.3
Rec. RD [10] 55.8 89.3 93.1 87.7 02.4 04.6 71.0 02.3 60.4
DiAD [16] 78.8 95.5 94.3 87.8 09.2 16.7 57.8 09.1 62.9
ViTAD [40] 54.1 88.1 93.2 90.1 03.1 06.2 73.7 03.2 60.9
InvAD [42] 65.8 92.6 93.2 91.8 06.2 11.6 80.0 06.2 65.6
InvAD-lite [42] 60.1 90.5 93.1 90.1 04.1 08.3 76.8 04.3 63.0
MambaAD [15] 56.3 89.5 93.1 87.3 02.5 04.4 69.4 02.2 60.3
Hybrid UniAD [35] 58.4 89.7 93.1 89.5 03.2 06.8 74.7 03.5 62.0
RD++ [31] 53.9 89.4 93.2 87.6 02.4 04.4 69.8 02.3 59.9
DesTSeg [44] 53.3 89.8 93.1 51.5 02.2 03.2 36.0 01.6 51.0
10Sheep Aug. DRAEM [37] 36.8 81.6 92.7 47.2 00.6 01.3 20.8 00.7 43.9
SimpleNet [27] 44.9 85.1 92.7 91.3 02.8 05.2 76.4 02.7 59.1
Emb. CFA [21] 66.4 92.3 92.7 71.9 02.9 07.4 50.2 03.8 58.5
CFLOW-AD [14] 63.0 90.1 92.9 91.9 03.9 06.8 79.2 03.5 63.7
PyramidalFlow [22] 58.6 90.6 92.7 90.6 02.2 05.0 75.9 02.6 62.0
Rec. RD [10] 62.6 91.8 92.7 92.5 05.1 08.8 84.1 04.6 65.0
DiAD [16] 87.4 97.3 94.3 93.6 04.5 08.8 76.6 04.6 66.1
ViTAD [40] 65.0 91.1 92.7 93.4 04.4 08.7 84.7 04.5 65.4
InvAD [42] 84.5 96.8 93.5 94.3 08.1 13.9 87.0 07.5 71.2
InvAD-lite [42] 80.5 96.2 93.9 93.7 07.0 10.7 85.7 05.7 69.7
MambaAD [15] 66.0 93.0 92.7 92.6 05.2 09.4 83.9 05.0 65.8
Hybrid UniAD [35] 65.9 92.4 92.9 93.6 05.6 09.9 84.2 05.2 66.0
RD++ [31] 70.4 94.2 92.7 92.6 06.5 10.6 84.1 05.6 67.1
DesTSeg [44] 71.1 94.1 92.9 54.8 05.9 08.2 43.4 04.3 57.1
11Pig Aug. DRAEM [37] 52.4 89.4 94.0 47.2 00.5 01.1 22.2 00.5 48.2
SimpleNet [27] 50.2 88.7 94.0 89.8 02.0 04.2 73.3 02.1 60.0
Emb. CFA [21] 57.1 91.4 94.0 84.7 09.7 17.7 56.1 09.7 61.4
CFLOW-AD [14] 52.6 90.1 94.0 92.4 06.5 05.6 74.5 02.9 61.8
PyramidalFlow [22] 50.9 89.4 94.0 91.3 02.6 05.7 71.5 02.9 60.4
Rec. RD [10] 59.3 92.0 94.0 94.1 06.0 11.7 83.2 06.2 65.3
DiAD [16] 53.1 92.4 94.0 93.9 09.2 16.3 71.5 08.9 61.5
ViTAD [40] 58.8 91.6 94.5 95.7 08.1 15.6 83.0 08.5 66.1
InvAD [42] 72.2 94.9 94.6 96.8 17.6 25.1 87.6 14.3 72.0
InvAD-lite [42] 66.3 93.6 94.2 96.1 13.6 21.2 85.8 11.8 69.4
MambaAD [15] 63.5 92.9 94.0 95.1 08.8 15.5 83.8 08.4 67.1
Hybrid UniAD [35] 52.6 88.9 94.0 94.5 05.1 09.6 81.2 05.0 63.0
RD++ [31] 60.8 92.4 94.0 94.5 06.5 12.7 83.6 06.8 65.9
DesTSeg [44] 65.5 93.0 94.0 62.3 08.1 13.1 48.0 07.0 58.5
12Zalika Aug. DRAEM [37] 50.4 87.0 92.9 40.9 00.9 02.2 17.0 01.1 46.0
SimpleNet [27] 50.4 86.8 92.9 84.1 02.5 05.0 57.9 02.5 57.0
Emb. CFA [21] 56.2 90.5 92.9 70.8 04.3 08.7 47.5 04.5 56.3
CFLOW-AD [14] 53.0 88.4 92.9 86.0 04.4 07.0 55.4 03.6 58.2
PyramidalFlow [22] 42.1 85.2 92.9 83.4 02.3 04.7 63.7 02.4 56.0
Rec. RD [10] 56.8 91.0 92.9 87.3 05.0 08.3 69.2 04.3 61.3
DiAD [16] 34.5 82.3 92.9 85.4 03.5 06.1 43.2 03.2 49.7
ViTAD [40] 56.4 91.0 92.9 87.4 04.8 08.6 68.2 04.5 61.2
InvAD [42] 61.2 92.2 92.9 90.2 09.4 15.5 75.4 08.4 64.9
InvAD-lite [42] 58.6 91.2 92.9 88.9 07.1 11.7 72.6 06.2 63.0
MambaAD [15] 58.6 91.5 92.9 86.8 04.7 08.0 68.0 04.2 61.4
Hybrid UniAD [35] 50.8 88.9 92.9 86.8 03.8 06.7 67.9 03.4 59.4
RD++ [31] 52.7 90.3 92.9 87.3 05.5 08.8 69.1 04.6 60.7
DesTSeg [44] 50.7 88.9 92.9 52.8 04.7 06.0 38.8 03.1 51.5
Table A78: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
13Pheonix Aug. DRAEM [37] 48.9 89.6 94.3 51.2 00.6 01.6 27.8 00.8 49.0
SimpleNet [27] 49.2 89.3 94.3 82.2 01.3 02.8 52.3 01.4 56.1
Emb. CFA [21] 56.9 91.0 94.4 66.2 01.6 04.6 50.0 02.4 55.7
CFLOW-AD [14] 52.5 90.2 94.4 82.3 01.5 02.9 57.6 01.5 57.6
PyramidalFlow [22] 44.6 86.1 94.3 78.0 01.1 02.5 61.6 01.2 55.4
Rec. RD [10] 52.1 90.1 94.4 83.5 02.2 03.3 69.3 01.7 59.2
DiAD [16] 39.7 87.8 94.3 83.6 01.5 03.0 49.9 01.5 51.4
ViTAD [40] 53.3 90.5 94.3 84.6 01.7 03.4 68.0 01.8 59.4
InvAD [42] 64.2 92.8 94.6 87.5 03.7 08.2 77.4 04.3 64.0
InvAD-lite [42] 56.3 91.9 94.5 86.0 03.1 04.6 74.2 02.3 61.4
MambaAD [15] 57.5 92.1 94.4 83.4 02.7 03.4 69.8 01.7 60.6
Hybrid UniAD [35] 55.3 90.6 94.3 85.2 01.8 03.6 68.1 01.8 59.9
RD++ [31] 56.6 91.9 94.4 84.1 01.9 03.7 70.3 01.9 60.5
DesTSeg [44] 56.9 92.0 94.3 52.9 02.3 04.8 35.7 02.5 52.5
14Elephant Aug. DRAEM [37] 44.6 87.2 93.9 50.3 01.0 01.9 19.3 00.9 46.7
SimpleNet [27] 48.0 87.8 93.9 71.8 01.8 04.0 49.8 02.0 54.2
Emb. CFA [21] 57.4 91.5 93.9 54.1 01.8 04.1 39.0 02.1 52.8
CFLOW-AD [14] 55.8 91.3 93.9 68.1 02.2 03.5 49.6 01.8 55.6
PyramidalFlow [22] 47.0 88.0 93.9 63.8 01.4 03.2 51.1 01.6 53.1
Rec. RD [10] 58.9 91.7 93.9 69.8 02.8 06.1 64.1 03.2 58.6
DiAD [16] 65.8 94.4 93.9 74.3 02.0 03.9 45.9 02.0 54.3
ViTAD [40] 57.0 91.9 93.9 70.7 03.1 07.2 64.2 03.7 58.6
InvAD [42] 66.7 93.9 94.0 72.5 04.6 11.5 68.5 06.1 62.1
InvAD-lite [42] 65.9 93.9 93.9 70.8 03.6 08.2 65.8 04.3 60.8
MambaAD [15] 63.3 93.4 93.9 69.6 02.8 06.3 62.8 03.2 59.5
Hybrid UniAD [35] 58.2 91.5 93.9 69.5 02.2 05.1 62.2 02.6 58.0
RD++ [31] 58.7 92.3 94.0 70.2 02.9 06.5 63.3 03.3 58.7
DesTSeg [44] 63.7 92.9 94.0 54.5 04.9 08.0 50.1 04.2 56.5
15Parrot Aug. DRAEM [37] 52.1 86.0 91.8 48.6 00.8 01.8 18.7 00.9 47.1
SimpleNet [27] 49.4 82.6 91.8 83.3 02.4 04.8 62.1 02.5 56.4
Emb. CFA [21] 57.1 89.2 91.8 65.9 02.3 04.8 48.2 02.4 54.8
CFLOW-AD [14] 53.0 87.0 92.0 83.4 03.0 05.3 60.2 02.7 57.7
PyramidalFlow [22] 59.2 86.3 92.2 80.0 02.2 04.7 59.7 02.4 57.9
Rec. RD [10] 50.7 86.7 92.0 84.0 03.2 06.9 67.1 03.6 58.4
DiAD [16] 53.1 88.9 91.8 83.4 02.5 05.8 63.9 03.0 55.6
ViTAD [40] 49.8 86.9 91.8 83.4 03.1 06.7 66.4 03.5 58.0
InvAD [42] 53.9 88.6 92.0 86.6 06.6 14.5 74.0 07.8 61.8
InvAD-lite [42] 50.0 87.3 91.8 85.4 04.6 09.6 69.8 05.0 59.4
MambaAD [15] 49.9 86.8 91.8 83.9 03.2 06.3 65.1 03.3 57.9
Hybrid UniAD [35] 46.9 86.0 91.8 85.5 03.5 06.5 69.9 03.4 58.1
RD++ [31] 53.9 88.1 92.0 84.1 03.1 06.7 66.5 03.5 59.0
DesTSeg [44] 54.4 88.4 91.8 52.7 04.9 06.0 25.8 03.1 50.3
16Cat Aug. DRAEM [37] 53.8 90.8 94.3 51.5 00.5 01.6 27.3 00.8 49.9
SimpleNet [27] 54.2 90.1 94.3 93.2 02.5 05.0 79.3 02.6 62.3
Emb. CFA [21] 53.5 90.3 94.4 71.8 01.4 04.9 42.5 02.5 54.8
CFLOW-AD [14] 54.5 90.6 94.3 93.1 02.4 04.8 80.3 02.5 62.5
PyramidalFlow [22] 46.7 87.7 94.4 91.7 01.7 04.1 74.2 02.1 59.6
Rec. RD [10] 49.3 88.8 94.3 93.6 03.0 06.1 84.3 03.1 62.1
DiAD [16] 63.5 93.3 94.3 93.9 02.9 05.6 82.0 02.9 62.2
ViTAD [40] 57.2 91.7 94.3 93.9 03.2 06.4 84.6 03.3 64.0
InvAD [42] 57.8 92.4 94.3 94.4 05.1 10.7 86.2 05.6 65.3
InvAD-lite [42] 56.8 91.4 94.3 94.3 03.9 08.0 85.6 04.2 64.4
MambaAD [15] 53.7 90.3 94.3 93.9 03.2 06.5 84.5 03.4 63.2
Hybrid UniAD [35] 52.8 89.8 94.3 94.1 03.0 06.4 83.3 03.3 62.8
RD++ [31] 54.2 91.3 94.3 93.5 03.3 06.1 84.6 03.1 63.4
DesTSeg [44] 54.8 90.7 94.3 57.7 01.6 04.5 45.0 02.3 53.6
Table A79: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
17Scorpion Aug. DRAEM [37] 48.2 86.3 92.9 50.6 00.5 01.2 21.5 00.6 47.1
SimpleNet [27] 52.6 88.2 92.9 83.8 01.5 03.0 55.5 01.5 56.9
Emb. CFA [21] 63.8 92.1 93.1 76.6 03.2 06.8 46.9 03.5 58.2
CFLOW-AD [14] 62.4 91.7 92.9 89.5 03.3 04.8 67.4 02.5 61.8
PyramidalFlow [22] 55.3 86.5 92.9 86.4 02.0 05.1 58.1 02.6 58.1
Rec. RD [10] 62.9 91.2 92.9 91.4 04.7 07.4 81.9 03.8 64.3
DiAD [16] 70.9 94.4 93.9 87.4 02.1 04.1 56.6 02.1 58.5
ViTAD [40] 59.1 89.8 93.1 91.1 03.3 07.2 77.2 03.8 62.7
InvAD [42] 80.9 96.6 93.4 93.7 07.2 10.2 85.4 05.4 69.7
InvAD-lite [42] 74.5 95.0 93.0 93.4 06.1 08.5 84.6 04.4 67.8
MambaAD [15] 65.1 92.4 92.9 92.7 04.8 07.3 82.3 03.8 65.1
Hybrid UniAD [35] 70.1 93.4 92.9 92.2 05.7 09.1 81.1 04.8 66.2
RD++ [31] 64.2 92.4 92.9 91.6 04.8 06.9 81.6 03.6 64.7
DesTSeg [44] 70.0 93.8 93.1 62.6 04.3 08.0 48.8 04.2 58.3
18Obesobeso Aug. DRAEM [37] 50.7 89.7 94.1 51.4 00.9 01.9 24.9 01.0 49.0
SimpleNet [27] 49.3 89.8 94.1 90.9 03.5 06.5 72.7 03.3 60.6
Emb. CFA [21] 61.4 92.8 94.1 78.8 12.4 18.1 60.4 10.0 62.6
CFLOW-AD [14] 62.0 91.9 94.1 92.6 06.0 07.8 79.6 04.0 64.6
PyramidalFlow [22] 62.9 93.1 94.1 93.5 05.5 10.6 79.8 05.6 65.4
Rec. RD [10] 63.0 93.4 94.1 93.4 10.7 19.2 87.0 10.6 68.0
DiAD [16] 92.2 99.0 95.2 95.8 13.2 21.9 84.5 12.3 71.7
ViTAD [40] 66.9 94.4 94.1 94.2 16.5 24.1 88.1 13.7 70.4
InvAD [42] 83.7 97.5 95.1 94.9 16.6 25.8 90.4 14.8 74.5
InvAD-lite [42] 73.4 95.7 94.2 93.7 12.2 20.2 87.9 11.2 70.6
MambaAD [15] 69.6 95.1 94.2 93.1 11.2 21.6 86.5 12.1 69.7
Hybrid UniAD [35] 71.7 94.8 94.2 93.1 09.5 15.8 84.8 08.6 68.8
RD++ [31] 67.0 93.2 94.2 93.1 07.9 15.6 86.0 08.4 67.7
DesTSeg [44] 71.3 95.4 94.1 63.7 18.5 23.7 53.0 13.4 63.3
19Bear Aug. DRAEM [37] 52.3 88.6 93.3 48.7 00.6 01.3 19.1 00.6 47.7
SimpleNet [27] 50.9 87.4 93.3 85.0 02.0 03.9 65.0 02.0 58.1
Emb. CFA [21] 56.1 90.7 93.5 78.6 04.2 08.8 50.1 04.6 57.8
CFLOW-AD [14] 61.1 91.7 93.3 88.9 04.0 06.3 64.4 03.3 61.5
PyramidalFlow [22] 53.6 89.5 93.3 87.3 03.1 05.7 61.6 02.9 59.1
Rec. RD [10] 61.5 92.1 93.6 90.8 05.6 10.0 76.1 05.3 64.0
DiAD [16] 64.3 93.7 93.3 90.1 05.1 13.0 65.0 07.0 60.6
ViTAD [40] 60.1 90.8 93.5 91.1 05.3 11.1 75.0 05.9 63.5
InvAD [42] 71.0 94.5 93.8 92.5 11.7 18.3 81.2 10.1 68.7
InvAD-lite [42] 67.9 93.4 93.4 92.2 11.2 17.0 78.8 09.3 67.3
MambaAD [15] 63.5 92.3 93.5 91.2 07.2 12.7 75.9 06.8 64.9
Hybrid UniAD [35] 61.7 91.9 93.6 90.6 05.0 09.4 79.4 04.9 64.2
RD++ [31] 64.9 92.9 93.5 90.9 06.5 11.2 76.6 05.9 65.0
DesTSeg [44] 68.3 93.5 93.5 60.4 06.8 11.7 52.0 06.2 58.9
20Puppy Aug. DRAEM [37] 51.7 86.3 92.7 47.5 00.5 01.0 17.2 00.5 46.7
SimpleNet [27] 55.6 88.5 92.7 83.1 01.3 02.6 55.6 01.3 57.3
Emb. CFA [21] 50.2 88.6 92.9 68.5 01.9 05.0 42.2 02.6 53.3
CFLOW-AD [14] 59.0 90.3 92.7 84.9 02.0 04.4 56.5 02.2 58.8
PyramidalFlow [22] 52.4 87.6 92.9 84.8 01.9 04.9 59.9 02.5 57.8
Rec. RD [10] 50.9 89.0 92.7 87.0 03.5 05.6 66.1 02.9 59.0
DiAD [16] 46.2 88.9 92.7 83.4 01.4 02.8 52.7 01.4 52.6
ViTAD [40] 55.4 88.2 93.0 87.5 02.4 04.8 66.3 02.4 59.6
InvAD [42] 61.9 92.0 92.9 89.7 05.7 09.7 73.8 05.1 63.5
InvAD-lite [42] 55.2 90.0 92.7 87.9 03.7 05.9 69.1 03.0 60.5
MambaAD [15] 48.8 87.3 92.7 86.6 02.3 04.9 64.6 02.5 57.9
Hybrid UniAD [35] 56.7 90.2 92.7 87.2 04.5 10.8 70.9 05.7 61.6
RD++ [31] 59.3 91.4 93.0 86.9 03.5 06.9 65.4 03.6 61.0
DesTSeg [44] 57.5 91.4 92.9 55.2 11.9 11.7 37.9 06.2 54.9
Avg Aug. DRAEM [37] 48.8 88.1 93.8 47.9 00.7 01.5 20.0 00.8 47.2
SimpleNet [27] 51.5 88.6 93.8 86.1 02.3 04.5 63.9 02.3 58.6
Emb. CFA [21] 55.7 91.1 93.8 71.8 03.7 07.7 45.3 04.1 56.2
CFLOW-AD [14] 56.9 90.9 93.8 86.9 03.4 05.1 63.7 02.6 60.2
PyramidalFlow [22] 49.0 88.1 93.8 84.4 02.4 05.0 62.7 02.6 57.8
Rec. RD [10] 57.9 91.3 93.8 88.3 04.4 08.2 73.7 04.3 62.3
DiAD [16] 65.6 93.4 94.1 87.7 04.4 08.8 61.4 04.7 59.4
ViTAD [40] 59.0 91.4 93.8 89.0 04.8 09.3 73.6 04.9 62.8
InvAD [42] 70.4 94.5 94.0 90.7 08.9 15.3 79.1 08.4 67.4
InvAD-lite [42] 65.3 93.4 93.9 89.7 06.6 11.4 76.4 06.1 65.1
MambaAD [15] 61.4 92.4 93.8 88.3 04.8 09.0 73.0 04.8 63.2
Hybrid UniAD [35] 59.4 91.5 93.8 89.0 04.4 08.4 74.6 04.4 62.8
RD++ [31] 58.6 91.7 93.8 88.3 04.4 08.1 73.2 04.2 62.4
DesTSeg [44] 59.2 91.9 93.8 56.8 05.7 08.6 41.8 04.5 54.9
Appendix V Detailed Quantitative Results on Uni-Medical Dataset Under 300 epochs
Table A80: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
Method Image-level Pixel-level mAU-PRO mIoU-max mAD
[…]
brain Aug. DRAEM [37] 53.0 84.2 90.6 28.7 02.0 06.2 05.7 03.2 43.3
SimpleNet [27] 77.7 94.1 90.7 93.7 36.4 39.3 71.1 24.4 73.8
Emb. CFA [21] 67.0 90.8 90.6 45.2 06.2 07.4 07.5 03.8 49.7
CFLOW-AD [14] 85.8 96.2 92.1 94.5 55.8 17.8 73.6 09.8 75.9
PyramidalFlow [22] 66.6 89.6 90.6 87.0 12.4 18.9 54.6 10.4 62.7
Rec. RD [10] 85.0 95.4 92.1 96.4 45.6 49.0 82.3 32.5 79.6
DiAD [16] 93.6 98.2 95.2 95.3 37.3 42.3 80.4 26.9 77.5
ViTAD [40] 89.0 97.3 92.9 97.6 60.2 58.0 83.6 40.9 84.0
InvAD [42] 89.5 97.2 92.5 97.6 56.5 57.4 85.6 40.3 83.7
InvAD-lite [42] 84.4 95.5 91.7 96.3 44.8 48.1 81.9 31.7 79.2
MambaAD [15] 91.9 98.0 93.5 97.6 56.6 58.0 86.4 40.8 84.6
Hybrid UniAD [35] 90.2 97.6 92.9 97.6 58.8 57.2 83.3 40.0 83.9
RD++ [31] 85.8 96.0 91.6 96.9 48.3 52.2 83.8 35.3 80.7
DesTSeg [44] 80.4 93.8 91.2 82.1 49.3 51.0 55.4 34.2 74.0
liver Aug. DRAEM [37] 56.2 47.3 64.0 26.2 00.5 02.4 09.8 01.2 32.8
SimpleNet [27] 52.5 44.7 62.3 96.8 08.5 13.7 86.9 07.4 52.3
Emb. CFA [21] 37.1 35.0 61.2 14.0 00.2 00.8 00.7 00.4 24.2
CFLOW-AD [14] 63.8 54.0 63.5 96.0 02.9 05.7 87.8 02.9 54.3
PyramidalFlow [22] 65.7 56.1 65.7 82.0 01.7 02.3 41.4 01.2 47.2
Rec. RD [10] 55.4 45.4 64.1 96.5 05.3 10.0 90.6 05.3 52.8
DiAD [16] 65.8 60.1 63.9 96.9 07.6 13.9 91.5 07.5 57.1
ViTAD [40] 63.5 54.5 65.4 97.8 11.1 18.9 92.5 10.4 58.1
InvAD [42] 62.5 52.1 65.2 96.8 06.8 12.6 92.3 06.7 56.0
InvAD-lite [42] 56.4 46.1 63.5 96.1 04.8 09.3 89.7 04.9 52.7
MambaAD [15] 61.1 50.9 63.8 96.8 08.3 15.5 91.7 08.4 55.8
Hybrid UniAD [35] 64.7 51.5 64.9 97.1 07.9 14.1 92.8 07.6 56.7
RD++ [31] 55.5 44.6 64.1 96.4 05.3 09.9 90.9 05.2 52.7
DesTSeg [44] 74.8 72.4 66.0 87.6 38.6 41.8 84.5 26.4 67.1
retinal Aug. DRAEM [37] 66.1 57.9 60.4 42.5 03.3 09.6 11.9 05.1 39.1
SimpleNet [27] 87.3 87.1 77.1 94.2 62.8 58.0 79.6 40.9 78.7
Emb. CFA [21] 64.6 62.6 59.5 55.2 08.1 11.5 18.6 06.1 42.8
CFLOW-AD [14] 87.4 85.5 76.3 95.6 60.0 55.5 83.7 38.4 78.4
PyramidalFlow [22] 47.7 42.8 59.5 77.7 12.2 23.4 44.7 13.3 44.8
Rec. RD [10] 88.8 86.7 77.7 96.4 65.7 60.6 86.6 43.4 80.9
DiAD [16] 89.2 87.5 79.0 95.7 62.7 58.3 84.3 41.2 79.5
ViTAD [40] 92.0 90.0 82.0 95.5 69.3 63.8 83.3 46.8 83.0
InvAD [42] 92.1 90.6 82.1 97.2 72.9 66.5 89.0 49.8 84.8
InvAD-lite [42] 92.1 90.2 82.8 96.1 66.5 61.8 85.7 44.7 82.9
MambaAD [15] 93.8 87.9 86.6 95.9 64.2 64.7 85.5 47.8 83.5
Hybrid UniAD [35] 86.2 80.7 74.8 94.9 50.4 52.7 81.2 35.8 75.2
RD++ [31] 90.3 87.7 79.9 96.7 69.5 63.8 87.5 46.9 82.7
DesTSeg [44] 90.9 88.9 81.9 80.0 64.0 58.7 58.5 41.6 76.3
Avg Aug. DRAEM [37] 58.5 63.2 71.7 32.5 02.0 06.1 09.1 03.2 38.4
SimpleNet [27] 72.5 75.3 76.7 94.9 35.9 37.0 79.2 24.2 68.3
Emb. CFA [21] 56.3 62.8 70.4 38.1 04.8 06.6 08.9 03.4 38.9
CFLOW-AD [14] 79.0 78.5 77.3 95.4 39.6 26.3 81.7 17.0 69.5
PyramidalFlow [22] 60.0 62.8 71.9 82.2 08.8 14.9 46.9 08.3 51.5
Rec. RD [10] 76.4 75.8 77.9 96.4 38.9 39.8 86.5 27.0 71.1
DiAD [16] 82.9 81.9 79.4 96.0 35.9 38.2 85.4 25.2 71.4
ViTAD [40] 81.5 80.6 80.1 97.0 46.8 46.9 86.5 32.7 75.0
InvAD [42] 81.3 80.0 79.9 97.2 45.4 45.5 89.0 32.3 74.8
InvAD-lite [42] 77.6 77.2 79.3 96.2 38.7 39.7 85.7 27.1 71.6
MambaAD [15] 82.3 78.9 81.3 96.7 43.0 46.1 87.9 32.4 74.6
Hybrid UniAD [35] 80.4 76.6 77.5 96.5 39.0 41.3 85.8 27.8 71.9
RD++ [31] 77.2 76.1 78.5 96.7 41.0 41.9 87.4 29.1 72.0
DesTSeg [44] 82.0 85.0 79.7 83.2 50.6 50.5 66.1 34.1 72.4

### v2

@@
 Broader Impacts. The open-sourcing ADer can accelerate the development of new VAD technology for the open-source community and become a valuable resource for practitioners in the field.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The abstract and introduction have clearly included the motivations, important assumptions, and contributions made in the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The authors have discussed the limitations of the work at the end of the Conclusion section.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The authors have clearly provided the full set of assumptions and complete proofs in the Method section.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The authors have presented all the experimental details in the Implementation Details section. Also, the codes are submitted as Supplementary Material.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The authors have submitted all the codes in Supplementary Material which provided sufficient instructions to faithfully reproduce the main experimental results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The authors have presented all the training and test details in the Implementation Details section. Also, the codes are submitted as Supplementary Material.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [N/A]
-  34.
-Justification: Due to the substantial workload and the high stability of most methods, with errors being nearly zero, this study does not use error bars, following the approach of previous RD [10].
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: The authors have included the sufficient information on the computer resources in the Implementation Details and Ablation and Analysis sections.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The authors conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: The authors have explained the broader impacts of the work at the end of the Conclusion section.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The authors have cited the original paper that produced the code package or dataset.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The authors have submitted the details of the code/model which includes details about training, license, limitations, etc.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-Overview
-The supplementary material presents the following sections to strengthen the main manuscript:
-  • Sec. A shows the average quantitative results on various datasets under 100 epochs.
-  • Sec. B shows the average quantitative results on various datasets under 300 epochs.
-  • Sec. C shows the specific quantitative results for each class under 100 epochs on MVTec AD [3] dataset.
-  • Sec. D shows the specific quantitative results for each class under 100 epochs on MVTec AD 3D [6] dataset.
-  • Sec. E shows the specific quantitative results for each class under 100 epochs on MVTec LOCO-AD [5] dataset.
-  • Sec. F shows the specific quantitative results for each class under 100 epochs on VisA [47] dataset.
-  • Sec. G shows the specific quantitative results for each class under 100 epochs on BTAD [28] dataset.
-  • Sec. H shows the specific quantitative results for each class under 100 epochs on MPDD [20] dataset.
-  • Sec. I shows the specific quantitative results for each class under 100 epochs on MAD_Real [46] dataset.
-  • Sec. J shows the specific quantitative results for each class under 100 epochs on MAD_Sim [46] dataset.
-  • Sec. K shows the specific quantitative results for each class under 100 epochs on Uni-Medical [40] dataset.
-  • Sec. L shows the specific quantitative results for each class under 100 epochs on Real-IAD [32] dataset.
-  • Sec. M shows the specific quantitative results for each class under 100 epochs on COCO-AD [42] dataset.
-  • Sec. N shows the specific quantitative results for each class under 300 epochs on MVTec AD [3] dataset.
-  • Sec. O shows the specific quantitative results for each class under 300 epochs on MVTec AD 3D [6] dataset.
-  • Sec. P shows the specific quantitative results for each class under 300 epochs on MVTec LOCO-AD [5] dataset.
-  • Sec. Q shows the specific quantitative results for each class under 300 epochs on VisA [47] dataset.
-  • Sec. R shows the specific quantitative results for each class under 300 epochs on BTAD [28] dataset.
-  • Sec. S shows the specific quantitative results for each class under 300 epochs on MPDD [20] dataset.
-  • Sec. T shows the specific quantitative results for each class under 300 epochs on MAD_Real [46] dataset.
-  • Sec. U shows the specific quantitative results for each class under 300 epochs on MAD_Sim [46] dataset.
-  • Sec. V shows the specific quantitative results for each class under 300 epochs on Uni-Medical [40] dataset.
 Appendix A Detailed Quantitative Results on All Dataset Under 100 epochs